### v1

# SIC-POVMs and orders of real quadratic fields

Abstract.
We consider the problem of counting and classifying symmetric informationally complete positive operator-valued measures (SICs or SIC-POVMs), that is, sets of $d^{2}$ equiangular lines in $\mathbb{C}^{d}$. For $4\leq d\leq 90$, we show the number of known equivalence classes of Weyl–Heisenberg covariant SICs in dimension $d$ equals the cardinality of the ideal class monoid of (not necessarily invertible) ideal classes in the real quadratic order of discriminant $(d+1)(d-3)$. Equivalently, this is the number of $\mathbf{GL}_{2}(\mathbb{Z})$ conjugacy classes in $\mathbf{SL}_{2}(\mathbb{Z})$ of trace $d-1$. We conjecture the equality extends to all $d\geq 4$. We prove that this conjecture implies more that one equivalence class of Weyl–Heisenberg covariant SICs for every $d>22$.
Additionally, we refine the “class field hypothesis” of Appleby, Flammia, McConnell, and Yard [10] to predict the exact class field generated by the ratios of vector entries for the equiangular lines defining a Weyl–Heisenberg covariant SIC. The class fields conjecturally associated to SICs in dimension $d$ come with a natural partial order under inclusion; we show that the natural inclusions of these fields are strict, except in a small family of cases.
Key words and phrases:
symmetric informationally complete positive operator-valued measure (SIC-POVM), complex equianglular lines, equiangular tight frame (ETF), real quadratic field, non-maximal order, binary quadratic form, ideal class group, explicit class field theory, Stark conjectures
1. Introduction
This paper concerns provable and conjectural properties of SIC-POVMs, which are ensembles of quantum measurements described by $d^{2}$ equiangular complex lines in $\mathbb{C}^{d}$. These properties are number-theoretic and relate to class numbers and class fields attached to orders of associated real quadratic number fields.
A set $S$ of complex lines through the origin in $\mathbb{C}^{d}$, written $S=\{\mathbb{C}\mathbf{v}_{1},\ldots,\mathbb{C}\mathbf{v}_{n}\}\subset\mathbb{P}^{d-1}(\mathbb{C})$, with $\mathbf{v}_{j}$ unit vectors, is equiangular if all the pairwise Hermitian inner products $\langle\mathbf{v}_{i},\mathbf{v}_{j}\rangle$ for $i\neq j$ have the same absolute value $|\langle\mathbf{v}_{i},\mathbf{v}_{j}\rangle|=\alpha\geq 0$. In 1975, Delsarte, Goethals, and Seidel [19] proved an upper bound of $n\leq d^{2}$ on the size of any set of complex equiangular lines in $\mathbb{C}^{d}$. They showed that if the bound $n=d^{2}$ is attained, then necessarily $\alpha=\frac{1}{d+1}$.
A symmetric informationally complete positive operator-valued measure or SIC-POVM (henceforth SIC) is a set of quantum measurements that is equivalent to a (maximal) set of $d^{2}$ complex equiangular lines in $\mathbb{C}^{d}$. Specifically, for a set of $d^{2}$ equiangular lines $\{\mathbb{C}\mathbf{v}_{1},\ldots,\mathbb{C}\mathbf{v}_{d^{2}}\}$ represented by unit column vectors $\mathbf{v}_{j}\in\mathbb{C}^{d}$, the associated SIC is the set of normalized rank one Hermitian projection operators ${\mathbf{\Pi}}[S]:=\left\{\frac{1}{d}\Pi_{1},\ldots,\frac{1}{d}\Pi_{d^{2}}\right\}$ where $\Pi_{j}=\mathbf{v}_{j}\mathbf{v}_{j}^{\dagger}$. (Vectors $\mathbf{v}$ are column vectors, $\mathbf{v}^{\dagger}$ is the conjugate transpose of $\mathbf{v}$, and the Hermitian inner product is $\langle\mathbf{v}_{i},\mathbf{v}_{j}\rangle=\mathbf{v}_{i}^{\dagger}\mathbf{v}_{j}$, following conventions from physics. Here, $\mathbf{v}_{i}^{\dagger}\mathbf{v}_{i}=1.$) The SIC condition is equivalent to the following $d^{4}$ identities on the traces of products of the projection operators:
$\operatorname{Tr}(\Pi_{i}\Pi_{j})$ =$=\begin{cases}1&\mbox{if }i=j,\\ \frac{1}{d+1}&\mbox{if }i\neq j,\end{cases}$ for $1\leq i,j\leq d^{2}$. (1.1)
We will use the term “SIC” interchangeably to mean either a “set of $d^{2}$ (scaled) Hermitian projection matrices ${\mathbf{\Pi}}(S)$ forming a SIC-POVM” or a “set of $d^{2}$ complex equiangular lines $S$ in $\mathbb{C}^{d}$.” When more precision is needed, we term the latter a line-SIC.
The name SIC-POVM was introduced in a 2004 paper of Renes, Blume-Kohout, Scott, and Caves [36], whose study was motivated by quantum state tomography, the problem of reconstructing a quantum state from a series of measurements performed on identical states. Scott showed that SICs are optimal measurement ensembles for tomography [37]; see also [26]. SICs have found further applications to quantum information processing [14, 42, 45, 18], quantum foundations [25], compressed sensing for radar [28], and classical phase retrieval [21].
The object of this paper is to extend the range of conjectures about the structure of SICs proposed by Appleby, Flammia, McConnell, and Yard [9]. Their conjectures relate Weyl–Heisenberg SICs in $\mathbb{C}^{d}$ (defined in Section 2) for $d\geq 4$ to class field theory over real quadratic fields. The ratios of entries of each of the $d^{2}$ vectors in such SICs lie in a number field that is an abelian extension of $K=\mathbb{Q}(\sqrt{\Delta_{d}})$ with $\Delta_{d}=(d+1)(d-3)$. Class field theory classifies the abelian extensions of a given number field $K$ in terms of ray class groups, which encode the arithmetic of ideals of the ring $\mathcal{O}_{K}$ of algebraic integers of $K$.
The number fields associated to individual SICs in dimension $d$ may vary with the SIC, and the subclass of SICs to which the most precise conjecture of [9] applies consists of those giving rise to minimal number field that occurs for a fixed $d$. Appleby, Flammia, McConnell, and Yard conjecture that this minimal field is a specific ray class field of $\mathbb{Q}(\sqrt{\Delta_{d}})$; see 1.1. The corresponding SICs comprise the minimal multiplet, as defined in Section 2. It is one of several multiplets, which are finite sets of geometric equivalence classes of SICs making up a single orbit of a Galois group action on geometric equivalence classes.
The current paper presents a refinement of these conjectures that accounts for all the remaining multiplets of Weyl–Heisenberg SICs, those having larger associated number fields. For $d\geq 4$ the refined conjecture says that each multiplet is indexed by a particular order $\mathcal{O}^{\prime}$ of a real quadratic number field $K_{d}=\mathbb{Q}(\sqrt{\Delta_{d}})$ intermediate between $\mathcal{O}=\mathbb{Z}[\varepsilon_{d}]$ with $\varepsilon_{d}=\frac{1}{2}(d-1+\sqrt{\Delta_{d}})$ and the maximal order $\mathcal{O}_{K_{d}}$. Here an order of a $K_{d}$ is a subring of $K_{d}$ whose additive group is free of rank $[K_{d}:\mathbb{Q}]=2$.
The number field associated to the $\mathcal{O}^{\prime}$-multiplet is conjectured to be a certain ray class field associated to the order $\mathcal{O}^{\prime}$. The definition of ray class fields associated to orders originates in the authors’ previous work [33]. Inclusions of orders induce reverse inclusion of the corresponding ray class fields, so the minimal multiplet corresponds to the maximal order.
1.1. SICs
Most work on SICs follows Zauner’s 1999 PhD thesis [48, 49] on quantum designs. A few sporadic examples (such as the Hoggar lines in dimension $8$ [29, 30]) were known prior to Zauner’s work. Zauner conjectured, based on his numerical investigations, that SICs exist in every dimension $d$, and moreover, that a SIC in $\mathbb{C}^{d}$ could be found as the orbit of a single vector under the Weyl–Heisenberg group [48, 49]; the latter objects are called Weyl–Heisenberg covariant SICs (or just Weyl–Heisenberg SICs) and are defined and discussed in Section 2. Weyl–Heisenberg SICs are the main topic of this paper. (An extensive treatment of the history of SICs is given by Fuchs, Hoang, and Stacy [24].)
Over the past 25 years, a large set of data has been computed about SICs in particular dimensions—sometimes approximately and sometime exactly—resulting in extensive datasets [39, 38, 22]. SICs have been rigorously constructed in a finite set of dimensions, including Weyl–Heisenberg SICs in all dimensions $1\leq d\leq 53$ and the Hoggar lines for $d=8$ (which are not Weyl–Heisenberg covariant). Beyond these exact constructions, numerical computations up to $d=90$ suggest there are only finitely many (equivalence classes of) Weyl–Heisenberg SICs in each dimension $d\neq 3$. It is known that $d=3$ has infinitely many equivalence classes of Weyl–Heisenberg SICs, and we will not consider it further.
The current numerical data is believed to account for all Weyl–Heisenberg SICs of dimension $4\leq d\leq 90$. There is a geometric action moving these SICs around by the normalizer of the Weyl–Heisenberg group inside the extended unitary group (which is a finite group whose order varies erratically with $d$, while the Weyl–Heisenberg group itself has order $d^{3}$.) In addition, since the (normalized) generating vectors of Weyl–Heisenberg SICs are (empirically) given by algebraic numbers, there is a Galois action, which sometimes moves SICs to new SICs and sometimes moves them to non-SICs.
The process of understanding this data has led to formulation of a series of interlocking hypotheses. The two most basic conjectures, which are both unsolved, are:
  1. (1)
    the existence of Weyl–Heisenberg SICs in all dimensions (conjectured by Zauner);
  2. (2)
    the finiteness of the set of (equivalence classes of) Weyl–Heisenberg SICs in every fixed dimension $d\neq 3$.
The finiteness assertion has been a “folklore” conjecture since about 2004 [2] and was observed informally up to dimension $d=50$ by Scott and Grassl (who state they are confident their list is complete for $d\leq 50$ in [39, Sec. 4]) and up to $d=90$ by Fuchs, Hoang, and Stacy [24, Sec. 6]. It was explicitly formulated as a conjecture in the first author’s PhD thesis [31, Conj. V.12]. A closely related conjecture was formulated by Waldron [46, Sec. 14.27].
The finiteness assertion would imply fiducial vectors of Weyl–Heisenberg SICs can be chosen to have all entries algebraic numbers. Further study of the algebraic numbers in existing examples has led to surprising discoveries.
1.2. Connection of SICs to algebraic number theory and class fields
Recent work has connected SICs to explicit class field theory over real quadratic number fields. Appleby, Yadsan-Appleby, and Zauner [4] discovered in 2013 that in all known examples, the algebraic number field $\mathbb{E}$ generated by the entries of the projections ${\mathbf{\Pi}}[S]$ contained the quadratic field $K_{d}=\mathbb{Q}(\sqrt{(d+1)(d-3)})$, and Galois automorphisms often mapped SICs to new SICs. Moreover, they observed that $\mathbb{E}$ was an abelian extension of $K_{d}$ in all known examples.
Such abelian Galois extensions are subfields of ray class fields over $K_{d}$. In 2016, Appleby, Flammia, McConnell, and Yard [10, 9] made explicit predictions relating the number fields generated by Weyl–Heisenberg SICs to ray class fields. Based on extensive data given in [10, Prop. 1], they formulated the following conjecture [10, Conj. 2].
Conjecture 1.1 (Appleby, Flammia, McConnell and Yard).
Let $d\geq 4$, $\Delta_{d}=(d+1)(d-3)$, and $K_{d}=\mathbb{Q}(\sqrt{\Delta_{d}})$.
  1. (1)
    At least one Weyl–Heisenberg line-SIC $S$ exists in dimension $d$ having SIC field $F^{\rm vec}\!\left(S\right)$ equal to $H_{d^{\prime}\infty_{1}\infty_{2}}$, the ray class field of $K_{d}$ having ray class modulus $d^{\prime}\infty_{1}\infty_{2}$, where $d^{\prime}=d$ if $d$ is odd, and $d^{\prime}=2d$ if $d$ is even.
  2. (2)
    Every Weyl–Heisenberg line-SIC $S$ in dimension $d$ has SIC-field $F^{\rm vec}\!\left(S\right)$ that is a finite Galois extension of $\mathbb{Q}$ containing $H_{d^{\prime}\infty_{1}\infty_{2}}$.
The technical definition of the SIC-field $F^{\rm vec}\!\left(S\right)$ appearing in 1.1 is given in Section 3.1. Ray class fields are discussed in Section 8.1.
Appleby, Flammia, McConnell, and Yard verified assertion (1) of 1.1 for
$d\in\{4,5,6,7,8,9,10,11,12,13,14,15,16,17,18,19,20,21,24,28,30,35,39,48\}.$ (1.2)
They verified assertion (2) of the conjecture for all currently known Weyl–Heisenberg SICs in these dimensions. They also verified that the fields $F^{\rm vec}\!\left(S\right)$ are all abelian over $K_{d}$ in the known cases. (The list of SICs used is believed to be complete in these dimensions.)
[…]
The origin of the discriminant $\Delta_{d}=(d+1)(d-3)$ appearing in the conjecture is not completely understood, but there are some tantalizing hints. The appearance of the factor $d+1$ dividing $\Delta_{d}$ seems related to the fact that in the inner products of distinct unit vectors in a SIC are of absolute value $\frac{1}{\sqrt{d+1}}$. The appearance of $d-3$ in the conjecture accounts for $d=3$ being special in that $\Delta_{3}=0$. The values $d=1,2$ have $\Delta_{d}<0$, so $K_{d}=\mathbb{Q}(\sqrt{\Delta_{d}})$ is an imaginary quadratic field, while for $d\geq 4$, $\Delta_{d}$ specifies a (nontrivial) real quadratic field. The occurrence of a small unit $\varepsilon_{d}=\frac{1}{2}(d-1+\sqrt{\Delta_{d}})$ (as compared to $\Delta_{d}$, which need not be a fundamental discriminant!) is certainly relevant. As noted by previously by the first author [32, p. 8], skipping repeated fields (e.g., $K_{8}=K_{4}$), the sequence $K_{4},K_{5},K_{6},K_{7},K_{9},K_{10},\ldots$ is precisely the ordering of all real quadratic fields by the size of the smallest nontrivial totally positive unit (which will necessarily be equal to $\varepsilon_{d}$).
1.3. Main results
In this paper, we refine 1.1 to conjectures that specify the number fields associated to all the Weyl–Heisenberg SICs in each dimension $d\geq 4$ and which count the number of such SICs associated to each field.
More precisely, we connect both SICs and fields to certain orders of the real quadratic number field $K_{d}=\mathbb{Q}(\sqrt{\Delta_{d}})$, with $\Delta_{d}=(d+1)(d-3)$. Write $\Delta_{d}=f^{2}\Delta_{0}$, where $\Delta_{0}$ is the fundamental discriminant $K_{d}$, and $f\geq 1$. The orders that arise are exactly those $\mathcal{O}^{\prime}$ satisfying $\mathbb{Z}[\varepsilon_{d}]\subseteq\mathcal{O}^{\prime}\subseteq\mathcal{O}_{K}$ for $\varepsilon_{d}=\frac{1}{2}(d-1+\sqrt{\Delta_{d}})$. These orders, arranged by set inclusion, form a distributive lattice, isomorphic to that of the positive divisors of $f$ under divisibility.
We organize Weyl–Heisenberg SICs into equivalence classes. For a line-SIC $S$,
  1. (1)
    the geometric equivalence class $[S]$ is the orbit of $S$ under the action of the projective extended Clifford group $\operatorname{PEC}(d)$ (see 2.9), and
  2. (2)
    the Galois multiplet $[[S]]$ is the set of geometric equivalence classes that are permuted with each other under a certain Galois action (see 4.11).
In Section 3, we define several fields attached to a line-SIC $S=\{\mathbb{C}\mathbf{v}_{k}:1\leq k\leq d^{2}\}$. These fields are finitely generated extensions of the rational numbers $\mathbb{Q}$.
  1. (1)
    The ratio SIC-field $F^{\rm vec}\!\left(S\right)$ (called the “SIC field” in [10]) is the field generated by all the ratios $\frac{v_{ki}}{v_{kj}}$ of distinct entries ($i\neq j$, $v_{kj}\neq 0$) of the vectors $\mathbf{v}_{k}$ with $\mathbf{v}_{k}\in S$.
  2. (2)
    The projection SIC-field $F^{\rm proj}\!\left(S\right)$ is the field generated by all the matrix entires of the Hermitian projection operators $\Pi_{i}=\mathbf{v}_{i}\mathbf{v}_{i}^{\dagger}$ of the $d^{2}$ unit vectors $\{{\mathbf{v}}_{i}:1\leq i\leq d^{2}\}$ defining $S$. It is shown to be the compositum of $F^{\rm vec}\!\left(S\right)$ with its complex conjugate field $\overline{F^{\rm vec}\!\left(S\right)}$ (in 3.3).
  3. (3)
    The extended projection SIC-field $F^{\rm eproj}\!\left(S\right)=F^{\rm proj}\!\left(S\right)\!(-e^{\frac{\pi i}{d}})$ is obtained by adjoining the root of unity $-e^{\frac{\pi i}{d}}$. For Weyl–Heisenberg SICs, we show $F^{\rm eproj}\!\left(S\right)$ is an invariant of the geometric equivalence class $[S]$ (in 3.8), and we show the index $[F^{\rm eproj}\!\left(S\right):F^{\rm proj}\!\left(S\right)]\in\{1,2\}$ and is $1$ if $d$ is odd (in 3.6).
  4. (4)
    For SICs defined over a finite extension of $\mathbb{Q}$, we define the Galois projection SIC-field $F^{\rm Gproj}\!\left(S\right)$ to be the normal closure of $F^{\rm eproj}\!\left(S\right)$, which is an invariant of the multiplet $[[S]]$. We show that the subgroup of $\operatorname{Gal}(F^{\rm Gproj}\!\left(S\right)/\mathbb{Q})$ consisting of elements that commute with complex conjugation sends SICs to other SICs with the same Galois projection SIC-field (in 4.8).
We have the field inclusions
$F^{\rm vec}\!\left(S\right)\subseteq F^{\rm proj}\!\left(S\right)\subseteq F^{\rm eproj}\!\left(S\right)\subseteq F^{\rm Gproj}\!\left(S\right).$ (1.3)
In 4.9, we show that 1.1 implies that
$F^{\rm vec}\!\left(S\right)=F^{\rm proj}\!\left(S\right)=F^{\rm eproj}\!\left(S\right)=F^{\rm Gproj}\!\left(S\right),$ (1.4)
whence $F^{\rm vec}\!\left(S\right)$ is a finite Galois extension of $\mathbb{Q}$.
The refinements of 1.1 formulated in this paper posit that Weil–Heisenberg SICs in dimension $d\neq 3$ occur in multiplets indexed by the intermediate orders $\mathcal{O}^{\prime}$ of $K_{d}$ the lie between $\mathcal{O}=\mathbb{Z}[\varepsilon_{d}]$ and $\mathcal{O}_{K_{d}}$. Each of these orders is in turn indexed by its conductor $f^{\prime}$, a positive integer such that $\operatorname{disc}(\mathcal{O}^{\prime})=(f^{\prime})^{2}\Delta_{0}$ where $\Delta_{0}=\operatorname{disc}(\mathcal{O}_{K_{d}})$. The conductors appearing are the positive divisors of the conductor $f$ of $\mathcal{O}$.
Conjecture 1.2 (Order-to-Multiplet Conjecture).
Fix a positive integer $d\neq 3$, let $\Delta=\Delta_{d}=(d+1)(d-3)$, and write $\Delta_{d}=f^{2}\Delta_{0}$ for $\Delta_{0}=\operatorname{disc}(\mathcal{O}_{K_{d}})$ and some positive integer $f$. Then there is a bijection $\mathcal{M}$ between conductors of intermediate orders and Galois multiplets,
[…]
having the following properties.
  • (1)
    The number of geometric equivalence classes $[S]$ in the multiplet $\mathcal{M}(f^{\prime})=[[S]]$ is $\left|\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})\right|$, the class number of the invertible class group of the order $\mathcal{O}^{\prime}$ of discriminant $(f^{\prime})^{2}\Delta_{0}$.
  • (2)
    If $f_{1}\mid f_{2}\mid f$, $[S_{1}]\in\mathcal{M}(f_{1})$, and $[S_{2}]\in\mathcal{M}(f_{2})$, then $F^{\rm vec}\!\left(S_{1}\right)\subseteq F^{\rm vec}\!\left(S_{2}\right)$.
A mildly strengthened version of this conjecture is stated as 7.2. The field inclusion property (2) puts a strong restriction on the map $\mathcal{M}$. Modulo the next conjecture, it determines $\mathcal{M}$ entirely for a fraction $\frac{47}{48}$ of all dimensions $d$ (in the sense of 9.10).
The following conjecture is an extension of 1.1 using the notion of ray class fields of an order $\mathcal{O}^{\prime}$ introduced in [33] and reviewed in Section 8.1.
Conjecture 1.3 (Ray Class Fields of Orders Conjecture).
Let $d\geq 4$. Then for any Weyl–Heisenberg line-SIC $S$ in dimension $d$ with associated order $\mathcal{O}^{\prime}$ having discriminant $(f^{\prime})^{2}\Delta_{0}$ (that is, with $\mathcal{M}(f^{\prime})=[[S]]$),
$F^{\rm vec}\!\left(S\right)=H_{d^{\prime}\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}^{\prime}}$ (1.6)
where $H_{d^{\prime}\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}^{\prime}}$ is the ray class field having level datum $(\mathcal{O}^{\prime};d^{\prime}\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\})$, $d^{\prime}=2d$ if $d$ is even, and $d^{\prime}=d$ if $d$ is odd.
The special case $f^{\prime}=1$ corresponds to 1.1, where the order $\mathcal{O}^{\prime}$ is the maximal order. The field inclusions asserted in 1.1(2) follow from 1.2. Appleby, Flammia, McConnell, and Yard [9] call the associated Galois multiplet $[[S]]$ the minimal multiplet, as it has the smallest associated field $F^{\rm vec}\!\left(S\right)$ under inclusion. 8.4 gives a further refinement of 1.3; these conjectures refine and revise [31, Conj. V.17].
The paper describes numerical checks of predictions of these conjectures against numerical data. The data for $4\leq d\leq 90$ exhibits patterns consistent with grouping SICs into multiplets indexed by orders. We formulate several narrower conjectures that make numerical predictions that can be checked against existing data and give tables of comparisons.
We also derive some consequences of the refined conjectures. In Section 6, we show they imply that the number of distinct geometric equivalence classes $[S]$ will be larger than $1$ for all $d\geq 23$ and will grow like $d^{1+o(1)}$ as $d\to\infty$. (See 6.3 and 6.4.)
1.4. Notation
  • Boldface roman upper case letters represent continuous groups, e.g., $\mathbf{EU}(d)$.
  • Boldface roman lower case letters represent column vectors in $\mathbb{C}^{d}$. For a complex vector, $\mathbf{w}^{\dagger}:=\overline{\mathbf{w}}^{T}$, and similarly for matrices.
  • Upper case roman letters represent $d\times d$ complex matrices (e.g., $X$, $Z$); they also appear in names of finite matrix groups (e.g., $\operatorname{WH}(d)$) and sets (e.g., $\operatorname{WHSIC}(d)$).
  • Matrices and vectors have rows and columns numbered from 00 to $d-1$. We follow the text of Waldron [46], which identifies indices with $\mathbb{Z}/d\mathbb{Z}$.
  • Greek letters generally represent algebraic numbers (e.g., $\xi_{d}=-e^{\frac{\pi i}{d}}$ and $\zeta_{d}=e^{\frac{2\pi i}{d}}$).
  • However, the letter $\sigma$ is exclusively used for Galois automorphisms, with $\sigma_{c}$ denoting complex conjugation. Galois group actions are treated either as acting on scalars or as acting entrywise on vectors or matrices.
  • The notation $S$ is used for a ($d$-dimensional) line-SIC; it is a set of $d^{2}$ complex lines.
  • The notation $[S]$ is used for a geometric equivalence class of SICs. (However, $[\mathbf{v}]$ denotes the complex line $\mathbb{C}\mathbf{v}$, when $\mathbf{v}$ is a vector.)
  • We use the convention that $d^{\prime}$ means $d^{\prime}=d$ if $d$ is odd and $d^{\prime}=2d$ if $d$ is even. The number $\xi_{d}=-e^{\frac{\pi i}{d}}$ is a $d^{\prime}$-th root of unity.
  • We write $\Delta_{d}=(d+1)(d-3)$ for the (often non-fundamental) discriminant associated to dimension $d$, $K_{d}=\mathbb{Q}(\sqrt{\Delta_{d}})$ for the associated field, and $\varepsilon_{d}=\frac{1}{2}(d-1+\sqrt{\Delta_{d}})$ for the associated unit.
2. Weyl–Heisenberg SICs
In this section, we defined and give background on Weyl–Heisenberg (covariant) SICs, which satisfy a particular group symmetry property.
2.1. Weyl–Heisenberg SICs
Specifically, Weyl–Heisenberg line-SICs will be defined as the orbit of a single line under the action of the Weyl-Heisenberg group, a finite subgroup of the group of $d\times d$ unitary matrices.
Definition 2.1.
The Weyl–Heisenberg group is $\operatorname{WH}(d)=\{\xi^{r}X^{p}Z^{q}:p,q,r\in\mathbb{Z}\}$, where $\xi=-e^{\frac{\pi i}{d}}$, $\zeta=\xi^{2}=e^{\frac{2\pi i}{d}}$, and
$X$ $=\left(\begin{array}[]{ccccc}0&0&\cdots&0&1\\ 1&0&\cdots&0&0\\ 0&1&\cdots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&1&0\end{array}\right),$ $Z$ $=\left(\begin{array}[]{ccccc}1&0&0&\cdots&0\\ 0&\zeta&0&\cdots&0\\ 0&0&\zeta^{2}&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\cdots&\zeta^{d-1}\end{array}\right),$ (2.11)
The projective Weyl–Heisenberg group $\operatorname{PWH}(d)$ is defined to be $\operatorname{WH}(d)$ modulo its scalar matrix subgroup $Z(\operatorname{WH}(d))=\{\xi^{j}I:\,j\geq 0\}$, where $I$ is the $d\times d$ identity matrix.
It is easily checked that $ZX=\zeta XZ$. It follows that $\operatorname{WH}(d)$ is a group. Its order is $d^{3}$ if $d$ is odd and $2d^{3}$ if $d$ is even, noting that $\xi$ is a $d$-th root of unity for odd $d$ and a $2d$-th root of unity for even $d$. The center $\operatorname{Z}(\operatorname{WH}(d))=\langle\xi I\rangle$ of $\operatorname{WH}(d)$ is given by scalar matrices, and the quotient group $\operatorname{PWH}(d)\cong\operatorname{WH}(d)/\operatorname{Z}(\operatorname{WH}(d))\cong\left(\mathbb{Z}/d\mathbb{Z}\right)^{2}$.
Definition 2.2.
For $\mathbf{w}\in\mathbb{C}^{d}$, its Weyl–Heisenberg orbit is
$\operatorname{orb}_{\operatorname{WH}}(\mathbf{w}):=\{H\mathbf{w}:\,H\in\operatorname{WH}(d)\}.$ (2.12)
Its associated set of complex lines, denoted $\operatorname{WHL}(\mathbf{w})$ is
$\operatorname{WHL}(\mathbf{w}):=\{[\mathbf{w}]:\,\mathbf{w}\in\operatorname{orb}_{\operatorname{WH}}(\mathbf{w})\}=\{[H\mathbf{w}]:H\in\operatorname{PWH}(d)\}.$ (2.13)
For any $\mathbf{w}$, $\operatorname{WHL}(\mathbf{w})$ has cardinality at most $d^{2}$, and for generic $\mathbf{w}$, exactly $d^{2}$.
Definition 2.3.
A Weyl–Heisenberg covariant SIC (or simply a Weyl–Heisenberg SIC) is any $S(\mathbf{v}):=\operatorname{WHL}(\mathbf{v})$ for which ${\mathbf{\Pi}}(S)$ is a $d$-dimensional SIC. We denote the set of all Weyl–Heisenberg SICs in dimension $d$ as $\operatorname{WHSIC}(d)$.
We call a generating vector $\mathbf{v}$ of a Weyl–Heisenberg SIC a fiducial vector. Such a vector is well-defined up to multiplication by a complex scalar $c\in\mathbb{C}^{\times}$. Each SIC has a choice of $d^{2}$ fiducial vectors, up to scalars. We call two Weyl–Heisenberg SICs scalar-equivalent if they have a common fiducial vector, up to multiplication by a nonzero scalar.
Conjecture 2.4 (Strong Zauner Conjecture).
For each dimension $d\geq 1$, there exists at least one Weyl–Heisenberg SIC.
This conjecture is currently known, by explicit constructions to hold for finitely many dimensions $d$. At present it includes dimensions $1\leq d\leq 53$ and a extensive number of higher dimensions.
Conjecture 2.5 (Weyl–Heisenberg Boundedness Conjecture).
For each dimension $d\geq 1$, except $d=3$, there are finitely many scalar-equivalence classes of Weyl–Heisenberg SICs.
There are a few small dimensions where this conjecture has been proved rigorously, including $d\in\{1,2,4,5\}$. In higher dimensions, up to at least $d=90$, it is supported by non-rigorous evidence given by numerical optimization routines repeatedly returning the same SICs [39, 38]. In dimension $3$, there is a one-parameter family of inequivalent Weyl–Heisenberg SICs giving uncountably many distinct scalar-equivalence classes of SICs.
A closely related conjecture is stated in Waldron [46, Sec. 14.27]: The (real) algebraic variety of (normalized Weyl–Heisenberg) SIC fiducials is zero-dimensional, except for $d=3$, where it is one-dimensional. This conjecture implies 2.5; it is equivalent if one takes “dimension” to mean “real dimension” (i.e., dimension of the Zariski closure) and is a stronger conjecture if one takes “dimension” to mean “dimension of the complex variety defined by the same equations.” Resolving 2.5 or Waldron’s conjecture in general (and specifically even for $n=6$) is an interesting open problem in optimization and real algebraic geometry.
2.2. Geometric equivalence classes of Weyl–Heisenberg SICs
There are two types of geometric operators acting on $\mathbb{C}^{d}$ that send SICs to SICs. Firstly, if $U$ is a unitary matrix in $\mathbf{U}(d)=\{U\in\mathbf{GL}(\mathbb{C}^{d}):UU^{\dagger}=1\}$ and $S=\{\mathbb{C}\mathbf{v}_{1},\ldots,\mathbb{C}\mathbf{v}_{d^{2}}\}$ is a SIC, then so is $US=\{\mathbb{C}U\mathbf{v}_{1},\ldots,\mathbb{C}U\mathbf{v}_{d^{2}}\}$. (This is a unitary equivalence.) Secondly, if $C_{d}$ is the complex conjugation operator acting pointwise on $\mathbb{C}^{d}$, so that $C_{d}\mathbf{v}:=\overline{\mathbf{v}}$, then $C_{d}$ also preserves the SIC property, as does any “antiunitary” operator of the form $C_{d}U$. These operators may be collected together to form the extended unitary group.
Definition 2.6.
The extended unitary group is $\mathbf{EU}(d):=\mathbf{U}(d)\sqcup C_{d}\mathbf{U}(d)$. The projective extended unitary group $\mathbf{PEU}(d)$ is defined to be $\mathbf{EU}(d)$ modulo its scalar matrix subgroup $\{zI:z\in\mathbb{C},\left|z\right|=1\}$.
All known SICs are group covariant, meaning that they are the orbit of a vector under the action of a finite subgroup of the unitary group.
Definition 2.7.
A SIC $S$ is group covariant for a subgroup $G$ of $\mathbf{U}(d)$ if there exists a vector $v\in\mathbb{C}^{d}$, called a fiducial vector, such that $S=\{\mathbb{C}Av:A\in G\}$.
If $\mathbf{v}$ is a fiducial vector for a Weyl–Heisenberg covariant SIC and $U\in\mathbf{EU}(d)$, then the vector $U\mathbf{v}$ will be a fiducial vector for a SIC covariant for the group $G^{\prime}=U\operatorname{WH}(d)U^{-1}$. All currently known SICs are unitary-equivalent to a Weyl–Heisenberg group $\operatorname{WH}(d)$ covariant SIC, with a single exception, the Hoggar lines in $\mathbb{C}^{8}$ [29, 30]. The Hoggar lines are group covariant for the tensor product $\operatorname{WH}(2)\otimes\operatorname{WH}(2)\otimes\operatorname{WH}(2)$. (The group $\operatorname{WH}(2)\otimes\operatorname{WH}(2)\otimes\operatorname{WH}(2)$ may itself be thought of as a generalized (Weyl)-Heisenberg group.)
The extended unitaries $U\in\mathbf{EU}(d)$ that preserve the Weyl–Heisenberg covariance property are precisely those for which $U\operatorname{WH}(d)U^{-1}=\operatorname{WH}(d)$, that is, elements of the extended Clifford group, defined as follows.
Definition 2.8.
We define several closely related groups.
  • (1)
    The Clifford group $\operatorname{C}(d)$ is the normalizer of $\operatorname{WH}(d)$ in $\mathbf{U}(d)$.
  • (2)
    The projective Clifford group $\operatorname{PC}(d)$ is the quotient group of $\operatorname{C}(d)$ modulo its scalar matrix subgroup.
  • (3)
    The extended Clifford group $\operatorname{EC}(d)$ is the normalizer of $\operatorname{WH}(d)$ in $\mathbf{EU}(d)$.
  • (4)
    The projective extended Clifford group $\operatorname{PEC}(d)$ is the quotient of the extended Clifford group $\operatorname{EC}(d)$ modulo its scalar matrix subgroup.
The Clifford group $\operatorname{C}(d)$ is infinite because it contains a copy of $\mathbf{U}(1)$ as scalar matrices $zI$ with $\left|z\right|=1$; thus, $\operatorname{EC}(d)$ is infinite. However, $\operatorname{PC}(d)$ and $\operatorname{PEC}(d)$ are finite groups. (The extended Clifford $\operatorname{EC}(d)$ group is studied in connection with SICs by Appleby [1].)
Definition 2.9.
The set of geometric equivalence classes of Weyl–Heisenberg SICs is the set of equivalence classes of $\operatorname{WHSIC}(d)$ under the equivalence relation $[S({\mathbf{v}})]\sim[S(U\mathbf{v})]$ whenever $U\in\operatorname{PEC}(d)$ (equivalently, whenever $U\in\operatorname{EC}(d)$.) This set of equivalence classes is denoted $\operatorname{WHSIC}(d)/\operatorname{PEC}(d)$. (It is not a group.) We denote the geometric equivalence class of a Weyl–Heisenberg SIC $S$ by $[S]$.
3. Fields generated by SICs
In this section, we define several number fields attached to arbitrary SICs. Many of these fields have been studied before, but they are often discussed within the context of assuming various conjectures about SICs. Here, we focus on relationships between “SIC fields” that can be proven rigorously, either in general (Section 3.1) or in the case of Weyl–Heisenberg SICs (Section 3.2).
In Section 4, we will focus on SICs expressible in terms of algebraic numbers and discuss some conditional results about fields generated by SICs in that setting.
3.1. Fields generated by general SICs
The fields we will study here are each associated to an individual SIC $S$ and depend only on data computed from the orthogonal projections $\{\Pi_{i}:1\leq i\leq d^{2}\}$ defining the SIC.
Definition 3.1.
Let ${\mathbf{\Pi}}(S)=\{\frac{1}{d}\Pi_{1},\ldots,\frac{1}{d}\Pi_{d^{2}}\}$ be an arbitrary SIC in dimension $d$, with each $\Pi_{i}=\mathbf{v}_{i}\mathbf{v}_{i}^{\dagger}$ for a unit vector $\mathbf{v}_{i}$. (Thus each $\mathbf{v}_{i}$ is a unit fiducial vector, unique up to multiplication by a unit scalar $e^{i\theta}$.)
  • (1)
    The ratio SIC-field $F^{\rm vec}\!\left(S\right)$ is the field extension of $\mathbb{Q}$ generated by all the ratios $v_{ij}/v_{ik}$ (for $v_{ik}\neq 0$) of the entries of the $d^{2}$ fiducial vectors $\mathbf{v}_{i}$ of $S$.
  • (2)
    The triple product SIC-field $F^{\rm trip}\!\left(S\right)$ is the field extension of $\mathbb{Q}$ generated by all the “triple products” $\operatorname{Tr}(\Pi_{i}\Pi_{j}\Pi_{k})$.
  • (3)
    The unitary invariant SIC-field $F^{\rm inv}\!\left(S\right)$ is the field extension of $\mathbb{Q}$ generated by all the unitary invariants $\operatorname{Tr}(\Pi_{i_{1}}\Pi_{i_{2}}\cdots\Pi_{i_{n}})$ for $n\in\mathbb{N}$.
  • (4)
    The projection SIC-field $F^{\rm proj}\!\left(S\right)$ is the field extension of $\mathbb{Q}$ generated by all the entries of the $d^{2}$ Hermitian projection matrices $\Pi_{j}$.
  • (5)
    The extended projection SIC-field $F^{\rm eproj}\!\left(S\right):=F^{\rm proj}\!\left(S\right)\!(\xi_{d})$, adjoining $\xi_{d}=-e^{\frac{\pi i}{d}}$.
A priori, these fields could be transcendental extensions of $\mathbb{Q}$, and this does occur in dimension $d=3$.
Remark 3.2.
These fields have previously been studied with various names and notations, which we now review.
  1. (1)
    The field $F^{\rm vec}\!\left(S\right)$ is the SIC field $\mathbb{Q}[S]$ attached to a general SIC $S$ by Appleby, Flammia, McConnell, and Yard [10, p. 212 bottom], as used in Thm. 1 and Conj. 2 of that paper. By definition, this field is independent of scaling factors in the $\mathbf{v}_{i}$. It is determined by ${\mathbf{\Pi}}(S)$ because the ratios
    $\frac{v_{ij}}{v_{ik}}=\frac{v_{ij}\overline{v}_{ik}}{v_{ik}\overline{v}_{ik}}=\frac{(\Pi_{i})_{jk}}{(\Pi_{i})_{kk}}.$ (3.1)
    We may rescale each of the $d^{2}$ vectors in $S$ to give $S=\{\mathbb{C}{\mathbf{w}}_{i}:1\leq i\leq d^{2}\}$ so that all entries of each ${\mathbf{w}}_{i}$ belong to $F^{\rm vec}\!\left(S\right)$. A given ${\mathbf{v}}_{i}=\left(v_{i0},v_{i1},\cdots,v_{i(d-1)}\right)^{T}$ has at least one nonzero entry, say $v_{ik}$, and we may set ${\mathbf{w}}_{i}=\frac{1}{v_{ik}}{\mathbf{v}}_{i}$, whence $w_{ij}=v_{ij}/v_{ik}\in F^{\rm vec}\!\left(S\right)$. The vectors ${\mathbf{w}}_{i}$ obtained this way may not be unit vectors.
    Prop. 3 of [10] states that, for a Weyl–Heisenberg SIC, the ratio SIC-field is the same for any SIC $S^{\prime}\in[S]$. The presented proof outline [10, Sec. 7.1] appears to assume some conjectural equivalences and characterizations of other related fields.
  2. (2)
    Concerning the triple product SIC-field $F^{\rm trip}\!\left(S\right)$, the fact that a SIC is determined up to unitary equivalence by its triple products was proved in 2011 by Appleby, Flammia, and Fuchs [3, Thm. 3]. We show unconditionally that $F^{\rm trip}\!\left(S\right)$ is an invariant of the geometric $\operatorname{PEC}(d)$-orbit of a Weyl–Heisenberg SIC $S$; see 3.4.
  3. (3)
    The unitary invariant SIC-field $F^{\rm inv}\!\left(S\right)$ is generated by the full set of projective unitary invariants of a general finite set of lines in $\mathbb{C}^{d}$; see [15] and [46, Ch. 8, Thm. 8.2]. We show in 3.3 that $F^{\rm trip}\!\left(S\right)=F^{\rm inv}\!\left(S\right)$.
  4. (4)
    The projection SIC-field $F^{\rm proj}\!\left(S\right)$, in the special case that $S$ is a Weyl–Heisenberg SIC, coincides with the field $L$ appearing in [10, Sec. 7], defined as the smallest subfield of $\mathbb{C}$ generated by the entries $v_{j}\overline{v}_{k}$ of unit vectors generating $\Pi_{\mathbf{v}}$ taken over all $d^{2}$ projectors $\Pi_{\mathbf{v}}$ of a Weyl–Heisenberg SIC.
  5. (5)
    The extended projection SIC-field $F^{\rm eproj}\!\left(S\right)$ is the SIC field $\mathbb{E}$ as defined by Appleby, Flammia, McConnell, and Yard [9, Sec. 4] for any SIC projector $S$. This definition of the SIC field is also used in [7]. Appleby, Flammia, McConnell, and Yard [9] state (without proof) that the field $\mathbb{E}$ is well-defined on the $\operatorname{PEC}(d)$-orbit $[S]$ of a Weyl–Heisenberg SIC $S$. We prove so unconditionally in 3.8.
The next result establishes some inclusion relations among these fields.
Proposition 3.3.
Let $S$ be an arbitrary line-SIC in dimension $d$. The fields associated to $S$ are related in the following ways.
  • (1)
    $F^{\rm proj}\!\left(S\right)=F^{\rm vec}\!\left(S\right)\overline{F^{\rm vec}\!\left(S\right)}$ (the compositum of $F^{\rm vec}\!\left(S\right)$ and its complex conjugate field). In particular, $F^{\rm proj}\!\left(S\right)$ is invariant under complex conjugation.
  • (2)
    $F^{\rm trip}\!\left(S\right)=F^{\rm inv}\!\left(S\right)\subseteq F^{\rm proj}\!\left(S\right)$.
Proof.
Write ${\mathbf{\Pi}}(S)=\{\frac{1}{d}\Pi_{1},\ldots,\frac{1}{d}\Pi_{d^{2}}\}$.
Proof of (1): Write $\Pi_{i}=\mathbf{v}_{i}\mathbf{v}_{i}^{\dagger}$, where $\mathbf{v}_{i}$ has entries $v_{ij}$ and $\left|\mathbf{v}_{i}\right|^{2}=1$. The nonzero entries of $\Pi_{i}$ are of the form $v_{ij}\overline{v}_{ik}$ for $v_{ij}\neq 0$, $v_{ik}\neq 0$ These entries can be expressed as
$v_{ij}\overline{v}_{ik}=\frac{v_{ij}\overline{v}_{ik}}{\left|\mathbf{v}_{i}\right|^{2}}=\frac{1}{\sum_{\ell=1}^{d}(v_{i\ell}/v_{ij})\overline{(v_{i\ell}/v_{ik})}}\in F^{\rm vec}\!\left(S\right)\overline{F^{\rm vec}\!\left(S\right)}.$ (3.2)
Thus, $F^{\rm proj}\!\left(S\right)\subseteq F^{\rm vec}\!\left(S\right)\overline{F^{\rm vec}\!\left(S\right)}$. To show the reverse inclusion, note that for $v_{ik}\neq 0$,
$v_{ij}/v_{ik}=\frac{v_{ij}\overline{v}_{ik}}{v_{ik}\overline{v}_{ik}}\mbox{ and }\overline{(v_{ij}/v_{ik})}=\frac{v_{ik}\overline{v}_{ij}}{v_{ik}\overline{v}_{ik}}.$ (3.3)
Proof of (2): First we show the equality $F^{\rm trip}\!\left(S\right)=F^{\rm inv}\!\left(S\right)$. Clearly $F^{\rm trip}\!\left(S\right)\subseteq F^{\rm inv}\!\left(S\right)$. To show the reverse inclusion, note that the $\Pi_{i}$ form a basis for the space of $d\times d$ complex matrices, and express
$\Pi_{i}\Pi_{j}=\sum_{\ell}\alpha_{ij}^{\ell}\Pi_{\ell}$ (3.4)
for structure constants $\alpha_{ij}^{\ell}\in\mathbb{C}$. We have
$\sum_{\ell}\alpha_{ij}^{\ell}=\operatorname{Tr}(\Pi_{i}\Pi_{j})=\frac{\delta_{ij}d+1}{d+1},\mbox{ where }\delta_{ij}=\begin{cases}1,&\mbox{if }i=j,\\ 0,&\mbox{if }i\neq j.\end{cases}$ (3.5)
Moreover, the triple products are related to the structure constants by
$\operatorname{Tr}(\Pi_{i}\Pi_{j}\Pi_{k})$ $=\sum_{\ell}\alpha_{ij}^{\ell}\operatorname{Tr}(\Pi_{k}\Pi_{\ell})$ (3.6)
$=\frac{d}{d+1}\alpha_{ij}^{k}+\frac{1}{d+1}\sum_{\ell}\alpha_{ij}^{\ell}$ (3.7)
$=\frac{d}{d+1}\alpha_{ij}^{k}+\frac{\delta_{ij}d+1}{(d+1)^{2}}.$ (3.8)
Thus, $\alpha_{ij}^{k}=\frac{d+1}{d}\operatorname{Tr}(\Pi_{i}\Pi_{j}\Pi_{k})-\frac{d(\delta_{ij}d+1)}{d+1}$, and the field generated over $\mathbb{Q}$ by the structure constants is $F^{\rm trip}\!\left(S\right)$. Moreover, using (3.4) repeatedly, we can express any product $\Pi_{i_{1}}\Pi_{i_{2}}\cdots\Pi_{i_{n}}$ in the $\left\{\Pi_{\ell}\right\}_{\ell}$-basis with coefficients that are integer polynomials in the structure constants. Thus, since $\operatorname{Tr}(\Pi_{\ell})=1$, $\operatorname{Tr}(\Pi_{i_{1}}\Pi_{i_{2}}\cdots\Pi_{i_{n}})$ is also an integer polynomials in the structure constants and is thus contained in $F^{\rm trip}\!\left(S\right)$.
Finally, the inclusion $F^{\rm inv}\!\left(S\right)\subseteq F^{\rm proj}\!\left(S\right)$ of (2) follows because $\operatorname{Tr}(\Pi_{i_{1}}\Pi_{i_{2}}\cdots\Pi_{i_{n}})$ is an integer polynomial in the entries of the various $\Pi_{j}$. ∎
Proposition 3.4.
Let $S$ be an arbitrary line-SIC in dimension $d$. Then the triple product field satisfies the invariance property $F^{\rm trip}\!\left(US\right)=F^{\rm trip}\!\left(S\right)$ for any $U\in\mathbf{EU}(d)$.
Proof.
Let $S=\{\mathbb{C}{\mathbf{v}}_{i}:1\leq i\leq d^{2}\}$ be an arbitrary line-SIC, and set ${\mathbf{\Pi}}(S)=\{\frac{1}{d}\Pi_{1},\ldots,\frac{1}{d}\Pi_{d^{2}}\}$, with $\Pi_{i}={\mathbf{v}}_{i}{\mathbf{v}}_{i}^{\dagger}$. Then ${\mathbf{\Pi}}(US)=\{\frac{1}{d}\Pi_{1}^{\prime},\ldots,\frac{1}{d}\Pi_{d^{2}}^{\prime}\}$. Write $U=CV$, where $C$ is either the identity or complex conjugation, and $V\in\mathbf{U}(d)$; then either $\Pi_{i}^{\prime}=V^{-1}\Pi_{i}V$, or $\Pi_{i}^{\prime}=V^{-1}\Pi_{i}^{\top}V$. In the first case, $\operatorname{Tr}(\Pi_{i}^{\prime}\Pi_{j}^{\prime}\Pi_{k}^{\prime})=\operatorname{Tr}(V^{-1}\Pi_{i}\Pi_{j}\Pi_{k}V)=\operatorname{Tr}(\Pi_{i}\Pi_{j}\Pi_{k})$. In the second case, $\operatorname{Tr}(\Pi_{i}^{\prime}\Pi_{j}^{\prime}\Pi_{k}^{\prime})=\operatorname{Tr}\!\left(V^{-1}\Pi_{i}^{\top}\Pi_{j}^{\top}\Pi_{k}^{\top}V\right)=\operatorname{Tr}\!\left((\Pi_{k}\Pi_{j}\Pi_{i})^{\top}\right)=\operatorname{Tr}(\Pi_{k}\Pi_{j}\Pi_{i})$. So $F^{\rm trip}\!\left(US\right)=F^{\rm trip}\!\left(S\right)$ in both cases. ∎
We derive some unconditional implications when these fields give Galois extensions of $\mathbb{Q}$.
Lemma 3.5.
Let $S$ be an arbitrary line-SIC in dimension $d$.
  • (1)
    If $F^{\rm vec}\!\left(S\right)$ is invariant under complex conjugation $\sigma_{c}$, then $F^{\rm vec}\!\left(S\right)=F^{\rm proj}\!\left(S\right).$ Thus, if $F^{\rm vec}\!\left(S\right)$ is a Galois extension of $\mathbb{Q}$, then $F^{\rm proj}\!\left(S\right)$ is a Galois extension of $\mathbb{Q}$.
  • (2)
    If $F^{\rm proj}\!\left(S\right)$ is a Galois extension of $\mathbb{Q}$, then $F^{\rm eproj}\!\left(S\right)$ is a Galois extension of $\mathbb{Q}$.
Proof.
Proof of (1): The implication $F^{\rm vec}\!\left(S\right)=F^{\rm proj}\!\left(S\right)$ follows from 3.3(1).
Proof of (2): If $K=F^{\rm proj}\!\left(S\right)$ is Galois over $\mathbb{Q}$, then $F^{\rm eproj}\!\left(S\right)$ is the compositum $KL$, with $L=\mathbb{Q}(e^{\frac{\pi i}{d}})$ an abelian extension of $\mathbb{Q}$, so the compositum is a Galois extension of $\mathbb{Q}$. ∎
3.2. Fields generated by Weyl–Heisenberg SICs
We now derive some properties of the field $F^{\rm vec}\!\left(S\right)$, $F^{\rm trip}\!\left(S\right)$ and $F^{\rm proj}\!\left(S\right)$ in the special case of Weyl–Heisenberg SICs.
Lemma 3.6.
Let $S$ be a Weyl–Heisenberg line-SIC $S=S([{\mathbf{v}}])$ of dimension $d$. Then:
  • (1)
    $F^{\rm proj}\!\left(S\right)$ contains a $d$-th root of unity $\zeta_{d}=e^{\frac{2\pi i}{d}}$.
  • (2)
    The index $[F^{\rm eproj}\!\left(S\right):F^{\rm proj}\!\left(S\right)]\in\{1,2\}$. It is $1$ whenever $d$ is odd.
Proof.
Proof of (1): Let ${\mathbf{v}}=(v_{0},v_{1},\ldots,v_{d-1})^{\top}$ be a fiducial vector for $S$, and let $\Pi={\mathbf{v}}{\mathbf{v}}^{\dagger}$. It has $(1,0)$-entry $(\Pi)_{10}=v_{1}\overline{v}_{0}\in F^{\rm proj}\!\left(S\right)$. Now the modulation matrix $Z\in\operatorname{WH}(d)$ gives another SIC vector $Z{\mathbf{v}}=(v_{0},e^{\frac{2\pi i}{d}}v_{1},\cdots,e^{\frac{2\pi i(d-1)}{d}}v_{d-1})$. The corresponding projection $\Pi^{\prime}:=Z{\mathbf{v}}(Z{\mathbf{v}})^{\dagger}=Z{\mathbf{v}}{\mathbf{v}}^{\dagger}Z^{-1}$ has $(\Pi^{\prime})_{1,0}=e^{\frac{2\pi i}{d}}v_{1}\overline{v}_{0}\in F^{\rm proj}\!\left(S\right)$. If $v_{1}\overline{v}_{0}\neq 0$, then we may conclude $e^{\frac{2\pi i}{d}}\in F^{\rm proj}\!\left(S\right)$.
We show every Weyl–Heisenberg SIC has a fiducial vector ${\mathbf{v}}$ with both $v_{0}$ and $v_{1}$ nonzero, so the argument applies. We first claim that every fiducial vector ${\mathbf{v}}$ has two consecutive entries $v_{i},v_{i+1}$ nonzero, allowing the wraparound product $v_{d-1}v_{0}$. Suppose not; then, the translation matrix $X\in\operatorname{WH}(d)$ gives a cyclic shift $X{\mathbf{v}}=(v_{1},v_{2},...,v_{d-1},v_{0})$, so $\langle{\mathbf{v}},X{\mathbf{v}}\rangle=\sum_{i=0}^{d-1}v_{i}\overline{v}_{i+1}=0$, contradicting the SIC property $\left|\langle{\mathbf{v}},X{\mathbf{v}}\rangle\right|=\frac{1}{d+1}$. The claim follows. If the given ${\mathbf{v}}$ has $v_{i}v_{i+1}\neq 0$, then ${\mathbf{v}}^{\prime}=X^{d-i}{\mathbf{v}}$ cyclically shifts entries so that $v_{0}^{\prime}v_{1}^{\prime}\neq 0$.
Proof of (2): By (1), $F^{\rm eproj}\!\left(S\right)=F^{\rm proj}\!\left(S\right)(e^{\frac{\pi i}{d}})$ is at most a quadratic extension of $F^{\rm proj}\!\left(S\right)$, since the field index $[\mathbb{Q}(e^{\frac{\pi i}{d}}):\mathbb{Q}(e^{\frac{2\pi i}{d}})]=2$ if $d$ is even and is $1$ if $d$ is odd. ∎
We now study the field $F^{\rm eproj}\!\left(S\right)$ on the $\operatorname{PEC}(d)$-orbit of $S$ and show it is a $\operatorname{PEC}(d)$-invariant. We recall a characterization of generators of the Clifford group.
Proposition 3.7.
The Clifford group $\operatorname{C}(d)$ is generated by all unitary scalar matrices $zI$ with $\left|z\right|=1$, together with
$X,Z,F,R$ (3.9)
where $X,Z$ are Weyl–Heisenberg group generators, $F$ is the discrete Fourier transform,
$F_{j,k}=\frac{1}{\sqrt{d}}e^{\frac{2\pi ijk}{d}}\,\mbox{ for }\,0\leq j,k\leq d-1.$ (3.10)
and $R$ is a diagonal matrix, with diagonal entries
$R_{j,j}=\xi_{d}^{j(j+d)}=(-1)^{jd}e^{\frac{\pi ij^{2}}{d}}\,\mbox{ for }\,0\leq j\leq d-1.$ (3.11)
Proof.
This is [46, Thm. 14.1]. ∎
Proposition 3.8.
Let $S$ be a Weyl–Heisenberg line-SIC $S=S([{\mathbf{v}}])$ of dimension $d$. For any $M\in\operatorname{EC}(d)$,
$F^{\rm eproj}\!\left(MS\right)=F^{\rm eproj}\!\left(S\right),$ (3.12)
where $MS$ denotes the Weyl–Heisenberg line-SIC having fiducial vector ${\mathbf{w}}:=M\mathbf{v}$. Moreover, $F^{\rm eproj}\!\left(S\right)$ is equal to the compositum of all the $F^{\rm proj}\!\left(MS\right)$ for $M\in\operatorname{EC}(d)$.
Proof.
The scalar matrices $zI$ with $\left|z\right|=1$ act trivially on projections, with $(zI\mathbf{v})(zI\mathbf{v})^{\dagger}=\mathbf{v}\mathbf{v}^{\dagger}$. The matrices $X$, $Z$, and $F$ all have entries in $\mathbb{Q}(\xi)$ (where $\xi=\xi_{d}=-e^{\frac{\pi i}{d}}$). The matrix $\sqrt{d}F$ also has entries in $\mathbb{Q}(\xi_{d})$, and $(F\mathbf{v})(F\mathbf{v})^{\dagger}=\frac{1}{d}(\sqrt{d}F\mathbf{v})(\sqrt{d}F\mathbf{v})^{\dagger}$. Therefore, writing any $M\in EC(d)$ as a product of the generators given in 3.7 shows that $(M\mathbf{v})(M\mathbf{v})^{\dagger}=M\mathbf{v}\mathbf{v}^{\dagger}M^{\dagger}$ has entries in $F^{\rm eproj}\!\left(S\right)$ (the compositum of $F^{\rm proj}\!\left(S\right)$ and $\mathbb{Q}(\xi)$). This argument shows that $F^{\rm proj}\!\left(MS\right)\subseteq F^{\rm eproj}\!\left(S\right)$, so $F^{\rm eproj}\!\left(MS\right)\subseteq F^{\rm eproj}\!\left(S\right)$. By considering $M^{-1}$, we also have $F^{\rm eproj}\!\left(S\right)=F^{\rm eproj}\!\left(M^{-1}MS\right)\subseteq F^{\rm eproj}\!\left(MS\right)$, so in fact $F^{\rm eproj}\!\left(MS\right)=F^{\rm eproj}\!\left(S\right)$.
To show that $F^{\rm eproj}\!\left(S\right)$ is equal to the compositum of all the $F^{\rm proj}\!\left(MS\right)$, it suffices to show that $\xi=-e^{-\frac{\pi i}{d}}$ occurs in the compositum of two fields $F^{\rm proj}\!\left(S\right)$ and $F^{\rm proj}\!\left(RS\right)$ where $R$ is the element of $\operatorname{C}(d)$ given in 3.7. Given a fiducial ${\mathbf{v}}$ generating $S=S([{\mathbf{v}}])$ having $v_{0}v_{1}\neq 0$, we consider the fiducial vector ${\mathbf{w}}=R{\mathbf{v}}$ of a geometrically equivalent SIC. It has entries $(R{\mathbf{v}})_{j}=\xi^{j(j+d)}v_{j}$ so $R{\mathbf{v}}_{0}=v_{0}$ and $R{\mathbf{v}}_{1}=-e^{\frac{\pi i}{d}}v_{1}$. Therefore $\Pi^{\prime}={\mathbf{w}}{\mathbf{w}}^{\dagger}$ has
$(\Pi^{\prime})_{1,0}=-e^{\frac{\pi i}{d}}v_{1}\overline{v}_{0}\in F^{\rm proj}\!\left(RS\right),$ (3.13)
while $v_{1}\overline{v}_{0}\in F^{\rm proj}\!\left(S\right)$. Now $v_{0}v_{1}\neq 0$ implies we may take their quotient in the compositum field and so get $-e^{\frac{\pi i}{d}}\in F^{\rm proj}\!\left(S\right)F^{\rm proj}\!\left(RS\right)$, as required. ∎
4. Algebraic SICs and the Galois action
We now focus on the subclass of algebraic SICs, for which the fields defined in Section 3.1 are algebraic number fields. In this paper, $\overline{\mathbb{Q}}$ is defined to be the algebraic closure of $\mathbb{Q}$ in $\mathbb{C}$, so there is a canonical choice of complex conjugation automorphism $\sigma_{c}\in\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$, $\sigma_{c}(z)=\overline{z}$. We show that there is a well-defined Galois action on algebraic SICs and that the property of $S$ being an algebraic SIC is preserved by the subgroup of all automorphisms $\sigma\in\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ that commute with $\sigma_{c}$ when restricted to the Galois closure of $F^{\rm proj}\!\left(S\right)$.
It is conjectured for dimensions $d\neq 3$ that all Weyl–Heisenberg SICs are algebraic. In Section 4.3, we define a combined geometric and Galois action on all algebraic Weyl–Heisenberg SICs and classify them into multiplets.
4.1. Algebraic SICs
We consider SICs generated by algebraic vectors.
Definition 4.1.
A complex line $\mathbb{C}{\mathbf{v}}$ generated by a nonzero vector ${\mathbf{v}}\in\mathbb{C}^{d}$ is a projective-algebraic line if all the ratios $v_{i}/v_{j}\in\overline{\mathbb{Q}}$ (for $v_{j}\neq 0$). (The definition is independent of the choice of ${\mathbf{v}}$.) A nonzero vector such that $\mathbb{C}\mathbf{v}$ is a projective-algebraic line is called a projective-algebraic vector.
An algebraic vector is some ${\mathbf{v}}\in{\overline{\mathbb{Q}}}^{d}$. For a projective-algebraic vector $\mathbf{v}$, we set its $\overline{\mathbb{Q}}$-equivalence class
$[[{{\mathbf{v}}}]]_{\operatorname{alg}}:=\{\lambda{\mathbf{v}}:\lambda\in\mathbb{C}\mbox{ and }\lambda{\mathbf{v}}\in\overline{\mathbb{Q}}^{d}\}.$ (4.1)
The next result implies that the algebraic class $[[{{\mathbf{v}}}]]_{\operatorname{alg}}$ of an algebraic complex line always includes algebraic unit vectors.
Lemma 4.2.
If ${\mathbf{v}}\in\mathbb{C}^{d}$ is a projective-algebraic vector, then there exists $\lambda\in\mathbb{C}^{\times}$ such that ${\mathbf{x}}=\lambda{\mathbf{v}}$ is algebraic and $\left|{\mathbf{x}}\right|=1$. The $\overline{\mathbb{Q}}$-equivalence class $[[{\mathbf{x}}]]_{\operatorname{alg}}$ is independent of $\lambda$.
Proof.
Let ${\mathbf{v}}$ have some $v_{i}\neq 0$ and set $\mu=\frac{1}{v_{i}}$. Then ${\mathbf{w}}:=\mu{\mathbf{v}}=\left(\frac{v_{0}}{v_{i}},\frac{v_{1}}{v_{i}},\ldots,\frac{v_{d-1}}{v_{i}}\right)\in\overline{\mathbb{Q}}^{d}$ by the projective-algebraicity of ${\mathbf{v}}$. Now the complex conjugates $\sigma_{c}\!\left(\frac{v_{k}}{v_{i}}\right)=\frac{\overline{v}_{k}}{\overline{v}_{i}}$ are algebraic numbers, hence so is $\left|\frac{v_{k}}{v_{i}}\right|^{2}$. Thus, $\left|{\mathbf{w}}\right|=\sqrt{\sum_{k}\left|\frac{v_{k}}{v_{i}}\right|^{2}}$ is a positive real algebraic number. We set $\lambda=\frac{\mu}{\left|{\mathbf{w}}\right|}$ and obtain the required ${\mathbf{x}}=\frac{{\mathbf{w}}}{\left|{\mathbf{w}}\right|}=\lambda{\mathbf{v}}$ having $\left|{\mathbf{x}}\right|=1$.
Finally, if $\lambda$ and $\lambda^{\prime}$ give two such ${\mathbf{x}},{\mathbf{x}}^{\prime}\!$, necessarily $\lambda^{\prime}/\lambda\in\overline{\mathbb{Q}}$. ∎
Definition 4.3.
A $d$-dimensional line-SIC $S=\{{\mathbf{v}}_{i}:1\leq i\leq d^{2}\}$ is algebraic if each of its $d^{2}$ vectors ${\mathbf{v}}_{i}$ are projective-algebraic.
Lemma 4.4.
Algebraic line-SICs satisfy the following properties.
  • (1)
[…]
  • (2)
    For an algebraic SIC the fields $F^{\rm vec}\!\left(S\right)$, $F^{\rm trip}\!\left(S\right)$ and $F^{\rm proj}\!\left(S\right)$ are algebraic number fields; that is, finite algebraic extensions of $\mathbb{Q}$.
  • (3)
    A Weyl–Heisenberg line-SIC with fiducial vector ${\mathbf{v}}$ is an algebraic SIC if and only if ${\mathbf{v}}$ is a projective-algebraic vector.
Proof.
Proof of (1): This follows using 4.2.
Proof of (2): That $F^{\rm vec}\!\left(S\right)$ is an algebraic number field follows directly from the definition of algebraic. Given a line-SIC specified by $d^{2}$ projective algebraic lines $\mathbb{C}{{\mathbf{v}}_{i}}$, using (1) we immediately have $F^{\rm trip}\!\left(S\right)$ being an algebraic number field. Furthermore, one may rescale all ${\mathbf{v}}_{i}$ to be algebraic unit vectors by 4.2. Now the associated projection matrices $\Pi_{{\mathbf{v}}}={\mathbf{v}}_{i}{\mathbf{v}}_{i}^{\dagger}$ have all entries in $\overline{\mathbb{Q}}$, hence $F^{\rm proj}\!\left(S\right)$ will be an algebraic number field.
Proof of (3): If the fiducial vector ${\mathbf{v}}$ of a Weyl–Heisenberg line-SIC $S$ is projective-algebraic, then by 4.2, ${\mathbf{v}}$ may be rescaled to be an algebraic unit vector ${\mathbf{w}}=\lambda{\mathbf{v}}$. The Weyl–Heisenberg group consists of matrices with algebraic entries, hence its action on ${\mathbf{w}}$ shows that all ${\mathbf{w}}_{i}$ for $1\leq i\leq d^{2}$ are algebraic, so $S$ is an algebraic line-SIC. The converse direction is immediate. ∎
It is conjectured that for each dimension $d\neq 3$, all Weyl–Heisenberg SICs are algebraic.
Conjecture 4.5 (Weyl–Heisenberg Algebraicity Conjecture).
For each dimension $d\geq 1$, except $d=3$, all Weyl–Heisenberg SICs are projective-algebraic.
This conjecture is equivalent to 2.5, which states that for all $d\neq 3$ the real projective algebraic variety $\operatorname{WHSIC}(d)$ is a finite set. It would follow by 4.4 that, if $S$ a Weyl–Heisenberg SIC, then $F^{\rm vec}\!\left(S\right)$, $F^{\rm trip}\!\left(S\right)$, and $F^{\rm proj}\!\left(S\right)$ are algebraic number fields.
4.2. Galois actions on algebraic SICs
There is a well-defined action of $\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ on algebraic complex lines which maps them to other algebraic complex lines. Given $\sigma\in\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ and ${\mathbf{v}}\in\overline{\mathbb{Q}}^{d}$ we send ${\mathbf{v}}\mathbb{C}\mapsto\sigma({\mathbf{v}})\mathbb{C}$. This action is well-defined on algebraic complex lines, since if $\lambda\in\overline{\mathbb{Q}}^{\times}$, then ${\mathbf{v}}\mathbb{C}=\lambda{\mathbf{v}}\mathbb{C}$ while
$\sigma(\lambda{\mathbf{v}})\mathbb{C}=\sigma(\lambda)\sigma({\mathbf{v}})\mathbb{C}=\sigma({\mathbf{v}})\mathbb{C}.$ (4.2)
Under this Galois action, the image under $\sigma$ of $d^{2}$ algebraic complex lines will be $d^{2}$ algebraic complex lines. If these algebraic lines form a line-SIC, then the image set of lines can be either the same SIC, a different SIC, or not a SIC at all.
For a particular SIC, there is a restricted set of Galois automorphisms (which may depend on the line-SIC) that map it to another line-SIC.
Definition 4.6.
Let $\sigma_{c}\in\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ denote complex conjugation and $E$ an algebraic Galois extension of $\mathbb{Q}$. Let $C(\sigma_{c}\!\mid\!E)$ be the centralizer of $\sigma_{c}$ restricted to $\operatorname{Gal}(E/\mathbb{Q})$, that is,
$C(\sigma_{c}\!\mid\!E):=\{\sigma\in\operatorname{Gal}(E/\mathbb{Q}):\,\sigma\sigma_{c}(\alpha)=\sigma_{c}\sigma(\alpha)\mbox{ for all }\alpha\in E\}.$ (4.3)
It is a subgroup of $\operatorname{Gal}(E/\mathbb{Q})$.
For a line-SIC $S$, we will show there is a Galois field $E$ (depending on $S$) for which $C(\sigma_{c}\!\mid\!E)$ is such a restricted set. However, neither the field $F^{\rm proj}\!\left(S\right)$ nor $F^{\rm eproj}\!\left(S\right)$ is known to always be a Galois extension of $\mathbb{Q}$, although there is no known counterexample for $d\neq 3$. Thus, we give a name to the Galois closure of $F^{\rm eproj}\!\left(S\right)$.
Definition 4.7.
Given an algebraic line-SIC $S$, let $F^{\rm Gproj}\!\left(S\right)$ denote the normal closure of $F^{\rm eproj}\!\left(S\right)$ over $\mathbb{Q}$, and call it the Galois-projection SIC field of $S$.
The next proposition shows that $C(\sigma_{c}\!\mid\!F^{\rm Gproj}\!\left(S\right))$ is a group of Galois automorphisms that preserve the property of being an algebraic line-SIC. For $d\geq 1$, the group $C(\sigma_{c}\!\mid\!F^{\rm Gproj}\!\left(S\right))$ is always nontrivial because it contains complex conjugation, which acts nontrivially because $F^{\rm proj}\!\left(S\right)$ cannot be contained in $\mathbb{R}$. (A set of real equiangular lines in $\mathbb{R}^{d}$ has cardinality at most $\frac{d(d+1)}{2}<d^{2}$.)
Proposition 4.8.
Let $S$ be an algebraic line-SIC. If $\sigma\in C(\sigma_{c}\!\mid\!F^{\rm Gproj}\!\left(S\right))$, then $\sigma(S)$ is also an algebraic line-SIC, and $F^{\rm Gproj}\!\left(\sigma(S)\right)=F^{\rm Gproj}\!\left(S\right)$.
Proof.
Write the line-SIC and (projection) SIC as
$S$ =$=\left\{\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{d^{2}}\right\},$ (4.4)
${\mathbf{\Pi}}(S)$ =$=\left\{\tfrac{1}{d}\Pi_{1},\tfrac{1}{d}\Pi_{2},\ldots,\tfrac{1}{d}\Pi_{d^{2}}\right\}.$ (4.5)
After rescaling if needed, all the entries of the $\mathbf{v}_{i}$ lie in $F^{\rm vec}\!\left(S\right)$, hence they lie in $F^{\rm proj}\!\left(S\right)$, so they lie in $F^{\rm Gproj}\!\left(S\right)$. The entries of the $\Pi_{i}$ also lie in $F^{\rm proj}\!\left(S\right)$, hence $F^{\rm Gproj}\!\left(S\right)$. Also,
$\sigma(\Pi_{i})=\sigma(\mathbf{v}_{i}\sigma_{c}(\mathbf{v}_{i})^{\top})=\sigma(\mathbf{v}_{i})\sigma(\sigma_{c}(\mathbf{v}_{i})^{\top})=\sigma(\mathbf{v}_{i})\sigma_{c}(\sigma(\mathbf{v}_{i})^{\top})$ (4.6)
is the Hermitian projection onto $\sigma(\mathbf{v}_{i})$, so
${\mathbf{\Pi}}(\sigma(S))=\left\{\tfrac{1}{d}\sigma(\Pi_{1}),\tfrac{1}{d}\sigma(\Pi_{2}),\ldots,\tfrac{1}{d}\sigma(\Pi_{d^{2}})\right\}.$ (4.7)
Moreover,
$\operatorname{Tr}(\sigma(\Pi_{i})\sigma(\Pi_{j}))=\sigma(\operatorname{Tr}(\Pi_{i}\Pi_{j}))=\sigma\!\left(\tfrac{\delta_{ij}d+1}{d+1}\right)=\tfrac{\delta_{ij}d+1}{d+1},$ (4.8)
so ${\mathbf{\Pi}}(\sigma(S))$ is a SIC. Finally, $F^{\rm Gproj}\!\left(\sigma(S)\right)\subseteq F^{\rm Gproj}\!\left(S\right)$, because the entries of $\Pi_{i}$ lie in $F^{\rm Gproj}\!\left(S\right)$, and $F^{\rm Gproj}\!\left(S\right)$ is a Galois field fixed by $\sigma$. It follows by symmetry that $F^{\rm Gproj}\!\left(S\right)=F^{\rm Gproj}\!\left(\sigma^{-1}(\sigma(S))\right)\subseteq F^{\rm Gproj}\!\left(\sigma(S)\right)$ and thus $F^{\rm Gproj}\!\left(\sigma(S)\right)=F^{\rm Gproj}\!\left(S\right)$. ∎
4.3. Galois multiplets and super-equivalence classes of Weyl–Heisenberg SICs
We now assemble $\operatorname{PEC}(d)$-orbits of Weyl–Heisenberg SICs in dimension $d$ into multiplets under the Galois action by the subgroup $C(\sigma_{c}\!\mid\!F^{\rm eproj}\!\left(S\right))$ of $\operatorname{Gal}(F^{\rm eproj}\!\left(S\right)/\mathbb{Q})$ that commutes with complex conjugation. The coarser equivalence classes obtained under the combined geometric and Galois actions (i.e., unions of the orbits in a multiplet) will called Galois super-equivalence classes.
The conjecture of Appleby, Flammia, McConnell, and Yard [10, Conj. 2] implies that many of the fields we have been studying are equivalent Galois extensions of $\mathbb{Q}$.
Lemma 4.9.
Assume 1.1 holds. Then in every dimension $d\geq 4$, for each Weyl–Heisenberg line-SIC $S$, the field $F^{\rm vec}\!\left(S\right)$ is a Galois extension of $\mathbb{Q}$, and
$F^{\rm vec}\!\left(S\right)=F^{\rm proj}\!\left(S\right)=F^{\rm eproj}\!\left(S\right)=F^{\rm Gproj}\!\left(S\right).$ (4.9)
Proof.
By assertion (2) of 1.1 $F^{\rm vec}\!\left(S\right)$ is a Galois extension of $\mathbb{Q}$. Consequently, $F^{\rm vec}\!\left(S\right)$ is preserved by complex conjugation, so $F^{\rm vec}\!\left(S\right)=F^{\rm proj}\!\left(S\right)$ holds by 3.5(1).
By assertion (2) of 1.1 $F^{\rm vec}\!\left(S\right)$ always contains the ray class field $H_{d^{\prime}\infty_{1}\infty_{2}}$ of $\mathbb{Q}(\Delta_{d})$. This class field must contain the class field of $\mathbb{Q}$ with conductor $d^{\prime}\infty$, which is $\mathbb{Q}(\zeta_{d^{\prime}})=\mathbb{Q}(\xi_{d})$. Therefore, $F^{\rm vec}\!\left(S\right)$ contains $F^{\rm proj}\!\left(S\right)(\xi_{d})=F^{\rm eproj}\!\left(S\right)$. Since $F^{\rm vec}\!\left(S\right)\subseteq F^{\rm proj}\!\left(S\right)\subseteq F^{\rm eproj}\!\left(S\right)$, the equality of those three fields follows. So $F^{\rm eproj}\!\left(S\right)$ is Galois and thus also equal to $F^{\rm Gproj}\!\left(S\right)$. ∎
We extract the following conjecture, of which all but the claim about $F^{\rm trip}\!\left(S\right)$ being Galois is already a consequence of 1.1.
Conjecture 4.10.
Fix a positive integer $d\neq 3$. For any $d$-dimensional Weyl–Heisenberg line SIC $S$, its $d^{2}$ vectors are always projective-algebraic. The associated fields $F^{\rm trip}\!\left(S\right)$ and $F^{\rm proj}\!\left(S\right)$ (and thus also $F^{\rm eproj}\!\left(S\right)$) are Galois extensions of $\mathbb{Q}$.
The evidence for a Galois action on Weyl–Heisenberg SICs was extensively detailed by Appleby, Yadsan-Appleby, and Zauner [4] in 2013.
The group $\operatorname{Gal}(\overline{\mathbb{Q}})/\mathbb{Q})$ acts on the elements of Weyl–Heisenberg group $\operatorname{WH}(d)$ coordinatewise on its matrices. Any automorphism $\sigma:\overline{\mathbb{Q}}\to\overline{\mathbb{Q}}$, sending a matrix $H\mapsto\sigma(H)$, acts as a homomorphism $\sigma(H_{1}H_{2})=\sigma(H_{1})\sigma(H_{2})$. This Galois action fixes the group $\operatorname{WH}(d)$ as a set, with any automorphism permuting its elements, since every Galois conjugate $\sigma(H)$ of a group element $H$ is also an element of $\operatorname{WH}(d)$.
By 3.8(2), $F^{\rm eproj}\!\left(S\right)$ is an invariant of the $\operatorname{PEC}(d)$ orbit of the Weyl–Heisenberg SIC $S$. If we assume 4.10, then $F^{\rm eproj}\!\left(S\right)$ is preserved by the Galois action by the subgroup $C(\sigma_{c}\!\mid\!F^{\rm eproj}\!\left(S\right))$ of $\operatorname{Gal}(F^{\rm eproj}\!\left(S\right)/\mathbb{Q})$. In other words, the extended projection field is invariant under both the geometric and the Galois actions.
We make the following definition.
Definition 4.11.
Assume 4.10 holds. For $d\geq 4$ the Galois multiplet of a $\operatorname{PEC}(d)$-equivalence class $[S]$ of Weyl–Heisenberg line-SICs is
$[[S]]=\{[\sigma(S)]:\sigma\in C(\sigma_{c}\!\mid\!F^{\rm eproj}\!\left(S\right))\},$ (4.10)
that is, the set of Galois equivalent orbits. We also define the equivalence relation $\approx_{GPEC}$, Galois super-equivalence, by
$S_{1}\approx_{GPEC}S_{2}\iff[[S_{1}]]=[[S_{2}]],$ (4.11)
and term the equivalence class of a SIC $S$ under $\approx_{GPEC}$ (that is, the union of the elements of $[[S]]$) the Galois super-equivalance class of $S$.
Remark 4.12.
Assuming 1.1 holds, this definition coincides with that of “Galois multiplet” in [9].
We summarize the properties noted above in the following proposition.
Proposition 4.13.
Assume 4.10 holds. Then, the set $\operatorname{WHSIC}(d)$ is divided into a finite number of Galois super-equivalence classes under the equivalence relation $\approx_{GPEC}$. Each Galois super-equivalence class is a finite union of (geometric) $\operatorname{PEC}(d)$-equivalence classes.
Galois multiplets are studied in Sections 7 and 8 using data and more refined conjectures. The multiplets in a fixed dimension often have different sizes.
5. A conjecture on counting geometric equivalence classes of Weyl–Heisenberg SICs
In this section, we formulate a conjecture relating the number of SICs in dimensions $d\geq 4$ to class numbers of certain real quadratic orders. We present numerical evidence.
5.1. Class groups and class monoids of orders
Let $K$ be a number field. An order in $K$ is any subring $\mathcal{O}$ including $1$ that is finitely additively generated and whose rational span $\mathbb{Q}\mathcal{O}=K$. There is a unique maximal order $\mathcal{O}_{K}$ of $K$ that contains all the others; $\mathcal{O}_{K}$ coincides with the set of algebraic integers in $K$. The maximal order is a Dedekind domain, while every non-maximal order is not a Dedekind domain. In particular, a non-maximal order will have non-invertible nonzero ideals. However, the set of invertible ideals of a non-maximal order behaves much like set of nonzero ideals in a Dedekind domain.
Recall that a fractional ideal of a (commutative) integral domain $\mathcal{O}$ is a finitely generated $\mathcal{O}$-module $\mathfrak{m}$ contained in its quotient field $K$. A fractional ideal $\mathfrak{a}$ is invertible if there exists a fractional ideal $\mathfrak{b}$ with $\mathfrak{a}\mathfrak{b}=\mathcal{O}$. Dedekind domains are characterized by the property that all nonzero fractional ideals are invertible.
The class group of an order, defined in terms of invertible fractional ideals, generalizes the usual notion of the class group of (the maximal order of) a number field.
Definition 5.1.
The class group (also called the Picard group or ring class group) of an order $\mathcal{O}$ is
$\operatorname{Cl}(\mathcal{O}):=\frac{\{\mbox{invertible fractional ideals of }\mathcal{O}\}}{\{\mbox{principal fractional ideals of }\mathcal{O}\}}.$ (5.1)
If one instead considers all nonzero ideals, one obtains a monoid (that is, a semigroup with identity) rather than a group.
Definition 5.2.
The class monoid of an order $\mathcal{O}$ is
$\operatorname{Clm}(\mathcal{O}):=\frac{\{\mbox{nonzero fractional ideals of }\mathcal{O}\}}{\{\mbox{principal fractional ideals of }\mathcal{O}\}}.$ (5.2)
For a more in-depth discussion of the nuances of orders, see [33] and [35, Ch. I, Sec. 12].
If $K$ is a quadratic field of discriminant $\Delta_{0}$, it has a unique suborder of discriminant $f^{2}\Delta_{0}$ for every positive integer $f$. This order is given explicitly as
$\mathcal{O}_{\!\Delta}:=\mathbb{Z}\!\left[\frac{\Delta+\sqrt{\Delta}}{2}\right]=\mathbb{Z}+\frac{\Delta+\sqrt{\Delta}}{2}\mathbb{Z}=\mathbb{Z}+f\frac{\Delta_{0}+\sqrt{\Delta_{0}}}{2}\mathbb{Z}.$ (5.3)
For the SIC problem, we are specifically interested in the case when $\Delta=\Delta_{d}=(d+1)(d-3)$ for some integer $d$. In this case, the class monoid $\operatorname{Clm}(\mathcal{O}_{\!\Delta_{d}})$ may be understood through bijections with other sets of number-theoretic interest.
Theorem 5.3.
Suppose $d\in\mathbb{Z}\setminus\{-1,3\}$, and let $\Delta=(d+1)(d-3)$. Write $\Delta_{d}=f^{2}\Delta_{0}$ for some $f\in\mathbb{N}$ and a fundamental discriminant $\Delta_{0}$. Then there are explicit, canonical bijections between the sets
  • (0)
    $\operatorname{Clm}(\mathcal{O}_{\Delta})$;
  • (1)
    the disjoint union $\bigsqcup_{f^{\prime}|f}\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})$;
  • (2)
    the set $\mathcal{Q}(\Delta)/\mathbf{GL}_{2}(\mathbb{Z})$, where $\mathcal{Q}(\Delta)$ is the set of binary quadratic forms of discriminant $\Delta$, and ${A}=\left(\begin{smallmatrix}s&t\\ u&v\end{smallmatrix}\right)\in\mathbf{GL}_{2}(\mathbb{Z})$ acts on the right by “twisted precomposition” $Q^{{A}}(x,y)=(\det{{A}})Q(sx+ty,ux+vy)$; and
  • (3)
    the set $\mathcal{C}_{d-1}=\{\{{R}^{-1}{A}{R}:{R}\in\mathbf{GL}_{2}(\mathbb{Z})\}:{A}\in\mathbf{SL}_{2}(\mathbb{Z}),\operatorname{Tr}({A})=d-1\}$ of $\mathbf{GL}_{2}(\mathbb{Z})$-conjugacy class of elements of $\mathbf{SL}_{2}(\mathbb{Z})$ of trace $d-1$.
Proof.
The bijection between (0) and (1) holds for quadratic orders in general. Specicially, by [33, Prop. A.12 and Prop. A.4], $\operatorname{Clm}(\mathcal{O}_{\Delta})=\operatorname{Clm}^{\flat}(\mathcal{O}_{\Delta})\cong\bigsqcup_{f^{\prime}|f}\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})$.
The bijection between (0) and (2) also holds for quadratic orders in general (and with appropriate modifications, nondegenerate quadratic rings) and is known as twisted Gauss composition. This is proven as [47, Thm. 1.2]; for an introduction to Gauss composition, see [16, Ch. 5].
We now exhibit a map $\phi:\mathcal{Q}(\Delta)\to\mathbf{GL}_{2}(\mathbb{Z})$ that induces a bijection between (2) and (3). For a quadratic form $Q(x,y)=ax^{2}+bxy+cy^{2}$ of discriminant $\Delta=(d+1)(d-3)$, define $\phi(Q):=\left(\begin{smallmatrix}\frac{1}{2}(b+d-1)&-a\\ c&\frac{1}{2}(-b+d-1)\end{smallmatrix}\right)$. Note that $\det(\phi(Q))=\frac{1}{4}(-b^{2}+(d-1)^{2})+ac=\frac{1}{4}(-\Delta+(d-1)^{2})=1$ and $\operatorname{Tr}(\phi(Q))=d-1$. It is straightforward to check that $\phi(Q^{R})={R}^{-1}\phi(Q){R}$ for ${R}\in\mathbf{GL}_{2}(\mathbb{Z})$. Thus, $\phi$ defines a map $\overline{\phi}:\mathcal{Q}(\Delta)/\mathbf{GL}_{2}(\mathbb{Z})\to\mathcal{C}_{d-1}$. In the other direction, for ${A}=\left(\begin{smallmatrix}s&t\\ u&v\end{smallmatrix}\right)\in SL_{2}(\mathbb{Z})$, define the quadratic form $(\psi({A}))(x,y):=-tx^{2}+(s-v)xy+uy^{2}$. It is straightforward to check that $\operatorname{disc}(\psi({A}))=\Delta$ and $\psi({R}^{-1}{A}{R})=\psi({A})^{R}$, so $\psi$ defines a map $\overline{\psi}:\mathcal{C}_{d-1}\to\mathcal{Q}(\Delta)/\mathbf{GL}_{2}(\mathbb{Z})$. A direct calculation shows $\overline{\phi}\circ\overline{\psi}=\operatorname{id}$ and $\overline{\psi}\circ\overline{\phi}=\operatorname{id}$. ∎
5.2. Statement of the Geometric Count Conjecture
The following conjecture counts the number of geometric equivalence classes of Weyl–Heisenberg SICs in dimension $d$.
Conjecture 5.4 (Geometric Count Conjecture).
Fix a positive integer $d\neq 3$, and let $\Delta=(d+1)(d-3)$ and $\mathcal{O}_{\Delta}$ be the quadratic order of discriminant $\Delta$. Then, the number of geometric equivalence classes $[S]$ of Weyl–Heisenberg covariant SICs is
$\left|{\rm WHSIC}(d)/\operatorname{PEC}(d)\right|=\left|\operatorname{Clm}(\mathcal{O}_{\Delta})\right|.$ (5.4)
By 5.3, the quantity $\left|\operatorname{Clm}(\mathcal{O}_{\Delta})\right|$ is equal to
  • (1)
    the sum of (ring) class numbers $\sum_{f^{\prime}|f}h_{(f^{\prime})^{2}\Delta_{0}}$, where $h_{(f^{\prime})^{2}\Delta_{0}}=\left|\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})\right|$;
  • (2)
    $\left|\mathcal{Q}(\Delta)/\mathbf{GL}_{2}(\mathbb{Z})\right|$, the number of twisted $\mathbf{GL}_{2}(\mathbb{Z})$-classes of binary quadratic forms of discriminant $\Delta$;
  • (3)
    $\left|\mathcal{C}_{d-1}\right|$, the number of $\mathbf{GL}_{2}(\mathbb{Z})$-conjugacy class of elements of $\mathbf{SL}_{2}(\mathbb{Z})$ of trace $d-1$.
5.3. Verification of cases of the Geometric Count Conjecture
We have verified that the two sets in the Geometric Count conjecture have the same size for dimensions $d\leq 90$, $d\neq 3$ under the assumption that the published lists of Weyl–Heisenberg SICs in these dimensions are complete.
Scott [38] gives a putatively complete list of numerical SICs in all dimensions $d\leq 90$, building on earlier work of Scott and Grassl [39]. Scott and Grassl find SIC fiducials by numerically minimizing a frame potential—a function whose global minima are Weyl–Heisenberg SIC fiducials. By repeating their search with a large number of seed vectors, they establish a high degree of numerical confidence in their results but do not rigorously rule out the existence of other SICs.
The dimensions $d\in\{1,2\}$ are special cases—the only dimensions for which $\Delta<0$, or equivalently, the conjugacy classes in $\mathcal{C}_{d-1}$ are elliptic. For $d=1$, there is exactly one SIC ($\mathbb{C}^{1}$ as a line in $\mathbb{C}^{1}$), and it is (trivially) Weyl–Heisenberg covariant. The associated order $\mathcal{O}_{-4}=\mathbb{Z}[i]$ is a maximal order of class number one, so the conjecture is true. For $d=2$, there is again a unique SIC up to unitary transformations (a regular tetrahedron on the Riemann sphere $\mathbb{P}^{1}(\mathbb{C})$), and it may be chosen to be Weyl–Heisenberg covariant. The associated order $\mathcal{O}_{-3}=\mathbb{Z}[e^{2\pi i/3}]$ is a maximal order of class number one, so the conjecture is true.
5.4 omits the case $d=3$, where $\Delta=0$, where associated field $K_{\!\Delta}=\mathbb{Q}$ has $\mathcal{O}_{\!\Delta}=\operatorname{Z}$. In this case, there is a continuous one-parameter family of Weyl–Heisenberg covariant SICs [39]. For $d=3$ the classification (3) above is the set $\mathcal{C}_{2}$ of parabolic conjugacy classes of $\mathbf{SL}_{2}(\mathbb{Z})$, which is infinite.
For $4\leq d\leq 90$ we computed $\left|\operatorname{Clm}(\mathcal{O}_{\Delta})\right|$ using Magma [17] by computing the equivalent quantity $\left|\mathcal{Q}(\Delta)/\mathbf{GL}_{2}(\mathbb{Z})\right|$. Write $\Delta=(d+1)(d-3)=f^{2}\Delta_{0}$ for a fundamental discriminant $\Delta_{0}$. For each $f^{\prime}\mid f$, we computed the set of reduced primitive binary quadratic forms of discriminant $(f^{\prime})^{2}\Delta_{0}$ using the Magma function ReducedForms(), written by David Kohel. This function produces a list of $\mathbf{SL}_{2}(\mathbb{Z})$-classes of primitive binary quadratic forms in which a twisted $\mathbf{GL}_{2}(\mathbb{Z})$-class is always represented either by a single form $ax^{2}+bxy+cy^{2}$ with $a>0$ or a pair of forms $\{ax^{2}+bxy+cy^{2},-ax^{2}+bxy-cy^{2}\}$. Removing forms with $a<0$ produced a list containing one representative for each primitive twisted $\mathbf{GL}_{2}(\mathbb{Z})$-class. Multiplying by $f/f^{\prime}$ and joining the lists gave a list of representatives of all (not necessarily primitive) classes in $\mathcal{Q}(\Delta)/\mathbf{GL}_{2}(\mathbb{Z})$.
We observe that for each $d$ ($4\leq d\leq 90$) that the size of
$\operatorname{Clm}(\mathcal{O}_{\Delta})=\sum_{f^{\prime}|f}\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})$ (5.5)
is equal to the sum of the corresponding values in the “#” column of Scott’s table beginning on page 3 of [38], which give the number of elements of ${\rm WHSIC}(d)/\operatorname{PEC}(d)$. This is agreement is shown in Table 1. The table also includes the sizes of the ring class groups $\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})$ (equivalently, the number of twisted $\mathbf{GL}_{2}(\mathbb{Z})$-classes of primitive binary quadratic forms with discriminant $(f^{\prime})^{2}\Delta_{0}$).
$d$ $\Delta_{0}$ $f^{\prime}$ $h_{(f^{\prime})^{2}\Delta_{0}}$#\##
$1$ $-4$ $1$ $1$ $1$
$2$ $-3$ $1$ $1$ $1$
$3$ 00 $\ast$ $\ast$ $\infty$
$4$ $5$ $1$ $1$ $1$
$5$ $12$ $1$ $1$ $1$
$6$ $21$ $1$ $1$ $1$
$7$ $8$ $1,2$ $1,1$ $2$
$8$ $5$ $1,3$ $1,1$ $2$
$9$ $60$ $1$ $2$ $2$
$10$ $77$ $1$ $1$ $1$
$11$ $24$ $1,2$ $1,2$ $3$
$12$ $31$,3131,31 , 3 $1,1$ $2$
$13$ $140$ $1$ $2$ $2$
$14$ $165$ $1$ $2$ $2$
$15$ $21$,2,41241,2,41 , 2 , 4 $1,1,2$ $4$
$16$ $221$ $1$ $2$ $2$
$17$ $28$ $1,3$ $1,2$ $3$
$18$ $285$ $1$ $2$ $2$
$19$ $5$ $1,2,4,8$ $1,1,1,2$ $5$
$20$ $357$ $1$ $2$ $2$
$21$ $44$ $1,3$ $1,4$ $5$
$22$ $437$ $1$ $1$ $1$
$23$ $201$,2121,21 , 2 $2,4$ $6$
$24$ $21$ $1,5$ $1,2$ $3$
$25$ $572$ $1$ $2$ $2$
$26$ $69$ $1,3$ $1,3$ $4$
$27$ $681$,2121,21 , 2 $2,4$ $6$
$28$ $29$ $1,5$ $1,2$ $3$
$29$ $780$ $1$ $4$ $4$
$30$ $93$ $1,3$ $1,3$ $4$
$31$ $56$ $1,2,4$ $1,2,4$ $7$
$32$ $957$ $1$ $2$ $2$
$34$ $1020$ $1$ $4$ $4$
$34$ $1085$ $1$ $2$ $2$
$35$ $8$ $1,2,3,4,6,12$ $1,1,1,1,2,4$ $10$
$36$ $1221$ $1$ $4$ $4$
$37$ $1292$ $1$ $4$ $4$
$38$ $1365$ $1$ $4$ $4$
$39$ $40$ $1,2,3,6$ $2,2,2,4$ $10$
$40$ $1517$ $1$ $2$ $2$
$41$ $1596$ $1$ $8$ $8$
$42$ $1677$ $1$ $4$ $4$
$43$ $440$ $1,2$ $2,4$ $6$
$44$ $205$ $1,3$ $2,4$ $6$
$45$ $1932$ $1$ $4$ $4$
 $d$ $\Delta_{0}$ $f^{\prime}$ $h_{(f^{\prime})^{2}\Delta_{0}}$#\## $46$ $2021$ $1$ $3$ $3$ $47$ $33$ $1,2,4,8$ $1,1,2,4$ $8$ $48$ $5$ $1,3,7,21$ $1,1,1,4$ $7$ $49$ $92$ $1,5$ $1,6$ $7$ $50$ $2397$ $1$ $2$ $2$ $51$ $561$,2,41241,2,41 , 2 , 4 $2,4,8$ $14$ $52$ $53$ $1,7$ $1,3$ $4$ $53$ $21$,3,$5,1$ 3 , 5 , 15 $1,1,2,6$ $10$ $54$ $2805$ $1$ $4$ $4$ $55$ $728$ $1,2$ $2,4$ $6$ $56$ $3021$ $1$ $6$ $6$ $57$ $348$ $1,3$ $2,6$ $8$ $58$ $3245$ $1$ $4$ $4$ $59$ $840$ $1,2$ $4,8$ $12$ $60$ $3477$ $1$ $4$ $4$ $61$ $3596$ $1$ $6$ $6$ $62$ $413$ $1,3$ $1,4$ $5$ $63$ $60$ $1,2,4,8$ $2,2,4,8$ $16$ $64$ $3965$ $1$ $4$ $4$ $65$ $4092$ $1$ $8$ $8$ $66$ $469$ $1,3$ $3,6$ $9$ $67$ $71$,2,4,8,161248161,2,4,8,161 , 2 , 4 , 8 , 16 $1,1,1,2,4$ $9$ $68$ $4485$ $1$ $4$ $4$ $69$ $4620$ $1$ $8$ $8$ $70$ $4757$ $1$ $5$ $5$ $71$ $361$,2,3,612361,2,3,61 , 2 , 3 , 6 $2,4,4,8$ $18$ $72$ $5037$ $1$ $4$ $4$ $73$ $5180$ $1$ $4$ $4$ $74$ $213$ $1,5$ $1,6$ $7$ $75$ $521$,2,3,612361,2,3,61 , 2 , 3 , 6 $1,2,4,8$ $15$ $76$ $5621$ $1$ $6$ $6$ $77$ $5772$ $1$ $8$ $8$ $78$ $237$ $1,5$ $1,6$ $7$ $79$ $380$ $1,2,4$ $2,4,8$ $14$ $80$ $77$ $1,3,9$ $1,2,6$ $9$ $81$ $6396$ $1$ $12$ $12$ $82$ $6557$ $1$ $3$ $3$ $83$ $051$,2,4,812481,2,4,81 , 2 , 4 , 8 $2,2,4,8$ $16$ $84$ $85$ $1,3,9$ $2,2,6$ $10$ $85$ $7052$ $1$ $4$ $4$ $86$ $7221$ $1$ $10$ $10$ $87$ $8481$,2121,21 , 2 $4,8$ $12$ $88$ $7565$ $1$ $4$ $4$ $89$ $860$ $1,3$ $2,8$ $10$ $90$ $7917$ $1$ $4$ $4$
Table 1. Comparison of ring class numbers and counts of SICs for $d\leq 90$. Here $(d+1)(d-3)=f^{2}\Delta_{0}$ for a fund. disc. $\Delta_{0}$, $f^{\prime}$ runs over divisors of $f$, $h_{(f^{\prime})^{2}\Delta_{0}}=\left|\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})\right|$, and ##\## counts W–H SICs of Scott [38] in dim. $d$.
6. Theorems on class monoids of real quadratic fields
In this section, we prove rigorous results about the behavior of $\operatorname{Clm}(\mathcal{O})$ for general orders. These results have conditional consequences for SICs via 5.4.
6.1. Growth of the number of SICs in dimension $d$
Going beyond Zauner’s conjecture, one would like to know not only that SICs exist in every dimension $d$, but also how many there are. The connection between SICs and class numbers gives us a precise prediction for the number of SICs up to equivalence (excluding any SICs that are not Weyl–Heisenberg covariant, to which our conjectures are agnostic). In this section, we use known results on the growth of class numbers to derive conditional results on the number of SICs as $d\to\infty$.
The first growth result characterizes those dimensions in which our conjectures predict a unique Weyl–Heisenberg SIC, using a theorem of Byeon, Kim, and Lee [13].
Proposition 6.1.
One has $\left|\operatorname{Clm}(\mathcal{O}_{\!\Delta_{d}})\right|>1$ for all $d>22$. The values of $d\neq 3$ with $\left|\operatorname{Clm}(\mathcal{O}_{\!\Delta_{d}})\right|=1$ are $d\in\{1,2,4,5,6,10,22\}$.
Proof.
Set $\Delta=\Delta_{d}=(d+1)(d-3)$ and $s(d):=\left|\operatorname{Clm}(\mathcal{O}_{\!\Delta})\right|$; we are trying to determine when $s(d)=1$. By 5.4,
$s(d)=\sum_{f^{\prime}|f}h_{(f^{\prime})^{2}\Delta_{0}}.$ (6.1)
where $\Delta=f^{2}\Delta_{0}$ and $\Delta_{0}$ is fundamental.
If $f>1$, then $s(d)>1$.
If $f=1$, then $\Delta=\Delta_{0}$ is fundamental. Write $\Delta=(d-1)^{2}-4=n^{2}-4$, where $n:=d-1$. Byeon, Kim, and Lee solved the class number 1 problem for real quadratic fields with fundamental discriminant of the form $n^{2}-4$, a result previously known as Mollin’s conjecture [13, Thm 1.2]; whenever $n>21$ (so $d>22$), $s(d)=h_{\Delta}>1$. The list of dimensions where $s(d)=1$ follows from a finite calculation of class numbers. ∎
Corollary 6.2.
Assuming 5.4, there is more than one $\operatorname{PEC}(d)$-class of Weyl–Heisenberg SIC in dimension $d$ for all $d>22$. The dimensions $d$ with a unique $\operatorname{PEC}(d)$-class of Weyl–Heisenberg SIC are $d\in\{1,2,4,5,6,10,22\}$.
Proof.
The corollary follows from 6.1. ∎
The second growth result is an asymptotic formula for the size of the ray class monoid of $\mathcal{O}_{\!\Delta_{d}}$, which is conjectured to be the equal to the number of extended Clifford orbits of SICs in dimension $d$. The implied constant in this formula is ineffective because it ultimately relies on a Siegel bound involving a possible Siegel zero of a Dirichlet $L$-function. We are grateful to Steven Flammia for independently finding numerical evidence for this asymptotic [23].
Theorem 6.3.
As $d\to\infty$, and $\Delta_{d}=(d+1)(d-3)$ the size of the class monoid of $\mathcal{O}_{\!\Delta_{d}}$ obeys the asymptotic formula
$\log\left|\operatorname{Clm}(\mathcal{O}_{\!\Delta_{d}})\right|=\log d+o(\log d).$ (6.2)
Proof.
The proof uses the Brauer–Siegel theorem, which is an (ineffective) estimate on the growth of the product of the class number and the regulator. The regulator is always negligibly small in our regime. The subtlety of the proof is entirely in accounting for the class numbers of non-maximal orders, which are not included in the Brauer–Siegel theorem.
As in the proof of 6.1, set $s(d):=\left|\operatorname{Clm}(\mathcal{O}_{\Delta_{d}})\right|$. By 5.3,
$s(d)=\sum_{f^{\prime}|f}h_{(f^{\prime})^{2}\Delta_{0}}$ (6.3)
where $\Delta=\Delta_{d}=(d+1)(d-3)=f^{2}\Delta_{0}$ and $\Delta_{0}$ is fundamental. Also, $h_{(f^{\prime})^{2}\Delta_{0}}\leq h_{\Delta}$ for all $f^{\prime}\mid f$. It follows that
$h_{\Delta}\leq s(d)\leq\sigma_{0}(f)h_{\Delta},$ (6.4)
where $\sigma_{0}(f)$ is the number of positive integer divisors of $f$. Moreover,
$\log\sigma_{0}(f)\ll\frac{\log f}{\log\log f}\ll\frac{\log d}{\log\log d}$ (6.5)
by [34, Thm. 2.11]. Thus,
$\log s(d)=\log h_{\Delta}+O\!\left(\frac{\log d}{\log\log d}\right).$ (6.6)
We then use the formula for the class number of an order given as [35, Thm. 12.12]:
$h_{\Delta}=\frac{h_{\!\Delta_{0}}}{[\mathcal{O}_{\!\Delta_{0}}^{\times}:\mathcal{O}_{\!\Delta}^{\times}]}\frac{\left|\left(\mathcal{O}_{\!\Delta_{0}}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\right|}{\left|\left(\mathcal{O}_{\Delta}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\right|}.$ (6.7)
We will estimate the size of each factor in this formula. For any real nonsquare discriminant $\Delta^{\prime}$, let $\varepsilon_{\!\Delta^{\prime}}$ be the smallest unit of $\mathcal{O}_{\!\Delta^{\prime}}$ such that $\varepsilon_{\!\Delta^{\prime}}>1$, and let $\varepsilon_{\!\Delta^{\prime}}^{+}$ be the smallest such totally positive unit. The Brauer–Siegel theorem [12]—specifically, Siegel’s part, concerning real quadratic fields [40]— gives the asymptotic formula
$\frac{\log\!\left(h_{\!\Delta_{0}}\operatorname{Reg}(\mathbb{Q}(\Delta_{0}))\right)}{\log{\Delta_{0}}}\sim 1\mbox{ as }\Delta_{0}\to\infty.$ (6.8)
Moreover,
$[\mathcal{O}_{\!\Delta_{0}}^{\times}:\mathcal{O}_{\!\Delta}^{\times}]=\frac{\log\varepsilon_{\!\Delta}}{\log\varepsilon_{\!\Delta_{0}}}=\frac{\log\varepsilon_{\!\Delta}}{\log\operatorname{Reg}(\mathbb{Q}(\Delta_{0}))}.$ (6.9)
Thus, (6.8) may be rewritten as
$\log\!\left(\frac{h_{\!\Delta_{0}}}{[\mathcal{O}_{\!\Delta_{0}}^{\times}:\mathcal{O}_{\!\Delta}^{\times}]}\log\varepsilon_{\Delta}\right)\sim\log{\Delta_{0}}\mbox{ as }\Delta_{0}\to\infty.$ (6.10)
Fix some $\epsilon>0$. Then, there exists some natural number $N$ such that, whenever $\Delta_{0}>N$,
$(1-\epsilon)\log\Delta_{0}-\log\log\varepsilon_{\Delta}<\log\!\left(\frac{h_{\!\Delta_{0}}}{[\mathcal{O}_{\!\Delta_{0}}^{\times}:\mathcal{O}_{\!\Delta}^{\times}]}\right)<(1+\epsilon)\log\Delta_{0}-\log\log\varepsilon_{\Delta}.$ (6.11)
For $\Delta_{0}\leq N$, there is some constant $B$ such that
$B-\log\log\varepsilon_{\Delta}\leq\log\!\left(\frac{h_{\!\Delta_{0}}}{[\mathcal{O}_{\!\Delta_{0}}^{\times}:\mathcal{O}_{\!\Delta}^{\times}]}\right)\leq B-\log\log\varepsilon_{\Delta}.$ (6.12)
We may bound $\left|\left(\mathcal{O}_{\!\Delta_{0}}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\right|$ above by $f^{2}$ and below by
$\left|\left(\mathcal{O}_{\!\Delta_{0}}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\right|$ $=f^{2}\prod_{\mathfrak{p}|f}\left(1-\frac{1}{\operatorname{Nm}(\mathfrak{p})}\right)\geq f^{2}\prod_{p|f}\left(1-\frac{1}{p}\right)^{2}=\varphi(f)^{2},$ (6.13)
where the first product is over prime ideals of $\mathcal{O}_{\!\Delta_{0}}$ and the second is over rational primes. Since $\sqrt{\Delta}=f\Delta_{0}\equiv 0\,\left(\operatorname{mod}\,f\mathcal{O}_{\!\Delta_{0}}\right)$, the multiplicative group $\left(\mathcal{O}_{\Delta}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\cong\left(\mathbb{Z}/f\mathbb{Z}\right)^{\times}$, so $\left|\left(\mathcal{O}_{\Delta}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\right|=\varphi(f)$. Thus,
$\varphi(f)\leq\frac{\left|\left(\mathcal{O}_{\!\Delta_{0}}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\right|}{\left|\left(\mathcal{O}_{\Delta}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\right|}\leq\frac{f^{2}}{\varphi(f)}.$ (6.14)
Using (on both sides) the inequality $\varphi(f)>\frac{f}{\ell(f)}$ where $\ell(f)$ is a real-valued function satisfying $\ell(n)>1$ for all $n$ and $\ell(n)=e^{\gamma}\log\log n$ as $n\to\infty$ [34, Thm. 2.9], we have
$\frac{f}{\ell(f)}<\frac{\left|\left(\mathcal{O}_{\!\Delta_{0}}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\right|}{\left|\left(\mathcal{O}_{\Delta}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\right|}\leq f\ell(f),$ (6.15)
which after taking logarithms becomes
$\log f-\log\ell(f)<\log\!\left(\frac{\left|\left(\mathcal{O}_{\!\Delta_{0}}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\right|}{\left|\left(\mathcal{O}_{\Delta}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\right|}\right)<\log f+\log\ell(f).$ (6.16)
Using (6.7) and combining (6.11) and (6.16), for $\Delta_{0}>N$, we obtain
$\left|\log h_{\Delta}-\left(\log\sqrt{\Delta}_{0}+\log f\right)\right|<\epsilon\log\sqrt{\Delta}_{0}+\left|\log\log\varepsilon_{\Delta}\right|+\log\ell(f).$ (6.17)
On the left-hand side, $\log\sqrt{\Delta}_{0}+\log f=\log\sqrt{\Delta}\sim\log d$ as $d\to\infty$. On the right-hand side, as $d\to\infty$, $\log\sqrt{\Delta_{0}}<\log\sqrt{\Delta}\ll\log d$, while
$\log\log\varepsilon_{\Delta}\leq\log\log\varepsilon_{\Delta}^{+}=\log\log\frac{d-1+\sqrt{\Delta}}{2}\ll\log\log d,$ (6.18)
and $\log\ell(f)\ll\log\log\log d$. Thus,
$\left|\log h_{\Delta}-\log d\right|\ll\epsilon\log d\mbox{ as }d\to\infty.$ (6.19)
In the case when $\Delta_{0}\leq N$, combining (6.11) and (6.17), we instead obtain
$\left|\log h_{\Delta}-\left(\log\sqrt{\Delta}_{0}+\log f\right)\right|<B+\left|\log\log\varepsilon_{\Delta}\right|+\log\ell(f),$ (6.20)
which by a similar argument also implies
$\left|\log h_{\Delta}-\log d\right|\ll\epsilon\log d\mbox{ as }d\to\infty.$ (6.21)
Combining (6.19) and (6.21) with (6.6) proves the theorem in all cases. ∎
Corollary 6.4.
Assuming 5.4, as $d\to\infty$,
$\log\left|{\rm WHSIC}(d)/\operatorname{PEC}(d)\right|=\log d+o(\log d).$ (6.22)
Proof.
Follows from 6.3. ∎
7. Conjectures on counting Galois multiplets of Weyl–Heisenberg SICs
Recall that 4.10 implies that there are a finite number of well-defined Galois multiplets of SICs in dimension $d$, each consisting of a finite number of geometric equivalence classes; we will now refine this prediction to a precise count in every dimension. Our Multiplet Count Conjecture (7.1) predicts the number of multiplets in dimension $d$. The Order-to-Multiplet Conjecture (7.2) predicts the number of geometric equivalence classes in each multiplet.
7.1. Statement of the Multiplet Count Conjecture
[…]
Conjecture 7.1 (Multiplet Count Conjecture).
Fix a positive integer $d\neq 3$. Let $\Delta=\Delta_{d}=(d+1)(d-3)$ and $K=\mathbb{Q}(\sqrt{\Delta})$. The number of Galois multiplets of Weyl-Heisenberg SICs in dimension $d$ equals the number of quadratic orders $\mathcal{O}^{\prime}$ with $\mathcal{O}_{\!\Delta}\subseteq\mathcal{O}^{\prime}\subseteq\mathcal{O}_{K}$. That is,
$\left|{\rm WHSIC}(d)/\approx_{GPEC}\right|=\left|\{\mathcal{O}^{\prime}:\,\mathcal{O}_{\!\Delta}\subseteq\mathcal{O}^{\prime}\subseteq\mathcal{O}_{K}\}\right|.$ (7.1)
(Note that, if we write $\Delta=f^{2}\Delta_{0}$ for a fundamental discriminant $\Delta_{0}$, with $f>0$, then
|$\left|\{\mathcal{O}^{\prime}:\,\mathcal{O}_{\!\Delta}\subseteq\mathcal{O}^{\prime}\subseteq\mathcal{O}_{K}\}\right|=\sigma_{0}(f),$ (7.2)
where $\sigma_{0}(f)$ denotes the number of positive divisors of $f$.)
The content of 7.1 is that the multiplets of Weyl–Heisenberg SICs are in one-to-one correspondence with orders between $\mathcal{O}$ and $\mathcal{O}_{K}$. The fact that the number of such orders is simply $\sigma_{0}(f)$ uses the fact that an order in a quadratic field is determined uniquely by its discriminant, which can be $(f^{\prime})^{2}\Delta_{0}$ for a fundamental discriminant $\Delta_{0}$ and a positive integer $f^{\prime}$. The order $\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}}$ contains $\mathcal{O}_{f^{2}\Delta_{0}}=\mathcal{O}_{\!\Delta}$ if and only if $f^{\prime}\mid f$. (These facts are not true for general number fields.)
7.2. Statement of the Extended Order-to-Multiplet Conjecture
We next formulate a conjecture positing a correspondence between orders and multiplets with certain properties. The number of geometric equivalence classes $[S]$ inside a Galois multiplet $[[S]]$ should be equal to the class number of the corresponding orders, and the fields $F^{\rm trip}\!\left(S\right)$ and $F^{\rm proj}\!\left(S\right)$ associated to different multiplets are related in a natural way. (Recall that we have already shown these fields to be are well-defined on $\operatorname{PEC}(d)$-equivalence classes, and they are well-defined on Galois super-equivalence classes under the assumption of 4.10.)
Conjecture 7.2 (Extended Order-to-Multiplet Conjecture).
Fix a positive integer $d\neq 3$, let $\Delta=\Delta_{d}=(d+1)(d-3)$, and write $\Delta=f^{2}\Delta_{0}$ for a fundamental discrimiant $\Delta_{0}$ and some positive integer $f$. Then there is a bijection $\mathcal{M}$,
{positive divisors of $\{\mbox{positive divisors of $f$}\}\xrightarrow{\mathcal{M}}\{[[S]]:S\mbox{ a Weyl--Heisenberg line-SIC in }\mathbb{C}^{d}\}$ (7.3)
having the following properties:
  • (1)
    The number of geometric equivalence classes $[S]$ in the multiplet $\mathcal{M}(f^{\prime})=[[S]]$ is $\left|\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})\right|$;
  • (2)
    If $f_{1}\mid f_{2}\mid f$, $[S_{1}]\in\mathcal{M}(f_{1})$, and $[S_{2}]\in\mathcal{M}(f_{2})$, then $F^{\rm trip}\!\left(S_{1}\right)\subseteq F^{\rm trip}\!\left(S_{2}\right)$ and $F^{\rm vec}\!\left(S_{1}\right)=F^{\rm proj}\!\left(S_{1}\right)\subseteq F^{\rm vec}\!\left(S_{2}\right)=F^{\rm proj}\!\left(S_{2}\right)$.
Part (1) of this conjecture is consistent with 5.4. The strengthening of this conjecture compared with 1.2 is that part (2) asks for additional compatibility under inclusion of the number fields $F^{\rm trip}\!\left(S\right)$ and $F^{\rm proj}\!\left(S\right)$ with divisibility of the associated conductors $f^{\prime}$ (i.e., reverse inclusion of the associated quadratic orders). Assuming 1.1, we have $F^{\rm vec}\!\left(S\right)=F^{\rm proj}\!\left(S\right)$.
This conjecture can be tested by study of numerical invariants. One may attempt to reverse-engineer the map $\mathcal{M}$ from numerical data.
  1. (1)
    One first verifies that the number of multiplets matches the number of orders (7.1).
  2. (2)
    One can compute all the class numbers of the orders $\mathcal{O}$, and one can numerically compute the number of $\operatorname{PEC}$-equivalence classes in each multiplet. These multisets of numbers must be identical for a map to exist.
  3. (3)
    If all the class numbers of the orders $\mathcal{O}^{\prime}$ are pairwise distinct, then the putative map $\mathcal{M}$ is uniquely specified. Otherwise, the map exists if and only if the number of $\operatorname{PEC}(d)$-orbits in the different Galois equivalence classes are a permutation of these class numbers.
  4. (4)
    The reverse-inclusion requirement of the triple product and projection SIC-fields as the orders vary provides a further test of the conjecture.
The inclusion relations between the different fields gives further constraints on existence of such a map. One can study weaker invariants, the degrees of the fields, which are easier to compute. A necessary condition for inclusion of fields is a divisibility relation between their field degrees over $\mathbb{Q}$.
Ambiguity may remain. A case where there would be unresolvable ambiguity using numerical invariants is any case where there are two distinct multiplet $[[S_{1}]]$ and $[[S_{2}]]$ having
  1. (1)
    identical sizes $\left|[[S_{1}]]\right|=\left|[[S_{2}]]\right|$,
  2. (2)
    identical triple product fields $F^{\rm trip}\!\left(S_{1}\right)=F^{\rm trip}\!\left(S_{2}\right)$,
  3. (3)
    identical projection fields $F^{\rm proj}\!\left(S_{1}\right)=F^{\rm proj}\!\left(S_{2}\right)$.
Such an ambiguity can occur, as discovered by Markus Grassl for $d=47$ and $d=67$; see Section 9.3.
In Section 8, we present a refined 8.4 about the fields $F^{\rm trip}\!\left(S\right)$ and $F^{\rm proj}\!\left(S\right)$, identifying them with certain ray class fields of the orders.
7.3. Verification of cases of the Multiplet Count Conjecture
We have verified that the two sets in the Multiplet Count Conjecture (7.1) have the same size for dimensions $d\leq 90$ ($d\neq 3$) under the assumption that the published lists of Weyl–Heisenberg SICs in these dimensions are complete.
Scott [38] gives a putatively complete list of numerical SICs in all dimensions $d\leq 90$, building on earlier work of Scott and Grassl [39]. Scott and Grassl find SIC fiducials by numerically minimizing a frame potential—a function whose global minima are SICs if SICs exist. Their search is constrained to Weyl–Heisenberg covariant sets of vectors. By repeating their search with a large number of seed vectors, they establish a high degree of numerical confidence in their results but do not rigorously rule out the existence of other SICs (Weyl–Heisenberg covaraint or otherwise).
7.4. Evidence for the Extended Order-to-Multiplet Conjecture
Some evidence for the Extended Order-to-Multiplet Conjecture (7.2) is given in Table 3 at the end of the paper, in dimensions $d\leq 15$. In the dimensions where there is more than one multiplet ($d\in\{7,8,11,12,15\}$), a conjectural correspondence between divisors $f^{\prime}\mid f$ and Galois multiplets (written using Scott and Grassl’s orbit labels) has been given. Under this correspondence, note that for divisors $f_{1},f_{2}$ of $f$ such that $f_{1}\mid f_{2}$, the corresponding degrees $\left.[F^{\rm proj}\!\left(S_{1}\right):K]\middle|[F^{\rm proj}\!\left(S_{2}\right):K]\right.$, and similarly for the triple product fields in the cases that have been computed. The table also supports the stronger 8.5, which implies 7.2.
8. Conjectures on precise fields of definition of Weyl–Heisenberg SICs
In this section, we give a refinement of both 5.4 and of the “class field hypothesis” of Appleby, Flammia, McConnell, and Yard [10]. The refined conjecture predicts the triple product field and the projector field associated to each multiplet.
8.1. Ray class fields of orders
The fields attached to a Weyl–Heisenberg SIC are believed to be abelian extensions of $K=\mathbb{Q}(\sqrt{\Delta_{d}})$. Class field theory gives an abstract description of abelian extensions of $K$, viewed with respect to the ideals of the maximal order $\mathcal{O}_{K}$. 7.2 implies that a Weyl–Heisenberg SIC can unambiguously be assigned an order $\mathcal{O}$ of $K$ with $\mathbb{Z}[\varepsilon_{d}]\subseteq\mathcal{O}\subseteq\mathcal{O}_{K}$. There is a long history of work which supplies a notion of class field theory attached to orders of number fields. In particular, the ring class field of an order $\mathcal{O}$ has Galois group over $K$ isomorphic to $\operatorname{Cl}(\mathcal{O})$ and generalizes the Hilbert class field, which is the maximal unramified extension of $K$ and has class group $\operatorname{Cl}(\mathcal{O}_{K})$.
In [33], the authors modify the standard constructions of class field theory to describe a distinguished collection of ray class fields of orders $H^{\mathcal{O}}_{\mathfrak{m},\Sigma}$, each specified by a level datum $(\mathcal{O};\mathfrak{m},\Sigma)$, where $\mathcal{O}$ is a (possibly non-maximal) order of $K$, $\mathfrak{m}$ is an integral $\mathcal{O}$-ideal, and $\Sigma$ is any subset of real places of $K$. In the case that $K$ is a real quadratic field, there are two infinite places, so $\Sigma$ is one of $\emptyset$, $\{\infty_{1}\}$, $\{\infty_{2}\}$, or $\{\infty_{1},\infty_{2}\}$. For the datum $(\mathcal{O};\mathcal{O},\emptyset)$, the associated field $H^{\mathcal{O}}_{\mathcal{O},\emptyset}$ is the ring class field associated to the order. For the datum $(\mathcal{O}_{K};\mathfrak{m},\Sigma)$, the associated field $H^{\mathcal{O}_{K}}_{\mathfrak{m},\Sigma}$ is the standard (Takagi) ray class field $H_{\mathfrak{m},\Sigma}$. For any fixed order $\mathcal{O}$, any abelian extension of $K$ as a subfield of some $H^{\mathcal{O}}_{\mathfrak{m},\Sigma}$.
We recall [33, Thm. 1.1]:
Theorem 8.1.
Let $K$ be a number field and $(\mathcal{O};\mathfrak{m},\Sigma)$ a level datum for $K$. There exists a unique abelian Galois extension $H_{\mathfrak{m},\Sigma}^{\mathcal{O}}/K$ with the property that a prime ideal $\mathfrak{p}$ of $\mathcal{O}_{K}$ that is coprime to the quotient ideal $\left(\mathfrak{m}:\mathcal{O}_{K}\right)$ splits completely in $H_{\mathfrak{m},\Sigma}^{\mathcal{O}}/K$ if and only if $\mathfrak{p}\cap\mathcal{O}=\pi\mathcal{O}$, a principal prime $\mathcal{O}$-ideal having $\pi\in\mathcal{O}$ with $\pi\equiv 1\,\left(\operatorname{mod}\,\mathfrak{m}\right)$ and $\rho(\pi)>0$ for $\rho\in\Sigma$.
To understand 8.1, recall that for any integral ideal $\mathfrak{m}$ of $\mathcal{O}$, the quotient ideal $\left(\mathfrak{m}:\mathcal{O}_{K}\right)$ is the largest ideal of $\mathcal{O}_{K}$ contained in $\mathfrak{m}$. An important invariant of an order $\mathcal{O}$ of an algebraic number field $K$ is its (absolute) conductor $\mathfrak{f}=\mathfrak{f}(\mathcal{O})=\left(\mathcal{O}:\mathcal{O}_{K}\right)$, which is the (set-theoretically) largest integral ideal $\mathfrak{f}\subseteq\mathcal{O}$ of $\mathcal{O}$ that is also an ideal of the maximal order $\mathcal{O}_{K}$. The conductor ideal $\mathfrak{f}(\mathcal{O})$ encodes information on all the non-invertible ideals in the order $\mathcal{O}$. The quotient ideal $\left(\mathfrak{m}:\mathcal{O}_{K}\right)$ always satisfies the inclusions (as $\mathcal{O}_{K}$-ideals)
$\mathfrak{f}(\mathcal{O})\mathfrak{m}\subseteq\left(\mathfrak{m}:\mathcal{O}_{K}\right)\subseteq\mathfrak{f}(\mathcal{O})\cap\mathfrak{m}\mathcal{O}_{K};$ (8.1)
in particular, $\left(\mathfrak{m}:\mathcal{O}_{K}\right)\subseteq\mathfrak{f}(\mathcal{O})$. The three ideals in (8.1) have the same prime divisors: A prime ideal $\mathfrak{p}$ of $\mathcal{O}_{K}$ such that $\mathfrak{p}\supseteq\mathfrak{f}(\mathcal{O})\mathfrak{m}=\mathfrak{f}(\mathcal{O})\mathfrak{m}\mathcal{O}_{K}$ satisfies either $\mathfrak{p}\supseteq\mathfrak{f}(\mathcal{O})$ or $\mathfrak{p}\supseteq\mathfrak{m}\mathcal{O}_{K}$, and thus, $\mathfrak{p}\supseteq\mathfrak{f}(\mathcal{O})\cap\mathfrak{m}\mathcal{O}_{K}$.
There is a Galois correspondence between ray class groups of an order and ray class fields of an order. The ray class group of an order $\operatorname{Cl}_{\mathfrak{m},\Sigma}(\mathcal{O})$ for level datum $(\mathcal{O};\mathfrak{m},\Sigma)$ is defined as a quotient of certain groups of invertible fractional ideals of the order $\mathcal{O}$. The ray class field $H_{\mathfrak{m},\Sigma}^{\mathcal{O}}$ of an order $\mathcal{O}$ is associated to an appropriate $\operatorname{Cl}_{\mathfrak{m},\Sigma}(\mathcal{O})$ in such a way that $\operatorname{Gal}(H_{\mathfrak{m},\Sigma}^{\mathcal{O}}/K)\simeq\operatorname{Cl}_{\mathfrak{m},\Sigma}(\mathcal{O})$ as abelian groups. Moreover, if $(\mathcal{O};\mathfrak{m},\Sigma)$ and $(\mathcal{O}^{\prime};\mathfrak{m}^{\prime},\Sigma^{\prime})$ are level data satisfying $\mathcal{O}\subseteq\mathcal{O}^{\prime}$, $\mathfrak{m}\mathcal{O}^{\prime}\subseteq\mathfrak{m}^{\prime}$, and $\Sigma\supseteq\Sigma^{\prime}$, then there is a natural quotient map $\operatorname{Cl}_{\mathfrak{m},\Sigma}(\mathcal{O})\twoheadrightarrow\operatorname{Cl}_{\mathfrak{m}^{\prime},\Sigma^{\prime}}(\mathcal{O}^{\prime})$ and a corresponding natural field inclusion $H_{\mathfrak{m}^{\prime},\Sigma^{\prime}}^{\mathcal{O}^{\prime}}\subseteq H_{\mathfrak{m},\Sigma}^{\mathcal{O}}$.
We recall [33, Thm. 1.3]:
Theorem 8.2.
For an order $\mathcal{O}$ in a number field $K$ and any level datum $(\mathcal{O};\mathfrak{m},\Sigma)$ with associated ray class field $H_{0}:=H_{\mathfrak{m},\Sigma}^{\mathcal{O}}$, there is an isomorphism $\operatorname{Art}_{\mathcal{O}}:\operatorname{Cl}_{\mathfrak{m},\Sigma}(\mathcal{O})\to\operatorname{Gal}(H_{0}/K)$, uniquely determined by its behavior on prime ideals $\mathfrak{p}$ of $\mathcal{O}$ that are coprime to $\mathfrak{f}(\mathcal{O})\cap\mathfrak{m}$, having the property that
$\operatorname{Art}_{\mathcal{O}}([\mathfrak{p}])(\alpha)\equiv\alpha^{q}\,\left(\operatorname{mod}\,\mathfrak{P}\right),$ (8.2)
where $\mathfrak{P}$ is any prime of $\mathcal{O}_{H_{0}}$ lying over $\mathfrak{p}\mathcal{O}_{K}$, and $q=p^{j}$ is the number of elements in the finite field $\mathcal{O}/\mathfrak{p}$. For any (not necessarily prime) ideal $\mathfrak{a}$ of $\mathcal{O}$ coprime to $\mathfrak{f}(\mathcal{O})\cap\mathfrak{m}$,
$\operatorname{Art}_{\mathcal{O}}([\mathfrak{a}])=\left.\operatorname{Art}([\mathfrak{a}\mathcal{O}_{K}])\right|_{H_{0}},$ (8.3)
where $\operatorname{Art}:\operatorname{Cl}_{\left(\mathfrak{m}:\mathcal{O}_{K}\right),\Sigma}(\mathcal{O}_{K})\to\operatorname{Gal}(H_{1}/K)$ is the usual Artin map in class field theory, with $H_{1}=H_{\left(\mathfrak{m}:\mathcal{O}_{K}\right),\Sigma}^{\mathcal{O}_{K}}$ being a (Takagi) ray class field for the maximal order $\mathcal{O}_{K}$, and $H_{0}\subseteq H_{1}$.
In 8.2, the set of prime ideals coprime to $\mathfrak{f}(\mathcal{O})\cap\mathfrak{m}$ includes all but finitely many of the prime ideals of $\mathcal{O}$. The Artin map is fully defined by (8.2), which specifies the map on prime ideals of $\mathcal{O}$ coprime to $\mathfrak{f}(\mathcal{O})\cap\mathfrak{m}$, because it turns out that any ray ideal class in $\operatorname{Cl}_{\mathfrak{m},\Sigma}(\mathcal{O})$ has a representative that is coprime to $\mathfrak{f}(\mathcal{O})\cap\mathfrak{m}$ (see [33, Lem. 4.12]), and ideals of $\mathcal{O}$ that are coprime to $\mathfrak{f}$ enjoy unique factorization into prime ideals (even though arbitrary nonzero ideals of $\mathcal{O}$ do not). When $\mathcal{O}=\mathcal{O}_{K}$, (8.2) specializes to the standard definition of the Artin map in the case of a Takagi ray class field.
For a real quadratic field $K$, the infinite places in $\Sigma$ relate to the notions of “wide” and “narrow” class fields. If $\mathfrak{m}=\mathcal{O}$, the (wide) ring class field is $H_{\mathcal{O},\emptyset}^{\mathcal{O}}=H_{\mathcal{O},\{\infty_{1}\}}^{\mathcal{O}}=H_{\mathcal{O},\{\infty_{2}\}}^{\mathcal{O}}$, and the narrow ring class field is $H_{\mathcal{O},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}}$. The narrow ring class field is either equal to the wide ring class field or is a degree $2$ extension thereof. More generally, there are four ray class fields associated to the ideal $\mathfrak{m}$ as well as a fifth subfield $H_{\mathfrak{m},+}^{\mathcal{O}}$ of $H_{\mathfrak{m},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}}$.
[…]
Each of the field inclusions in this diagram is either trivial or degree $2$. In the “nondegenerate” case when all have degree $2$, the Galois group $\operatorname{Gal}\!\left(H_{\mathfrak{m},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}}/H_{\mathfrak{m},\emptyset}^{\mathcal{O}}\right)\cong\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$. The fields $H_{\mathfrak{m},\emptyset}^{\mathcal{O}}$, $H_{\mathfrak{m},+}^{\mathcal{O}}$, and $H_{\mathfrak{m},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}}$ are all Galois over $\mathbb{Q}$, whereas in the nondegenerate case $H_{\mathfrak{m},\{\infty_{1}\}}^{\mathcal{O}}$ and $H_{\mathfrak{m},\{\infty_{2}\}}^{\mathcal{O}}$ are not. The field $H_{\mathfrak{m},\emptyset}^{\mathcal{O}}$ is totally real, $H_{\mathfrak{m},+}^{\mathcal{O}}$ is (in the nondegenerate case) totally complex and CM, and $H_{\mathfrak{m},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}}$ is (in the nondegenerate case) totally complex and not CM.
8.2. Statement of the Extended Ray Class Fields of Orders Conjecture
We conjecture an identification of fields of triple product SIC-fields and projection SIC-fields with certain class fields of orders. We recall the fact that orders of a quadratic field $K$ are completely specified by their discriminants $\Delta^{\prime}=(f^{\prime})^{2}\Delta_{0}$, where $\Delta_{0}$ is the field discriminant of $K$ (a fundamental discriminant; see the beginning of Section 7.1) and $f^{\prime}$ is a positive integer.
Definition 8.3.
Fix a positive integer $d\geq 4$, write $\Delta=(d+1)(d-3)=f^{2}\Delta_{0}$ for a fundamental discriminant $\Delta_{0}$, and let $f^{\prime}$ be a positive divisor of $f$. Let $\Delta^{\prime}=(f^{\prime})^{2}\Delta_{0}$ and $\mathcal{O}^{\prime}=\mathcal{O}_{\Delta^{\prime}}$. We introduce two families of class fields parametrized by $d$ and $f^{\prime}$:
$E_{d,f^{\prime}}$ $=H_{d\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}^{\prime}};$ (8.5)
$\tilde{E}_{d,f^{\prime}}$ $=H_{d^{\prime}\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}^{\prime}}.$ (8.6)
Here $d^{\prime}=d$ if $d$ is odd and $d^{\prime}=2d$ if $d$ is even.
Note that $E_{d,f^{\prime}}=\tilde{E}_{d,f^{\prime}}$ if $d$ is odd, and $E_{d,f^{\prime}}\subseteq\tilde{E}_{d,f^{\prime}}$ if $d$ is even. Specifically, when $d$ is even, the index $[\tilde{E}_{d,f^{\prime}}:E_{d,f^{\prime}}]=\left|\operatorname{Cl}_{2d\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})/\operatorname{Cl}_{d\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})\right|=2$ by 10.2.
Conjecture 8.4 (Extended Ray Class Fields of Orders Conjecture).
Fix a positive integer $d\geq 4$, write $\Delta=(d+1)(d-3)=f^{2}\Delta_{0}$ for a fundamental discriminant $\Delta_{0}$, and let $f^{\prime}\mid f$, $f^{\prime}>0$.
Let :{positive divisors of $\mathcal{M}:\{\mbox{positive divisors of $f$}\}\to\{[[S]]:S\mbox{ a Weyl--Heisenberg line-SIC in }\mathbb{C}^{d}\}$ be any function satisfying the conditions of 7.2. If $S$ is a Weyl–Heisenberg SIC with $[S]\in\mathcal{M}(f^{\prime})$, then
$F^{\rm trip}\!\left(S\right)$ $=E_{d,f^{\prime}}$ and $F^{\rm vec}\!\left(S\right)=F^{\rm proj}\!\left(S\right)$ $=\tilde{E}_{d,f^{\prime}}.$ (8.7)
8.4 implies that $F^{\rm proj}\!\left(S\right)=F^{\rm eproj}\!\left(S\right)$ for any Weyl–Heisenberg SIC $S$. This is because $\tilde{E}_{d,f^{\prime}}=H^{\mathcal{O}^{\prime}}_{d^{\prime}\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\}}$ contains the ray class field of $\mathbb{Q}$ with level datum $(\mathbb{Z};d^{\prime}\mathbb{Z},\{\infty\})$, which is simply the field $\mathbb{Q}(\zeta_{d^{\prime}})=\mathbb{Q}(\xi_{d})$ of $d^{\prime}$-th roots of unity; thus, $E_{d,f^{\prime}}(\xi_{d})=\tilde{E}_{d,f^{\prime}}$.
8.4 is a natural extension of 1.1 of Appleby, Flammia, McConnell, and Yard. Under the additional (likely true) assumption that $F^{\rm vec}\!\left(S\right)$ is closed under complex conjugation (so that $F^{\rm proj}\!\left(S\right)=F^{\rm vec}\!\left(S\right)$), their conjecture becomes equivalent to the statement that $F^{\rm proj}\!\left(S\right)=\tilde{E}_{d,1}$ for any $S$ in the minimal multiplet.
8.3. Statement of the Ideal-Class-to-SIC Conjecture
We now formulate a conjecture that refines 5.4, 7.1, and 8.4 by insisting on compatibility with the Artin map. To state this conjecture, we decompose
$\operatorname{Clm}(\mathcal{O}_{\!\Delta})=\bigsqcup_{f^{\prime}|f}\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})$ (8.8)
by means of 5.3.
Conjecture 8.5 (Ideal-Class-to-SIC Conjecture).
Fix a positive integer $d>3$, let $\Delta=(d+1)(d-3)$, and write $\Delta=f^{2}\Delta_{0}$ where $\Delta_{0}$ is a fundamental discriminant of a quadratic field. Then there is a bijection
$\bigsqcup_{f^{\prime}|f}\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})\xrightarrow{\Omega}\operatorname{WHSIC}(d)/\operatorname{PEC}(d)$ (8.9)
under which an ideal class $\mathcal{A}\in\operatorname{Cl}(\mathcal{O}^{\prime})$ with $\mathcal{O}^{\prime}=\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}}$ maps to a geometric equivalence class $\Omega(\mathcal{A})=[S_{\mathcal{A}}]$ with associated fields $F^{\rm trip}\!\left(S_{\mathcal{A}}\right)=E_{d,f^{\prime}}$ and $F^{\rm proj}\!\left(S_{\mathcal{A}}\right)=\tilde{E}_{d,f^{\prime}}$.
Furthermore, $\Omega$ can be chosen to be compatible with the Artin map (defined in 8.2)
$Art_{\mathcal{O}^{\prime}}:\operatorname{Cl}_{d^{\prime}\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\}}(\mathcal{O}^{\prime})\to\operatorname{Gal}(\tilde{E}_{d,f^{\prime}}/\mathbb{Q}(\Delta_{0})),$ (8.10)
in the sense that
$[S_{\bar{\mathcal{A}}\bar{\mathfrak{B}}}]=[(\operatorname{Art}_{\mathcal{O}^{\prime}}(\mathcal{A}))(S_{\bar{\mathfrak{B}}})]\mbox{ for any }\mathcal{A},\mathfrak{B}\in\operatorname{Cl}_{d^{\prime}\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\}}(\mathcal{O}^{\prime})$ (8.11)
and their images $\bar{\mathcal{A}},\bar{\mathfrak{B}}\in\operatorname{Cl}(\mathcal{O}^{\prime})$ under the natural quotient map.
The required compatibility with the Artin map implies that there is exactly one Galois super-equivalence class of Weyl–Heisenberg SICs—equivalently, one Galois multiplet—for each order $\mathcal{O}$ between $\mathcal{O}_{\!\Delta_{0}}$ and $\mathcal{O}_{\!\Delta}$, which is 7.1.
We formulate this hypothesis because its truth would provide a unifying principle explaining the structure of all the other conjectures; we present no mechanism to construct a candidate map $\Omega$. 8.5 would define a map $\mathcal{M}$ of the type given in 7.2 by means of the following commutative diagram,
${\displaystyle\bigsqcup_{f^{\prime}|f}\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})}$${\operatorname{WHSIC}(d)/\operatorname{PEC}(d)}${positive divisors of ${\{\mbox{positive divisors of $f$}\}}$${\{[[S]]:S\mbox{ a line-SIC in }\mathbb{C}^{d}\}}$$\scriptstyle{\Omega}$$\scriptstyle{\mathcal{M}}$ (8.12)
where the downward maps send $\mathcal{A}\in\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})$ to $f^{\prime}$ and $[S]$ to $[[S]]$, respectively.
8.5 is compatible with conjectures of the first author [32] and other authors [6, 11] connecting SICs to the Stark conjectures for real quadratic fields [41], particularly as formulated by Tate [43, 44], where the Stark units carry a Galois action compatible with the Artin map.
9. Theorems on degeneration and nondegeneration of inclusions of ray class fields of orders
In this section, we prove unconditional results about when the fields $E_{d,f^{\prime}}$ introduced in 8.3 can coincide, while fixing $d$ and varying $f^{\prime}$. We prove similar results about some closely related fields. These results allows us to explain some observed phenomena about SICs and their associated number fields.
The fact that two distinct multiplets $[[S_{1}]]\neq[[S_{2}]]$ can have the same associated field $F^{\rm proj}\!\left(S_{1}\right)=F^{\rm proj}\!\left(S_{2}\right)$ was discovered experimentally by Markus Grassl [27]. This observation came as a surprise, as experimental evidence in low dimensions had pointed to multiplets being determined by the associated fields.
We determine exactly when there is an equality of class fields $E_{d,f^{\prime}}=E_{d,f^{\prime\prime}}$ (equivalently, $\tilde{E}_{d,f^{\prime}}=\tilde{E}_{d,f^{\prime\prime}}$) in the case when $f^{\prime\prime}\mid f^{\prime}$. We do not rule out such an equality for $f^{\prime}$, $f^{\prime\prime}$ incomparable in the divisibility order. (Some such cases could be ruled out by considering ramification properties of primes in these fields as extensions of $K$.) Nonetheless, our results explain the examples discovered by Grassl.
We characterize all cases where an equality $E_{d,f^{\prime}}=E_{d,f^{\prime\prime}}$ happens for some distinct $f^{\prime},f^{\prime\prime}$ satisfying $f^{\prime\prime}\mid f^{\prime}\mid f$ (with $(d+1)(d-3)=f^{2}\Delta_{0}$), purely in terms of the dimension $d$. We show that the set of such $d$ has a limiting natural density of $\frac{1}{48}$.
8.4 then predicts that, in such dimensions $d$, there will be two distinct SIC multiplets that generate the same field. This assertion is experimentally testable. A calculation of this criterion for $d\leq 500$ is compared with Grassl’s experimental data for occurrence of equality of associated fields.
9.1. Preliminary results on class groups and class fields of orders
We recall results from [33] needed to relate different ray class fields to each other and determine the relative Galois group (including whether or not it is the trivial group). We first recall a result on when there is an inclusion between ray class fields of two orders $\mathcal{O}\subseteq\mathcal{O}^{\prime}$, which is [33, Thm. 6.5].
Theorem 9.1.
For two orders $\mathcal{O}\subseteq\mathcal{O}^{\prime}$ in a number field $K$ and any level datum $(\mathcal{O};\mathfrak{m},\Sigma)$ for $\mathcal{O}$, there are inclusions of ray class fields $H_{\mathfrak{m}\mathcal{O}^{\prime},\Sigma}^{\mathcal{O}^{\prime}}\subseteq H_{\mathfrak{m},\Sigma}^{\mathcal{O}}\subseteq H_{\left(\mathfrak{m}:\mathcal{O}^{\prime}\right),\Sigma}^{\mathcal{O}^{\prime}}$. (Here, the extended ideal $\mathfrak{m}\mathcal{O}^{\prime}$ is the smallest $\mathcal{O}^{\prime}$-ideal containing $\mathfrak{m}$, and the quotient ideal $\left(\mathfrak{m}:\mathcal{O}^{\prime}\right)$ is the maximal $\mathcal{O}^{\prime}$-ideal contained in $\mathfrak{m}$.)
To determine cases of equality $H_{\mathfrak{m}\mathcal{O}^{\prime},\Sigma}^{\mathcal{O}^{\prime}}=H_{\mathfrak{m},\Sigma}^{\mathcal{O}}$, we will convert the problem via the Artin isomorphism (8.2) to a calculation of isomorphisms between ray class groups in the two orders.
We next recall an exact sequence important for the calculation of ray class groups of orders and maps between them. The exact sequence relates class groups of two different orders in terms of the behavior of various subgroups of local and global unit groups of the two orders. In order to write down the exact sequence, we first need the following definition and conventions concerning the unit groups.
Definition 9.2.
For a commutative ring with unity $R$ and an ideal $I$ of $R$, define the group
$\operatorname{U}_{I}(R):=\{\alpha\in R^{\times}:\alpha\equiv 1\,\left(\operatorname{mod}\,I\right)\}=(1+I)\cap R^{\times}.$ (9.1)
If in addition $R$ has real embeddings and $\Sigma$ is a subset of the real embeddings of $R$, define
$\operatorname{U}_{I,\Sigma}(R):=\{\alpha\in R^{\times}:\alpha\equiv 1\,\left(\operatorname{mod}\,I\right)\mbox{ and }\rho(\alpha)>0\mbox{ for }\rho\in\Sigma\}.$ (9.2)
We also use the following extension of this notation: If there is an map $\phi:R_{1}\to R_{2}$ implicit in the discussion, and $I$ is an ideal of $R_{1}$, then we let $\operatorname{U}_{I}(R_{2}):=\operatorname{U}_{\phi(I)R_{2}}(R_{2})$ and $\operatorname{U}_{I,\Sigma}(R_{2}):=\operatorname{U}_{\phi(I)R_{2},\Sigma}(R_{2})$. The follow theorem is [33, Thm. 5.4].
Theorem 9.3.
Let $K$ be a number field and $\mathcal{O}\subseteq\mathcal{O}^{\prime}\subseteq\mathcal{O}_{K}$ be orders of $K$, and let $\mathfrak{f}=\mathfrak{f}_{\mathcal{O}^{\prime}}(\mathcal{O})$. Let $\mathfrak{m}$ be an ideal of $\mathcal{O}$, $\mathfrak{m}^{\prime}$ an ideal of $\mathcal{O}^{\prime}$ such that $\mathfrak{m}\mathcal{O}^{\prime}\subseteq\mathfrak{m}^{\prime}$, and $\Sigma^{\prime}\subseteq\Sigma\subseteq\{\mbox{embeddings }K\hookrightarrow\mathbb{R}\}$. Let ${\mathfrak{d}}$ be any $\mathcal{O}^{\prime}$-ideal such that ${\mathfrak{d}}\subseteq\left(\mathfrak{m}:\mathcal{O}^{\prime}\right)$. With the $\operatorname{U}$-groups defined as in 9.2, we have the following exact sequence.
$1\to\frac{\operatorname{U}_{\mathfrak{m}^{\prime},\Sigma^{\prime}}\!\left(\mathcal{O}^{\prime}\right)}{\operatorname{U}_{\mathfrak{m},\Sigma}\!\left(\mathcal{O}\right)}\to\frac{\operatorname{U}_{\mathfrak{m}^{\prime}}\!\left(\mathcal{O}^{\prime}/{\mathfrak{d}}\right)}{\operatorname{U}_{\mathfrak{m}}\!\left(\mathcal{O}/{\mathfrak{d}}\right)}\times\{\pm 1\}^{\left|\Sigma\setminus\Sigma^{\prime}\right|}\to\operatorname{Cl}_{\mathfrak{m},\Sigma}(\mathcal{O})\to\operatorname{Cl}_{\mathfrak{m}^{\prime},\Sigma^{\prime}}(\mathcal{O}^{\prime})\to 1$ (9.3)
9.2. Nondegeneracy of certain ray class fields in a fixed order $\mathcal{O}$ varying infinite places
In this subsection, we show nondegeneracy of the inclusion of subfields of $E_{d,f^{\prime}}=H_{d\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}^{\prime}}$ obtained by dropping one or both of the infinite places from the level datum. It follows that $E_{d,f^{\prime}}$ is totally complex and non-CM. The field $H_{d\mathcal{O}^{\prime},\{\infty_{1}\}}^{\mathcal{O}^{\prime}}$ appears in [10] (with the notation $\mathfrak{R}_{1}$) as a field generated by SIC overlaps after making appropriate choices.
Nondegeneracy will be deduced from the following lemma. This lemma is very special to the family of ray class groups we are considering, and it is crucial to all the calculations. To not interrupt the exposition, we defer the proof to Appendix A.
Lemma 9.4.
Let $d\geq 4$, $\Delta=(d+1)(d-3)$, and $\varepsilon_{d}=\frac{d-1+\sqrt{\Delta}}{2}$. Let $\Sigma$ be a subset of the real embedding of $\mathbb{Q}(\sqrt{\Delta})$. Fix an order $\mathcal{O}$ satisfying
$\mathcal{O}_{\!\Delta}:=\mathbb{Z}[\varepsilon_{d}]\subseteq\mathcal{O}\subseteq\mathcal{O}_{\mathbb{Q}(\sqrt{\Delta})}.$ (9.4)
Then,
$\operatorname{U}_{d\mathcal{O},\Sigma}(\mathcal{O})=\langle\varepsilon_{d}^{3}\rangle.$ (9.5)
The following result shows nondegeneracy of class groups whilst varying the infinite places. It generalizes [32, Lem. 5.3].
Theorem 9.5.
Let $d\geq 4$, $\Delta=(d+1)(d-3)$, and $\varepsilon_{d}=\frac{d-1+\sqrt{\Delta}}{2}$. For every order
$\mathbb{Z}[\varepsilon_{d}]\subseteq\mathcal{O}\subseteq\mathcal{O}_{\mathbb{Q}(\sqrt{\Delta})},$ (9.6)
all the maps in the diagram
${\operatorname{Cl}_{d\mathcal{O},\{\infty_{1},\infty_{2}\}}(\mathcal{O})}$${\operatorname{Cl}_{d\mathcal{O},\{\infty_{1}\}}(\mathcal{O})}$${\operatorname{Cl}_{d\mathcal{O},\{\infty_{2}\}}(\mathcal{O})}$${\operatorname{Cl}_{d\mathcal{O},\emptyset}(\mathcal{O})}$ (9.7)
are surjections with kernel $\mathbb{Z}/2\mathbb{Z}$.
Proof.
We apply the exact sequence of 9.3, with $\mathcal{O}^{\prime}=\mathcal{O}$, $\mathfrak{m}=\mathfrak{m}^{\prime}=d\mathcal{O}$, $\Sigma^{\prime}\subseteq\Sigma$, and ${\mathfrak{d}}=(d\mathcal{O}:\mathcal{O})=d\mathcal{O}$. It says that there is an exact sequence
$1\to\frac{\operatorname{U}_{d\mathcal{O},\Sigma^{\prime}}\!\left(\mathcal{O}\right)}{\operatorname{U}_{d\mathcal{O},\Sigma}\!\left(\mathcal{O}\right)}\to\frac{\operatorname{U}_{d\mathcal{O}}\!\left(\mathcal{O}/d\mathcal{O}\right)}{\operatorname{U}_{d\mathcal{O}}\!\left(\mathcal{O}/d\mathcal{O}\right)}\times\{\pm 1\}^{\left|\Sigma\setminus\Sigma^{\prime}\right|}\to\operatorname{Cl}_{d\mathcal{O},\Sigma}(\mathcal{O})\to\operatorname{Cl}_{d\mathcal{O},\Sigma^{\prime}}(\mathcal{O})\to 1$ (9.8)
9.4 shows for any order $\mathcal{O}\subseteq\mathcal{O}_{\mathbb{Q}(\sqrt{\Delta})}$ containing $\varepsilon_{d}$ that $\frac{\operatorname{U}_{d\mathcal{O},\Sigma^{\prime}}\!\left(\mathcal{O}\right)}{\operatorname{U}_{d\mathcal{O},\Sigma}\!\left(\mathcal{O}\right)}=\frac{\langle\varepsilon_{d}^{3}\rangle}{\langle\varepsilon_{d}^{3}\rangle}=1$. We deduce that
$1\to\{\pm 1\}^{\left|\Sigma\setminus\Sigma^{\prime}\right|}\to\operatorname{Cl}_{d\mathcal{O},\Sigma}(\mathcal{O})\to\operatorname{Cl}_{d\mathcal{O},\Sigma^{\prime}}(\mathcal{O})\to 1$ (9.9)
It follows that the class groups $\operatorname{Cl}_{(d)}(\mathcal{O})$, $\operatorname{Cl}_{(d)\infty_{1}}(\mathcal{O})$, $\operatorname{Cl}_{(d)\infty_{2}}(\mathcal{O})$, and $\operatorname{Cl}_{(d)\infty_{1}\infty_{2}}(\mathcal{O})$ are all distinct and have the relationships given in the theorem statement. ∎
Remark 9.6.
By 8.2, we obtain for any order $\mathcal{O}$ satisfying $\mathbb{Z}[\varepsilon_{d}]\subseteq\mathcal{O}\subseteq\mathcal{O}_{\mathbb{Q}(\sqrt{\Delta})}$ the following inclusions of ray class fields of orders,
[…]
where each edge in the diagram denotes an extension of degree 2. We also know that $\operatorname{Gal}(H_{d\mathcal{O},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}}/H_{d\mathcal{O},\emptyset}^{\mathcal{O}})\cong\{\pm 1\}^{2}\cong(\mathbb{Z}/2\mathbb{Z})^{2}$.
9.3. Cases of equality in ray class groups and ray class fields of two orders $\mathcal{O}\subsetneq\mathcal{O}^{\prime}$
In certain cases, the ray class group of an order does not change the under extension of the order, and consequently, the correpsonding ray class fields are identical. This phenomenon can be understood using the exact sequence given in 9.3: The rightmost two groups in the exact sequence are isomorphic if and only if the leftmost two groups in the sequence are isomorphic. We quantify exactly when this degeneration phenomenon occurs for the level data appearing in our conjectures about SICs.
Theorem 9.7.
Let $d\geq 4$, $\Delta=(d+1)(d-3)$, and $\varepsilon_{d}=\frac{d-1+\sqrt{\Delta}}{2}$. Write $\Delta=f^{2}\Delta_{0}$, where $\Delta_{0}$ is a fundamental discriminant. Let $\Sigma^{\prime}\subseteq\Sigma$ be subsets of $\{\infty_{1},\infty_{2}\}$. For every containment of orders
$\mathbb{Z}[\varepsilon_{d}]\subseteq\mathcal{O}\subseteq\mathcal{O}^{\prime}\subseteq\mathcal{O}_{\mathbb{Q}(\sqrt{\Delta})},$ (9.11)
the surjective map
$\operatorname{Cl}_{d\mathcal{O},\Sigma}(\mathcal{O})\to\operatorname{Cl}_{d\mathcal{O}^{\prime},\Sigma^{\prime}}(\mathcal{O}^{\prime})$ (9.12)
induced by extension of ideals is an isomorphism if and only if $\Sigma=\Sigma^{\prime}$ and either:
  • (1)
    $\mathcal{O}=\mathcal{O}^{\prime}$, or
  • (2)
[…]
Proof.
Let $g=[\mathcal{O}^{\prime}:\mathcal{O}]$ be the relative index of the orders and let $\Delta^{\prime}=\operatorname{disc}{\mathcal{O}^{\prime}}$, so that the two orders have the following expressions as $\mathbb{Z}$-modules:
$\mathcal{O}^{\prime}$ $=\mathbb{Z}+\frac{\Delta^{\prime}+\sqrt{\Delta^{\prime}}}{2}$ and $\mathcal{O}$ $=\mathbb{Z}+g\frac{\Delta^{\prime}+\sqrt{\Delta^{\prime}}}{2}.$ (9.13)
A direct computation shows that the conductor ideal $\left(\mathcal{O}:\mathcal{O}^{\prime}\right)=g\mathcal{O}^{\prime}$, so the colon ideal $\left(d\mathcal{O}:\mathcal{O}^{\prime}\right)=dg\mathcal{O}^{\prime}$. We apply the exact sequence of 9.3, choosing $\mathfrak{m}=d\mathcal{O}$, $\mathfrak{m}^{\prime}=d\mathcal{O}^{\prime}$, $\Sigma^{\prime}\subseteq\Sigma$, and ${\mathfrak{d}}=dg\mathcal{O}^{\prime}$, to obtain
$1\to\frac{\operatorname{U}_{d\mathcal{O}^{\prime},\Sigma^{\prime}}(\mathcal{O}^{\prime})}{\operatorname{U}_{d\mathcal{O},\Sigma}(\mathcal{O})}\to\frac{\operatorname{U}_{d\mathcal{O}^{\prime}}\!\left(\mathcal{O}^{\prime}/dg\mathcal{O}^{\prime}\right)}{\operatorname{U}_{d\mathcal{O}}\!\left(\mathcal{O}/dg\mathcal{O}^{\prime}\right)}\times\{\pm 1\}^{\left|\Sigma\setminus\Sigma^{\prime}\right|}\to\operatorname{Cl}_{d\mathcal{O},\Sigma}(\mathcal{O})\to\operatorname{Cl}_{d\mathcal{O}^{\prime},\Sigma^{\prime}}(\mathcal{O}^{\prime})\to 1$ (9.14)
By 9.4, $\operatorname{U}_{d\mathcal{O}^{\prime},\Sigma^{\prime}}(\mathcal{O}^{\prime})=\operatorname{U}_{d\mathcal{O},\Sigma}(\mathcal{O})=\langle\varepsilon_{d}^{3}\rangle$, so the leftmost term in the sequence is the trivial group. Thus, the rightmost map (which is the surjection induced by extension of ideals mentioned in the theorem statement) is an isomorphism if and only if the group $\frac{\operatorname{U}_{d\mathcal{O}^{\prime}}\!\left(\mathcal{O}^{\prime}/df\mathcal{O}^{\prime}\right)}{\operatorname{U}_{d\mathcal{O}}\!\left(\mathcal{O}/df\mathcal{O}^{\prime}\right)}\times\{\pm 1\}^{\left|\Sigma\setminus\Sigma^{\prime}\right|}=1$. We compute the order of this group as
$\left|\frac{\operatorname{U}_{d\mathcal{O}^{\prime}}\!\left(\mathcal{O}^{\prime}/dg\mathcal{O}^{\prime}\right)}{\operatorname{U}_{d\mathcal{O}}\!\left(\mathcal{O}/dg\mathcal{O}^{\prime}\right)}\times\{\pm 1\}^{\left|\Sigma\setminus\Sigma^{\prime}\right|}\right|=\left|{\frac{\operatorname{U}_{d\mathcal{O}^{\prime}}\!\left(\mathcal{O}^{\prime}/dg\mathcal{O}^{\prime}\right)}{\operatorname{U}_{d\mathcal{O}}\!\left(\mathcal{O}/dg\mathcal{O}^{\prime}\right)}}\right|\cdot\left|\{\pm 1\}^{\left|\Sigma\setminus\Sigma^{\prime}\right|}\right|=\left|{\frac{\operatorname{U}_{d\mathcal{O}^{\prime}}\!\left(\mathcal{O}^{\prime}/dg\mathcal{O}^{\prime}\right)}{\operatorname{U}_{d\mathcal{O}}\!\left(\mathcal{O}/dfg\mathcal{O}^{\prime}\right)}}\right|\cdot 2^{\left|\Sigma\setminus\Sigma^{\prime}\right|}.$ (9.15)
Thus, the map is an isomorphism if and only if $\operatorname{U}_{d\mathcal{O}^{\prime}}\!\left(\mathcal{O}^{\prime}/dg\mathcal{O}^{\prime}\right)=\operatorname{U}_{d\mathcal{O}}\!\left(\mathcal{O}/dg\mathcal{O}^{\prime}\right)$ and $\Sigma=\Sigma^{\prime}$. We now determine the cases when the former happens.
We wish to compare the group $U^{\prime}:=\operatorname{U}_{d\mathcal{O}^{\prime}}\!\left(\mathcal{O}^{\prime}/dg\mathcal{O}^{\prime}\right)$ to its subgroup $U:=\operatorname{U}_{d\mathcal{O}}\!\left(\mathcal{O}/dg\mathcal{O}^{\prime}\right)$ and determine when they are equal. They are clearly equal when $g=1$, giving (1); from now on, we assume $g>1$. The structure of the proof is as follows:
  • (A)
    If $g$ is an odd prime, we will show that $U^{\prime}\neq U$.
  • (B)
    If $g=2$, we will characterize exactly when $U^{\prime}=U$, giving (2).
  • (C)
    If $g=4$, we will use (B) to show that $U^{\prime}\neq U$ by considering intermediate orders between $\mathcal{O}$ and $\mathcal{O}^{\prime}$.
  • (D)
    We will use (A) and (C) to conclude that $U^{\prime}\neq U$ for all $g>2$ by considering intermediate orders between $\mathcal{O}$ and $\mathcal{O}^{\prime}$.
Let $\omega=\frac{\Delta^{\prime}+\sqrt{\Delta^{\prime}}}{2}$, so $\mathcal{O}^{\prime}=\mathbb{Z}+\omega\mathbb{Z}$ and $\mathcal{O}=\mathbb{Z}+g\omega\mathbb{Z}$. Let
$M^{\prime}$ =$=\{\beta\in\mathcal{O}^{\prime}/dg\mathcal{O}^{\prime}:\beta\equiv 1\,\left(\operatorname{mod}\,d\mathcal{O}^{\prime}\right)\}\mbox{ and}$ (9.16)
$M$ =$=\{\beta\in\mathcal{O}/dg\mathcal{O}^{\prime}:\beta\equiv 1\,\left(\operatorname{mod}\,d\mathcal{O}\right)\}.$ (9.17)
Then $U^{\prime}$ consists of the invertible elements of $M^{\prime}$, and $U$ consists of the invertible elements of $M$, that is, the elements of each that are coprime to $dg$.
The elements of $M^{\prime}$ and $M$ may be written down as follows, where $\overline{\alpha}$ stands for the equivalence class $\,\left(\operatorname{mod}\,dg\mathcal{O}^{\prime}\right)$ represented by $\alpha$.
$M^{\prime}$ =$=\left\{\overline{1+dm+dn\omega}:0\leq m,n<g\right\},\mbox{ and}$ (9.18)
$M$ =$=\left\{\overline{1+dm}:0\leq m<g\right\}.$ (9.19)
Case (A). Suppose $g=p$ is an odd prime. For $n\in\mathbb{Z}$ we have, using the norm in $\mathcal{O}_{K}$,
$\operatorname{Nm}(1+dn\omega)$ $=\left(1+dn\frac{\Delta^{\prime}+\sqrt{\Delta^{\prime}}}{2}\right)\left(1+dn\frac{\Delta^{\prime}-\sqrt{\Delta^{\prime}}}{2}\right)$ (9.20)
$=1+dn\Delta^{\prime}+d^{2}n^{2}\frac{(\Delta^{\prime})^{2}-\Delta^{\prime}}{4}.$ (9.21)
The quadratic polynomial $F(x)=1+d\Delta^{\prime}x+d^{2}\frac{(\Delta^{\prime})^{2}-\Delta^{\prime}}{4}x^{2}$ has at most two roots in $\mathbb{Z}/p\mathbb{Z}$, and $F(0)=1$ (so $x=0$ is not a root), so for $p\geq 5$ there exists some $n\in\{1,\ldots,p-1\}$ such that $\operatorname{Nm}(1+dn\omega)=F(n)\not\equiv 0\,\left(\operatorname{mod}\,p\right)$.
[…]
Finally $\operatorname{Nm}(1+dn\omega)\equiv 1\,\left(\operatorname{mod}\,d\right)$, so we conclude $\operatorname{Nm}(1+dn\omega)$ is a unit modulo $dg$, and thus $\overline{1+dn\omega}\in U^{\prime}\setminus U$, so $U\subsetneq U^{\prime}$.
Case (B). Suppose $g=2$. The elements of $M^{\prime}$ and $M$ may be written down as follows.
$M^{\prime}$ =$=\{\overline{1},\overline{1+d},\overline{1+d\omega},\overline{1+d+d\omega}\},\mbox{ and}$ (9.22)
$M$ =$=\{\overline{1},\overline{1+d}\}.$ (9.23)
Clearly, every element of $M^{\prime}$ and $M$ are coprime to $d$. If $d$ is even, then they are all coprime to $2d$, so $M^{\prime}=U^{\prime}$, $M=U$, and $U^{\prime}\neq U$. Now suppose that $d$ is odd; then $U=\{\overline{1}\}$. We compute
$\operatorname{Nm}(1+d\omega)=\begin{cases}1\,\left(\operatorname{mod}\,2d\right)&\mbox{if }\Delta^{\prime}\equiv 0\,\left(\operatorname{mod}\,8\right),\\ 1+d\,\left(\operatorname{mod}\,2d\right)&\mbox{if }\Delta^{\prime}\equiv 1\,\left(\operatorname{mod}\,8\right),\\ 1+d\,\left(\operatorname{mod}\,2d\right)&\mbox{if }\Delta^{\prime}\equiv 4\,\left(\operatorname{mod}\,8\right),\\ 1\,\left(\operatorname{mod}\,2d\right)&\mbox{if }\Delta^{\prime}\equiv 5\,\left(\operatorname{mod}\,8\right).\end{cases}$ (9.24)
and
$\operatorname{Nm}(1+d+d\omega)=\begin{cases}1+d\,\left(\operatorname{mod}\,2d\right)&\mbox{if }\Delta^{\prime}\equiv 0\,\left(\operatorname{mod}\,8\right),\\ 1+d\,\left(\operatorname{mod}\,2d\right)&\mbox{if }\Delta^{\prime}\equiv 1\,\left(\operatorname{mod}\,8\right),\\ 1\,\left(\operatorname{mod}\,2d\right)&\mbox{if }\Delta^{\prime}\equiv 4\,\left(\operatorname{mod}\,8\right),\\ 1\,\left(\operatorname{mod}\,2d\right)&\mbox{if }\Delta^{\prime}\equiv 5\,\left(\operatorname{mod}\,8\right).\end{cases}$ (9.25)
Thus,
$U^{\prime}=\begin{cases}\{\overline{1},\overline{1+d\omega}\}&\mbox{if }\Delta^{\prime}\equiv 0\,\left(\operatorname{mod}\,8\right),\\ \{\overline{1}\}&\mbox{if }\Delta^{\prime}\equiv 1\,\left(\operatorname{mod}\,8\right),\\ \{\overline{1},\overline{1+d+d\omega}\}&\mbox{if }\Delta^{\prime}\equiv 4\,\left(\operatorname{mod}\,8\right),\\ \{\overline{1},\overline{1+d\omega},\overline{1+d+d\omega}\}&\mbox{if }\Delta^{\prime}\equiv 5\,\left(\operatorname{mod}\,8\right).\end{cases}$ (9.26)
[…]
Case (C). Suppose $g=4$. Consider the intermediate order $\mathcal{O}_{1}=\mathbb{Z}+2\omega\mathbb{Z}$. We have
$[\operatorname{Cl}_{d,\Sigma}(\mathcal{O}):\operatorname{Cl}_{d,\Sigma}(\mathcal{O}^{\prime})]=[\operatorname{Cl}_{d,\Sigma}(\mathcal{O}):\operatorname{Cl}_{d,\Sigma}(\mathcal{O}_{1})][\operatorname{Cl}_{d,\Sigma}(\mathcal{O}_{1}):\operatorname{Cl}_{d,\Sigma}(\mathcal{O}^{\prime})],$ (9.27)
but $\operatorname{disc}(\mathcal{O}_{1})$ is even, and thus $[\operatorname{Cl}_{d,\Sigma}(\mathcal{O}):\operatorname{Cl}_{d,\Sigma}(\mathcal{O}_{1})]=2$ by (B). Thus, $\operatorname{Cl}_{d,\Sigma}(\mathcal{O})\neq\operatorname{Cl}_{d,\Sigma}(\mathcal{O}^{\prime})$ in this case (i.e., $U^{\prime}\neq U$).
Case (D). Now consider any $g>2$. Write $g=h_{1}h_{2}$ where either $h_{1}=4$ or $h_{1}$ is an odd prime. There is then an intermediate order $\mathcal{O}_{1}=\mathbb{Z}+h_{1}\omega\mathbb{Z}$ with $[\mathcal{O}^{\prime}:\mathcal{O}_{1}]=h_{1}$, and by (A) and (C), $[\operatorname{Cl}_{d,\Sigma}(\mathcal{O}_{1}):\operatorname{Cl}_{d,\Sigma}(\mathcal{O}^{\prime})]>1$. Thus, as in (C), $\operatorname{Cl}_{d,\Sigma}(\mathcal{O})\neq\operatorname{Cl}_{d,\Sigma}(\mathcal{O}^{\prime})$ (i.e., $U^{\prime}\neq U$). ∎
We obtain the following criterion for equality of the associated ray class fields.
Theorem 9.8.
The set of $d\geq 4$ such that, for $\Delta=(d+1)(d-3)$ and $\varepsilon_{d}=\frac{d-1+\sqrt{\Delta}}{2}$, there exist distinct orders $\mathcal{O}\subseteq\mathcal{O}^{\prime}$ with
$\mathbb{Z}[\varepsilon_{d}]\subseteq\mathcal{O}\subseteq\mathcal{O}^{\prime}\subseteq\mathcal{O}_{\mathbb{Q}(\sqrt{\Delta})},$ (9.28)
[…]
Proof.
[…]
The Galois correspondence gives an isomorphism
$\operatorname{Gal}\!\left(H_{d\mathcal{O},\Sigma}^{\mathcal{O}}/H_{d\mathcal{O}^{\prime},\Sigma}^{\mathcal{O}^{\prime}}\right)\cong\ker\!\left(\operatorname{Cl}_{d\mathcal{O},\Sigma}(\mathcal{O})\to\operatorname{Cl}_{d\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})\right).$ (9.29)
Thus, $H_{d\mathcal{O},\Sigma}^{\mathcal{O}}=H_{d\mathcal{O}^{\prime},\Sigma}^{\mathcal{O}^{\prime}}$ if and only if the (always surjective) map $\operatorname{Cl}_{d\mathcal{O},\Sigma}(\mathcal{O})\to\operatorname{Cl}_{d\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})$ is an isomorphism. The corollary now follows. ∎
We now summarize the degeneration phenomenon in the context and notation of the fields $E_{d,f^{\prime}}$, $\tilde{E}_{d,f^{\prime}}$ appearing in our conjectures.
Corollary 9.9.
Let $d\geq 4$. Write $\Delta=(d+1)(d-3)=f^{2}\Delta_{0}$ for $\Delta_{0}$ a fundamental discriminant, and suppose $f^{\prime},f^{\prime\prime}$ are positive integers satisfying $f^{\prime\prime}\mid f^{\prime}\mid f$ and $f^{\prime}\neq f^{\prime\prime}$. The following are equivalent:
  • (1)
    rad(Δ)1(mod $\operatorname{rad}(\Delta)\equiv 1\,\left(\operatorname{mod}\,8\right)$ and $f^{\prime}=2f^{\prime\prime}$.
  • (2)
[…]
  • (3)
    $E_{d,f^{\prime}}=E_{d,f^{\prime\prime}}$
  • (4)
    $\tilde{E}_{d,f^{\prime}}=\tilde{E}_{d,f^{\prime\prime}}$
Proof.
The equivalence of (1), (2), and (3) follows immediately from 9.7 and the proof of 9.8, taking $\mathcal{O}=\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}}$ and $\mathcal{O}^{\prime}=\mathcal{O}_{(f^{\prime\prime})^{2}\Delta_{0}}$.
The equivalence of (3) and (4) follows from the fact that
$\tilde{E}_{d,f^{\prime}}^{\prime}$ $=E_{d,f^{\prime}}$ and $\tilde{E}_{d,f^{\prime\prime}}$ $=E_{d,f^{\prime\prime}}$ (9.30)
[…]
9.4. Rarity of degeneration for ray class fields of orders
We determine the density of $d$ for which the criterion of 9.7 holds.
Proposition 9.10.
We have
[…]
Proof.
[…]
Let $\mathcal{D}=\bigcup_{k\geq 0}{\mathcal{D}}_{k}$ where
$\mathcal{D}_{k}=\{d\geq 4:2^{2k}\mid\mid\Delta_{1}\mbox{ and }\Delta_{1}\equiv 1\,\left(\operatorname{mod}\,8\right)\}.$ (9.32)
We show that each $\mathcal{D}_{k}$ has a natural density
${\bf d}_{k}:=\lim_{n\to\infty}\frac{1}{N}\#\{d\in\mathcal{D}_{k}:\,d\leq N\}$ (9.33)
by showing it is a finite union of complete arithmetic progressions. The analysis below gives
${\bf d_{0}=\bf d_{1}=\bf d_{2}=\bf d_{3}=\bf d_{4}=\bf d_{5}}=0$ (9.34)
and ${\bf d_{k}}=\frac{2}{2^{2k}}$ for $k\geq 6$.
  1. (1)
[…]
  2. (2)
[…]
    1. (a)
[…]
    2. (b)
[…]
    3. (c)
[…]
    4. (d)
      The pattern continues for higher $k$, yielding a density set $\frac{1}{2^{2k+3}}$ set at level $k$.
  3. (3)
[…]
    1. (a)
[…]
    2. (b)
      The pattern continues for higher $k$, yielding a density set $\frac{1}{2^{2k+3}}$ set at level $k$.
[…]
$\sum_{k=1}^{\infty}\frac{1}{2^{5+2k}}=\frac{1}{96}.$
[…]
Remark 9.11.
The rarity of degeneration is special to the case of SIC fields. Degeneration is common in ring class towers over real quadratic fields; indeed, for any nonsquare $\Delta>0$, any prime $p$, and any fixed modulus $\mathfrak{m}$, it may be shown that the tower of fields
$H_{\mathfrak{m}}^{\mathcal{O}_{0}}\subseteq H_{\mathfrak{m}}^{\mathcal{O}_{1}}\subseteq H_{\mathfrak{m}}^{\mathcal{O}_{2}}\subseteq\cdots\subseteq H_{\mathfrak{m}}^{\mathcal{O}_{n}}\subseteq\cdots$ (9.35)
associated to the orders $\mathcal{O}_{n}=\mathbb{Z}[p^{n}\sqrt{\Delta}]$ is eventually constant. The lack of degeneration for inclusions of SIC fields seems due to the particular properties of the unit group of the base field modulo $d$. The tendency of SIC fields to be large, and “aligned SIC” towers incorporating SIC fields in different dimensions [5] to be noncyclotomic (and thus necessarily “horizontal,” in the sense of introducing infinitely many primes into both the degree and the modulus), holds for similar reasons.
9.5. Experimental testing of degeneration for ray class fields of orders
One can compare the prediction of 9.8 with experimental data on algebraic SICs, which can detect equality of SIC-fields for SICs in different Galois multiplets (which conjecturally correspond to different orders.) Only three cases of equality have been found experimentally, all by Marcus Grassl in unpublished work [27]. Those occur in dimensions
$d\in\{47,67,259\},$ (9.36)
and in each case, there is a secondary multiplet having the same projection SIC-field as the minimal multiplet (conjecturally corresponding to $\mathcal{O}^{\prime}=\mathcal{O}_{K}$ or $f^{\prime}=1$). Algebraic SICs have not been computed in every dimension up to $d=259$, but rather, Grassl has computed them is certain dimensions (including $d=259$) where the projection field has small enough expected degree to permit the calculation.
It is much easier to write down a list of dimensions $d$ and corresponding order conductors $f^{\prime},2f^{\prime}$ where our conjectures predict degeneration of SIC-fields, that is, where the conditions of 9.8 are satisfied. For $d\leq 500$, equality of ray class fields occurs for
$d\in\{47,67,83,175,211,259,303,339,431,447,467\}.$ (9.37)
This data is consistent with Grassl’s data for SIC-fields. For every such $d$, the conductor pair $(1,2)$ degenerates (i.e., gives the same field ray class field in the manner of 9.8), and an additional conductor pair degenerates in three of these dimensions. The data on ray class field degeneration for $d\leq 500$ is given in Table 2.
$d$ $\Delta_{0}$ $f^{\prime},2f^{\prime}$ $h$ $[E_{d,f^{\prime}}:\mathbb{Q}(\Delta_{0})]$
$47$ $33$ $1,2$ $1$ $1472$
$67$ $71$,2121,21 , 2 $1$ $1452$
$83$ $051$,2121,21 , 2 $2$ $9184$
$175$ $473$ $1,2$ $3$ $43200$
$211$ $689$ $1,2$ $4$ $117600$
$259$ $65$ $1,2$ $2$ $31104$
$303$ $57$ $1,2$ $1$ $40800$
 $d$ $\Delta_{0}$ $f^{\prime},2f^{\prime}$ $h$ $[E_{d,f^{\prime}}:\mathbb{Q}(\Delta_{0})]$ $303$ $75$,105105,105 , 10 $6$ $244800$ $339$ $7851$,2121,21 , 2 $8$ $408576$ $431$ $321$ $1,2$ $3$ $371520$ $431$ $213$,6363,63 , 6 $9$ $1114560$ $447$ $777$ $1,2$ $4$ $355200$ $467$ $377$ $1,2$ $2$ $290784$ $467$ $773$,6363,63 , 6 $8$ $1163136$
Table 2. SIC dims. $d\leq 500$ with an equality of ray class fields of orders $E_{d,f^{\prime}}=E_{d,f^{\prime\prime}}$. The columns show the fund. disc. $\Delta_{0}$, each pair of conductors $(f^{\prime},f^{\prime\prime})=(f^{\prime},2f^{\prime})$ exhibiting field equality, the order class number $h=h_{(f^{\prime})^{2}\Delta_{0}}=h_{(2f^{\prime})^{2}\Delta_{0}}$, and the ray class field extension degree $[E_{d,f^{\prime}}:\mathbb{Q}(\Delta_{0})]=[E_{d,2f^{\prime}}:\mathbb{Q}(\Delta_{0})]$.
10. Comparison of degrees of ray class fields of orders and degrees of SIC-fields
We give formulas for the degrees of the ray class fields of orders $E_{d,f^{\prime}}$ and $\tilde{E}_{d,f^{\prime}}$. These degrees conjecturally coincide with the SIC fields $F^{\rm trip}\!\left(S\right)$ and $F^{\rm proj}\!\left(S\right)$ for those Weyl–Heisenberg SICs $S$ assigned to these orders in 8.4, and we make a numerical comparison in Section 11.
Definition 10.1.
If $K$ is a field and $n\in\mathbb{N}$, define the number field totient function by
$\varphi_{K}(n)=\left|\left(\mathcal{O}_{K}/n\mathcal{O}_{K}\right)^{\times}\right|.$ (10.1)
This function is a natural generalization of Euler’s totient function $\varphi(n)=\varphi_{\mathbb{Q}}(n)$. It is easily seen to be multiplicative by the Chinese remainder theorem. It may be further generalized to $\varphi_{\mathcal{O}}(\mathfrak{a}):=\left|(\mathcal{O}/\mathfrak{a})^{\times}\right|$, for any order $\mathcal{O}$ and any nonzero $\mathcal{O}$-ideal $\mathfrak{a}$, which we do not need here.
Theorem 10.2.
Let $d\geq 4$ and let $\Delta=(d+1)(d-3)=f^{2}\Delta_{0}$ for some fundamental discriminant $\Delta_{0}$. Let $K=\mathbb{Q}(\sqrt{\Delta})$. Let $\varepsilon_{d}=\frac{(d-1)+\sqrt{\Delta}}{2}$, and let $\mathcal{O}^{\prime}$ be a $K$-order such that $\mathbb{Z}[\varepsilon_{d}]\subseteq\mathcal{O}^{\prime}\subseteq\mathcal{O}_{K}$, that is, with $\operatorname{disc}(\mathcal{O}^{\prime})=(f^{\prime})^{2}\Delta_{0}$ for some $f^{\prime}\mid f$. Then
$\left|\operatorname{Cl}_{d\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})\right|=\begin{cases}\dfrac{2^{\left|\Sigma\right|}h_{K}\operatorname{Reg}_{K}\varphi_{K}(df^{\prime})}{9\log(\varepsilon_{d})\varphi(f^{\prime})},&\mbox{if}d\equiv 3\,\left(\operatorname{mod}\,9\right)\mbox{ and }3\mid f^{\prime},\vspace{5pt}\\ \dfrac{2^{\left|\Sigma\right|}h_{K}\operatorname{Reg}_{K}\varphi_{K}(df^{\prime})}{6\log(\varepsilon_{d})\varphi(f^{\prime})},&\mbox{otherwise}.\end{cases}$ (10.2)
If $d$ is even, then $\left|\operatorname{Cl}_{2d\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})\right|=2\left|\operatorname{Cl}_{d\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})\right|$.
Proof.
Let $\tilde{d}=d$ if $d$ is odd and $\tilde{d}\in\{d,2d\}$ if $d$ is even. By [33, Thm. 5.6], the cardinality of the ray class group is
$\left|\operatorname{Cl}_{\tilde{d}\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})\right|=\frac{h_{K}}{\left[\mathcal{O}_{K}^{\times}:\operatorname{U}_{\tilde{d}\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})\right]}\cdot\frac{2^{\left|\Sigma\right|}\left|\left(\mathcal{O}_{K}\middle/\left(\tilde{d}\mathcal{O}^{\prime}:\mathcal{O}_{K}\right)\right)^{\times}\right|}{\left|\operatorname{U}_{\tilde{d}\mathcal{O}^{\prime}}\!\left(\mathcal{O}^{\prime}\middle/\left(\tilde{d}\mathcal{O}^{\prime}:\mathcal{O}_{K}\right)\right)\right|}.$ (10.3)
By 9.4, $\operatorname{U}_{d\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})=\langle\varepsilon_{d}^{3}\rangle$. On the other hand, when $\tilde{d}=2d$,
[…]
$\left[\mathcal{O}_{K}^{\times}:\operatorname{U}_{\tilde{d}\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})\right]=\begin{cases}\left[\langle-1,\varepsilon_{K}\rangle:\langle\varepsilon_{d}^{3}\rangle\right]=\dfrac{2\log(\varepsilon_{d}^{3})}{\log(\varepsilon_{K})}=\dfrac{6\log(\varepsilon_{d})}{\operatorname{Reg}_{K}},&\mbox{if }\tilde{d}=d,\vspace{5pt}\\ \left[\langle-1,\varepsilon_{K}\rangle:\langle\varepsilon_{d}^{6}\rangle\right]=\dfrac{2\log(\varepsilon_{d}^{6})}{\log(\varepsilon_{K})}=\dfrac{12\log(\varepsilon_{d})}{\operatorname{Reg}_{K}},&\mbox{if }\tilde{d}=2d.\end{cases}$ (10.5)
The quotient ideal appearing in (10.3) is $\left(\tilde{d}\mathcal{O}^{\prime}:\mathcal{O}_{K}\right)=\tilde{d}\left(\mathcal{O}^{\prime}:\mathcal{O}_{K}\right)=\tilde{d}f^{\prime}\mathcal{O}_{K}$. Thus,
$\left|\left(\mathcal{O}_{K}\middle/\left(\tilde{d}\mathcal{O}^{\prime}:\mathcal{O}_{K}\right)\right)^{\times}\right|=\left|\left(\mathcal{O}_{K}\middle/\tilde{d}f^{\prime}\mathcal{O}_{K}\right)^{\times}\right|=\varphi_{K}(\tilde{d}f^{\prime}).$ (10.6)
Now we must compute the size of $\operatorname{U}_{\tilde{d}\mathcal{O}^{\prime}}\!\left(\mathcal{O}^{\prime}\middle/\left(\tilde{d}\mathcal{O}^{\prime}:\mathcal{O}_{K}\right)\right)=\operatorname{U}_{\tilde{d}\mathcal{O}^{\prime}}\!\left(\mathcal{O}^{\prime}\middle/\tilde{d}f^{\prime}\mathcal{O}_{K}\right)$. Write $\mathcal{O}_{K}=\mathbb{Z}+\omega\mathbb{Z}$ and $\mathcal{O}^{\prime}=\mathbb{Z}+f^{\prime}\omega\mathbb{Z}$. Thus, $\mathcal{O}^{\prime}/\tilde{d}f^{\prime}\mathcal{O}_{K}=\frac{\mathbb{Z}+f^{\prime}\omega\mathbb{Z}}{\tilde{d}f^{\prime}\mathbb{Z}+\tilde{d}f^{\prime}\omega\mathbb{Z}}$ and $\tilde{d}\mathcal{O}^{\prime}=\tilde{d}\mathbb{Z}+\tilde{d}f^{\prime}\omega\mathbb{Z}$, so the multiplicative group
$\operatorname{U}_{\tilde{d}\mathcal{O}^{\prime}}\!\left(\mathcal{O}^{\prime}\middle/\left(\tilde{d}\mathcal{O}^{\prime}:\mathcal{O}_{K}\right)\right)$ =$=\left\{u\in\frac{\mathbb{Z}+f^{\prime}\omega\mathbb{Z}}{\tilde{d}f^{\prime}\mathbb{Z}+\tilde{d}f^{\prime}\omega\mathbb{Z}}:u\equiv 1\,\left(\operatorname{mod}\,\tilde{d}\mathbb{Z}+\tilde{d}f^{\prime}\omega\mathbb{Z}\right)\right\}$ (10.7)
$\cong\left\{u\in\mathbb{Z}/\tilde{d}f^{\prime}\mathbb{Z}:u\equiv 1\,\left(\operatorname{mod}\,\tilde{d}\right)\right\}.$ (10.8)
[…]
$\left|\operatorname{U}_{\tilde{d}\mathcal{O}^{\prime}}\!\left(\mathcal{O}^{\prime}\middle/\left(\tilde{d}\mathcal{O}^{\prime}:\mathcal{O}_{K}\right)\right)\right|=\begin{cases}\frac{3}{2}\varphi(f^{\prime}),&\mbox{if }d\equiv 3\,\left(\operatorname{mod}\,9\right)\mbox{ and }3\mid f^{\prime},\\ \varphi(f^{\prime}),&\mbox{otherwise}.\end{cases}$ (10.9)
In the case $\tilde{d}=d$, plugging in (10.5), (10.6), and (10.9) to (10.3) yields (10.2). In the case $\tilde{d}=2d$ (with $d$ even), we have $\varphi_{K}(\tilde{d}f^{\prime})=4\varphi_{K}(df^{\prime})$, and again using plugging in (10.5), (10.6), and (10.9) to (10.3) and comparing to (10.2), we see that $\left|\operatorname{Cl}_{2d\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})\right|=2\left|\operatorname{Cl}_{d\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})\right|$. ∎
To use 10.2 for computation, the following elementary lemma is helpful.
Lemma 10.3.
Let $K$ be a quadratic field of discriminant $\Delta_{0}$. Then
$\varphi_{K}(n)=n^{2}\prod_{p|n}\left(1-\frac{1}{p}\right)\left(1-\left(\frac{\Delta_{0}}{p}\right)\frac{1}{p}\right),$ (10.10)
where the product is over primes dividing $n$, and $\left(\frac{\Delta_{0}}{p}\right)$ is the Kronecker symbol.
Proof.
The function $\varphi_{K}(n)$ is multiplicative by the Chinese remainder theorem. It is computed on prime powers by splitting into cases based on the splitting behavior of $p$ in $K$, which is measured by the Kronecker symbol. A straightforward calculation yields (10.10). ∎
Remark 10.4.
Applying the Brauer–Siegel formula to 10.2, one obtains (for $\mathbb{Z}[\varepsilon_{d}]\subseteq\mathcal{O}^{\prime}\subseteq\mathcal{O}_{\mathbb{Q}(\sqrt{\Delta_{0}})}$, $\Delta_{0}=\operatorname{disc}(\mathbb{Q}(\varepsilon_{d}))$, $\operatorname{disc}(\mathcal{O}^{\prime})=(f^{\prime})^{2}\Delta_{0}$) the asymptotic estimate
$\log\left|\operatorname{Cl}_{d\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})\right|\sim\frac{1}{2}\log((f^{\prime})^{2}\Delta_{0})+2\log(d)\mbox{ as }d\to\infty.$ (10.11)
In particular,
$d^{2+o(1)}\ll\left|\operatorname{Cl}_{d\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})\right|\ll d^{3+o(1)},$ (10.12)
with both extremes being achieved for infinitely many pairs $(d,f^{\prime})$.
11. Numerical evidence for the conjectures
We present formal numerical evidence for the conjectures, based on the published tables of Scott and Grassl [39, 38], a paper of Appleby, Chien, Flammia, and Waldron [7], and Flammia’s online tables [22]. The tables of Scott and Grassl contain putatively complete lists of $\operatorname{EC}(d)$-orbits of Weyl–Heisenberg covariant SICs as numerical approximations for $d\leq 90$ and as exact algebraic solutions for $d\leq 14$. The paper [7] produced putatively complete lists of $\operatorname{EC}(d)$-orbits as exact algebraic solutions for all $d\leq 21$. (All of these works also produced many SICs in higher dimensions, without producing a complete list.)
  • For dimensions $d\leq 90$, we check that the size of $\operatorname{Clm}(\mathbb{Z}[\varepsilon_{d}])$ is equal to the number of known $\operatorname{PEC}(d)$-orbits (or geometric equivalence classes) of Weyl–Heisenberg covariant SICs in dimension $d$, as per 5.4. This data was tabulated in Table 1 in Section 5.3.
  • For dimensions $d\leq 15$, we check that the known $\operatorname{PEC}(d)$-orbits and may be assigned to orders $\mathbb{Z}[\varepsilon_{d}]\subseteq\mathcal{O}^{\prime}\subseteq\mathcal{O}_{K}$ in such a way that $\left|\operatorname{Cl}(\mathcal{O}^{\prime})\right|$ orbits are assigned to $\mathcal{O}^{\prime}$, and those orbits form a Galois multiplet, as per 7.2. In addition, the degrees of the class fields $E_{d,f^{\prime}}$ and $\tilde{E}_{d,f^{\prime}}$ over $K=\mathbb{Q}(\sqrt{\Delta_{d}})$ match the degrees of $F^{\rm trip}\!\left(S\right)$ and $F^{\rm proj}\!\left(S\right)$, respectively, as per 8.4. The data is tabulated in Table 3 below.
Isomorphisms of fields have not been checked except in a few small dimensions, due to computational difficulties.
Other sources of evidence also point toward the truth of these conjectures but have not been formally tabulated here. The connection to Stark units supports 8.5, including the compatibility with the Artin map [32, 6, 11]. In all cases tested (including all $d\leq 15$), the cardinality of the subgroup of $\operatorname{PEC}(d)$ fixing the line $\mathbb{C}v$ spanned by a fiducial vector $v$ (called ${\rm S}$ in [39, Table I]) equals the cardinality of the image of corresponding order unit group $(\mathcal{O}^{\prime})^{\times}$ in $(\mathcal{O}^{\prime}/d\mathcal{O}^{\prime})^{\times}$. Finally, work in preparation of Appleby, Flammia, and Kopp will present further evidence and refinements for the conjectures [8].
Orbit labels in the “multiplet” column are based on Flammia’s online table [22], following [39, 38, 7], specifically using the labels for exact SICs, which differ in some cases from Scott and Grassl’s labels for numerical SICs. We recall the definitions of the ray class fields appearing in the table:
  1. (1)
    $E_{d,f^{\prime}}:=H_{d\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}^{\prime}}$, the ray class field having level datum $(\mathcal{O}^{\prime};d\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\})$, where $(d+1)(d-3)=f^{2}\Delta_{0}$ for a fundamental discrimiant $\Delta_{0}$, $f^{\prime}$ is a positive divisor of $f$, and $\mathcal{O}^{\prime}$ is the quadratic order of discriminant $(f^{\prime})^{2}\Delta_{0}$.
  2. (2)
    $\tilde{E}_{d,f^{\prime}}:=H_{d^{\prime}\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}^{\prime}}$, where $d^{\prime}=d$ if $d$ is odd and $d^{\prime}=2d$ if $d$ is even.
8.4 predicts that $F^{\rm trip}\!\left(S\right)=E_{d,f^{\prime}}$ and $F^{\rm proj}\!\left(S\right)=\tilde{E}_{d,f^{\prime}}$.
$d$ $\mathcal{O}^{\prime}$ $f^{\prime}$ $h_{\mathcal{O}^{\prime}}$ $[E_{d,f^{\prime}}:K]$ $[\tilde{E}_{d,f^{\prime}}:K]$ multiplet $[F^{\rm trip}\!\left(S\right):K]$ $[F^{\rm proj}\!\left(S\right):K]$
$4$ $\mathbb{Z}[\frac{1+\sqrt{5}}{2}]$ $1$ $1$ $4$ $8$ 4a $4$ $8$
$5$ $\mathbb{Z}[\sqrt{3}]$ $1$ $1$ $16$ $16$ 5a $16$ $16$
$6$ $\mathbb{Z}[\frac{1+\sqrt{21}}{2}]$ $1$ $1$ $12$ $24$ 6a $12$ $24$
$7$ $\mathbb{Z}[\sqrt{2}]$ $1$ $1$ $12$ $12$ 7b $12$ $12$
$\mathbb{Z}[2\sqrt{2}]$ $2$ $1$ $24$ $24$ 7a $24$ $24$
$8$ $\mathbb{Z}[\frac{1+\sqrt{5}}{2}]$ $1$ $1$ $8$ $16$ 8b $8$ $16$
$\mathbb{Z}[\frac{1+3\sqrt{5}}{2}]$ $3$ $1$ $32$ $64$ 8a $32$ $64$
$9$ $\mathbb{Z}[\sqrt{15}]$ $1$ $2$ $72$ $72$ 9ab $72$ $72$
$10$ $\mathbb{Z}[\frac{1+\sqrt{77}}{2}]$ $1$ $1$ $48$ $96$ 10a $48$ $96$
$11$ $\mathbb{Z}[\sqrt{6}]$ $1$ $1$ $80$ $80$ 11c $\ast$ $80$
$\mathbb{Z}[2\sqrt{6}]$ $2$ $2$ $160$ $160$ 11ab $\ast$ $160$
$12$ $\mathbb{Z}[\frac{1+\sqrt{13}}{2}]$ $1$ $1$ $16$ $32$ 12b $\ast$ $32$
$\mathbb{Z}[\frac{1+3\sqrt{13}}{2}]$ $3$ $1$ $48$ $96$ 12a $\ast$ $96$
$13$ $\mathbb{Z}[\sqrt{35}]$ $1$ $2$ $192$ $192$ 13ab $\ast$ $192$
$14$ $\mathbb{Z}[\frac{1+\sqrt{165}}{2}]$ $1$ $2$ $144$ $288$ 14ab $\ast$ $288$
$15$ $\mathbb{Z}[\sqrt{2}]$ $1$ $1$ $48$ $48$ 15d $\ast$ $48$
$\mathbb{Z}[2\sqrt{2}]$ $2$ $1$ $96$ $96$ 15b $\ast$ $96$
$\mathbb{Z}[4\sqrt{2}]$ $4$ $2$ $192$ $192$ 15ac $\ast$ $192$
Table 3. Table of $d$, orders $\mathcal{O}^{\prime}=\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}}$ with $\Delta=(d+1)(d-3)=f^{2}\Delta_{0}$, class numbers $h_{\mathcal{O}^{\prime}}$, degrees of fields $E_{d,f^{\prime}}$ and $\tilde{E}_{d,f^{\prime}}$, and index of fields $[F^{\rm trip}\!\left(S\right):K]$ and $[F^{\rm proj}\!\left(S\right):K]$. Entries denoted by $\ast$ have not been computed.
12. Concluding Remarks
This paper considered fields that are geometric invariants of a SIC. It can be useful in practice to define SICs using data in smaller subfields that are not geometric invariants. One is the field generated by the entries of a single fiducial projector $\Pi$. That field can be smaller than the full projector field by nearly a factor of $d$ is some special cases, and this is extremely useful for computing exact SICs in some large dimensions that would otherwise be inaccessible to numerical investigation [6, 11]. Another is the field generated by the SIC overlaps $\operatorname{Tr}(\Pi D_{p})$ (or else the closely related SIC overlap phases $\sqrt{d+1}\operatorname{Tr}(\Pi D_{p})$), where $D_{p}$ are the displacement operators.
The overlap field (and the overlap phase field) appear to be ray class fields with ramification at just one of the infinite places $\infty_{1}$, provided one chooses a “strongly centered” fiducial projector; see [9, 20]. That phenomenon relates to the connection to Stark units, which live in a ray class field with ramification at the other infinite place $\infty_{2}$ and are (at least in some cases) Galois conjugate to squares of overlap phases.
Work in preparation by Appleby, Flammia, and Kopp [8] will make the classification scheme suggested by this paper explicit, giving a conjectural construction of each presumptive equivalence class of Weyl–Heisenberg covariant SICs in dimension $d$. Specifically, an explicit construction of a candidate for a map $\Omega$ satisfying the conditions of 8.5 (our strongest conjecture) will be given.
13. Acknowledgements
Work of the first author was supported by NSF DMS grant #2302514. Work of the second author was supported by NSF DMS grant #1701576.
We are grateful to the developers of Magma and in particular David Kohel, whose software package reduced_forms (and whose package class_number, on which reduced_forms relies) we used to compute class numbers of real quadratic orders.
We thank Markus Grassl for finding numerical examples of degeneration of SIC-fields and sharing those examples with us. Professor Grassl’s insights prompted the first author to correct an originally incorrect statement and proof of 9.7.
We thank Marcus Appleby, Steven Flammia, and Gary McConnell for valuable conversations. We also acknowledge the contributions of Steve Donnelly, who independently connected SICs to orders in real quadratic number fields and shared his insights with Marcus Appleby, Steven Flammia, and Markus Grassl through informal discussions.
Appendix A Proof of 9.4
We will now prove 9.4, which we restate as A.1.
Lemma A.1.
Let $d\geq 4$, $\Delta=(d+1)(d-3)$, and $\varepsilon_{d}=\frac{1}{2}(d-1+\sqrt{\Delta})$. Let $\Sigma$ be a subset of the two real embeddings of $\mathbb{Q}(\sqrt{\Delta})$. Fix an order $\mathcal{O}$ satisfying
$\mathcal{O}_{\!\Delta}:=\mathbb{Z}[\varepsilon_{d}]\subseteq\mathcal{O}\subseteq\mathcal{O}_{\mathbb{Q}(\sqrt{\Delta})}.$ (A.1)
Then,
$\operatorname{U}_{d\mathcal{O},\Sigma}(\mathcal{O})=\langle\varepsilon_{d}^{3}\rangle.$ (A.2)
Proof.
Let $K=\mathbb{Q}(\sqrt{\Delta})$. Write $\Delta=f^{2}D$, $f\geq 1$, where $D$ is squarefree, so $K=\mathbb{Q}(\sqrt{D})$ and
$\mathcal{O}_{K}=\left\{\frac{1}{2}(a+b\sqrt{D}):(a,b)\in\mathbb{Z}^{2},\,a\equiv b\,\left(\operatorname{mod}\,2\right),\,a\equiv 0\,\left(\operatorname{mod}\,2\right)\mbox{ if }D\not\equiv 1\,\left(\operatorname{mod}\,4\right)\right\}.$ (A.3)
For the datum $(\mathfrak{m},\Sigma)$ consisting of an ideal $\mathfrak{m}$ of the order $\mathcal{O}$ in $K$ and a set of real places $\Sigma$ of $K$, recall the definition of the unit subgroup
$\operatorname{U}_{\mathfrak{m},\Sigma}(\mathcal{O}):=\left\{\alpha\in\mathcal{O}^{\times}:\alpha\equiv 1\,\left(\operatorname{mod}\,\mathfrak{m}\right),\,\sigma(\alpha)>0\mbox{ for }\sigma\in\Sigma\right\}.$ (A.4)
For A.1, the unit group of interest is $\operatorname{U}_{d\mathcal{O},\Sigma}(\mathcal{O})$. The unit $\varepsilon_{d}$ is totally positive, and
$\varepsilon_{d}^{3}=\frac{1}{2}\left((d-1)(d^{2}-2d-2)+d(d-2)f\sqrt{D}\right)\equiv 1\,\left(\operatorname{mod}\,d\mathcal{O}_{\!\Delta}\right).$ (A.5)
Thus, we have the containment relations
$\langle\varepsilon_{d}^{3}\rangle\subseteq\operatorname{U}_{d\mathcal{O}_{\!\Delta},\{\infty_{1},\infty_{2}\}}\subseteq\operatorname{U}_{d\mathcal{O},\Sigma}(\mathcal{O})\subseteq\operatorname{U}_{d\mathcal{O}_{K}}(\mathcal{O}_{K}).$ (A.6)
These relations show it suffices to prove (A.2) for the special case $\mathcal{O}=\mathcal{O}_{K}$ with $\Sigma=\emptyset$.
The proof studies the minimal unit $\eta>1\in\mathcal{O}_{K}$ such that
$\eta=\pm 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right).$ (A.7)
[…]
Claim 1. We have $\operatorname{Nm}(\eta)=\eta\eta^{\prime}=1$. Consequently $\eta$ is totally positive, and $10<\eta^{\prime}<$ 1.
Proof. We know $\eta\eta^{\prime}=\pm 1$. By hypothesis, $\eta=\pm 1+d\omega$ for some $\omega\in\mathcal{O}_{K}$, and taking algebraic conjugates yields $\eta^{\prime}=\pm 1+d\omega^{\prime}$. Then $\eta\eta^{\prime}=1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$, which forces $\eta\eta^{\prime}=1$. Total positivity and the bound for $\eta^{\prime}$ follow, proving the claim.
Claim 2. For some integer $k\geq 1$, $\eta^{k}=\varepsilon_{d}^{3}.$
Proof. We have $1<\eta\leq\varepsilon_{d}^{3}$, so for some $k\geq 1$, there holds $\eta^{k-1}<\varepsilon_{d}^{3}\leq\eta^{k}$. Then $\omega:=\eta^{k}\varepsilon_{d}^{-3}$ is a unit, $1\leq\omega<\eta$, and $\omega\equiv\pm 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$; hence by minimality of $\eta>1$, we must have $\omega=1$, proving the claim.
Claim 3. $\eta^{k}=\varepsilon_{d}^{3}$ for $k\geq 3$ cannot occur.
Proof. If $k=3$, then $\eta^{3}=\varepsilon_{d}^{3}$, which, since $\eta>1$ forces $\eta=\varepsilon_{d}$. We know $\eta\equiv\pm 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$ and obtain a contradiction by showing that $\varepsilon_{d}\not\equiv\pm 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$. To show the latter, we argue by contradiction and suppose $\varepsilon_{d}\equiv\pm 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$. We know $\varepsilon_{d}\varepsilon_{d}^{\prime}=\eta\eta^{\prime}=1$, which then holds $\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$ and thus implies $\varepsilon_{d}^{\prime}\equiv\pm 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$; hence, $\varepsilon_{d}+\varepsilon_{d}^{\prime}\equiv\pm 2\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$. However, $\varepsilon_{d}+\varepsilon_{d}^{\prime}=\frac{1}{2}(d-1+\sqrt{\Delta})+\frac{1}{2}(d-1+\sqrt{\Delta})=d-1\equiv-1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$, giving a contradiction for $d\geq 4$. Thus $\varepsilon_{d}\not\equiv\pm 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$ holds, and $\eta\neq\varepsilon_{d}$.
If $k\geq 4$, then $\eta=(\varepsilon_{d})^{3/k}<\varepsilon_{d}$ (taking the positive $k$-th root.) Now Claim 1 gives (since $x+\frac{1}{x}$ is an increasing function for $x\geq 1$) that
$11<\eta+\eta^{\prime}=\operatorname{Tr}(\eta)<\varepsilon_{d}+\varepsilon_{d}^{\prime}=\operatorname{Tr}(\varepsilon_{d})=d-$ 1 . (A.8)
The trace is an integer, and
$\operatorname{Tr}(\eta)$ $\equiv\eta+\eta^{\prime}\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)\equiv\pm 2\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right).$ (A.9)
Since $d\geq 4$ and $\operatorname{Tr}(\eta)\in\mathbb{Z}$, the bounds on $\operatorname{Tr}(\eta)$ give
$\operatorname{Tr}(\eta)=\begin{cases}2&\mbox{if }\eta\equiv 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right),\\ d-2&\mbox{if }\eta\equiv-1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right).\end{cases}$ (A.10)
We have two cases.
  • Case 1: Suppose $\operatorname{Tr}(\eta)=2$. Then $\eta+\eta^{\prime}=2$ and $\eta\eta^{\prime}=1$, so $\eta$ is a root of the polynomial $x^{2}-2x+1=(x-1)^{2}$, so $\eta=1$, contradicting $\eta>1$.
  • Case 2: Suppose $\operatorname{Tr}(\eta)=d-2$. We may assume $d\geq 5$, since the case $d=4$ is covered by Case 1. Now
    $d-3<\eta=\operatorname{Tr}(\eta)-\eta^{\prime}<\varepsilon_{d}<\operatorname{Tr}(\varepsilon_{d})=d-1.$ (A.11)
    We have for $d\geq 6$ that
    $(d-3)^{3}<\eta^{3}<\varepsilon_{d}^{3}<(d-1)^{3}<(d-3)^{5}<\eta^{5},$ (A.12)
    since $(d-1)^{3}<(d-3)^{5}$ for $d\geq 6$. Thus $\eta^{k}=\varepsilon_{d}^{3}$ cannot hold. For $d=5$, we have $\eta+\eta^{\prime}=3$ and $\eta\eta^{\prime}=1$, so $\eta$ is a root of the polynomial $x^{3}-3x+1$, so $\eta=\frac{3+\sqrt{5}}{2}$. However, for $d=5$, $\Delta=12$, so $\varepsilon_{5}\in\mathbb{Q}(\sqrt{3})$ cannot be an integer power of $\eta$.
This completes the proof of Claim 3. It remains to rule out $k=2$, where $\varepsilon_{d}^{3}$ is the square of a unit $\tilde{\eta}$. We first classify all $d$ where such a unit exists.
Claim 4. The set of all $d\geq 4$ having a unit $\tilde{\eta}\in\mathcal{O}_{K}$ that satisfies $\tilde{\eta}^{2}=\varepsilon_{d}^{3}$ is:
  1. (1)
    $d=a^{2}-1$ for any $a\geq 3$. Then, $\tilde{\eta}=\pm\eta\varepsilon_{d}$, where
    $\eta=\varepsilon_{a+1}=\frac{a+\sqrt{a^{2}-4}}{2}$ (A.13)
    has $\operatorname{Nm}(\eta)=+1$. Here $\sqrt{\Delta}=a\sqrt{d-3}=a\sqrt{a^{2}-4}$.
  2. (2)
    $d=a^{2}+3$ for any $a\geq 1$. Then, $\tilde{\eta}=\pm\eta\varepsilon_{d}$, where $\eta=\frac{a+\sqrt{a^{2}+4}}{2}$ has $\operatorname{Nm}(\eta)=-1.$ Here $\sqrt{\Delta}=a\sqrt{d+1}=a\sqrt{a^{2}+4}$.
[…]
$\eta^{2}=\frac{1}{4}\left((a^{2}+b^{2}D)+2ab\sqrt{D}\right),$ (A.14)
while $\varepsilon_{d}=\frac{1}{2}(d-1+f\sqrt{D})$. To obtain a solution, we must have
$ab$ $=f;$ (A.15)
$a^{2}+b^{2}D$ $=2(d-1).$ (A.16)
We have $a\neq 0$ since $f\neq 0$, so the first equation gives $b^{2}=\frac{f^{2}}{a^{2}}$, and the second equation can be rewritten (since $\Delta=Df^{2}$) as
$a^{2}+\frac{\Delta}{a^{2}}=2(d-1).$ (A.17)
Introducing the new variable $y=a^{2}$, we find
$y^{2}-2(d-1)y+(d+1)(d-3)=(y-(d+1))(y-(d-3))=0,$ (A.18)
which yields two cases. The first is $y=a^{2}=d+1$, which gives solutions $d=a^{2}-1$ for $a\geq 3$, to make $d\geq 4$. The second is $y=a^{2}-3$, which gives solutions $d=a^{2}+3$ for $a\geq 1$.
  • Case 1: Suppose $d=a^{2}+1$ (so $a\geq 3$). Then $\Delta=a^{2}(a^{2}-4)$, $\sqrt{\Delta}=a\sqrt{a^{2}-4}$, and $a^{2}-4=Db^{2}$. Now
    $\eta$ $=\varepsilon_{a+1}=\frac{1}{2}(a+\sqrt{a^{2}-4});$ (A.19)
    $\varepsilon_{d}$ $=\eta^{2}=\frac{1}{2}(a^{2}-2+a\sqrt{a^{2}-4}).$ (A.20)
    Here $\operatorname{Nm}(\eta)=+1$, and the unit $\tilde{\eta}=\eta\varepsilon_{d}>1$ has $\operatorname{Nm}(\tilde{\eta})=1$.
  • Case 2: Suppose $d=a^{2}-3$, ($a\geq 1$). Then $\Delta=a^{2}(a^{2}+4)$ and $\sqrt{\Delta}=a\sqrt{a^{2}+4}$. So
    $\eta$ $=\frac{1}{2}(a+\sqrt{a^{2}+4});$ (A.21)
    $\varepsilon_{d}$ $=\eta^{2}=\frac{1}{2}(a^{2}+2+a\sqrt{a^{2}+4}).$ (A.22)
    Here $\operatorname{Nm}(\eta)=-1$, and the unit $\tilde{\eta}=\eta\varepsilon_{d}>1$ has $\operatorname{Nm}(\tilde{\eta})=-1$.
This proves Claim 4. To complete the proof, we must rule out the case $k=2$.
Claim 5. $\eta^{k}=\varepsilon_{d}^{3}$ for $k=2$ does not occur.
We must show that all the $\tilde{\eta}$ described by Claim 4 have
$\tilde{\eta}\not\equiv\pm 1\,\left(\operatorname{mod}\,(a^{2}-1)\mathcal{O}_{K}\right).$ (A.23)
The Claim 4 solutions for $d=a^{2}+1$ are ruled out by Claim 1, since $\operatorname{Nm}(\tilde{\eta})=-1$. To rule out the Claim 4 solutions having $d=a^{2}-1$, note that these solutions $\tilde{\eta}$ with $\tilde{\eta}^{2}=\varepsilon_{d}^{3}$ are given by $\tilde{\eta}=e_{a+1}e_{a^{2}-1}=e_{a+1}^{3}$ with $\varepsilon_{a+1}=\frac{1}{2}(a+\sqrt{a^{2}-4})$. We have
$\tilde{\eta}=\frac{1}{2}\left(a(a^{2}-3)+(a^{2}-1)\sqrt{a^{2}-4}\right).$ (A.24)
Now $\Delta=(d+1)(d-3)=a^{2}(a^{2}-4)$; hence, $\Delta=f^{2}D$ has $D\mid(a^{2}-4)$ and $K=\mathbb{Q}(\sqrt{a^{2}-4})$. We assert for $d=a^{2}-1$ that
$\tilde{\eta}\equiv-a\,\left(\operatorname{mod}\,(a^{2}-1)\mathcal{O}_{K}\right).$ (A.25)
The assertion implies $a\pm 1\not\in d\mathcal{O}_{K}$ holds for $d=a^{2}-1$, with $d\geq 3$, which implies Claim 5. It is proved in two cases.
[…]
    $\tilde{\eta}$ $=\frac{1}{2}(a(a^{2}-3)+(a^{2}-1)\sqrt{a^{2}-4}))$ (A.26)
    $\equiv\frac{1}{2}a(a^{2}-3)-\frac{1}{2}(a^{2}-1)$ $\,\left(\operatorname{mod}\,(a^{2}-1)\mathcal{O}_{K}\right)$ (A.27)
    $\equiv\frac{a-1}{2}(a^{2}-1)-a$ $\,\left(\operatorname{mod}\,(a^{2}-1)\mathcal{O}_{K}\right)$ (A.28)
    $\equiv-a$ $\,\left(\operatorname{mod}\,(a^{2}-1)\mathcal{O}_{K}\right).$ (A.29)
  • Case 2. Suppose $a$ is even. Let $a=2a^{\prime}$. Then, $\Delta=a^{2}\Delta_{0}$ with $\Delta_{0}=a^{2}-4=4((a^{\prime})^{2}-1)$, so $K=\mathbb{Q}(\sqrt{(a^{\prime})^{2}-1})$. Now
    $\tilde{\eta}$ $=\frac{a}{2}(a^{2}-3)+(a^{2}-1)\sqrt{(a^{\prime})^{2}-1}$ (A.30)
    $\equiv\frac{a}{2}(a^{2}-1)-a$ $\,\left(\operatorname{mod}\,(a^{2}-1)\mathcal{O}_{K}\right)$ (A.31)
    $\equiv-a$ $\,\left(\operatorname{mod}\,(a^{2}-1)\mathcal{O}_{K}\right).$ (A.32)
We have proved the assertion, hence Claim 5, which completes the proof. ∎

### v3

@@
 Abstract.
-We consider the problem of counting and classifying symmetric informationally complete positive operator-valued measures (SICs or SIC-POVMs), that is, sets of $d^{2}$ equiangular lines in $\mathbb{C}^{d}$.
-For $4\leq d\leq 90$, we show the number of known equivalence classes of Weyl–Heisenberg covariant SICs in dimension $d$ equals the cardinality of the ideal class monoid of (not necessarily invertible) ideal classes in the real quadratic order of discriminant $(d+1)(d-3)$.
-Equivalently, this is the number of $\mathbf{GL}_{2}(\mathbb{Z})$ conjugacy classes in $\mathbf{SL}_{2}(\mathbb{Z})$ of trace $d-1$.
-We conjecture the equality extends to all $d\geq 4$.
-We prove that this conjecture implies more that one equivalence class of Weyl–Heisenberg covariant SICs for every $d>22$.
-Additionally, we refine the “class field hypothesis” of Appleby, Flammia, McConnell, and Yard [10] to predict the exact class field generated by the ratios of vector entries for the equiangular lines defining a Weyl–Heisenberg covariant SIC.
-The class fields conjecturally associated to SICs in dimension $d$ come with a natural partial order under inclusion; we show that the natural inclusions of these fields are strict, except in a small family of cases.
+This paper concerns SIC-POVMs and their relationship to class field theory.
+SIC-POVMs are generalized quantum measurements (POVMs) described by $d^{2}$ equiangular complex lines through the origin in $\mathbb{C}^{d}$.
+Weyl–Heisenberg covariant SICs are those SIC-POVMs described by the orbit a single vector under a finite Weyl–Heisenberg group $\operatorname{WH}(d)$.
+We relate known data on the structure and classification of Weyl–Heisenberg covariant SICs in low dimensions to arithmetic data attached to certain orders of real quadratic fields.
+For $4\leq d\leq 90$, we show the number of known geometric equivalence classes of Weyl–Heisenberg covariant SICs in dimension $d$ equals the cardinality of the ideal class monoid of (not necessarily invertible) ideal classes in the real quadratic order $\mathcal{O}_{\Delta}$ of discriminant $\Delta=(d+1)(d-3)$; we conjecture the equality extends to all $d\geq 4$.
+We prove that this conjecture implies the existence of more than one geometric equivalence class of Weyl–Heisenberg covariant SICs for every $d>22$.
+We conjecture that Galois multiplets of SICs may be put in one-to-one correspondence with the over-orders $\mathcal{O}^{\prime}$ of $\mathcal{O}_{\Delta}$ in such a way that the number of geometric classes in the multiplet equals the ring class number of $\mathcal{O}^{\prime}$.
+We test that conjecture against known data on exact SICs in low dimensions.
+We refine the “class field hypothesis” of Appleby, Flammia, McConnell, and Yard [9] to predict the exact class field over $\mathbb{Q}(\sqrt{(d+1)(d-3)})$ generated by the ratios of vector entries for the equiangular lines defining a Weyl–Heisenberg SIC.
+The refined conjectures use a recently developed class field theory for orders of number fields [34].
+The refined class fields assigned to over-orders $\mathcal{O}^{\prime}$ have a natural partial order under inclusion; the inclusions of these class fields fail to be strict in some cases.
+We characterize such cases and give a table of them for $d<500$.
 Key words and phrases:
-symmetric informationally complete positive operator-valued measure (SIC-POVM), complex equianglular lines, equiangular tight frame (ETF), real quadratic field, non-maximal order, binary quadratic form, ideal class group, explicit class field theory, Stark conjectures
+SIC-POVM, complex equiangular lines, equiangular tight frame (ETF), real quadratic field, non-maximal order, ideal class group, explicit class field theory
+2020 Mathematics Subject Classification:
+11R37 (primary), 11R29, 11R65, 81P15, 81P18, 81R05, 42C15
 1. Introduction
-This paper concerns provable and conjectural properties of SIC-POVMs, which are ensembles of quantum measurements described by $d^{2}$ equiangular complex lines in $\mathbb{C}^{d}$.
-These properties are number-theoretic and relate to class numbers and class fields attached to orders of associated real quadratic number fields.
-A set $S$ of complex lines through the origin in $\mathbb{C}^{d}$, written $S=\{\mathbb{C}\mathbf{v}_{1},\ldots,\mathbb{C}\mathbf{v}_{n}\}\subset\mathbb{P}^{d-1}(\mathbb{C})$, with $\mathbf{v}_{j}$ unit vectors, is equiangular if all the pairwise Hermitian inner products $\langle\mathbf{v}_{i},\mathbf{v}_{j}\rangle$ for $i\neq j$ have the same absolute value $|\langle\mathbf{v}_{i},\mathbf{v}_{j}\rangle|=\alpha\geq 0$.
-In 1975, Delsarte, Goethals, and Seidel [19] proved an upper bound of $n\leq d^{2}$ on the size of any set of complex equiangular lines in $\mathbb{C}^{d}$.
-They showed that if the bound $n=d^{2}$ is attained, then necessarily $\alpha=\frac{1}{d+1}$.
-A symmetric informationally complete positive operator-valued measure or SIC-POVM (henceforth SIC) is a set of quantum measurements that is equivalent to a (maximal) set of $d^{2}$ complex equiangular lines in $\mathbb{C}^{d}$.
-Specifically, for a set of $d^{2}$ equiangular lines $\{\mathbb{C}\mathbf{v}_{1},\ldots,\mathbb{C}\mathbf{v}_{d^{2}}\}$ represented by unit column vectors $\mathbf{v}_{j}\in\mathbb{C}^{d}$, the associated SIC is the set of normalized rank one Hermitian projection operators ${\mathbf{\Pi}}[S]:=\left\{\frac{1}{d}\Pi_{1},\ldots,\frac{1}{d}\Pi_{d^{2}}\right\}$ where $\Pi_{j}=\mathbf{v}_{j}\mathbf{v}_{j}^{\dagger}$.
-(Vectors $\mathbf{v}$ are column vectors, $\mathbf{v}^{\dagger}$ is the conjugate transpose of $\mathbf{v}$, and the Hermitian inner product is $\langle\mathbf{v}_{i},\mathbf{v}_{j}\rangle=\mathbf{v}_{i}^{\dagger}\mathbf{v}_{j}$, following conventions from physics.
-Here, $\mathbf{v}_{i}^{\dagger}\mathbf{v}_{i}=1.$) The SIC condition is equivalent to the following $d^{4}$ identities on the traces of products of the projection operators:
-$\operatorname{Tr}(\Pi_{i}\Pi_{j})$ =$=\begin{cases}1&\mbox{if }i=j,\\ \frac{1}{d+1}&\mbox{if }i\neq j,\end{cases}$ for $1\leq i,j\leq d^{2}$. (1.1)
-We will use the term “SIC” interchangeably to mean either a “set of $d^{2}$ (scaled) Hermitian projection matrices ${\mathbf{\Pi}}(S)$ forming a SIC-POVM” or a “set of $d^{2}$ complex equiangular lines $S$ in $\mathbb{C}^{d}$.” When more precision is needed, we term the latter a line-SIC.
-The name SIC-POVM was introduced in a 2004 paper of Renes, Blume-Kohout, Scott, and Caves [36], whose study was motivated by quantum state tomography, the problem of reconstructing a quantum state from a series of measurements performed on identical states.
-Scott showed that SICs are optimal measurement ensembles for tomography [37]; see also [26].
+SIC-POVMs are generalized quantum measurements (POVMs) described by $d^{2}$ equiangular complex lines through the origin in $\mathbb{C}^{d}$.
+Sets of $d^{2}$ equiangular complex lines in $\mathbb{C}^{d}$ are extremal objects in combinatorial design theory; they give equiangular tight frames [46, Ch.
+12–14].
+The name SIC-POVM (abbreviated ‘SIC’) was introduced in a 2004 paper of Renes, Blume-Kohout, Scott, and Caves [38], whose study was motivated by quantum state tomography, the problem of reconstructing a quantum state from a series of measurements performed on identical states.
+Scott showed that SICs are optimal measurement ensembles for quantum tomography [39]; see also [25].
 SICs have found further applications to quantum information processing [13, 44, 45, 17], quantum foundations [24], compressed sensing for radar [29], and classical phase retrieval [21].
-The object of this paper is to extend the range of conjectures about the structure of SICs proposed by Appleby, Flammia, McConnell, and Yard [9].
-Their conjectures relate Weyl–Heisenberg SICs in $\mathbb{C}^{d}$ (defined in Section 2) for $d\geq 4$ to class field theory over real quadratic fields.
-The ratios of entries of each of the $d^{2}$ vectors in such SICs lie in a number field that is an abelian extension of $K=\mathbb{Q}(\sqrt{\Delta_{d}})$ with $\Delta_{d}=(d+1)(d-3)$.
+At present SICs have been constructed for finitely many dimensions $d$, and conjecturally they exist for all $d$.
+The objects of this paper are to extend the range of number-theoretic conjectures about the existence and structure of SICs proposed by Appleby, Flammia, McConnell, and Yard [8, 9] and to test the refined conjectures against empirical data.
+We also draw consequences of the refined conjectures for counting and classifying equivalence classes of SICs.
+The conjectures of [9] relate Weyl–Heisenberg SICs in $\mathbb{C}^{d}$ (defined in Section 2) for $d\geq 4$ to class field theory over real quadratic fields.
+The ratios of entries of each of the $d^{2}$ vectors in the currently known SICs lie in a number field that is an abelian extension of $K=\mathbb{Q}(\sqrt{\Delta_{d}})$ with $\Delta_{d}=(d+1)(d-3)$.
 Class field theory classifies the abelian extensions of a given number field $K$ in terms of ray class groups, which encode the arithmetic of ideals of the ring $\mathcal{O}_{K}$ of algebraic integers of $K$.
-The number fields associated to individual SICs in dimension $d$ may vary with the SIC, and the subclass of SICs to which the most precise conjecture of [9] applies consists of those giving rise to minimal number field that occurs for a fixed $d$.
+The number fields associated to individual SICs in dimension $d$ may vary with the SIC, and the subclass of SICs to which the most precise conjecture of [9] applies consists of those giving rise to the minimal number field that occurs for a fixed $d$.
 Appleby, Flammia, McConnell, and Yard conjecture that this minimal field is a specific ray class field of $\mathbb{Q}(\sqrt{\Delta_{d}})$; see 1.1.
 The corresponding SICs comprise the minimal multiplet, as defined in Section 2.
-It is one of several multiplets, which are finite sets of geometric equivalence classes of SICs making up a single orbit of a Galois group action on geometric equivalence classes.
-The current paper presents a refinement of these conjectures that accounts for all the remaining multiplets of Weyl–Heisenberg SICs, those having larger associated number fields.
-For $d\geq 4$ the refined conjecture says that each multiplet is indexed by a particular order $\mathcal{O}^{\prime}$ of a real quadratic number field $K_{d}=\mathbb{Q}(\sqrt{\Delta_{d}})$ intermediate between $\mathcal{O}=\mathbb{Z}[\varepsilon_{d}]$ with $\varepsilon_{d}=\frac{1}{2}(d-1+\sqrt{\Delta_{d}})$ and the maximal order $\mathcal{O}_{K_{d}}$.
-Here an order of a $K_{d}$ is a subring of $K_{d}$ whose additive group is free of rank $[K_{d}:\mathbb{Q}]=2$.
-The number field associated to the $\mathcal{O}^{\prime}$-multiplet is conjectured to be a certain ray class field associated to the order $\mathcal{O}^{\prime}$.
-The definition of ray class fields associated to orders originates in the authors’ previous work [33].
+It is one of several Galois multiplets, which are finite sets of geometric equivalence classes of SICs making up a single orbit of a Galois group action.
+The current paper formulates a refinement of these conjectures that accounts for all the remaining Galois multiplets of Weyl–Heisenberg SICs, those having larger associated number fields.
+For $d\geq 4$ the refined conjecture asserts that the multiplets can be indexed bijectively by the the orders $\mathcal{O}^{\prime}$ of the real quadratic number field $K_{d}=\mathbb{Q}(\sqrt{\Delta_{d}})$ intermediate between the order $\mathcal{O}=\mathbb{Z}[\varepsilon_{d}]$ with $\varepsilon_{d}=\frac{1}{2}(d-1+\sqrt{\Delta_{d}})$ and the maximal order $\mathcal{O}_{K_{d}}$ of all algebraic integers in $K_{d}$.
+Here an order $\mathcal{O}$ of $K_{d}$ is a subring of $K_{d}$ which contains $1$ and is of finite index in $\mathcal{O}_{K}$.
+In addition, the number field associated to the $\mathcal{O}^{\prime}$-multiplet is conjectured to be a certain ray class field associated to the order $\mathcal{O}^{\prime}$.
+The notion of ray class fields associated to orders is an extension of the usual class field theory (which is the special case of the maximal order), variants of which were developed (independently) by Campagna and Pengo [12] and the authors [34].
 Inclusions of orders induce reverse inclusion of the corresponding ray class fields, so the minimal multiplet corresponds to the maximal order.
 1.1. SICs
+A set $S$ of complex lines through the origin in $\mathbb{C}^{d}$, written $S=\{\mathbb{C}\mathbf{v}_{1},\ldots,\mathbb{C}\mathbf{v}_{n}\}\subset\mathbb{P}^{d-1}(\mathbb{C})$ with $\mathbf{v}_{j}$ unit vectors, is equiangular if all the pairwise Hermitian inner products $\langle\mathbf{v}_{i},\mathbf{v}_{j}\rangle$ for $i\neq j$ have the same absolute value $\left|\langle\mathbf{v}_{i},\mathbf{v}_{j}\rangle\right|=\alpha\geq 0$.
+In 1975, Delsarte, Goethals, and Seidel [18] proved an upper bound of $n\leq d^{2}$ on the size of any set of complex equiangular lines in $\mathbb{C}^{d}$.
+They showed that if the bound $n=d^{2}$ is attained, then necessarily $\alpha=\frac{1}{\sqrt{d+1}}$.
+A symmetric informationally complete positive operator-valued measure or SIC-POVM (henceforth SIC) is a set of normalized projections onto a (maximal) set of $d^{2}$ complex equiangular lines in $\mathbb{C}^{d}$.
+For $d^{2}$ equiangular lines $\{\mathbb{C}\mathbf{v}_{1},\ldots,\mathbb{C}\mathbf{v}_{d^{2}}\}$ represented by unit column vectors $\mathbf{v}_{j}\in\mathbb{C}^{d}$, the associated SIC is the set of normalized rank one Hermitian projection operators ${\mathbf{\Pi}}(S):=\left\{\frac{1}{d}\Pi_{1},\ldots,\frac{1}{d}\Pi_{d^{2}}\right\}$ where $\Pi_{j}=\mathbf{v}_{j}\mathbf{v}_{j}^{\dagger}$.
+(Vectors $\mathbf{v}$ are column vectors, $\mathbf{v}^{\dagger}$ is the conjugate transpose of $\mathbf{v}$, and the Hermitian inner product is $\langle\mathbf{v}_{i},\mathbf{v}_{j}\rangle=\mathbf{v}_{i}^{\dagger}\mathbf{v}_{j}$, following the quantum mechanics convention established by Dirac [19, Sec.
+1.6].
+Here, $\mathbf{v}_{i}^{\dagger}\mathbf{v}_{i}=1$.) The SIC condition is equivalent to the following $d^{4}$ identities on the traces of products of the projections:
+$\operatorname{Tr}(\Pi_{i}\Pi_{j})$ =$=\begin{cases}1&\mbox{if }i=j,\\ \frac{1}{d+1}&\mbox{if }i\neq j,\end{cases}$ for $1\leq i,j\leq d^{2}$. (1.1)
+We will use the term “SIC” interchangeably to mean either a “set of $d^{2}$ (scaled) Hermitian projection matrices ${\mathbf{\Pi}}(S)$ forming a SIC-POVM” or a “set of $d^{2}$ complex equiangular lines $S$ in $\mathbb{C}^{d}$.” When more specificity is needed, we term the latter a line-SIC.
 Most work on SICs follows Zauner’s 1999 PhD thesis [48, 49] on quantum designs.
@@
 Zauner conjectured, based on his numerical investigations, that SICs exist in every dimension $d$, and moreover, that a SIC in $\mathbb{C}^{d}$ could be found as the orbit of a single vector under the Weyl–Heisenberg group [48, 49]; the latter objects are called Weyl–Heisenberg covariant SICs (or just Weyl–Heisenberg SICs) and are defined and discussed in Section 2.
+A generating vector ${\mathbf{v}}\in\mathbb{C}^{d}$ of such an orbit is called a fiducial vector.
 Weyl–Heisenberg SICs are the main topic of this paper.
 (An extensive treatment of the history of SICs is given by Fuchs, Hoang, and Stacy [23].)
-Over the past 25 years, a large set of data has been computed about SICs in particular dimensions—sometimes approximately and sometime exactly—resulting in extensive datasets [39, 38, 22].
-SICs have been rigorously constructed in a finite set of dimensions, including Weyl–Heisenberg SICs in all dimensions $1\leq d\leq 53$ and the Hoggar lines for $d=8$ (which are not Weyl–Heisenberg covariant).
-Beyond these exact constructions, numerical computations up to $d=90$ suggest there are only finitely many (equivalence classes of) Weyl–Heisenberg SICs in each dimension $d\neq 3$.
-It is known that $d=3$ has infinitely many equivalence classes of Weyl–Heisenberg SICs, and we will not consider it further.
-The current numerical data is believed to account for all Weyl–Heisenberg SICs of dimension $4\leq d\leq 90$.
-There is a geometric action moving these SICs around by the normalizer of the Weyl–Heisenberg group inside the extended unitary group (which is a finite group whose order varies erratically with $d$, while the Weyl–Heisenberg group itself has order $d^{3}$.) In addition, since the (normalized) generating vectors of Weyl–Heisenberg SICs are (empirically) given by algebraic numbers, there is a Galois action, which sometimes moves SICs to new SICs and sometimes moves them to non-SICs.
+Over the past 25 years, a large data set has been computed about SICs in particular dimensions—sometimes approximately and sometime exactly—resulting in extensive datasets [41, 40, 22].
+SICs have been rigorously constructed in a finite set of dimensions, including Weyl–Heisenberg SICs in all dimensions $1\leq d\leq 53$ and the Hoggar lines for $d=8$.
+Beyond these exact constructions, numerical computations in dimensions including every $d\leq 90$ suggest there are only finitely many (equivalence classes of) Weyl–Heisenberg SICs in each dimension $d\neq 3$.
+It is known that $d=3$ has infinitely many equivalence classes of Weyl–Heisenberg SICs, and we will not treat $d=3$ here.
+The current numerical data is believed to account for all Weyl–Heisenberg SICs of dimension $d\leq 90$.
+There is a geometric action moving these SICs around by the normalizer of the Weyl–Heisenberg group inside the extended unitary group (which is a finite group whose order varies erratically with $d$, while the Weyl–Heisenberg group itself has order $d^{3}$ or $2d^{3}$, the latter occurring for even $d$.) In addition, since the (normalized) generating vectors of Weyl–Heisenberg SICs can be (empirically) given by algebraic numbers, there is a Galois action, which sometimes moves SICs to new SICs and sometimes moves them to non-SICs.
 The process of understanding this data has led to formulation of a series of interlocking hypotheses. The two most basic conjectures, which are both unsolved, are:
@@
   2. (2)
-    the finiteness of the set of (equivalence classes of) Weyl–Heisenberg SICs in every fixed dimension $d\neq 3$.
-The finiteness assertion has been a “folklore” conjecture since about 2004 [2] and was observed informally up to dimension $d=50$ by Scott and Grassl (who state they are confident their list is complete for $d\leq 50$ in [39, Sec.
-4]) and up to $d=90$ by Fuchs, Hoang, and Stacy [24, Sec.
-6].
-It was explicitly formulated as a conjecture in the first author’s PhD thesis [31, Conj.
+    the finiteness of the set of Weyl–Heisenberg line-SICs in every fixed dimension $d\neq 3$.
+The finiteness assertion has been a “folklore” conjecture since about 2004.
+It was observed empirically under heuristics up to dimension $d=7$ by Renes, Blume-Kohout, Scott, and Caves [38, Tab.
+I and Sec.
+V], up to dimension $d=50$ by Scott and Grassl (who state they are confident their list is complete for $d\leq 50$ in [41, Sec.
+4]), and up to $d=90$ by Fuchs, Hoang, and Stacy [23, Sec.
+6] using data of Scott [40].
+It was explicitly formulated as a conjecture in the first author’s PhD thesis [32, Conj.
 V.12].
@@
 The finiteness assertion would imply fiducial vectors of Weyl–Heisenberg SICs can be chosen to have all entries algebraic numbers.
-Further study of the algebraic numbers in existing examples has led to surprising discoveries.
+Further study of these algebraic numbers in existing examples has led to surprising discoveries.
 1.2. Connection of SICs to algebraic number theory and class fields
 Recent work has connected SICs to explicit class field theory over real quadratic number fields.
-Appleby, Yadsan-Appleby, and Zauner [4] discovered in 2013 that in all known examples, the algebraic number field $\mathbb{E}$ generated by the entries of the projections ${\mathbf{\Pi}}[S]$ contained the quadratic field $K_{d}=\mathbb{Q}(\sqrt{(d+1)(d-3)})$, and Galois automorphisms often mapped SICs to new SICs.
-Moreover, they observed that $\mathbb{E}$ was an abelian extension of $K_{d}$ in all known examples.
-Such abelian Galois extensions are subfields of ray class fields over $K_{d}$.
+Appleby, Yadsan-Appleby, and Zauner [4] discovered in 2013 that in all known examples, the algebraic number field $\mathbb{E}$ generated by the entries of the projections ${\mathbf{\Pi}}(S)$ contains the quadratic field $K_{d}=\mathbb{Q}(\sqrt{(d+1)(d-3)})$, and Galois automorphisms often map SICs to new SICs.
+Moreover, they observed in all known examples (for $d\neq 3$) that $\mathbb{E}$ is an abelian extension of $K_{d}$.
 In 2016, Appleby, Flammia, McConnell, and Yard [9, 8] made explicit predictions relating the number fields generated by Weyl–Heisenberg SICs to ray class fields.
-Based on extensive data given in [10, Prop.
-1], they formulated the following conjecture [10, Conj.
+Based on extensive data given in [9, Prop.
+1], they formulated the following conjecture [9, Conj.
 2].
@@
   2. (2)
-    Every Weyl–Heisenberg line-SIC $S$ in dimension $d$ has SIC-field $F^{\rm vec}\!\left(S\right)$ that is a finite Galois extension of $\mathbb{Q}$ containing $H_{d^{\prime}\infty_{1}\infty_{2}}$.
-The technical definition of the SIC-field $F^{\rm vec}\!\left(S\right)$ appearing in 1.1 is given in Section 3.1. Ray class fields are discussed in Section 8.1.
+    Every Weyl–Heisenberg line-SIC $S$ in dimension $d$ has SIC field $F^{\rm vec}\!\left(S\right)$ that is a finite Galois extension of $\mathbb{Q}$ containing $H_{d^{\prime}\infty_{1}\infty_{2}}$.
+The technical definition of the SIC field $F^{\rm vec}\!\left(S\right)$ appearing in 1.1 is given in Section 3.1. Ray class fields are discussed in Section 6.1.
 Appleby, Flammia, McConnell, and Yard verified assertion (1) of 1.1 for
 $d\in\{4,5,6,7,8,9,10,11,12,13,14,15,16,17,18,19,20,21,24,28,30,35,39,48\}.$ (1.2)
-They verified assertion (2) of the conjecture for all currently known Weyl–Heisenberg SICs in these dimensions.
+They verified assertion (2) of the conjecture for all known Weyl–Heisenberg SICs in these dimensions.
 They also verified that the fields $F^{\rm vec}\!\left(S\right)$ are all abelian over $K_{d}$ in the known cases.
@@
 […]
-The origin of the discriminant $\Delta_{d}=(d+1)(d-3)$ appearing in the conjecture is not completely understood, but there are some tantalizing hints.
-The appearance of the factor $d+1$ dividing $\Delta_{d}$ seems related to the fact that in the inner products of distinct unit vectors in a SIC are of absolute value $\frac{1}{\sqrt{d+1}}$.
-The appearance of $d-3$ in the conjecture accounts for $d=3$ being special in that $\Delta_{3}=0$.
+The origin of the discriminant $\Delta_{d}=(d+1)(d-3)$ appearing in the conjecture is not completely understood.
+The appearance of the factor $d+1$ dividing $\Delta_{d}$ seems related to the fact that the inner products of distinct unit vectors in a SIC are of absolute value $\frac{1}{\sqrt{d+1}}$.
+The appearance of $d-3$ accounts for $d=3$ being special, in that $\Delta_{3}=0$.
 The values $d=1,2$ have $\Delta_{d}<0$, so $K_{d}=\mathbb{Q}(\sqrt{\Delta_{d}})$ is an imaginary quadratic field, while for $d\geq 4$, $\Delta_{d}$ specifies a (nontrivial) real quadratic field.
-The occurrence of a small unit $\varepsilon_{d}=\frac{1}{2}(d-1+\sqrt{\Delta_{d}})$ (as compared to $\Delta_{d}$, which need not be a fundamental discriminant!) is certainly relevant.
-As noted by previously by the first author [32, p.
-8], skipping repeated fields (e.g., $K_{8}=K_{4}$), the sequence $K_{4},K_{5},K_{6},K_{7},K_{9},K_{10},\ldots$ is precisely the ordering of all real quadratic fields by the size of the smallest nontrivial totally positive unit (which will necessarily be equal to $\varepsilon_{d}$).
+The occurrence of a small unit $\varepsilon_{d}=\frac{1}{2}(d-1+\sqrt{\Delta_{d}})$ (as compared in size to $\Delta_{d}$, which need not be a fundamental discriminant) is certainly relevant.
+The first author previously noted that, on skipping repeated fields (e.g., $K_{8}=K_{4}$), the sequence $K_{4},K_{5},K_{6},K_{7},K_{9},K_{10},\ldots$ gives precisely the ordering of all real quadratic fields by increasing size of the smallest nontrivial totally positive unit (necessarily equal to $\varepsilon_{d}$); see [33, p.
+13819].
 1.3. Main results
-In this paper, we refine 1.1 to conjectures that specify the number fields associated to all the Weyl–Heisenberg SICs in each dimension $d\geq 4$ and which count the number of such SICs associated to each field.
-More precisely, we connect both SICs and fields to certain orders of the real quadratic number field $K_{d}=\mathbb{Q}(\sqrt{\Delta_{d}})$, with $\Delta_{d}=(d+1)(d-3)$.
+In this paper, we refine 1.1 to conjectures that specify the number fields associated to all the Weyl–Heisenberg SICs in each dimension $d\geq 4$, and which also count the number of geometric equivalence classes of such SICs associated to each such field.
+More precisely, we connect both SICs and their associated fields to certain orders of the real quadratic number field $K_{d}=\mathbb{Q}(\sqrt{\Delta_{d}})$, with $\Delta_{d}=(d+1)(d-3)$.
 Write $\Delta_{d}=f^{2}\Delta_{0}$, where $\Delta_{0}$ is the fundamental discriminant $K_{d}$, and $f\geq 1$.
@@
 These orders, arranged by set inclusion, form a distributive lattice, isomorphic to that of the positive divisors of $f$ under divisibility.
-We organize Weyl–Heisenberg SICs into equivalence classes. For a line-SIC $S$,
+We organize Weyl–Heisenberg SICs into equivalence classes. For such a line-SIC $S$,
   1. (1)
-    the geometric equivalence class $[S]$ is the orbit of $S$ under the action of the projective extended Clifford group $\operatorname{PEC}(d)$ (see 2.9), and
+    the geometric equivalence class $[S]$ is the orbit of $S$ under the action of the projective extended Clifford group $\operatorname{PEC}(d)$ (by 2.9), and
   2. (2)
-    the Galois multiplet $[[S]]$ is the set of geometric equivalence classes that are permuted with each other under a certain Galois action (see 4.11).
-In Section 3, we define several fields attached to a line-SIC $S=\{\mathbb{C}\mathbf{v}_{k}:1\leq k\leq d^{2}\}$. These fields are finitely generated extensions of the rational numbers $\mathbb{Q}$.
+    the Galois multiplet $[[S]]$ is the set of geometric equivalence classes that are permuted with each other under a certain Galois action (by 4.9).
+In Sections 3 and 4, we define several fields attached to a line-SIC $S=\{\mathbb{C}\mathbf{v}_{k}:1\leq k\leq d^{2}\}$.
+These fields are finitely generated extensions of the rational numbers $\mathbb{Q}$.
   1. (1)
-    The ratio SIC-field $F^{\rm vec}\!\left(S\right)$ (called the “SIC field” in [10]) is the field generated by all the ratios $\frac{v_{ki}}{v_{kj}}$ of distinct entries ($i\neq j$, $v_{kj}\neq 0$) of the vectors $\mathbf{v}_{k}$ with $\mathbf{v}_{k}\in S$.
+    The ratio SIC field $F^{\rm vec}\!\left(S\right)$ (called the “SIC field” in [9]) is the field generated by all the ratios $\frac{v_{ki}}{v_{kj}}$ of distinct entries ($i\neq j$, $v_{kj}\neq 0$) of the vectors $\mathbf{v}_{k}$ with $\mathbf{v}_{k}\in S$.
   2. (2)
-    The projection SIC-field $F^{\rm proj}\!\left(S\right)$ is the field generated by all the matrix entires of the Hermitian projection operators $\Pi_{i}=\mathbf{v}_{i}\mathbf{v}_{i}^{\dagger}$ of the $d^{2}$ unit vectors $\{{\mathbf{v}}_{i}:1\leq i\leq d^{2}\}$ defining $S$.
-It is shown to be the compositum of $F^{\rm vec}\!\left(S\right)$ with its complex conjugate field $\overline{F^{\rm vec}\!\left(S\right)}$ (in 3.3).
+    The projection SIC field $F^{\rm proj}\!\left(S\right)$ is the field generated by the matrix entries of the Hermitian projection operators $\Pi_{i}=\mathbf{v}_{i}\mathbf{v}_{i}^{\dagger}$ of the $d^{2}$ unit vectors $\{{\mathbf{v}}_{i}:1\leq i\leq d^{2}\}$ defining $S$.
   3. (3)
-    The extended projection SIC-field $F^{\rm eproj}\!\left(S\right)=F^{\rm proj}\!\left(S\right)\!(-e^{\frac{\pi i}{d}})$ is obtained by adjoining the root of unity $-e^{\frac{\pi i}{d}}$.
-For Weyl–Heisenberg SICs, we show $F^{\rm eproj}\!\left(S\right)$ is an invariant of the geometric equivalence class $[S]$ (in 3.8), and we show the index $[F^{\rm eproj}\!\left(S\right):F^{\rm proj}\!\left(S\right)]\in\{1,2\}$ and is $1$ if $d$ is odd (in 3.6).
+    The extended projection SIC field $F^{\rm eproj}\!\left(S\right)=F^{\rm proj}\!\left(S\right)\!(\xi_{d})$ is obtained by adjoining the root of unity $\xi_{d}=-e^{\frac{\pi i}{d}}$.
   4. (4)
-    For SICs defined over a finite extension of $\mathbb{Q}$, we define the Galois projection SIC-field $F^{\rm Gproj}\!\left(S\right)$ to be the normal closure of $F^{\rm eproj}\!\left(S\right)$, which is an invariant of the multiplet $[[S]]$.
-We show that the subgroup of $\operatorname{Gal}(F^{\rm Gproj}\!\left(S\right)/\mathbb{Q})$ consisting of elements that commute with complex conjugation sends SICs to other SICs with the same Galois projection SIC-field (in 4.8).
-We have the field inclusions
+    For algebraic SICs (ones with $F^{\rm vec}\!\left(S\right)$ a finite extension of $\mathbb{Q}$), the Galois projection SIC field $F^{\rm Gproj}\!\left(S\right)$ is the normal closure of $F^{\rm eproj}\!\left(S\right)$.
+We have the following field inclusions, with the last being valid only for algebraic SICs:
 $F^{\rm vec}\!\left(S\right)\subseteq F^{\rm proj}\!\left(S\right)\subseteq F^{\rm eproj}\!\left(S\right)\subseteq F^{\rm Gproj}\!\left(S\right).$ (1.3)
-In 4.9, we show that 1.1 implies that
+The field $F^{\rm Gproj}\!\left(S\right)$ is an invariant of the Galois multiplet $[[S]]$ of an algebraic SIC $S$. In 4.8, we show that 1.1 implies the equalities
 $F^{\rm vec}\!\left(S\right)=F^{\rm proj}\!\left(S\right)=F^{\rm eproj}\!\left(S\right)=F^{\rm Gproj}\!\left(S\right),$ (1.4)
 whence $F^{\rm vec}\!\left(S\right)$ is a finite Galois extension of $\mathbb{Q}$.
-The refinements of 1.1 formulated in this paper posit that Weil–Heisenberg SICs in dimension $d\neq 3$ occur in multiplets indexed by the intermediate orders $\mathcal{O}^{\prime}$ of $K_{d}$ the lie between $\mathcal{O}=\mathbb{Z}[\varepsilon_{d}]$ and $\mathcal{O}_{K_{d}}$.
-Each of these orders is in turn indexed by its conductor $f^{\prime}$, a positive integer such that $\operatorname{disc}(\mathcal{O}^{\prime})=(f^{\prime})^{2}\Delta_{0}$ where $\Delta_{0}=\operatorname{disc}(\mathcal{O}_{K_{d}})$.
-The conductors appearing are the positive divisors of the conductor $f$ of $\mathcal{O}$.
+The refinements of 1.1 formulated in this paper assert that Weyl–Heisenberg SICs in dimension $d\neq 3$ occur in Galois multiplets indexed by the intermediate orders $\mathcal{O}^{\prime}$ of $K_{d}$ that lie between $\mathcal{O}_{\Delta_{d}}=\mathbb{Z}[\varepsilon_{d}]$ and $\mathcal{O}_{K_{d}}$.
+Each intermediate order is in turn labeled by its conductor $f^{\prime}$, a positive integer such that $\operatorname{disc}(\mathcal{O}^{\prime})=(f^{\prime})^{2}\Delta_{0}$ where $\Delta_{0}=\operatorname{disc}(\mathcal{O}_{K_{d}})$.
+Quadratic orders are uniquely specified by their discriminants, so we may write $\mathcal{O}^{\prime}=\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}}$.
+The conductors appearing are the positive divisors $f^{\prime}$ of the conductor $f$ of $\mathcal{O}$.
 Conjecture 1.2 (Order-to-Multiplet Conjecture).
-Fix a positive integer $d\neq 3$, let $\Delta=\Delta_{d}=(d+1)(d-3)$, and write $\Delta_{d}=f^{2}\Delta_{0}$ for $\Delta_{0}=\operatorname{disc}(\mathcal{O}_{K_{d}})$ and some positive integer $f$.
-Then there is a bijection $\mathcal{M}$ between conductors of intermediate orders and Galois multiplets,
+Fix a positive integer $d\neq 3$, let $\Delta=\Delta_{d}=(d+1)(d-3)$, and write $\Delta=f^{2}\Delta_{0}$ for $\Delta_{0}=\operatorname{disc}(\mathcal{O}_{K_{d}})$ and some positive integer $f$.
+Then there is a bijection $\mathcal{M}$ between intermediate orders $\mathcal{O}^{\prime}=\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}}$, labeled by their conductors $f^{\prime}$, and Galois multiplets,
 […]
@@
   • (1)
-    The number of geometric equivalence classes $[S]$ in the multiplet $\mathcal{M}(f^{\prime})=[[S]]$ is $\left|\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})\right|$, the class number of the invertible class group of the order $\mathcal{O}^{\prime}$ of discriminant $(f^{\prime})^{2}\Delta_{0}$.
+    The number of geometric equivalence classes $[S]$ in the multiplet $\mathcal{M}(f^{\prime})=[[S]]$ is the class number $h_{(f^{\prime})^{2}\Delta_{0}}:=\left|\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})\right|$ of the the order $\mathcal{O}^{\prime}$ of discriminant $(f^{\prime})^{2}\Delta_{0}$.
   • (2)
-    If $f_{1}\mid f_{2}\mid f$, $[S_{1}]\in\mathcal{M}(f_{1})$, and $[S_{2}]\in\mathcal{M}(f_{2})$, then $F^{\rm vec}\!\left(S_{1}\right)\subseteq F^{\rm vec}\!\left(S_{2}\right)$.
-A mildly strengthened version of this conjecture is stated as 7.2.
-The field inclusion property (2) puts a strong restriction on the map $\mathcal{M}$.
-Modulo the next conjecture, it determines $\mathcal{M}$ entirely for a fraction $\frac{47}{48}$ of all dimensions $d$ (in the sense of 9.10).
-The following conjecture is an extension of 1.1 using the notion of ray class fields of an order $\mathcal{O}^{\prime}$ introduced in [33] and reviewed in Section 8.1.
+    If $f_{1}\mid f_{2}$ and $f_{2}\mid f$, and $[S_{1}]\in\mathcal{M}(f_{1})$, and $[S_{2}]\in\mathcal{M}(f_{2})$, then $F^{\rm vec}\!\left(S_{1}\right)\subseteq F^{\rm vec}\!\left(S_{2}\right)$.
+The special case $f^{\prime}=1$ corresponds to 1.1, where the order $\mathcal{O}^{\prime}$ is the maximal order.
+Appleby, Flammia, McConnell, and Yard [8] call the associated Galois multiplet $[[S]]$ the minimal multiplet, as it has the smallest associated field $F^{\rm vec}\!\left(S\right)$ under inclusion.
+The field inclusions asserted in 1.1(2) imply the truth of 1.2(2) in the special case $f_{1}=1$.
+The following conjecture is an extension of both 1.1 and 1.2, which uses the notion of ray class fields of an order $\mathcal{O}^{\prime}$ introduced in [34] and reviewed in Section 6.1.
 Conjecture 1.3 (Ray Class Fields of Orders Conjecture).
 Let $d\geq 4$.
-Then for any Weyl–Heisenberg line-SIC $S$ in dimension $d$ with associated order $\mathcal{O}^{\prime}$ having discriminant $(f^{\prime})^{2}\Delta_{0}$ (that is, with $\mathcal{M}(f^{\prime})=[[S]]$),
-$F^{\rm vec}\!\left(S\right)=H_{d^{\prime}\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}^{\prime}}$ (1.6)
+Then the map $\mathcal{M}$ in 1.2 may be chosen so that, for any Weyl–Heisenberg line-SIC $S$ in dimension $d$ with associated order $\mathcal{O}^{\prime}$ having discriminant $(f^{\prime})^{2}\Delta_{0}$ (that is, with $\mathcal{M}(f^{\prime})=[[S]]$), one has
+$F^{\rm vec}\!\left(S\right)=F^{\rm proj}\!\left(S\right)=H_{d^{\prime}\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}^{\prime}},$ (1.6)
 where $H_{d^{\prime}\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}^{\prime}}$ is the ray class field having level datum $(\mathcal{O}^{\prime};d^{\prime}\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\})$, $d^{\prime}=2d$ if $d$ is even, and $d^{\prime}=d$ if $d$ is odd.
-The special case $f^{\prime}=1$ corresponds to 1.1, where the order $\mathcal{O}^{\prime}$ is the maximal order.
-The field inclusions asserted in 1.1(2) follow from 1.2.
-Appleby, Flammia, McConnell, and Yard [9] call the associated Galois multiplet $[[S]]$ the minimal multiplet, as it has the smallest associated field $F^{\rm vec}\!\left(S\right)$ under inclusion.
-8.4 gives a further refinement of 1.3; these conjectures refine and revise [31, Conj.
-V.17].
-The paper describes numerical checks of predictions of these conjectures against numerical data.
-The data for $4\leq d\leq 90$ exhibits patterns consistent with grouping SICs into multiplets indexed by orders.
-We formulate several narrower conjectures that make numerical predictions that can be checked against existing data and give tables of comparisons.
-We also derive some consequences of the refined conjectures.
-In Section 6, we show they imply that the number of distinct geometric equivalence classes $[S]$ will be larger than $1$ for all $d\geq 23$ and will grow like $d^{1+o(1)}$ as $d\to\infty$.
-(See 6.3 and 6.4.)
+The paper describes numerical checks of predictions of these conjectures against numerical data for SICs obtained in the datasets [41, 40, 22, 23].
+The numerical data include both exact Weyl–Heisenberg SICs given by algebraic fiducial vectors and numerical SICs given as rational vectors approximately satisfying the fiducial vector SIC relations (to many decimal places).
+The notion of geometric equivalence class makes sense for numerical SICs, because they have well-defined $\operatorname{PEC}(d)$-orbits; however, the notion of Galois multiplet is not well-defined for them.
+In consequence we cannot test 1.2 using numerical SICs, since its statement involves Galois multiplets.
+Instead, we extract a numerical prediction implied by 1.2 that is testable.
+This numerical prediction counts the total number of geometric equivalence classes of Weyl–Heisenberg SICs in dimension $d$ and is formulated as 5.4.
+It says that the total number of geometric equivalence classes of Weyl–Heisenberg SICs in dimension $d\geq 4$ is $\left|\operatorname{Clm}(\mathcal{O}_{\Delta_{d}})\right|$, the cardinality of the class monoid of the order $\mathcal{O}_{\Delta_{d}}$, as defined in Section 5.1.
+We formulate this prediction as 5.4, and we test it in 5.7 using both exact and numerical SICs.
+These tests are summarized by Table 5.1 and Table 5.2 for all dimensions $4\leq d\leq 90$.
+5.4 gives agreement in all these dimensions, assuming that the list of exact plus numerical SICs is complete.
+(The number $\left|\operatorname{Clm}(\mathcal{O}_{\Delta_{d}})\right|$ is equal to a sum of the ring class numbers $h_{(f^{\prime})^{2}\Delta_{0}}$ appearing in 1.2; see 5.3.)
+The existence of the bijection $\mathcal{M}$ asserted in 1.2 implies a numerical prediction concerning the number of Galois multiplets, predicting that it equals the number of quadratic orders between $\mathcal{O}_{\Delta_{d}}$ and $\mathcal{O}_{K_{d}}$, which is the number of over-orders of $\mathcal{O}_{\Delta_{d}}$ in $K=\mathbb{Q}(\sqrt{\Delta_{d}})$.
+This numerical prediction is formulated as 5.8 and tested in 5.9 against certain dimensions $d$ in which all known SICs are exact SICs.
+Supporting data is given in Table 5.1.
+In Section 6, we provide further tests of 1.2(2) and 1.3.
+We first summarize results on ray class fields of orders required to understand the conjectures.
+In 6.6, we numerically test the equality of degrees over $\mathbb{Q}$ of the ray class fields of orders in 1.3 to those of the fields $F^{\rm vec}\!\left(S\right)$.
+Tests are performed for exact SICs in dimensions $4\leq d\leq 15$ and $d=35$, and a supporting table of fields is given in Table 6.1.
+We display for $d=35$ in 5.10 and 6.7 a (unique) bijective map $\mathcal{M}$ of orders to Galois multiplets that fulfills both conditions (1) and (2) of 1.2.
+In Section 8 we show that 1.3 strongly restricts the map $\mathcal{M}$ and, according to results proved elsewhere [36], determines it uniquely for a fraction $\frac{47}{48}$ of dimensions $d$.
+We show in 8.2 that the set of $d$ for which there exist two distinct orders $\mathcal{O}^{\prime},\mathcal{O}^{\prime\prime}$ with identical $H_{d^{\prime}\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}^{\prime}}$ in 1.3 is the set of $d$ for which the squarefree part of $(d+1)(d-3)$ is 1(mod $1\,\left(\operatorname{mod}\,8\right)$.
+Then 8.3 computes the asymptotic density of this set to be $\frac{1}{48}$.
+We present Table 8.1, which lists all cases of equality of such orders occurring for $d\leq 500$.
+These equality cases include all cases of observed equality of SIC fields $d\in\{47,67,259\}$ by Grassl [26].
 1.4. Notation
   • Boldface roman upper case letters represent continuous groups, e.g., $\mathbf{EU}(d)$.
-  • Boldface roman lower case letters represent column vectors in $\mathbb{C}^{d}$. For a complex vector, $\mathbf{w}^{\dagger}:=\overline{\mathbf{w}}^{T}$, and similarly for matrices.
+  • Boldface roman lower case letters $\mathbf{v}=(v_{0},v_{1},\cdots v_{d-1})^{T}$ represent column vectors in $\mathbb{C}^{d}$.
+For a complex vector, $\mathbf{w}^{\dagger}:=\overline{\mathbf{w}}^{T}$, and similarly for matrices.
   • Upper case roman letters represent $d\times d$ complex matrices (e.g., $X$, $Z$); they also appear in names of finite matrix groups (e.g., $\operatorname{WH}(d)$) and sets (e.g., $\operatorname{WHSIC}(d)$).
@@
 2. Weyl–Heisenberg SICs
-In this section, we defined and give background on Weyl–Heisenberg (covariant) SICs, which satisfy a particular group symmetry property.
+In this section, we define and give background on Weyl–Heisenberg (covariant) SICs, which satisfy a particular group symmetry property.
 2.1. Weyl–Heisenberg SICs
-Specifically, Weyl–Heisenberg line-SICs will be defined as the orbit of a single line under the action of the Weyl-Heisenberg group, a finite subgroup of the group of $d\times d$ unitary matrices.
+Specifically, Weyl–Heisenberg line-SICs will be defined as the orbit of a single line under the action of the Weyl–Heisenberg group, a finite subgroup of the group of $d\times d$ unitary matrices.
 Definition 2.1.
@@
 $X$ $=\left(\begin{array}[]{ccccc}0&0&\cdots&0&1\\ 1&0&\cdots&0&0\\ 0&1&\cdots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&1&0\end{array}\right),$ $Z$ $=\left(\begin{array}[]{ccccc}1&0&0&\cdots&0\\ 0&\zeta&0&\cdots&0\\ 0&0&\zeta^{2}&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\cdots&\zeta^{d-1}\end{array}\right),$ (2.11)
-The projective Weyl–Heisenberg group $\operatorname{PWH}(d)$ is defined to be $\operatorname{WH}(d)$ modulo its scalar matrix subgroup $Z(\operatorname{WH}(d))=\{\xi^{j}I:\,j\geq 0\}$, where $I$ is the $d\times d$ identity matrix.
+The projective Weyl–Heisenberg group $\operatorname{PWH}(d)$ is defined to be $\operatorname{WH}(d)$ modulo its scalar matrix subgroup $Z(\operatorname{WH}(d))=\{\xi^{j}I:\,j\geq 0\}$, where $I$ is the $d\times d$ identity matrix, $X$ is called the cyclic shift matrix, and $Z$ is called the modulation matrix [46, Sec.
+14.5].
 It is easily checked that $ZX=\zeta XZ$.
@@
 $\operatorname{orb}_{\operatorname{WH}}(\mathbf{w}):=\{H\mathbf{w}:H\in\operatorname{WH}(d)\}.$ (2.12)
-Its associated set of complex lines, denoted $\operatorname{WHL}(\mathbf{w})$ is
+Its associated set of complex lines, denoted $\operatorname{WHL}(\mathbf{w})$, is
 $\operatorname{WHL}(\mathbf{w}):=\{[\mathbf{w}]:\mathbf{w}\in\operatorname{orb}_{\operatorname{WH}}(\mathbf{w})\}=\{[H\mathbf{w}]:H\in\operatorname{PWH}(d)\}.$ (2.13)
@@
 We call a generating vector $\mathbf{v}$ of a Weyl–Heisenberg SIC a fiducial vector.
-Such a vector is well-defined up to multiplication by a complex scalar $c\in\mathbb{C}^{\times}$.
+Such a vector is well-defined up to multiplication by a complex scalar $z\in\mathbb{C}^{\times}$.
 Each SIC has a choice of $d^{2}$ fiducial vectors, up to scalars.
-We call two Weyl–Heisenberg SICs scalar-equivalent if they have a common fiducial vector, up to multiplication by a nonzero scalar.
 Conjecture 2.4 (Strong Zauner Conjecture).
 For each dimension $d\geq 1$, there exists at least one Weyl–Heisenberg SIC.
-This conjecture is currently known, by explicit constructions to hold for finitely many dimensions $d$. At present it includes dimensions $1\leq d\leq 53$ and a extensive number of higher dimensions.
+This conjecture is currently known by explicit constructions to hold for finitely many dimensions $d$, including $1\leq d\leq 53$ and a extensive number of higher dimensions.
 Conjecture 2.5 (Weyl–Heisenberg Boundedness Conjecture).
-For each dimension $d\geq 1$, except $d=3$, there are finitely many scalar-equivalence classes of Weyl–Heisenberg SICs.
-There are a few small dimensions where this conjecture has been proved rigorously, including $d\in\{1,2,4,5\}$.
-In higher dimensions, up to at least $d=90$, it is supported by non-rigorous evidence given by numerical optimization routines repeatedly returning the same SICs [39, 38].
-In dimension $3$, there is a one-parameter family of inequivalent Weyl–Heisenberg SICs giving uncountably many distinct scalar-equivalence classes of SICs.
+For each dimension $d\geq 1$, except $d=3$, there are finitely many Weyl–Heisenberg SICs.
+There are a few small dimensions where this conjecture has been proved rigorously, including $d\in\{1,2,4,5,6\}$.
+The cases $d=1,2$ may be done by hand; the cases $d=4,5,6$ have been handled by computer using Gröbner bases [27, unpublished].
+In higher dimensions, up to at least $d=50$, it is supported by non-rigorous evidence given by numerical optimization routines repeatedly returning the same SICs [41, 40]; see Table 5.1.
+In dimension $3$, there is a one-parameter family of inequivalent Weyl–Heisenberg SICs.
 A closely related conjecture is stated in Waldron [46, Sec.
-14.27]: The (real) algebraic variety of (normalized Weyl–Heisenberg) SIC fiducials is zero-dimensional, except for $d=3$, where it is one-dimensional.
-This conjecture implies 2.5; it is equivalent if one takes “dimension” to mean “real dimension” (i.e., dimension of the Zariski closure) and is a stronger conjecture if one takes “dimension” to mean “dimension of the complex variety defined by the same equations.” Resolving 2.5 or Waldron’s conjecture in general (and specifically even for $n=6$) is an interesting open problem in optimization and real algebraic geometry.
+14.27] as follows: The algebraic variety of SIC fiducials is zero dimensional (and nonempty), except for $d=3$, where it is one dimensional.
+Resolving 2.5 or Waldron’s conjecture in general (and specifically even for $n=7$) is an interesting open problem in optimization and real algebraic geometry.
 2.2. Geometric equivalence classes of Weyl–Heisenberg SICs
 There are two types of geometric operators acting on $\mathbb{C}^{d}$ that send SICs to SICs.
-Firstly, if $U$ is a unitary matrix in $\mathbf{U}(d)=\{U\in\mathbf{GL}(\mathbb{C}^{d}):UU^{\dagger}=1\}$ and $S=\{\mathbb{C}\mathbf{v}_{1},\ldots,\mathbb{C}\mathbf{v}_{d^{2}}\}$ is a SIC, then so is $US=\{\mathbb{C}U\mathbf{v}_{1},\ldots,\mathbb{C}U\mathbf{v}_{d^{2}}\}$.
+Firstly, if $U$ is a unitary matrix in $\mathbf{U}(d)=\{U\in\mathbf{GL}(\mathbb{C}^{d}):UU^{\dagger}=1\}$, and $S=\{\mathbb{C}\mathbf{v}_{1},\ldots,\mathbb{C}\mathbf{v}_{d^{2}}\}$ is a SIC, then so is $US=\{\mathbb{C}U\mathbf{v}_{1},\ldots,\mathbb{C}U\mathbf{v}_{d^{2}}\}$.
 (This is a unitary equivalence.) Secondly, if $C_{d}$ is the complex conjugation operator acting pointwise on $\mathbb{C}^{d}$, so that $C_{d}\mathbf{v}:=\overline{\mathbf{v}}$, then $C_{d}$ also preserves the SIC property, as does any “antiunitary” operator of the form $C_{d}U$.
@@
 Definition 2.6.
-The extended unitary group is $\mathbf{EU}(d):=\mathbf{U}(d)\sqcup C_{d}\mathbf{U}(d)$.
-The projective extended unitary group $\mathbf{PEU}(d)$ is defined to be $\mathbf{EU}(d)$ modulo its scalar matrix subgroup $\{zI:z\in\mathbb{C},\left|z\right|=1\}$.
+The extended unitary group is $\mathbf{EU}(d):=\mathbf{U}(d)\sqcup C_{d}\mathbf{U}(d)$. The projective extended unitary group $\mathbf{PEU}(d):=\mathbf{EU}(d)/\{zI:z\in\mathbb{C},\left|z\right|=1\}$.
 All known SICs are group covariant, meaning that they are the orbit of a vector under the action of a finite subgroup of the unitary group.
@@
 The Hoggar lines are group covariant for the tensor product $\operatorname{WH}(2)\otimes\operatorname{WH}(2)\otimes\operatorname{WH}(2)$.
-(The group $\operatorname{WH}(2)\otimes\operatorname{WH}(2)\otimes\operatorname{WH}(2)$ may itself be thought of as a generalized (Weyl)-Heisenberg group.)
+(The group $\operatorname{WH}(2)\otimes\operatorname{WH}(2)\otimes\operatorname{WH}(2)$ may itself be thought of as a generalized (Weyl)–Heisenberg group.)
 The extended unitaries $U\in\mathbf{EU}(d)$ that preserve the Weyl–Heisenberg covariance property are precisely those for which $U\operatorname{WH}(d)U^{-1}=\operatorname{WH}(d)$, that is, elements of the extended Clifford group, defined as follows.
@@
 The Clifford group $\operatorname{C}(d)$ is infinite because it contains a copy of $\mathbf{U}(1)$ as scalar matrices $zI$ with $\left|z\right|=1$; thus, $\operatorname{EC}(d)$ is infinite.
-However, $\operatorname{PC}(d)$ and $\operatorname{PEC}(d)$ are finite groups.
+The groups $\operatorname{PC}(d)$ and $\operatorname{PEC}(d)$ are finite groups.
 (The extended Clifford $\operatorname{EC}(d)$ group is studied in connection with SICs by Appleby [1].)
@@
 In this section, we define several number fields attached to arbitrary SICs.
-Many of these fields have been studied before, but they are often discussed within the context of assuming various conjectures about SICs.
+Many of these fields have been studied before, often within the context of assuming various conjectures about SICs.
 Here, we focus on relationships between “SIC fields” that can be proven rigorously, either in general (Section 3.1) or in the case of Weyl–Heisenberg SICs (Section 3.2).
@@
 3.1. Fields generated by general SICs
-The fields we will study here are each associated to an individual SIC $S$ and depend only on data computed from the orthogonal projections $\{\Pi_{i}:1\leq i\leq d^{2}\}$ defining the SIC.
+The fields we study here are each associated to an individual SIC $S$ and depend only on data computed from the orthogonal projections $\{\Pi_{i}:1\leq i\leq d^{2}\}$ defining the SIC.
 Definition 3.1.
 Let ${\mathbf{\Pi}}(S)=\{\frac{1}{d}\Pi_{1},\ldots,\frac{1}{d}\Pi_{d^{2}}\}$ be an arbitrary SIC in dimension $d$, with each $\Pi_{i}=\mathbf{v}_{i}\mathbf{v}_{i}^{\dagger}$ for a unit vector $\mathbf{v}_{i}$.
-(Thus each $\mathbf{v}_{i}$ is a unit fiducial vector, unique up to multiplication by a unit scalar $e^{i\theta}$.)
+(Thus each $\mathbf{v}_{i}$ is unique up to multiplication by a unit scalar.)
   • (1)
-    The ratio SIC-field $F^{\rm vec}\!\left(S\right)$ is the field extension of $\mathbb{Q}$ generated by all the ratios $v_{ij}/v_{ik}$ (for $v_{ik}\neq 0$) of the entries of the $d^{2}$ fiducial vectors $\mathbf{v}_{i}$ of $S$.
+    The ratio SIC field $F^{\rm vec}\!\left(S\right)$ is the field extension of $\mathbb{Q}$ generated by all the ratios $v_{ij}/v_{ik}$ (for $v_{ik}\neq 0$) of the entries of the $d^{2}$ fiducial vectors $\mathbf{v}_{i}$ of $S$.
   • (2)
-    The triple product SIC-field $F^{\rm trip}\!\left(S\right)$ is the field extension of $\mathbb{Q}$ generated by all the “triple products” $\operatorname{Tr}(\Pi_{i}\Pi_{j}\Pi_{k})$.
+    The triple product SIC field $F^{\rm trip}\!\left(S\right)$ is the field extension of $\mathbb{Q}$ generated by all the triple products $\operatorname{Tr}(\Pi_{i}\Pi_{j}\Pi_{k})$ (also called $3$-vertex Bargmann invariants [46, Sec.
+8.2]).
   • (3)
-    The unitary invariant SIC-field $F^{\rm inv}\!\left(S\right)$ is the field extension of $\mathbb{Q}$ generated by all the unitary invariants $\operatorname{Tr}(\Pi_{i_{1}}\Pi_{i_{2}}\cdots\Pi_{i_{n}})$ for $n\in\mathbb{N}$.
+    The unitary invariant SIC field $F^{\rm inv}\!\left(S\right)$ is the field extension of $\mathbb{Q}$ generated by all the unitary invariants $\operatorname{Tr}(\Pi_{i_{1}}\Pi_{i_{2}}\cdots\Pi_{i_{n}})$ for $n\in\mathbb{N}$ (also called $n$-vertex Bargmann invariants [46, Sec.
+8.2]).
   • (4)
-    The projection SIC-field $F^{\rm proj}\!\left(S\right)$ is the field extension of $\mathbb{Q}$ generated by all the entries of the $d^{2}$ Hermitian projection matrices $\Pi_{j}$.
+    The projection SIC field $F^{\rm proj}\!\left(S\right)$ is the field extension of $\mathbb{Q}$ generated by all the entries of the $d^{2}$ Hermitian projection matrices $\Pi_{j}$.
   • (5)
-    The extended projection SIC-field $F^{\rm eproj}\!\left(S\right):=F^{\rm proj}\!\left(S\right)\!(\xi_{d})$, adjoining $\xi_{d}=-e^{\frac{\pi i}{d}}$.
-A priori, these fields could be transcendental extensions of $\mathbb{Q}$, and this does occur in dimension $d=3$.
+    The extended projection SIC field $F^{\rm eproj}\!\left(S\right):=F^{\rm proj}\!\left(S\right)\!(\xi_{d})$, adjoining $\xi_{d}=-e^{\frac{\pi i}{d}}$.
+A priori, these fields could be transcendental extensions of $\mathbb{Q}$, as happens in dimension $d=3$.
 Remark 3.2.
-These fields have previously been studied with various names and notations, which we now review.
+These fields have previously been studied with various names and notations.
   1. (1)
-    The field $F^{\rm vec}\!\left(S\right)$ is the SIC field $\mathbb{Q}[S]$ attached to a general SIC $S$ by Appleby, Flammia, McConnell, and Yard [10, p.
+    The field $F^{\rm vec}\!\left(S\right)$ is the SIC field $\mathbb{Q}[S]$ attached to a general SIC $S$ by Appleby, Flammia, McConnell, and Yard [9, p.
 212 bottom], as used in Thm.
@@
 By definition, this field is independent of scaling factors in the $\mathbf{v}_{i}$.
-It is determined by ${\mathbf{\Pi}}(S)$ because the ratios
+It is determined by ${\mathbf{\Pi}}(S)$, because the ratios
     $\frac{v_{ij}}{v_{ik}}=\frac{v_{ij}\overline{v}_{ik}}{v_{ik}\overline{v}_{ik}}=\frac{(\Pi_{i})_{jk}}{(\Pi_{i})_{kk}}.$ (3.1)
@@
 The vectors ${\mathbf{w}}_{i}$ obtained this way may not be unit vectors.
-    Prop.
-3 of [10] states that, for a Weyl–Heisenberg SIC, the ratio SIC-field is the same for any SIC $S^{\prime}\in[S]$.
-The presented proof outline [10, Sec.
+    It is stated as [9, Prop.
+3] that, for a Weyl–Heisenberg SIC, the ratio SIC field is the same for any SIC $S^{\prime}\in[S]$.
+The presented proof outline [9, Sec.
 7.1] appears to assume some conjectural equivalences and characterizations of other related fields.
   2. (2)
-    Concerning the triple product SIC-field $F^{\rm trip}\!\left(S\right)$, the fact that a SIC is determined up to unitary equivalence by its triple products was proved in 2011 by Appleby, Flammia, and Fuchs [3, Thm.
-3].
+    Concerning the triple product SIC field $F^{\rm trip}\!\left(S\right)$, the fact that a SIC is determined up to unitary equivalence by its triple products was proved in 2011 by Appleby, Flammia, and Fuchs [3, Thm.
+3]; see also [46, Cor.
+8.1].
 We show unconditionally that $F^{\rm trip}\!\left(S\right)$ is an invariant of the geometric $\operatorname{PEC}(d)$-orbit of a Weyl–Heisenberg SIC $S$; see 3.4.
   3. (3)
-    The unitary invariant SIC-field $F^{\rm inv}\!\left(S\right)$ is generated by the full set of projective unitary invariants of a general finite set of lines in $\mathbb{C}^{d}$; see [15] and [46, Ch.
+    The unitary invariant SIC field $F^{\rm inv}\!\left(S\right)$ is generated by the full set of projective unitary invariants of a general finite set of lines in $\mathbb{C}^{d}$; see [14] and [46, Ch.
 8, Thm.
@@
   4. (4)
-    The projection SIC-field $F^{\rm proj}\!\left(S\right)$, in the special case that $S$ is a Weyl–Heisenberg SIC, coincides with the field $L$ appearing in [10, Sec.
+    The projection SIC field $F^{\rm proj}\!\left(S\right)$, in the special case that $S$ is a Weyl–Heisenberg SIC, coincides with the field $L$ appearing in [9, Sec.
 7], defined as the smallest subfield of $\mathbb{C}$ generated by the entries $v_{j}\overline{v}_{k}$ of unit vectors generating $\Pi_{\mathbf{v}}$ taken over all $d^{2}$ projectors $\Pi_{\mathbf{v}}$ of a Weyl–Heisenberg SIC.
   5. (5)
-    The extended projection SIC-field $F^{\rm eproj}\!\left(S\right)$ is the SIC field $\mathbb{E}$ as defined by Appleby, Flammia, McConnell, and Yard [9, Sec.
+    The extended projection SIC field $F^{\rm eproj}\!\left(S\right)$ is the SIC field $\mathbb{E}$ as defined by Appleby, Flammia, McConnell, and Yard [8, Sec.
 4] for any SIC projector $S$.
-This definition of the SIC field is also used in [7].
-Appleby, Flammia, McConnell, and Yard [9] state (without proof) that the field $\mathbb{E}$ is well-defined on the $\operatorname{PEC}(d)$-orbit $[S]$ of a Weyl–Heisenberg SIC $S$.
-We prove so unconditionally in 3.8.
+This definition of the SIC field is also used in [6] and is treated in detail in [46, Sec.
+14.20].
+Appleby, Flammia, McConnell, and Yard [8] state that the field $\mathbb{E}$ is well-defined on the $\operatorname{PEC}(d)$-orbit $[S]$ of a Weyl–Heisenberg SIC $S$.
+We supply a proof in 3.8.
 The next result establishes some inclusion relations among these fields.
@@
 Write ${\mathbf{\Pi}}(S)=\{\frac{1}{d}\Pi_{1},\ldots,\frac{1}{d}\Pi_{d^{2}}\}$.
-Proof of (1): Write $\Pi_{i}=\mathbf{v}_{i}\mathbf{v}_{i}^{\dagger}$, where $\mathbf{v}_{i}$ has entries $v_{ij}$ and $\left|\mathbf{v}_{i}\right|^{2}=1$.
-The nonzero entries of $\Pi_{i}$ are of the form $v_{ij}\overline{v}_{ik}$ for $v_{ij}\neq 0$, $v_{ik}\neq 0$ These entries can be expressed as
-$v_{ij}\overline{v}_{ik}=\frac{v_{ij}\overline{v}_{ik}}{\left|\mathbf{v}_{i}\right|^{2}}=\frac{1}{\sum_{\ell=1}^{d}(v_{i\ell}/v_{ij})\overline{(v_{i\ell}/v_{ik})}}\in F^{\rm vec}\!\left(S\right)\overline{F^{\rm vec}\!\left(S\right)}.$ (3.2)
-Thus, $F^{\rm proj}\!\left(S\right)\subseteq F^{\rm vec}\!\left(S\right)\overline{F^{\rm vec}\!\left(S\right)}$. To show the reverse inclusion, note that for $v_{ik}\neq 0$,
-$v_{ij}/v_{ik}=\frac{v_{ij}\overline{v}_{ik}}{v_{ik}\overline{v}_{ik}}\mbox{ and }\overline{(v_{ij}/v_{ik})}=\frac{v_{ik}\overline{v}_{ij}}{v_{ik}\overline{v}_{ik}}.$ (3.3)
-Proof of (2): First we show the equality $F^{\rm trip}\!\left(S\right)=F^{\rm inv}\!\left(S\right)$.
+Proof of (1).
+Write $\Pi_{i}=\mathbf{v}_{i}\mathbf{v}_{i}^{\dagger}$, where $\mathbf{v}_{i}$ has entries $v_{im}$ and $\left|\mathbf{v}_{i}\right|^{2}=1$, for $1\leq i\leq d^{2}$.
+The entries of $\Pi_{i}$ are of the form $v_{im}\overline{v}_{in}$ for $10\leq m,n\leq d-$ 1 When both $v_{im}\neq 0$, $v_{in}\neq 0$, these entries can be expressed as
+$v_{im}\overline{v}_{in}=\frac{v_{im}\overline{v}_{in}}{\left|\mathbf{v}_{i}\right|^{2}}=\frac{1}{\sum_{p=0}^{d-1}(v_{ip}/v_{im})\overline{(v_{ip}/v_{in})}}\in F^{\rm vec}\!\left(S\right)\overline{F^{\rm vec}\!\left(S\right)},$ (3.2)
+where $10\leq m,n,p\leq d-$ 1.
+Thus, $F^{\rm proj}\!\left(S\right)\subseteq F^{\rm vec}\!\left(S\right)\overline{F^{\rm vec}\!\left(S\right)}$.
+To show the reverse inclusion, note that for $v_{ik}\neq 0$,
+$v_{im}/v_{in}=\frac{v_{im}\overline{v}_{in}}{v_{im}\overline{v}_{im}}\mbox{ and }\overline{(v_{im}/v_{in})}=\frac{v_{in}\overline{v}_{im}}{v_{in}\overline{v}_{in}}.$ (3.3)
+Proof of (2).
+First, the inclusion $F^{\rm inv}\!\left(S\right)\subseteq F^{\rm proj}\!\left(S\right)$ of (2) follows because $\operatorname{Tr}(\Pi_{i_{1}}\Pi_{i_{2}}\cdots\Pi_{i_{n}})$ is an integer polynomial in the entries of the various $\Pi_{j}$.
+Second, we show the equality $F^{\rm trip}\!\left(S\right)=F^{\rm inv}\!\left(S\right)$.
 Clearly $F^{\rm trip}\!\left(S\right)\subseteq F^{\rm inv}\!\left(S\right)$.
@@
 $\Pi_{i}\Pi_{j}=\sum_{\ell}\alpha_{ij}^{\ell}\Pi_{\ell}$ (3.4)
-for structure constants $\alpha_{ij}^{\ell}\in\mathbb{C}$. We have
+for structure constants $\alpha_{ij}^{\ell}\in\mathbb{C}$, having $1\leq i,j,\ell\leq d^{2}$. We have
 $\sum_{\ell}\alpha_{ij}^{\ell}=\operatorname{Tr}(\Pi_{i}\Pi_{j})=\frac{\delta_{ij}d+1}{d+1},\mbox{ where }\delta_{ij}=\begin{cases}1,&\mbox{if }i=j,\\ 0,&\mbox{if }i\neq j.\end{cases}$ (3.5)
@@
 Moreover, using (3.4) repeatedly, we can express any product $\Pi_{i_{1}}\Pi_{i_{2}}\cdots\Pi_{i_{n}}$ in the $\left\{\Pi_{\ell}\right\}_{\ell}$-basis with coefficients that are integer polynomials in the structure constants.
-Thus, since $\operatorname{Tr}(\Pi_{\ell})=1$, $\operatorname{Tr}(\Pi_{i_{1}}\Pi_{i_{2}}\cdots\Pi_{i_{n}})$ is also an integer polynomials in the structure constants and is thus contained in $F^{\rm trip}\!\left(S\right)$.
-Finally, the inclusion $F^{\rm inv}\!\left(S\right)\subseteq F^{\rm proj}\!\left(S\right)$ of (2) follows because $\operatorname{Tr}(\Pi_{i_{1}}\Pi_{i_{2}}\cdots\Pi_{i_{n}})$ is an integer polynomial in the entries of the various $\Pi_{j}$.
+Thus, since $\operatorname{Tr}(\Pi_{\ell})=1$, $\operatorname{Tr}(\Pi_{i_{1}}\Pi_{i_{2}}\cdots\Pi_{i_{n}})$ is also an integer polynomial in the structure constants and is thus contained in $F^{\rm trip}\!\left(S\right)$.
 ∎
@@
 Proof.
-Let $S=\{\mathbb{C}{\mathbf{v}}_{i}:1\leq i\leq d^{2}\}$ be an arbitrary line-SIC, and set ${\mathbf{\Pi}}(S)=\{\frac{1}{d}\Pi_{1},\ldots,\frac{1}{d}\Pi_{d^{2}}\}$, with $\Pi_{i}={\mathbf{v}}_{i}{\mathbf{v}}_{i}^{\dagger}$.
-Then ${\mathbf{\Pi}}(US)=\{\frac{1}{d}\Pi_{1}^{\prime},\ldots,\frac{1}{d}\Pi_{d^{2}}^{\prime}\}$.
-Write $U=CV$, where $C$ is either the identity or complex conjugation, and $V\in\mathbf{U}(d)$; then either $\Pi_{i}^{\prime}=V^{-1}\Pi_{i}V$, or $\Pi_{i}^{\prime}=V^{-1}\Pi_{i}^{\top}V$.
+Write $S=\{\mathbb{C}{\mathbf{v}}_{i}:1\leq i\leq d^{2}\}$, and set ${\mathbf{\Pi}}(S)=\{\frac{1}{d}\Pi_{1},\ldots,\frac{1}{d}\Pi_{d^{2}}\}$ with $\Pi_{i}={\mathbf{v}}_{i}{\mathbf{v}}_{i}^{\dagger}$.
+Then set ${\mathbf{\Pi}}(US)=\{\frac{1}{d}\Pi_{1}^{\prime},\ldots,\frac{1}{d}\Pi_{d^{2}}^{\prime}\}$.
+Write $U=CV$, where $C$ is identity or complex conjugation, and $V\in\mathbf{U}(d)$; then either $\Pi_{i}^{\prime}=V^{-1}\Pi_{i}V$, or $\Pi_{i}^{\prime}=V^{-1}\Pi_{i}^{\top}V$.
 In the first case, $\operatorname{Tr}(\Pi_{i}^{\prime}\Pi_{j}^{\prime}\Pi_{k}^{\prime})=\operatorname{Tr}(V^{-1}\Pi_{i}\Pi_{j}\Pi_{k}V)=\operatorname{Tr}(\Pi_{i}\Pi_{j}\Pi_{k})$.
-In the second case, $\operatorname{Tr}(\Pi_{i}^{\prime}\Pi_{j}^{\prime}\Pi_{k}^{\prime})=\operatorname{Tr}\!\left(V^{-1}\Pi_{i}^{\top}\Pi_{j}^{\top}\Pi_{k}^{\top}V\right)=\operatorname{Tr}\!\left((\Pi_{k}\Pi_{j}\Pi_{i})^{\top}\right)=\operatorname{Tr}(\Pi_{k}\Pi_{j}\Pi_{i})$.
-So $F^{\rm trip}\!\left(US\right)=F^{\rm trip}\!\left(S\right)$ in both cases.
-∎
+In the second case,
+$\operatorname{Tr}(\Pi_{i}^{\prime}\Pi_{j}^{\prime}\Pi_{k}^{\prime})=\operatorname{Tr}\!\left(V^{-1}\Pi_{i}^{\top}\Pi_{j}^{\top}\Pi_{k}^{\top}V\right)=\operatorname{Tr}\!\left((\Pi_{k}\Pi_{j}\Pi_{i})^{\top}\right)=\operatorname{Tr}(\Pi_{k}\Pi_{j}\Pi_{i}).$ (3.9)
+So $F^{\rm trip}\!\left(US\right)=F^{\rm trip}\!\left(S\right)$ in both cases. ∎
 We derive some unconditional implications when these fields give Galois extensions of $\mathbb{Q}$.
@@
 Proof.
-Proof of (1): The implication $F^{\rm vec}\!\left(S\right)=F^{\rm proj}\!\left(S\right)$ follows from 3.3(1).
-Proof of (2): If $K=F^{\rm proj}\!\left(S\right)$ is Galois over $\mathbb{Q}$, then $F^{\rm eproj}\!\left(S\right)$ is the compositum $KL$, with $L=\mathbb{Q}(e^{\frac{\pi i}{d}})$ an abelian extension of $\mathbb{Q}$, so the compositum is a Galois extension of $\mathbb{Q}$.
+The implication $F^{\rm vec}\!\left(S\right)=F^{\rm proj}\!\left(S\right)$ follows from 3.3(1).
+Moreover, if $K=F^{\rm proj}\!\left(S\right)$ is Galois over $\mathbb{Q}$, then $F^{\rm eproj}\!\left(S\right)$ is the compositum $KL$, with $L=\mathbb{Q}(e^{\frac{\pi i}{d}})$ an abelian extension of $\mathbb{Q}$, so the compositum is a Galois extension of $\mathbb{Q}$.
 ∎
 3.2. Fields generated by Weyl–Heisenberg SICs
-We now derive some properties of the field $F^{\rm vec}\!\left(S\right)$, $F^{\rm trip}\!\left(S\right)$ and $F^{\rm proj}\!\left(S\right)$ in the special case of Weyl–Heisenberg SICs.
-Lemma 3.6.
-Let $S$ be a Weyl–Heisenberg line-SIC $S=S([{\mathbf{v}}])$ of dimension $d$. Then:
+We now derive some properties of the fields $F^{\rm vec}\!\left(S\right)$, $F^{\rm trip}\!\left(S\right)$, $F^{\rm proj}\!\left(S\right)$, and $F^{\rm eproj}\!\left((\right)S)$ in the special case of Weyl–Heisenberg SICs.
+Proposition 3.6.
+Let $S$ be a Weyl–Heisenberg line-SIC $S=S([{\mathbf{v}}])$ of dimension $d\geq 2$. Then:
   • (1)
-    $F^{\rm proj}\!\left(S\right)$ contains a $d$-th root of unity $\zeta_{d}=e^{\frac{2\pi i}{d}}$.
+    The field $F^{\rm trip}\!\left(S\right)$ contains the $d$-th root of unity $\zeta_{d}=e^{\frac{2\pi i}{d}}$.
+  • (2)
+    The field $F^{\rm proj}\!\left(S\right)\subseteq L(S)=F^{\rm trip}\!\left(S\right)\!({\sqrt[d]{\beta_{1}},\sqrt[d]{\beta_{2}}})$, for some $\beta_{1},\beta_{2}\in F^{\rm trip}\!\left(S\right)$.
+$L(S)$ is an abelian extension of $F^{\rm trip}\!\left(S\right)$ whose degree divides $d^{2}$, hence $[F^{\rm proj}\!\left(S\right):F^{\rm trip}\!\left(S\right)]$ divides $d^{2}$.
   • (3)
@@
 Proof.
-Proof of (1): Let ${\mathbf{v}}=(v_{0},v_{1},\ldots,v_{d-1})^{\top}$ be a fiducial vector for $S$, and let $\Pi={\mathbf{v}}{\mathbf{v}}^{\dagger}$.
-It has $(1,0)$-entry $(\Pi)_{10}=v_{1}\overline{v}_{0}\in F^{\rm proj}\!\left(S\right)$.
-Now the modulation matrix $Z\in\operatorname{WH}(d)$ gives another SIC vector $Z{\mathbf{v}}=(v_{0},e^{\frac{2\pi i}{d}}v_{1},\cdots,e^{\frac{2\pi i(d-1)}{d}}v_{d-1})$.
-The corresponding projection $\Pi^{\prime}:=Z{\mathbf{v}}(Z{\mathbf{v}})^{\dagger}=Z{\mathbf{v}}{\mathbf{v}}^{\dagger}Z^{-1}$ has $(\Pi^{\prime})_{1,0}=e^{\frac{2\pi i}{d}}v_{1}\overline{v}_{0}\in F^{\rm proj}\!\left(S\right)$.
-If $v_{1}\overline{v}_{0}\neq 0$, then we may conclude $e^{\frac{2\pi i}{d}}\in F^{\rm proj}\!\left(S\right)$.
-We show every Weyl–Heisenberg SIC has a fiducial vector ${\mathbf{v}}$ with both $v_{0}$ and $v_{1}$ nonzero, so the argument applies.
-We first claim that every fiducial vector ${\mathbf{v}}$ has two consecutive entries $v_{i},v_{i+1}$ nonzero, allowing the wraparound product $v_{d-1}v_{0}$.
-Suppose not; then, the translation matrix $X\in\operatorname{WH}(d)$ gives a cyclic shift $X{\mathbf{v}}=(v_{1},v_{2},...,v_{d-1},v_{0})$, so $\langle{\mathbf{v}},X{\mathbf{v}}\rangle=\sum_{i=0}^{d-1}v_{i}\overline{v}_{i+1}=0$, contradicting the SIC property $\left|\langle{\mathbf{v}},X{\mathbf{v}}\rangle\right|=\frac{1}{d+1}$.
-The claim follows.
-If the given ${\mathbf{v}}$ has $v_{i}v_{i+1}\neq 0$, then ${\mathbf{v}}^{\prime}=X^{d-i}{\mathbf{v}}$ cyclically shifts entries so that $v_{0}^{\prime}v_{1}^{\prime}\neq 0$.
-Proof of (2): By (1), $F^{\rm eproj}\!\left(S\right)=F^{\rm proj}\!\left(S\right)(e^{\frac{\pi i}{d}})$ is at most a quadratic extension of $F^{\rm proj}\!\left(S\right)$, since the field index $[\mathbb{Q}(e^{\frac{\pi i}{d}}):\mathbb{Q}(e^{\frac{2\pi i}{d}})]=2$ if $d$ is even and is $1$ if $d$ is odd.
+Let ${\mathbf{v}}=(v_{0},v_{1},\ldots,v_{d-1})^{\top}$ be a fiducial vector for $S$, and let $\Pi={\mathbf{v}}{\mathbf{v}}^{\dagger}$.
+For $X,Z$ as in 2.1, let $D_{{\mathbf{p}}}=X^{p_{1}}Z^{p_{2}}$ Then, for $0\leq p_{1}<d$ and $0\leq p_{2}<d$, the vectors ${\mathbf{v}}_{{\mathbf{p}}}=X^{p_{1}}Z^{p_{2}}{\mathbf{v}}$ form a set of generating set of vectors of the SIC, all having $\left|\left|\hskip 1.7pt{\mathbf{v}}_{{\mathbf{p}}}\hskip 1.7pt\right|\right|^{2}=\left|\left|\hskip 1.7pt{\mathbf{v}}\hskip 1.7pt\right|\right|^{2}$.
+Here ${\mathbf{p}}=(p_{1},p_{2})\in(\mathbb{Z}/p\mathbb{Z})^{2}$, and we also write ${\mathbf{v}}_{i}$ with $i=(d-1)p_{1}+p_{2}$ having $1\leq i\leq d^{2}$.
+We note for column vectors ${\mathbf{w}}_{i}$ that
+$\langle{\mathbf{w}}_{1},{\mathbf{w}}_{2}\rangle\langle{\mathbf{w}}_{2},{\mathbf{v}}_{3}\rangle\langle{\mathbf{w}}_{3},{\mathbf{w}}_{1}\rangle$ $=\operatorname{Tr}\left({\mathbf{w}}_{1}^{\dagger}{\mathbf{w}}_{2}{\mathbf{w}}_{2}^{\dagger}{\mathbf{w}}_{3}{\mathbf{w}}_{3}^{\dagger}{\mathbf{w}}_{1}\right)$ (3.10)
+$=\operatorname{Tr}\left({\mathbf{w}}_{1}{\mathbf{w}}_{1}^{\dagger}{\mathbf{w}}_{2}{\mathbf{w}}_{2}^{\dagger}{\mathbf{w}}_{3}{\mathbf{w}}_{3}^{\dagger}{\mathbf{w}}_{3}\right)$ (3.11)
+$=\operatorname{Tr}\left(\Pi_{{\mathbf{w}}_{1}}\Pi_{{\mathbf{w}}_{2}}\Pi_{{\mathbf{w}}_{3}}\right)$ (3.12)
+is a triple product, and similarly for $r$-fold cyclic scalar products of ${\mathbf{w}}_{i}$.
+Proof of (1). The ratio of triple inner products
+$\frac{\langle{\mathbf{v}},X^{-1}{\mathbf{v}}\rangle\langle X^{-1}{\mathbf{v}},Z{\mathbf{v}}\rangle\langle Z{\mathbf{v}},{\mathbf{v}}\rangle}{\langle{\mathbf{v}},Z^{-1}{\mathbf{v}}\rangle\langle Z^{-1}{\mathbf{v}},X{\mathbf{v}}\rangle\langle X{\mathbf{v}},{\mathbf{v}}\rangle}$ $=\frac{\langle X{\mathbf{v}},{\mathbf{v}}\rangle\langle{\mathbf{v}},XZ{\mathbf{v}}\rangle\langle Z{\mathbf{v}},{\mathbf{v}}\rangle}{\langle Z{\mathbf{v}},{\mathbf{v}}\rangle\langle{\mathbf{v}},ZX{\mathbf{v}}\rangle\langle X{\mathbf{v}},{\mathbf{v}}\rangle}$ (3.13)
+$=\frac{\langle{\mathbf{v}},XZ{\mathbf{v}}\rangle}{\langle{\mathbf{v}},ZX{\mathbf{v}}\rangle}=\zeta_{d}.$ (3.14)
+Via the SIC relations $\left|\langle{\mathbf{v}}_{i},{\mathbf{v}}_{j}\rangle\right|=\frac{1}{d+1}\left|{\mathbf{v}}\right|^{2}$, if $i\neq j$, and $\left|\left|\hskip 1.7pt{\mathbf{v}}\hskip 1.7pt\right|\right|^{2}$ if $i=j$, all factors in these ratios are nonzero.
+Therefore $\zeta_{d}\in F^{\rm trip}\!\left(S\right)$.
+Proof of (2).
+Set $\nu_{\mathbf{p}}=\langle{\mathbf{v}},D_{\mathbf{p}}{\mathbf{v}}\rangle$.
+Set $\alpha_{1}:=\nu_{(1,0)}=\langle{\mathbf{v}},X{\mathbf{v}}\rangle$ and $\alpha_{2}=\nu_{(0,1)}=\langle{\mathbf{v}},Z{\mathbf{v}}\rangle$.
+Define the field $L(S):=F^{\rm trip}\!\left(S\right)(\alpha_{1},\alpha_{2})$.
+By (1), $\zeta_{d}\in F^{\rm trip}\!\left(S\right)$.
+To show $F^{\rm proj}\!\left(S\right)\subseteq L(S)$, it suffices =to show that all entries $(\Pi_{i})_{k,\ell}\in L(S)$ for $1\leq i\leq d^{2}$ and $0\leq k,\ell<d$.
+Set ${\bf e}_{1}=(1,0)$ and ${\bf e}_{2}=(0,1)$.
+Now for ${\mathbf{p}}=p_{1}{\bf e}_{1}+p_{2}{\bf e_{2}}$, the product $\nu_{{\bf e}_{1}}^{p_{1}}\nu_{{\bf e}_{2}}^{p_{2}}\nu_{-{\mathbf{p}}}$ is a $(p_{1}+p_{2}+1)$-fold product (up to multiplication by a $d$-th root of unity), and thus by 3.3(2), $\nu_{{\bf e}_{1}}^{p_{1}}\nu_{{\bf e}_{2}}^{p_{2}}\nu_{-{\mathbf{p}}}\in F^{\rm trip}\!\left(S\right)(\alpha_{1},\alpha_{2},\zeta_{d}).$ We conclude that all $\nu_{-{\mathbf{p}}}$ are in $F^{\rm trip}\!\left(S\right)(\alpha_{1},\alpha_{2},\zeta_{d})$.
+The $D_{\mathbf{p}}$ form a basis of the matrix algebra ${\rm Mat}_{d\times d}(\mathbb{Q}(e^{\frac{2\pi i}{d}}))$.
+In particular all the matrices $E_{k,\ell}$ for $10\leq k,\ell\leq d-$ 1 with entry $1$ in the $(k,\ell)$ position and 00 elsewhere are linear combinations of the $D_{p}$ over the field $\mathbb{Q}(e^{\frac{2\pi i}{d}})$.
+Thus $(\Pi_{1})_{k,\ell}=\langle{\mathbf{v}},E_{k,\ell}{\mathbf{v}}\rangle\in F^{\rm trip}\!\left(S\right)(\alpha_{1},\alpha_{2})=L(S)$.
+Now for $i=(d-1)p_{1}+p_{2}$ we have
+$(\Pi_{i})_{k,\ell}=\langle X^{p_{1}}Z^{p_{2}}{\mathbf{v}},E_{k,\ell}X^{p_{1}}Z^{p_{2}}{\mathbf{v}}\rangle=(\Pi_{1})_{k,\ell}$ (3.15)
+holding for all $1\leq i\leq p^{2}$, by unitarity. Hence $(\Pi_{i})_{k,\ell}\in L(S)$ for all $i,k,\ell$. We conclude that $F^{\rm proj}\!\left(S\right)\subseteq L(S)$.
+Finally, the product $\beta_{1}=(\alpha_{1})^{d}=\nu_{{\bf e}_{1}}^{d}=\nu_{{\bf e}_{1}}^{d}\nu_{0}$ is also a $(d+1)$-fold cyclic product and thus is in $F^{\rm inv}\!\left(S\right)=F^{\rm trip}\!\left(S\right)$, and similarly for $\beta_{2}=(\alpha_{2})^{d}=\nu_{{\bf e}_{2}}^{d}\in F^{\rm trip}\!\left(S\right)$.
+Thus $\alpha_{1}=\sqrt[d]{\beta_{1}}$ and $\alpha_{2}=\sqrt[d]{\beta_{2}}$ are both $d$-th roots of elements of $F^{\rm trip}\!\left(S\right)$.
+Since $\zeta_{d}\in F^{\rm trip}\!\left(S\right)$ we conclude adjoining these two elements to $F^{\rm trip}\!\left(S\right)$ gives an abelian extension $L(S)$ of $F^{\rm trip}\!\left(S\right)$, which is a Kummer extension whose degree divides $d^{2}$.
+The field $L(S)$ contains $F^{\rm proj}\!\left(S\right)$, and $F^{\rm trip}\!\left(S\right)\subseteq F^{\rm proj}\!\left(S\right)$, which implies $[F^{\rm proj}\!\left(S\right):F^{\rm trip}\!\left(S\right)]$ divides $d^{2}$.
+Proof of (3).
+By (1), since $F^{\rm trip}\!\left(S\right)\subseteq F^{\rm proj}\!\left(S\right)$, we have $\zeta_{d}\in F^{\rm proj}\!\left(S\right)$.
+Now $F^{\rm eproj}\!\left(S\right)=F^{\rm proj}\!\left(S\right)(e^{\frac{\pi i}{d}})$ is at most a quadratic extension of $F^{\rm proj}\!\left(S\right)$, since the field index $[\mathbb{Q}(e^{\frac{\pi i}{d}}):\mathbb{Q}(e^{\frac{2\pi i}{d}})]=2$ if $d$ is even and is $1$ if $d$ is odd.
 ∎
@@
 $X,Z,F,R$ (3.16)
-where $X,Z$ are Weyl–Heisenberg group generators, $F$ is the discrete Fourier transform,
-$F_{j,k}=\frac{1}{\sqrt{d}}e^{\frac{2\pi ijk}{d}}\,\mbox{ for }\,0\leq j,k\leq d-1.$ (3.10)
-and $R$ is a diagonal matrix, with diagonal entries
+where $X,Z$ are Weyl–Heisenberg group generators, $F$ is the discrete Fourier transform
+$F_{j,k}=\frac{1}{\sqrt{d}}e^{\frac{2\pi ijk}{d}}\,\mbox{ for }\,0\leq j,k\leq d-1,$ (3.17)
+and $R$ is a diagonal matrix with diagonal entries
 $R_{j,j}=\xi_{d}^{j(j+d)}=(-1)^{jd}e^{\frac{\pi ij^{2}}{d}}\,\mbox{ for }\,0\leq j\leq d-1.$ (3.18)
 Proof.
-This is [46, Thm. 14.1]. ∎
+This is [46, Thm. 14.1]. Every group element is a product of finitely many generators, with exactly one use of a unitary scalar matrix $zI$. ∎
 Proposition 3.8.
-Let $S$ be a Weyl–Heisenberg line-SIC $S=S([{\mathbf{v}}])$ of dimension $d$. For any $M\in\operatorname{EC}(d)$,
+Let $S$ be a Weyl–Heisenberg line-SIC $S=S([{\mathbf{v}}])$ of dimension $d$.
+  • (1)
+    For any $M\in\operatorname{EC}(d)$,
     $F^{\rm eproj}\!\left(MS\right)=F^{\rm eproj}\!\left(S\right),$ (3.19)
     where $MS$ denotes the Weyl–Heisenberg line-SIC having fiducial vector ${\mathbf{w}}:=M\mathbf{v}$.
-Moreover, $F^{\rm eproj}\!\left(S\right)$ is equal to the compositum of all the $F^{\rm proj}\!\left(MS\right)$ for $M\in\operatorname{EC}(d)$.
+  • (2)
+    $F^{\rm eproj}\!\left(S\right)$ is equal to the compositum of all the $F^{\rm proj}\!\left(MS\right)$ for $M\in\operatorname{EC}(d)$.
 Proof.
-The scalar matrices $zI$ with $\left|z\right|=1$ act trivially on projections, with $(zI\mathbf{v})(zI\mathbf{v})^{\dagger}=\mathbf{v}\mathbf{v}^{\dagger}$.
+Proof of (1).
+To show $F^{\rm eproj}\!\left(MS\right)=F^{\rm eproj}\!\left(S\right)$, note that scalar matrices $zI$ with $\left|z\right|=1$ act trivially on projections, with $(zI\mathbf{v})(zI\mathbf{v})^{\dagger}=\mathbf{v}\mathbf{v}^{\dagger}$.
 The matrices $X$, $Z$, and $F$ all have entries in $\mathbb{Q}(\xi)$ (where $\xi=\xi_{d}=-e^{\frac{\pi i}{d}}$).
-The matrix $\sqrt{d}F$ also has entries in $\mathbb{Q}(\xi_{d})$, and $(F\mathbf{v})(F\mathbf{v})^{\dagger}=\frac{1}{d}(\sqrt{d}F\mathbf{v})(\sqrt{d}F\mathbf{v})^{\dagger}$.
-Therefore, writing any $M\in EC(d)$ as a product of the generators given in 3.7 shows that $(M\mathbf{v})(M\mathbf{v})^{\dagger}=M\mathbf{v}\mathbf{v}^{\dagger}M^{\dagger}$ has entries in $F^{\rm eproj}\!\left(S\right)$ (the compositum of $F^{\rm proj}\!\left(S\right)$ and $\mathbb{Q}(\xi)$).
-This argument shows that $F^{\rm proj}\!\left(MS\right)\subseteq F^{\rm eproj}\!\left(S\right)$, so $F^{\rm eproj}\!\left(MS\right)\subseteq F^{\rm eproj}\!\left(S\right)$.
-By considering $M^{-1}$, we also have $F^{\rm eproj}\!\left(S\right)=F^{\rm eproj}\!\left(M^{-1}MS\right)\subseteq F^{\rm eproj}\!\left(MS\right)$, so in fact $F^{\rm eproj}\!\left(MS\right)=F^{\rm eproj}\!\left(S\right)$.
-To show that $F^{\rm eproj}\!\left(S\right)$ is equal to the compositum of all the $F^{\rm proj}\!\left(MS\right)$, it suffices to show that $\xi=-e^{-\frac{\pi i}{d}}$ occurs in the compositum of two fields $F^{\rm proj}\!\left(S\right)$ and $F^{\rm proj}\!\left(RS\right)$ where $R$ is the element of $\operatorname{C}(d)$ given in 3.7.
+The matrix $\sqrt{d}F$ also has entries in $\mathbb{Q}(\xi)$, and $(F\mathbf{v})(F\mathbf{v})^{\dagger}=\frac{1}{d}(\sqrt{d}F\mathbf{v})(\sqrt{d}F\mathbf{v})^{\dagger}$.
+Writing any $M\in EC(d)$ as a finite product of the generators given in 3.7 shows that $(M\mathbf{v})(M\mathbf{v})^{\dagger}=M\mathbf{v}\mathbf{v}^{\dagger}M^{\dagger}$ has entries in $F^{\rm eproj}\!\left(S\right)$ (the compositum of $F^{\rm proj}\!\left(S\right)$ and $\mathbb{Q}(\xi)$).
+This argument shows $F^{\rm proj}\!\left(MS\right)\subseteq F^{\rm eproj}\!\left(S\right)$, so $F^{\rm eproj}\!\left(MS\right)\subseteq F^{\rm eproj}\!\left(S\right)$.
+By considering $M^{-1}$, we also have $F^{\rm eproj}\!\left(S\right)=F^{\rm eproj}\!\left(M^{-1}MS\right)\subseteq F^{\rm eproj}\!\left(MS\right)$, so $F^{\rm eproj}\!\left(MS\right)=F^{\rm eproj}\!\left(S\right)$.
+Proof of (2).
+To show $F^{\rm eproj}\!\left(S\right)$ is equal to the compositum of all the $F^{\rm proj}\!\left(MS\right)$, it suffices to show that $\xi=-e^{-\frac{\pi i}{d}}$ occurs in the compositum of two fields $F^{\rm proj}\!\left(S\right)$ and $F^{\rm proj}\!\left(RS\right)$ where $R$ is the element of $\operatorname{C}(d)$ given in 3.7.
 Given a fiducial ${\mathbf{v}}$ generating $S=S([{\mathbf{v}}])$ having $v_{0}v_{1}\neq 0$, we consider the fiducial vector ${\mathbf{w}}=R{\mathbf{v}}$ of a geometrically equivalent SIC.
-It has entries $(R{\mathbf{v}})_{j}=\xi^{j(j+d)}v_{j}$ so $R{\mathbf{v}}_{0}=v_{0}$ and $R{\mathbf{v}}_{1}=-e^{\frac{\pi i}{d}}v_{1}$.
+It has entries $(R{\mathbf{v}})_{j}=\xi^{j(j+d)}v_{j}$ so $R{\mathbf{v}}_{0}=v_{0}$ and $R{\mathbf{v}}_{1}=\xi v_{1}$.
 Therefore $\Pi^{\prime}={\mathbf{w}}{\mathbf{w}}^{\dagger}$ has
-$(\Pi^{\prime})_{1,0}=-e^{\frac{\pi i}{d}}v_{1}\overline{v}_{0}\in F^{\rm proj}\!\left(RS\right),$ (3.13)
+$(\Pi^{\prime})_{1,0}=\xi v_{1}\overline{v}_{0}\in F^{\rm proj}\!\left(RS\right),$ (3.20)
 while $v_{1}\overline{v}_{0}\in F^{\rm proj}\!\left(S\right)$.
-Now $v_{0}v_{1}\neq 0$ implies we may take their quotient in the compositum field and so get $-e^{\frac{\pi i}{d}}\in F^{\rm proj}\!\left(S\right)F^{\rm proj}\!\left(RS\right)$, as required.
+Now $v_{0}v_{1}\neq 0$ implies we may take their quotient in the compositum field and so get $\xi_{d}\in F^{\rm proj}\!\left(S\right)F^{\rm proj}\!\left(RS\right)$, as required.
 ∎
 4. Algebraic SICs and the Galois action
-We now focus on the subclass of algebraic SICs, for which the fields defined in Section 3.1 are algebraic number fields.
-In this paper, $\overline{\mathbb{Q}}$ is defined to be the algebraic closure of $\mathbb{Q}$ in $\mathbb{C}$, so there is a canonical choice of complex conjugation automorphism $\sigma_{c}\in\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$, $\sigma_{c}(z)=\overline{z}$.
+We now treat the subclass of algebraic SICs, for which the fields defined in Section 3.1 are algebraic number fields.
+In this paper, $\overline{\mathbb{Q}}$ is defined to be the algebraic closure of $\mathbb{Q}$ inside $\mathbb{C}$, so there is a canonical choice of complex conjugation automorphism $\sigma_{c}\in\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$, $\sigma_{c}(z)=\overline{z}$.
 We show that there is a well-defined Galois action on algebraic SICs and that the property of $S$ being an algebraic SIC is preserved by the subgroup of all automorphisms $\sigma\in\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ that commute with $\sigma_{c}$ when restricted to the Galois closure of $F^{\rm proj}\!\left(S\right)$.
@@
 (The definition is independent of the choice of ${\mathbf{v}}$.) A nonzero vector such that $\mathbb{C}\mathbf{v}$ is a projective-algebraic line is called a projective-algebraic vector.
-An algebraic vector is some ${\mathbf{v}}\in{\overline{\mathbb{Q}}}^{d}$. For a projective-algebraic vector $\mathbf{v}$, we set its $\overline{\mathbb{Q}}$-equivalence class
-$[[{{\mathbf{v}}}]]_{\operatorname{alg}}:=\{\lambda{\mathbf{v}}:\lambda\in\mathbb{C}\mbox{ and }\lambda{\mathbf{v}}\in\overline{\mathbb{Q}}^{d}\}.$ (4.1)
+An algebraic vector is any ${\mathbf{v}}\in{\overline{\mathbb{Q}}}^{d}$. For a projective-algebraic vector $\mathbf{v}$, its $\overline{\mathbb{Q}}$-equivalence class is
+$[[{{\mathbf{v}}}]]_{\operatorname{alg}}:=\left\{\lambda{\mathbf{v}}:\lambda\in\mathbb{C}\mbox{ and }\lambda{\mathbf{v}}\in\overline{\mathbb{Q}}^{d}\right\}.$ (4.1)
 The next result implies that the algebraic class $[[{{\mathbf{v}}}]]_{\operatorname{alg}}$ of an algebraic complex line always includes algebraic unit vectors.
@@
 Thus, $\left|{\mathbf{w}}\right|=\sqrt{\sum_{k}\left|\frac{v_{k}}{v_{i}}\right|^{2}}$ is a positive real algebraic number.
-We set $\lambda=\frac{\mu}{\left|{\mathbf{w}}\right|}$ and obtain the required ${\mathbf{x}}=\frac{{\mathbf{w}}}{\left|{\mathbf{w}}\right|}=\lambda{\mathbf{v}}$ having $\left|{\mathbf{x}}\right|=1$.
-Finally, if $\lambda$ and $\lambda^{\prime}$ give two such ${\mathbf{x}},{\mathbf{x}}^{\prime}\!$, necessarily $\lambda^{\prime}/\lambda\in\overline{\mathbb{Q}}$. ∎
+We set $\lambda=\frac{\mu}{\left|{\mathbf{w}}\right|}$ and obtain the required ${\mathbf{x}}:=\lambda{\mathbf{v}}=\frac{{\mathbf{w}}}{\left|{\mathbf{w}}\right|}$, having $\left|{\mathbf{x}}\right|=1$.
+Finally, if $\lambda$ and $\lambda^{\prime}$ give two algebraic ${\mathbf{x}},{\mathbf{x}}^{\prime}\!$, necessarily $\lambda^{\prime}/\lambda\in\overline{\mathbb{Q}}$. ∎
 Definition 4.3.
-A $d$-dimensional line-SIC $S=\{{\mathbf{v}}_{i}:1\leq i\leq d^{2}\}$ is algebraic if each of its $d^{2}$ vectors ${\mathbf{v}}_{i}$ are projective-algebraic.
+A $d$-dimensional line-SIC $S=\{{\mathbf{v}}_{i}:1\leq i\leq d^{2}\}$ is algebraic if each of its $d^{2}$ vectors ${\mathbf{v}}_{i}$ are projective-algebraic (or equivalently, if $F^{\rm vec}\!\left(S\right)$ is a finite extension of $\mathbb{Q}$).
 Lemma 4.4.
@@
   • (1)
-[…]
+    A line-SIC is algebraic if and only if, for unit vectors ${\mathbf{v}}_{i}$, all the SIC projection matrices $\{\Pi_{{\mathbf{v}}_{i}}:1\leq i\leq d^{2}\}$ have algebraic number entries.
   • (2)
-    For an algebraic SIC the fields $F^{\rm vec}\!\left(S\right)$, $F^{\rm trip}\!\left(S\right)$ and $F^{\rm proj}\!\left(S\right)$ are algebraic number fields; that is, finite algebraic extensions of $\mathbb{Q}$.
+    For an algebraic SIC $S$, the fields $F^{\rm vec}\!\left(S\right)$, $F^{\rm trip}\!\left(S\right)$, $F^{\rm proj}\!\left(S\right)$ and $F^{\rm eproj}\!\left(S\right)$ are algebraic number fields, that is, finite algebraic extensions of $\mathbb{Q}$.
   • (3)
@@
 Proof.
-Proof of (1): This follows using 4.2.
-Proof of (2): That $F^{\rm vec}\!\left(S\right)$ is an algebraic number field follows directly from the definition of algebraic.
+Proof of (1). This follows using 4.2.
+Proof of (2).
+That $F^{\rm vec}\!\left(S\right)$ is an algebraic number field follows directly from the definition of algebraic.
 Given a line-SIC specified by $d^{2}$ projective algebraic lines $\mathbb{C}{{\mathbf{v}}_{i}}$, using (1) we immediately have $F^{\rm trip}\!\left(S\right)$ being an algebraic number field.
 Furthermore, one may rescale all ${\mathbf{v}}_{i}$ to be algebraic unit vectors by 4.2.
-Now the associated projection matrices $\Pi_{{\mathbf{v}}}={\mathbf{v}}_{i}{\mathbf{v}}_{i}^{\dagger}$ have all entries in $\overline{\mathbb{Q}}$, hence $F^{\rm proj}\!\left(S\right)$ will be an algebraic number field.
-Proof of (3): If the fiducial vector ${\mathbf{v}}$ of a Weyl–Heisenberg line-SIC $S$ is projective-algebraic, then by 4.2, ${\mathbf{v}}$ may be rescaled to be an algebraic unit vector ${\mathbf{w}}=\lambda{\mathbf{v}}$.
-The Weyl–Heisenberg group consists of matrices with algebraic entries, hence its action on ${\mathbf{w}}$ shows that all ${\mathbf{w}}_{i}$ for $1\leq i\leq d^{2}$ are algebraic, so $S$ is an algebraic line-SIC.
+Now the associated projection matrices $\Pi_{{\mathbf{v}}}={\mathbf{v}}_{i}{\mathbf{v}}_{i}^{\dagger}$ have all entries in $\overline{\mathbb{Q}}$, so $F^{\rm proj}\!\left(S\right)$ is an algebraic number field.
+Finally, $F^{\rm eproj}\!\left(S\right)$ is a finite extension of $F^{\rm proj}\!\left(S\right)$.
+Proof of (3).
+If the fiducial vector ${\mathbf{v}}$ of a Weyl–Heisenberg line-SIC $S$ is projective-algebraic, then by 4.2, ${\mathbf{v}}$ may be rescaled to be an algebraic unit vector ${\mathbf{w}}=\lambda{\mathbf{v}}$.
+The Weyl–Heisenberg group $\operatorname{WH}(d)$ consists of matrices with algebraic entries, hence its action on ${\mathbf{w}}$ shows that all ${\mathbf{w}}_{i}$ for $1\leq i\leq d^{2}$ are algebraic, so $S$ is an algebraic line-SIC.
 The converse direction is immediate.
 ∎
-It is conjectured that for each dimension $d\neq 3$, all Weyl–Heisenberg SICs are algebraic.
-Conjecture 4.5 (Weyl–Heisenberg Algebraicity Conjecture).
-For each dimension $d\geq 1$, except $d=3$, all Weyl–Heisenberg SICs are projective-algebraic.
-This conjecture is equivalent to 2.5, which states that for all $d\neq 3$ the real projective algebraic variety $\operatorname{WHSIC}(d)$ is a finite set.
-It would follow by 4.4 that, if $S$ a Weyl–Heisenberg SIC, then $F^{\rm vec}\!\left(S\right)$, $F^{\rm trip}\!\left(S\right)$, and $F^{\rm proj}\!\left(S\right)$ are algebraic number fields.
 4.2. Galois actions on algebraic SICs
 There is a well-defined action of $\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ on algebraic complex lines which maps them to other algebraic complex lines.
-Given $\sigma\in\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ and ${\mathbf{v}}\in\overline{\mathbb{Q}}^{d}$ we send ${\mathbf{v}}\mathbb{C}\mapsto\sigma({\mathbf{v}})\mathbb{C}$.
+Given $\sigma\in\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ and ${\mathbf{v}}\in\overline{\mathbb{Q}}^{d}$, we send ${\mathbf{v}}\mathbb{C}\mapsto\sigma({\mathbf{v}})\mathbb{C}$.
 This action is well-defined on algebraic complex lines, since if $\lambda\in\overline{\mathbb{Q}}^{\times}$, then ${\mathbf{v}}\mathbb{C}=\lambda{\mathbf{v}}\mathbb{C}$ while
@@
 Let $\sigma_{c}\in\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ denote complex conjugation and $E$ an algebraic Galois extension of $\mathbb{Q}$.
-Let $C(\sigma_{c}\!\mid\!E)$ be the centralizer of $\sigma_{c}$ restricted to $\operatorname{Gal}(E/\mathbb{Q})$, that is,
+Let $C(\sigma_{c}\!\mid\!E)$ denote the centralizer of $\sigma_{c}$ restricted to $\operatorname{Gal}(E/\mathbb{Q})$, that is,
 $C(\sigma_{c}\!\mid\!E):=\{\sigma\in\operatorname{Gal}(E/\mathbb{Q}):\,\sigma\sigma_{c}(\alpha)=\sigma_{c}\sigma(\alpha)\mbox{ for all }\alpha\in E\}.$ (4.3)
 It is a subgroup of $\operatorname{Gal}(E/\mathbb{Q})$.
-For a line-SIC $S$, we will show there is a Galois field $E$ (depending on $S$) for which $C(\sigma_{c}\!\mid\!E)$ is such a restricted set.
-However, neither the field $F^{\rm proj}\!\left(S\right)$ nor $F^{\rm eproj}\!\left(S\right)$ is known to always be a Galois extension of $\mathbb{Q}$, although there is no known counterexample for $d\neq 3$.
-Thus, we give a name to the Galois closure of $F^{\rm eproj}\!\left(S\right)$.
+For a algebraic line-SIC $S$, we will show there is a Galois field $E$ (depending on $S$) for which $C(\sigma_{c}\!\mid\!E)$ is such a restricted set.
+Since neither the field $F^{\rm proj}\!\left(S\right)$ nor $F^{\rm eproj}\!\left(S\right)$ is known to always be a Galois extension of $\mathbb{Q}$, we give a name to the Galois closure of $F^{\rm eproj}\!\left(S\right)$.
 Definition 4.6.
@@
 $S$ =$=\left\{\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{d^{2}}\right\},$ (4.4)
-${\mathbf{\Pi}}(S)$ =$=\left\{\tfrac{1}{d}\Pi_{1},\tfrac{1}{d}\Pi_{2},\ldots,\tfrac{1}{d}\Pi_{d^{2}}\right\}.$ (4.5)
-After rescaling if needed, all the entries of the $\mathbf{v}_{i}$ lie in $F^{\rm vec}\!\left(S\right)$, hence they lie in $F^{\rm proj}\!\left(S\right)$, so they lie in $F^{\rm Gproj}\!\left(S\right)$.
-The entries of the $\Pi_{i}$ also lie in $F^{\rm proj}\!\left(S\right)$, hence $F^{\rm Gproj}\!\left(S\right)$.
-Also,
+${\mathbf{\Pi}}(S)$ =$=\left\{\tfrac{1}{d}\Pi_{1},\tfrac{1}{d}\Pi_{2},\ldots,\tfrac{1}{d}\Pi_{d^{2}}\right\},$ (4.5)
+where $\mathbf{v}_{i}$ are unit vectors.
+By definition of $F^{\rm proj}\!\left(S\right)$, the entries of all the $\Pi_{i}$ lie in $F^{\rm proj}\!\left(S\right)$, so they lie in $F^{\rm Gproj}\!\left(S\right)$.
+Also, using the commutativity of $\sigma$ and $\sigma_{c}$,
 $\sigma(\Pi_{i})=\sigma(\mathbf{v}_{i}\sigma_{c}(\mathbf{v}_{i})^{\top})=\sigma(\mathbf{v}_{i})\sigma(\sigma_{c}(\mathbf{v}_{i})^{\top})=\sigma(\mathbf{v}_{i})\sigma_{c}(\sigma(\mathbf{v}_{i})^{\top})$ (4.6)
-is the Hermitian projection onto $\sigma(\mathbf{v}_{i})$, so
+is the Hermitian projection onto $\sigma(\mathbf{v}_{i})$, hence
 ${\mathbf{\Pi}}(\sigma(S))=\left\{\tfrac{1}{d}\sigma(\Pi_{1}),\tfrac{1}{d}\sigma(\Pi_{2}),\ldots,\tfrac{1}{d}\sigma(\Pi_{d^{2}})\right\}.$ (4.7)
@@
 $\operatorname{Tr}(\sigma(\Pi_{i})\sigma(\Pi_{j}))=\sigma(\operatorname{Tr}(\Pi_{i}\Pi_{j}))=\sigma\!\left(\tfrac{\delta_{ij}d+1}{d+1}\right)=\tfrac{\delta_{ij}d+1}{d+1},$ (4.8)
-so ${\mathbf{\Pi}}(\sigma(S))$ is a SIC.
-Finally, $F^{\rm Gproj}\!\left(\sigma(S)\right)\subseteq F^{\rm Gproj}\!\left(S\right)$, because the entries of $\Pi_{i}$ lie in $F^{\rm Gproj}\!\left(S\right)$, and $F^{\rm Gproj}\!\left(S\right)$ is a Galois field fixed by $\sigma$.
+certifying that ${\mathbf{\Pi}}(\sigma(S))$ is a SIC.
+Finally, $F^{\rm Gproj}\!\left(\sigma(S)\right)\subseteq F^{\rm Gproj}\!\left(S\right)$, because the entries of $\Pi_{i}$ lie in $F^{\rm Gproj}\!\left(S\right)$, and $F^{\rm Gproj}\!\left(S\right)$ is Galois over $\mathbb{Q}$, hence $\sigma(F^{\rm Gproj}\!\left(S\right)=F^{\rm Gproj}\!\left(S\right)$.
 It follows by symmetry that $F^{\rm Gproj}\!\left(S\right)=F^{\rm Gproj}\!\left(\sigma^{-1}(\sigma(S))\right)\subseteq F^{\rm Gproj}\!\left(\sigma(S)\right)$ and thus $F^{\rm Gproj}\!\left(\sigma(S)\right)=F^{\rm Gproj}\!\left(S\right)$.
@@
 4.3. Galois multiplets and super-equivalence classes of Weyl–Heisenberg SICs
-We now assemble $\operatorname{PEC}(d)$-orbits of Weyl–Heisenberg SICs in dimension $d$ into multiplets under the Galois action by the subgroup $C(\sigma_{c}\!\mid\!F^{\rm eproj}\!\left(S\right))$ of $\operatorname{Gal}(F^{\rm eproj}\!\left(S\right)/\mathbb{Q})$ that commutes with complex conjugation.
-The coarser equivalence classes obtained under the combined geometric and Galois actions (i.e., unions of the orbits in a multiplet) will called Galois super-equivalence classes.
-The conjecture of Appleby, Flammia, McConnell, and Yard [10, Conj. 2] implies that many of the fields we have been studying are equivalent Galois extensions of $\mathbb{Q}$.
+We now assemble $\operatorname{PEC}(d)$-orbits of Weyl–Heisenberg SICs in dimension $d$ into multiplets under the Galois action by the subgroup $C(\sigma_{c}\!\mid\!F^{\rm Gproj}\!\left(S\right))$ of $\operatorname{Gal}(F^{\rm Gproj}\!\left(S\right)/\mathbb{Q})$ that commutes with complex conjugation.
+The coarser equivalence classes obtained under the combined geometric and Galois actions (i.e., unions of the orbits in a multiplet) will be called Galois super-equivalence classes.
+The conjecture of Appleby, Flammia, McConnell, and Yard [9, Conj. 2] implies that many of the fields we have been studying are equivalent Galois extensions of $\mathbb{Q}$.
 Lemma 4.8.
-Assume 1.1 holds. Then in every dimension $d\geq 4$, for each Weyl–Heisenberg line-SIC $S$, the field $F^{\rm vec}\!\left(S\right)$ is a Galois extension of $\mathbb{Q}$, and
+Assume 1.1 holds. Then in every dimension $d\geq 4$, for each Weyl–Heisenberg line-SIC $S$, the field $F^{\rm vec}\!\left(S\right)$ is a finite Galois extension of $\mathbb{Q}$, and
 $F^{\rm vec}\!\left(S\right)=F^{\rm proj}\!\left(S\right)=F^{\rm eproj}\!\left(S\right)=F^{\rm Gproj}\!\left(S\right).$ (4.9)
 Proof.
-By assertion (2) of 1.1 $F^{\rm vec}\!\left(S\right)$ is a Galois extension of $\mathbb{Q}$.
+By 1.1(2), $F^{\rm vec}\!\left(S\right)$ is a Galois extension of $\mathbb{Q}$.
 Consequently, $F^{\rm vec}\!\left(S\right)$ is preserved by complex conjugation, so $F^{\rm vec}\!\left(S\right)=F^{\rm proj}\!\left(S\right)$ holds by 3.5(1).
-By assertion (2) of 1.1 $F^{\rm vec}\!\left(S\right)$ always contains the ray class field $H_{d^{\prime}\infty_{1}\infty_{2}}$ of $\mathbb{Q}(\Delta_{d})$.
+Also by 1.1(2), $F^{\rm vec}\!\left(S\right)$ always contains the ray class field $H_{d^{\prime}\infty_{1}\infty_{2}}$ of $\mathbb{Q}(\Delta_{d})$.
 This class field must contain the class field of $\mathbb{Q}$ with conductor $d^{\prime}\infty$, which is $\mathbb{Q}(\zeta_{d^{\prime}})=\mathbb{Q}(\xi_{d})$.
 Therefore, $F^{\rm vec}\!\left(S\right)$ contains $F^{\rm proj}\!\left(S\right)(\xi_{d})=F^{\rm eproj}\!\left(S\right)$.
-Since $F^{\rm vec}\!\left(S\right)\subseteq F^{\rm proj}\!\left(S\right)\subseteq F^{\rm eproj}\!\left(S\right)$, the equality of those three fields follows.
-So $F^{\rm eproj}\!\left(S\right)$ is Galois and thus also equal to $F^{\rm Gproj}\!\left(S\right)$.
+Since $F^{\rm vec}\!\left(S\right)\subseteq F^{\rm proj}\!\left(S\right)\subseteq F^{\rm eproj}\!\left(S\right)$, equality of those three fields follows.
+So $F^{\rm eproj}\!\left(S\right)$ is Galois and hence equal to $F^{\rm Gproj}\!\left(S\right)$.
 ∎
-We extract the following conjecture, of which all but the claim about $F^{\rm trip}\!\left(S\right)$ being Galois is already a consequence of 1.1.
-Conjecture 4.10.
-Fix a positive integer $d\neq 3$.
-For any $d$-dimensional Weyl–Heisenberg line SIC $S$, its $d^{2}$ vectors are always projective-algebraic.
-The associated fields $F^{\rm trip}\!\left(S\right)$ and $F^{\rm proj}\!\left(S\right)$ (and thus also $F^{\rm eproj}\!\left(S\right)$) are Galois extensions of $\mathbb{Q}$.
 The evidence for a Galois action on Weyl–Heisenberg SICs was extensively detailed by Appleby, Yadsan-Appleby, and Zauner [4] in 2013.
-The group $\operatorname{Gal}(\overline{\mathbb{Q}})/\mathbb{Q})$ acts on the elements of Weyl–Heisenberg group $\operatorname{WH}(d)$ coordinatewise on its matrices.
+The group $\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ acts on the elements of Weyl–Heisenberg group $\operatorname{WH}(d)$ coordinatewise on its matrices.
 Any automorphism $\sigma:\overline{\mathbb{Q}}\to\overline{\mathbb{Q}}$, sending a matrix $H\mapsto\sigma(H)$, acts as a homomorphism $\sigma(H_{1}H_{2})=\sigma(H_{1})\sigma(H_{2})$.
 This Galois action fixes the group $\operatorname{WH}(d)$ as a set, with any automorphism permuting its elements, since every Galois conjugate $\sigma(H)$ of a group element $H$ is also an element of $\operatorname{WH}(d)$.
-By 3.8(2), $F^{\rm eproj}\!\left(S\right)$ is an invariant of the $\operatorname{PEC}(d)$ orbit of the Weyl–Heisenberg SIC $S$.
-If we assume 4.10, then $F^{\rm eproj}\!\left(S\right)$ is preserved by the Galois action by the subgroup $C(\sigma_{c}\!\mid\!F^{\rm eproj}\!\left(S\right))$ of $\operatorname{Gal}(F^{\rm eproj}\!\left(S\right)/\mathbb{Q})$.
-In other words, the extended projection field is invariant under both the geometric and the Galois actions.
-We make the following definition.
+By 3.8(2), $F^{\rm eproj}\!\left(S\right)$ is an invariant of the $\operatorname{PEC}(d)$ orbit of the Weyl–Heisenberg SIC $S$, hence so is its Galois closure $F^{\rm Gproj}\!\left(S\right)$.
+Thus the field $F^{\rm Gproj}\!\left(S\right)$ is invariant under both the geometric and the Galois actions.
+The following definition coincides with “Galois multiplet” as defined in [8].
 Definition 4.9.
-Assume 4.10 holds. For $d\geq 4$ the Galois multiplet of a $\operatorname{PEC}(d)$-equivalence class $[S]$ of Weyl–Heisenberg line-SICs is
-$[[S]]=\{[\sigma(S)]:\sigma\in C(\sigma_{c}\!\mid\!F^{\rm eproj}\!\left(S\right))\},$ (4.10)
-that is, the set of Galois equivalent orbits. We also define the equivalence relation $\approx_{GPEC}$, Galois super-equivalence, by
-$S_{1}\approx_{GPEC}S_{2}\iff[[S_{1}]]=[[S_{2}]],$ (4.11)
-and term the equivalence class of a SIC $S$ under $\approx_{GPEC}$ (that is, the union of the elements of $[[S]]$) the Galois super-equivalance class of $S$.
-Remark 4.12.
-Assuming 1.1 holds, this definition coincides with that of “Galois multiplet” in [9].
+For $d\geq 4$, the Galois multiplet of a $\operatorname{PEC}(d)$-equivalence class $[S]$ of Weyl–Heisenberg line-SICs is
+$[[S]]=\{[\sigma(S)]:\sigma\in C(\sigma_{c}\!\mid\!F^{\rm Gproj}\!\left(S\right))\}.$ (4.10)
+We also define an associated equivalence relation $\approx_{GPEC}$, termed Galois super-equivalence, by
+$[S_{1}]\approx_{GPEC}[S_{2}]\iff[[S_{1}]]=[[S_{2}]],$ (4.11)
+and term the union of the elements $[S_{i}]$ of $[[S]]$ the Galois super-equivalance class of $S$.
 We summarize the properties noted above in the following proposition.
 Proposition 4.10.
-Assume 4.10 holds.
-Then, the set $\operatorname{WHSIC}(d)$ is divided into a finite number of Galois super-equivalence classes under the equivalence relation $\approx_{GPEC}$.
+Assume 1.1 holds.
+Then for $d\geq 4$ the set $\operatorname{WHSIC}(d)$ is partitioned into a finite number of Galois super-equivalence classes under the equivalence relation $\approx_{GPEC}$.
 Each Galois super-equivalence class is a finite union of (geometric) $\operatorname{PEC}(d)$-equivalence classes.
-Galois multiplets are studied in Sections 7 and 8 using data and more refined conjectures. The multiplets in a fixed dimension often have different sizes.
-5. A conjecture on counting geometric equivalence classes of Weyl–Heisenberg SICs
-In this section, we formulate a conjecture relating the number of SICs in dimensions $d\geq 4$ to class numbers of certain real quadratic orders. We present numerical evidence.
+Galois multiplets in a fixed dimension often have different sizes.
+5. Conjectures on counting geometric equivalence classes and Galois multiplets of Weyl–Heisenberg SICs
+In this section, we formulate two conjectures relating the number of SICs in dimensions $d\geq 4$ to class numbers of certain real quadratic orders.
+Both these conjectures are numerically testable, and they are consequences of 1.2.
 5.1. Class groups and class monoids of orders
@@
 $\operatorname{Clm}(\mathcal{O}):=\frac{\{\mbox{nonzero fractional ideals of }\mathcal{O}\}}{\{\mbox{principal fractional ideals of }\mathcal{O}\}}.$ (5.2)
-For a more in-depth discussion of the nuances of orders, see [33] and [35, Ch. I, Sec. 12].
+For a more in-depth discussion of the nuances of orders, see [34] and [37, Ch. I, Sec. 12].
 If $K$ is a quadratic field of discriminant $\Delta_{0}$, it has a unique suborder of discriminant $f^{2}\Delta_{0}$ for every positive integer $f$. This order is given explicitly as
 $\mathcal{O}_{\!\Delta}:=\mathbb{Z}\!\left[\frac{\Delta+\sqrt{\Delta}}{2}\right]=\mathbb{Z}+\frac{\Delta+\sqrt{\Delta}}{2}\mathbb{Z}=\mathbb{Z}+f\frac{\Delta_{0}+\sqrt{\Delta_{0}}}{2}\mathbb{Z}.$ (5.3)
-For the SIC problem, we are specifically interested in the case when $\Delta=\Delta_{d}=(d+1)(d-3)$ for some integer $d$.
+For the SIC problem, we specifically consider the case when $\Delta=\Delta_{d}=(d+1)(d-3)$ for some integer $d$.
 In this case, the class monoid $\operatorname{Clm}(\mathcal{O}_{\!\Delta_{d}})$ may be understood through bijections with other sets of number-theoretic interest.
@@
   • (2)
-    the set $\mathcal{Q}(\Delta)/\mathbf{GL}_{2}(\mathbb{Z})$, where $\mathcal{Q}(\Delta)$ is the set of binary quadratic forms of discriminant $\Delta$, and ${A}=\left(\begin{smallmatrix}s&t\\ u&v\end{smallmatrix}\right)\in\mathbf{GL}_{2}(\mathbb{Z})$ acts on the right by “twisted precomposition” $Q^{{A}}(x,y)=(\det{{A}})Q(sx+ty,ux+vy)$; and
+    the set $\mathcal{Q}(\Delta)/\mathbf{GL}_{2}(\mathbb{Z})$, where $\mathcal{Q}(\Delta)$ is the set of binary quadratic forms of discriminant $\Delta$, and ${A}=\left(\begin{smallmatrix}s&t\\ u&v\end{smallmatrix}\right)\in\mathbf{GL}_{2}(\mathbb{Z})$ acts on the right by $Q^{{A}}(x,y)=(\det{{A}})Q(sx+ty,ux+vy)$; and
   • (3)
@@
 The bijection between (0) and (1) holds for quadratic orders in general.
-Specicially, by [33, Prop.
-A.12 and Prop.
-A.4], $\operatorname{Clm}(\mathcal{O}_{\Delta})=\operatorname{Clm}^{\flat}(\mathcal{O}_{\Delta})\cong\bigsqcup_{f^{\prime}|f}\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})$.
+Specifically, by Halter-Koch [28, Sec.
+5.5], one has, as sets, $\operatorname{Clm}(\mathcal{O}_{\Delta})=\bigsqcup_{f^{\prime}|f}\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})$.
 The bijection between (0) and (2) also holds for quadratic orders in general (and with appropriate modifications, nondegenerate quadratic rings) and is known as twisted Gauss composition.
 This is proven as [47, Thm.
-1.2]; for an introduction to Gauss composition, see [16, Ch.
+1.2]; for an introduction to Gauss composition, see [15, Ch.
 5].
-We now exhibit a map $\phi:\mathcal{Q}(\Delta)\to\mathbf{GL}_{2}(\mathbb{Z})$ that induces a bijection between (2) and (3).
+We now exhibit a bijection between (2) and (3) which specifically holds only for $\Delta=(d+1)(d-3)$.
+It is defined by a map $\phi:\mathcal{Q}(\Delta)\to\mathbf{GL}_{2}(\mathbb{Z})$ that induces the bijection.
 For a quadratic form $Q(x,y)=ax^{2}+bxy+cy^{2}$ of discriminant $\Delta=(d+1)(d-3)$, define $\phi(Q):=\left(\begin{smallmatrix}\frac{1}{2}(b+d-1)&-a\\ c&\frac{1}{2}(-b+d-1)\end{smallmatrix}\right)$.
-Note that $\det(\phi(Q))=\frac{1}{4}(-b^{2}+(d-1)^{2})+ac=\frac{1}{4}(-\Delta+(d-1)^{2})=1$ and $\operatorname{Tr}(\phi(Q))=d-1$.
+Note that
+$\det(\phi(Q))=\frac{1}{4}(-b^{2}+(d-1)^{2})+ac=\frac{1}{4}(-\Delta+(d-1)^{2})=1$ (5.4)
+and $\operatorname{Tr}(\phi(Q))=d-1$.
 It is straightforward to check that $\phi(Q^{R})={R}^{-1}\phi(Q){R}$ for ${R}\in\mathbf{GL}_{2}(\mathbb{Z})$.
 Thus, $\phi$ defines a map $\overline{\phi}:\mathcal{Q}(\Delta)/\mathbf{GL}_{2}(\mathbb{Z})\to\mathcal{C}_{d-1}$.
-In the other direction, for ${A}=\left(\begin{smallmatrix}s&t\\ u&v\end{smallmatrix}\right)\in SL_{2}(\mathbb{Z})$, define the quadratic form $(\psi({A}))(x,y):=-tx^{2}+(s-v)xy+uy^{2}$.
+In the other direction, for ${A}=\left(\begin{smallmatrix}s&t\\ u&v\end{smallmatrix}\right)\in\mathbf{SL}_{2}(\mathbb{Z})$, define the quadratic form $(\psi({A}))(x,y):=-tx^{2}+(s-v)xy+uy^{2}$.
 It is straightforward to check that $\operatorname{disc}(\psi({A}))=\Delta$ and $\psi({R}^{-1}{A}{R})=\psi({A})^{R}$, so $\psi$ defines a map $\overline{\psi}:\mathcal{C}_{d-1}\to\mathcal{Q}(\Delta)/\mathbf{GL}_{2}(\mathbb{Z})$.
@@
 5.2. Statement of the Geometric Count Conjecture
-The following conjecture counts the number of geometric equivalence classes of Weyl–Heisenberg SICs in dimension $d$.
+The following conjecture counts the number of geometric equivalence classes of Weyl–Heisenberg SICs in dimension $d$. It is implied by 1.2.
 Conjecture 5.4 (Geometric Count Conjecture).
@@
     $\left|\mathcal{C}_{d-1}\right|$, the number of $\mathbf{GL}_{2}(\mathbb{Z})$-conjugacy class of elements of $\mathbf{SL}_{2}(\mathbb{Z})$ of trace $d-1$.
+5.4 immediately gives an upper bound for the total number of Weyl–Heisenberg line-SICs in dimensions $d\geq 4$.
+Proposition 5.5.
+5.4 implies that the number $\left|\operatorname{WHSIC}(d)\right|$ of Weyl–Heisenberg line-SICs in dimensions $d\geq 4$ satisfies
+$\left|\operatorname{WHSIC}(d)\right|\leq\frac{\left|\operatorname{PEC}(d)\right|}{d^{2}}\left|\operatorname{Clm}(\mathcal{O}_{\Delta_{d}})\right|.$ (5.6)
+Proof.
+The number of distinct Weyl–Heisenberg line-SICs in a $\operatorname{PEC}(d)$-equivalence class $[S]$ is $\frac{\left|\operatorname{PEC}(d)\right|}{\left|\operatorname{stab}(S)\right|}$, where $\operatorname{stab}(S)$ is the subgroup of $\operatorname{PEC}(d)$ stabilizing a line-SIC $S\in[S]$.
+The cardinality $\left|\operatorname{stab}(S)\right|$ is independent of the choice of $S\in[S]$.
+The Weyl–Heisenberg group $\operatorname{WH}(d)$ acts transitively on the set of $d^{2}$ complex lines in the $\operatorname{WH}(d)$-orbit of a fiducial vector, so $\operatorname{WH}(d)$ is a subgroup of $\operatorname{stab}(S)$.
+Thus $d^{2}$ divides $\left|\operatorname{stab}(S)\right|$.
+Hence we have $\frac{\left|\operatorname{PEC}(d)\right|}{\left|\operatorname{stab}(S)\right|}\leq\frac{\left|\operatorname{PEC}(d)\right|}{d^{2}}$.
+The result now follows from 5.4.
+∎
+Remark 5.6.
+Table 5.1 presents, in the column labeled $\left|\operatorname{WHSIC}(d)\right|$, empirical data for the number of observed (numerical) SICs for $4\leq d\leq 50$ from Scott and Grassl [41, Table 1].
+That table also observes that all known numerical Weyl–Heisenberg SICs of dimensions $4\leq d\leq 50$ have an additional stabilizer element of order $3$.
+If this statement were true in general, then we could strengthen 5.5 to conclude
+$\left|\operatorname{WHSIC}(d)\right|\leq\frac{1}{3}\frac{\left|\operatorname{PEC}(d)\right|}{d^{2}}\left|\operatorname{Clm}(\mathcal{O}_{\Delta_{d}})\right|$ (5.7)
+holds for all $d\geq 4$. Examination of Table 5.1 shows for known numerical SICs that (5.7) holds for $4\leq d\leq 50$ and that equality in (5.7) holds for
+$d\in\{5,13,17,23,25,27,29,37,41,43,47,49\}.$ (5.8)
 5.3. Verification of cases of the Geometric Count Conjecture
-We have verified that the two sets in the Geometric Count conjecture have the same size for dimensions $d\leq 90$, $d\neq 3$ under the assumption that the published lists of Weyl–Heisenberg SICs in these dimensions are complete.
-Scott [38] gives a putatively complete list of numerical SICs in all dimensions $d\leq 90$, building on earlier work of Scott and Grassl [39].
-Scott and Grassl find SIC fiducials by numerically minimizing a frame potential—a function whose global minima are Weyl–Heisenberg SIC fiducials.
-By repeating their search with a large number of seed vectors, they establish a high degree of numerical confidence in their results but do not rigorously rule out the existence of other SICs.
-The dimensions $d\in\{1,2\}$ are special cases—the only dimensions for which $\Delta<0$, or equivalently, the conjugacy classes in $\mathcal{C}_{d-1}$ are elliptic.
-For $d=1$, there is exactly one SIC ($\mathbb{C}^{1}$ as a line in $\mathbb{C}^{1}$), and it is (trivially) Weyl–Heisenberg covariant.
-The associated order $\mathcal{O}_{-4}=\mathbb{Z}[i]$ is a maximal order of class number one, so the conjecture is true.
-For $d=2$, there is again a unique SIC up to unitary transformations (a regular tetrahedron on the Riemann sphere $\mathbb{P}^{1}(\mathbb{C})$), and it may be chosen to be Weyl–Heisenberg covariant.
-The associated order $\mathcal{O}_{-3}=\mathbb{Z}[e^{2\pi i/3}]$ is a maximal order of class number one, so the conjecture is true.
-5.4 omits the case $d=3$, where $\Delta=0$, where associated field $K_{\!\Delta}=\mathbb{Q}$ has $\mathcal{O}_{\!\Delta}=\operatorname{Z}$.
-In this case, there is a continuous one-parameter family of Weyl–Heisenberg covariant SICs [39].
-For $d=3$ the classification (3) above is the set $\mathcal{C}_{2}$ of parabolic conjugacy classes of $\mathbf{SL}_{2}(\mathbb{Z})$, which is infinite.
-For $4\leq d\leq 90$ we computed $\left|\operatorname{Clm}(\mathcal{O}_{\Delta})\right|$ using Magma [17] by computing the equivalent quantity $\left|\mathcal{Q}(\Delta)/\mathbf{GL}_{2}(\mathbb{Z})\right|$.
-Write $\Delta=(d+1)(d-3)=f^{2}\Delta_{0}$ for a fundamental discriminant $\Delta_{0}$.
+We have verified that the two sets in the Geometric Count conjecture have the same size for dimensions $4\leq d\leq 90$, under the assumption that the published lists of Weyl–Heisenberg SICs in these dimensions, given in Scott [40], are complete.
+(The list builds on earlier work of Scott and Grassl [41].)
+Proposition 5.7 (Empirical Proposition 1).
+For dimensions $d$ with $\Delta=(d+1)(d-3)$, the total number of $\operatorname{PEC}(d)$ equivalence classes $[S]$ of numerical SICs equals $\left|\operatorname{Clm}(\mathcal{O}_{\Delta})\right|$ for $4\leq d\leq 50$ in the Scott and Grassl list [41], and for $51\leq d\leq 90$ in the Scott list [40].
+Proof.
+For $4\leq d\leq 90$, we computed $\left|\operatorname{Clm}(\mathcal{O}_{\Delta})\right|$ using Magma [16] by computing the equivalent quantity $\left|\mathcal{Q}(\Delta)/\mathbf{GL}_{2}(\mathbb{Z})\right|$.
+Write $\Delta=(d+1)(d-3)=f^{2}\Delta_{0}$ for a fundamental discriminant $\Delta_{0}$. Recall that for any quadratic order
+$\operatorname{Clm}(\mathcal{O}_{\Delta})=\bigsqcup_{f^{\prime}|f}\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}}).$ (5.9)
 For each $f^{\prime}\mid f$, we computed the set of reduced primitive binary quadratic forms of discriminant $(f^{\prime})^{2}\Delta_{0}$ using the Magma function ReducedForms(), written by David Kohel.
@@
 Multiplying by $f/f^{\prime}$ and joining the lists gave a list of representatives of all (not necessarily primitive) classes in $\mathcal{Q}(\Delta)/\mathbf{GL}_{2}(\mathbb{Z})$.
-We observe that for each $d$ ($4\leq d\leq 90$) that the size of
-$\operatorname{Clm}(\mathcal{O}_{\Delta})=\sum_{f^{\prime}|f}\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})$ (5.5)
-is equal to the sum of the corresponding values in the “#” column of Scott’s table beginning on page 3 of [38], which give the number of elements of ${\rm WHSIC}(d)/\operatorname{PEC}(d)$.
-This is agreement is shown in Table 1.
-The table also includes the sizes of the ring class groups $\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})$ (equivalently, the number of twisted $\mathbf{GL}_{2}(\mathbb{Z})$-classes of primitive binary quadratic forms with discriminant $(f^{\prime})^{2}\Delta_{0}$).
-$d$ $\Delta_{0}$ $f^{\prime}$ $h_{(f^{\prime})^{2}\Delta_{0}}$#\##
-$1$ $-4$ $1$ $1$ $1$
-$2$ $-3$ $1$ $1$ $1$
-$3$ 00 $\ast$ $\ast$ $\infty$
-$4$ $5$ $1$ $1$ $1$
-$5$ $12$ $1$ $1$ $1$
-$6$ $21$ $1$ $1$ $1$
-$7$ $8$ $1,2$ $1,1$ $2$
-$8$ $5$ $1,3$ $1,1$ $2$
-$9$ $60$ $1$ $2$ $2$
-$10$ $77$ $1$ $1$ $1$
-$11$ $24$ $1,2$ $1,2$ $3$
-$12$ $31$,3131,31 , 3 $1,1$ $2$
-$13$ $140$ $1$ $2$ $2$
-$14$ $165$ $1$ $2$ $2$
-$15$ $21$,2,41241,2,41 , 2 , 4 $1,1,2$ $4$
-$16$ $221$ $1$ $2$ $2$
-$17$ $28$ $1,3$ $1,2$ $3$
-$18$ $285$ $1$ $2$ $2$
-$19$ $5$ $1,2,4,8$ $1,1,1,2$ $5$
-$20$ $357$ $1$ $2$ $2$
-$21$ $44$ $1,3$ $1,4$ $5$
-$22$ $437$ $1$ $1$ $1$
-$23$ $201$,2121,21 , 2 $2,4$ $6$
-$24$ $21$ $1,5$ $1,2$ $3$
-$25$ $572$ $1$ $2$ $2$
-$26$ $69$ $1,3$ $1,3$ $4$
-$27$ $681$,2121,21 , 2 $2,4$ $6$
-$28$ $29$ $1,5$ $1,2$ $3$
-$29$ $780$ $1$ $4$ $4$
-$30$ $93$ $1,3$ $1,3$ $4$
-$31$ $56$ $1,2,4$ $1,2,4$ $7$
-$32$ $957$ $1$ $2$ $2$
-$34$ $1020$ $1$ $4$ $4$
-$34$ $1085$ $1$ $2$ $2$
-$35$ $8$ $1,2,3,4,6,12$ $1,1,1,1,2,4$ $10$
-$36$ $1221$ $1$ $4$ $4$
-$37$ $1292$ $1$ $4$ $4$
-$38$ $1365$ $1$ $4$ $4$
-$39$ $40$ $1,2,3,6$ $2,2,2,4$ $10$
-$40$ $1517$ $1$ $2$ $2$
-$41$ $1596$ $1$ $8$ $8$
-$42$ $1677$ $1$ $4$ $4$
-$43$ $440$ $1,2$ $2,4$ $6$
-$44$ $205$ $1,3$ $2,4$ $6$
-$45$ $1932$ $1$ $4$ $4$
- $d$ $\Delta_{0}$ $f^{\prime}$ $h_{(f^{\prime})^{2}\Delta_{0}}$#\## $46$ $2021$ $1$ $3$ $3$ $47$ $33$ $1,2,4,8$ $1,1,2,4$ $8$ $48$ $5$ $1,3,7,21$ $1,1,1,4$ $7$ $49$ $92$ $1,5$ $1,6$ $7$ $50$ $2397$ $1$ $2$ $2$ $51$ $561$,2,41241,2,41 , 2 , 4 $2,4,8$ $14$ $52$ $53$ $1,7$ $1,3$ $4$ $53$ $21$,3,$5,1$ 3 , 5 , 15 $1,1,2,6$ $10$ $54$ $2805$ $1$ $4$ $4$ $55$ $728$ $1,2$ $2,4$ $6$ $56$ $3021$ $1$ $6$ $6$ $57$ $348$ $1,3$ $2,6$ $8$ $58$ $3245$ $1$ $4$ $4$ $59$ $840$ $1,2$ $4,8$ $12$ $60$ $3477$ $1$ $4$ $4$ $61$ $3596$ $1$ $6$ $6$ $62$ $413$ $1,3$ $1,4$ $5$ $63$ $60$ $1,2,4,8$ $2,2,4,8$ $16$ $64$ $3965$ $1$ $4$ $4$ $65$ $4092$ $1$ $8$ $8$ $66$ $469$ $1,3$ $3,6$ $9$ $67$ $71$,2,4,8,161248161,2,4,8,161 , 2 , 4 , 8 , 16 $1,1,1,2,4$ $9$ $68$ $4485$ $1$ $4$ $4$ $69$ $4620$ $1$ $8$ $8$ $70$ $4757$ $1$ $5$ $5$ $71$ $361$,2,3,612361,2,3,61 , 2 , 3 , 6 $2,4,4,8$ $18$ $72$ $5037$ $1$ $4$ $4$ $73$ $5180$ $1$ $4$ $4$ $74$ $213$ $1,5$ $1,6$ $7$ $75$ $521$,2,3,612361,2,3,61 , 2 , 3 , 6 $1,2,4,8$ $15$ $76$ $5621$ $1$ $6$ $6$ $77$ $5772$ $1$ $8$ $8$ $78$ $237$ $1,5$ $1,6$ $7$ $79$ $380$ $1,2,4$ $2,4,8$ $14$ $80$ $77$ $1,3,9$ $1,2,6$ $9$ $81$ $6396$ $1$ $12$ $12$ $82$ $6557$ $1$ $3$ $3$ $83$ $051$,2,4,812481,2,4,81 , 2 , 4 , 8 $2,2,4,8$ $16$ $84$ $85$ $1,3,9$ $2,2,6$ $10$ $85$ $7052$ $1$ $4$ $4$ $86$ $7221$ $1$ $10$ $10$ $87$ $8481$,2121,21 , 2 $4,8$ $12$ $88$ $7565$ $1$ $4$ $4$ $89$ $860$ $1,3$ $2,8$ $10$ $90$ $7917$ $1$ $4$ $4$
-Table 1.
-Comparison of ring class numbers and counts of SICs for $d\leq 90$.
-Here $(d+1)(d-3)=f^{2}\Delta_{0}$ for a fund.
+The counts of quantities for the different $f^{\prime}$ are given are the unordered set of all ring class numbers $h_{(f^{\prime})^{2}\Delta_{0}}=\left|\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})\right|$.
+They are given in column labeled by $h_{(f^{\prime})^{2}\Delta_{0}}$ in Table 5.1.
+The total counts of class numbers are , are given in the column labeled by $\left|\operatorname{Clm}(\mathcal{O}_{\Delta})\right|$.
+Numerical Weyl–Heisenberg SICs are systems of $d^{2}$ complex lines that are the Weyl–Heisenberg orbit of a single numerical fiducial vector and that appears to have equal angles to $10$ decimal places.
+Numerical SICs are produced as output of an optimization algorithm.
+Empirical counts for the number of $\operatorname{PEC}(d)$-equivalence classes of numerical SICs were given by Scott and Grassl for $4\leq d\leq 50$ [41, Tab.
+1 and Tab.
+2] and by Scott [40, pp.
+3–6] for $51\leq d\leq 90$.
+The classes in the tables are denoted by dimension and letter, eg., $31a$.
+Tab.
+1 of [41] contains the subset of exact SICs, grouped into Galois multiplets, while Tab.
+2 of [41] gives a list of numerical SICs, not grouped into Galois multiplets.
+In both cases one can count the total number of $\operatorname{PEC}(d)$-equivalence classes.
+We also found agreement for all dimensions $51\leq d\leq 90$ of $\left|\operatorname{Clm}(\mathcal{O}_{\Delta_{d}})\right|$ given in Table 5.2 with the total count of $\operatorname{PEC}(d)$-classes of SICs in each dimension listed in [40, Appendix C].
+∎
+$d$$\frac{\left|\operatorname{PEC}(d)\right|}{d^{2}}h_{(f^{\prime})^{2}\Delta_{0}}$ $\Delta$ $\Delta_{0}$ $f^{\prime}$ $\sigma_{0}(f)$ $h_{(f^{\prime})^{2}\Delta_{0}}$ $\left|\operatorname{Clm}(\mathcal{O}_{\Delta})\right|$
+$1$ $-4$ $-4$ $1$ $1$ $1$ $1$
+$2$ $-3$ $-3$ $1$ $1$ $1$ $1$
+$3$ 00 00 $\mathbb{N}$infty $1$infty
+$4$ $5$ $5$ $1$ $1$ $1$ $1$
+$5$ $12$ $12$ $1$ $1$ $1$ $1$
+$6$ $21$ $21$ $1$ $1$ $1$ $1$
+$7$ $32$ $8$ $1,2$ $2$ $1,1$ $2$
+$8$ $45$ $5$ $1,3$ $2$ $1,1$ $2$
+$9$ $60$ $60$ $1$ $1$ $2$ $2$
+$10$ $77$ $77$ $1$ $1$ $1$ $1$
+$11$ $96$ $24$ $1,2$ $2$ $1,2$ $3$
+$12$ $117$ $31$,3131,31 , 3 $2$ $1,1$ $2$
+$13$ $140$ $140$ $1$ $1$ $2$ $2$
+$14$ $165$ $165$ $1$ $1$ $2$ $2$
+$15$ $192$ $21$,2,41241,2,41 , 2 , 4 $3$ $1,1,2$ $4$
+$16$ $221$ $221$ $1$ $1$ $2$ $2$
+$17$ $252$ $28$ $1,3$ $2$ $1,2$ $3$
+$18$ $285$ $285$ $1$ $1$ $2$ $2$
+$19$ $320$ $5$ $1,2,4,8$ $4$ $1,1,1,2$ $5$
+$20$ $357$ $357$ $1$ $1$ $2$ $2$
+$21$ $396$ $44$ $1,3$ $2$ $1,4$ $5$
+$22$ $437$ $437$ $1$ $1$ $1$ $1$
+$23$ $480$ $201$,2121,21 , 2 $2$ $2,4$ $6$
+$24$ $525$ $21$ $1,5$ $2$ $1,2$ $3$
+$25$ $572$ $572$ $1$ $1$ $2$ $2$
+$26$ $621$ $69$ $1,3$ $2$ $1,3$ $4$
+$27$ $672$ $681$,2121,21 , 2 $2$ $2,4$ $6$
+$28$ $725$ $29$ $1,5$ $2$ $1,2$ $3$
+$29$ $780$ $780$ $1$ $1$ $4$ $4$
+$30$ $837$ $93$ $1,3$ $2$ $1,3$ $4$
+$31$ $896$ $56$ $1,2,4$ $3$ $1,2,4$ $7$
+$32$ $957$ $957$ $1$ $1$ $2$ $2$
+$34$ $1020$ $1020$ $1$ $1$ $4$ $4$
+$34$ $1085$ $1085$ $1$ $1$ $2$ $2$
+$35$ $1152$ $8$ $1,2,3,4,6,12$ $6$ $1,1,1,1,2,4$ $10$
+$36$ $1221$ $1221$ $1$ $1$ $4$ $4$
+$37$ $1292$ $1292$ $1$ $1$ $4$ $4$
+$38$ $1365$ $1365$ $1$ $1$ $4$ $4$
+$39$ $1440$ $40$ $1,2,3,6$ $4$ $2,2,2,4$ $10$
+$40$ $1517$ $1517$ $1$ $1$ $2$ $2$
+$41$ $1596$ $1596$ $1$ $1$ $8$ $8$
+$42$ $1677$ $1677$ $1$ $1$ $4$ $4$
+$43$ $1760$ $440$ $1,2$ $2$ $2,4$ $6$
+$44$ $1845$ $205$ $1,3$ $2$ $2,4$ $6$
+$45$ $1932$ $1932$ $1$ $1$ $4$ $4$
+$46$ $2021$ $2021$ $1$ $1$ $3$ $3$
+$47$ $2112$ $33$ $1,2,4,8$ $4$ $1,1,2,4$ $8$
+$48$ $2205$ $5$ $1,3,7,21$ $4$ $1,1,1,4$ $7$
+$49$ $2300$ $92$ $1,5$ $2$ $1,6$ $7$
+$50$ $2397$ $2397$ $1$ $1$ $2$ $2$
+ $d$$\frac{\left|\operatorname{PEC}(d)\right|}{d^{2}}h_{(f^{\prime})^{2}\Delta_{0}}$ $\left|\operatorname{WHSIC}(d)\right|$ $\frac{\left|\operatorname{PEC}(d)\right|}{d^{2}}$#\## $1$ $1$ $1$ $1$ $2$ $2$ $12$ $1$ $3$infty $48$infty $4$ $16$ $96$ $1$ $5$ $80$ $240$ $1$ $6$ $96$ $288$ $1$ $7$ $336$ $672$ $2$ $8$ $320$ $768$ $2$ $9$ $864$ $1296$ $2$ $10$ $480$ $1440$ $1$ $11$ $2640$ $2640$ $3$ $12$ $1152$ $2304$ $2$ $13$ $2912$ $4368$ $2$ $14$ $2688$ $4032$ $2$ $15$ $6720$ $5760$ $4$ $16$ $4096$ $6144$ $2$ $17$ $9792$ $9792$ $3$ $18$ $5184$ $7776$ $2$ $19$ $16720$ $13680$ $5$ $20$ $7680$ $11520$ $2$ $21$ $26880$ $16128$ $5$ $22$ $5280$ $15840$ $1$ $23$ $48576$ $24288$ $6$ $24$ $15360$ $18432$ $3$ $25$ $20000$ $30000$ $2$ $26$ $34944$ $26208$ $4$ $27$ $69984$ $34992$ $6$ $28$ $26880$ $32256$ $3$ $29$ $64960$ $48720$ $4$ $30$ $46080$ $34560$ $4$ $31$ $138880$ $59520$ $7$ $32$ $32768$ $49152$ $2$ $33$ $84480$ $63360$ $4$ $34$ $39168$ $58752$ $2$ $35$ $235200$ $80640$ $10$ $36$ $82944$ $62208$ $4$ $37$ $134976$ $101232$ $4$ $38$ $109440$ $82080$ $4$ $39$ $314496$ $104832$ $10$ $40$ $61440$ $92160$ $2$ $41$ $367360$ $137360$ $8$ $42$ $129024$ $96768$ $4$ $43$ $317856$ $158928$ $6$ $44$ $253440$ $126720$ $6$ $45$ $207360$ $155520$ $4$ $46$ $145728$ $145728$ $3$ $47$ $553472$ $207552$ $8$ $48$ $276480$ $147456$ $7$ $49$ $537824$ $230496$ $7$ $50$ $120000$ $180000$ $2$
+Table 5.1.
+Comparison for $1\leq d\leq 50$ of ring class numbers with counts of SICs.
+The three rightmost columns are from [41, Table 1].
+They include empirical data for $\left|\operatorname{WHSIC}(d)\right|$, and column ##\## gives empirical data for $\left|\operatorname{WHSIC}(d)/\operatorname{PEC}(d)\right|$.
+The left columns give $\Delta=(d+1)(d-3)=f^{2}\Delta_{0}$ for a fund.
 disc.
-$\Delta_{0}$, $f^{\prime}$ runs over divisors of $f$, $h_{(f^{\prime})^{2}\Delta_{0}}=\left|\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})\right|$, and ##\## counts W–H SICs of Scott [38] in dim.
-$d$.
-6. Theorems on class monoids of real quadratic fields
-In this section, we prove rigorous results about the behavior of $\operatorname{Clm}(\mathcal{O})$ for general orders. These results have conditional consequences for SICs via 5.4.
-6.1. Growth of the number of SICs in dimension $d$
-Going beyond Zauner’s conjecture, one would like to know not only that SICs exist in every dimension $d$, but also how many there are.
-The connection between SICs and class numbers gives us a precise prediction for the number of SICs up to equivalence (excluding any SICs that are not Weyl–Heisenberg covariant, to which our conjectures are agnostic).
-In this section, we use known results on the growth of class numbers to derive conditional results on the number of SICs as $d\to\infty$.
-The first growth result characterizes those dimensions in which our conjectures predict a unique Weyl–Heisenberg SIC, using a theorem of Byeon, Kim, and Lee [13].
-Proposition 6.1.
-One has $\left|\operatorname{Clm}(\mathcal{O}_{\!\Delta_{d}})\right|>1$ for all $d>22$.
-The values of $d\neq 3$ with $\left|\operatorname{Clm}(\mathcal{O}_{\!\Delta_{d}})\right|=1$ are $d\in\{1,2,4,5,6,10,22\}$.
-Proof.
-Set $\Delta=\Delta_{d}=(d+1)(d-3)$ and $s(d):=\left|\operatorname{Clm}(\mathcal{O}_{\!\Delta})\right|$; we are trying to determine when $s(d)=1$. By 5.4,
-$s(d)=\sum_{f^{\prime}|f}h_{(f^{\prime})^{2}\Delta_{0}}.$ (6.1)
-where $\Delta=f^{2}\Delta_{0}$ and $\Delta_{0}$ is fundamental.
-If $f>1$, then $s(d)>1$.
-If $f=1$, then $\Delta=\Delta_{0}$ is fundamental.
-Write $\Delta=(d-1)^{2}-4=n^{2}-4$, where $n:=d-1$.
-Byeon, Kim, and Lee solved the class number 1 problem for real quadratic fields with fundamental discriminant of the form $n^{2}-4$, a result previously known as Mollin’s conjecture [13, Thm 1.2]; whenever $n>21$ (so $d>22$), $s(d)=h_{\Delta}>1$.
-The list of dimensions where $s(d)=1$ follows from a finite calculation of class numbers.
-∎
-Corollary 6.2.
-Assuming 5.4, there is more than one $\operatorname{PEC}(d)$-class of Weyl–Heisenberg SIC in dimension $d$ for all $d>22$.
-The dimensions $d$ with a unique $\operatorname{PEC}(d)$-class of Weyl–Heisenberg SIC are $d\in\{1,2,4,5,6,10,22\}$.
-Proof.
-The corollary follows from 6.1. ∎
-The second growth result is an asymptotic formula for the size of the ray class monoid of $\mathcal{O}_{\!\Delta_{d}}$, which is conjectured to be the equal to the number of extended Clifford orbits of SICs in dimension $d$.
-The implied constant in this formula is ineffective because it ultimately relies on a Siegel bound involving a possible Siegel zero of a Dirichlet $L$-function.
-We are grateful to Steven Flammia for independently finding numerical evidence for this asymptotic [23].
-Theorem 6.3.
-As $d\to\infty$, and $\Delta_{d}=(d+1)(d-3)$ the size of the class monoid of $\mathcal{O}_{\!\Delta_{d}}$ obeys the asymptotic formula
-$\log\left|\operatorname{Clm}(\mathcal{O}_{\!\Delta_{d}})\right|=\log d+o(\log d).$ (6.2)
-Proof.
-The proof uses the Brauer–Siegel theorem, which is an (ineffective) estimate on the growth of the product of the class number and the regulator.
-The regulator is always negligibly small in our regime.
-The subtlety of the proof is entirely in accounting for the class numbers of non-maximal orders, which are not included in the Brauer–Siegel theorem.
-As in the proof of 6.1, set $s(d):=\left|\operatorname{Clm}(\mathcal{O}_{\Delta_{d}})\right|$. By 5.3,
-$s(d)=\sum_{f^{\prime}|f}h_{(f^{\prime})^{2}\Delta_{0}}$ (6.3)
-where $\Delta=\Delta_{d}=(d+1)(d-3)=f^{2}\Delta_{0}$ and $\Delta_{0}$ is fundamental. Also, $h_{(f^{\prime})^{2}\Delta_{0}}\leq h_{\Delta}$ for all $f^{\prime}\mid f$. It follows that
-$h_{\Delta}\leq s(d)\leq\sigma_{0}(f)h_{\Delta},$ (6.4)
-where $\sigma_{0}(f)$ is the number of positive integer divisors of $f$. Moreover,
-$\log\sigma_{0}(f)\ll\frac{\log f}{\log\log f}\ll\frac{\log d}{\log\log d}$ (6.5)
-by [34, Thm. 2.11]. Thus,
-$\log s(d)=\log h_{\Delta}+O\!\left(\frac{\log d}{\log\log d}\right).$ (6.6)
-We then use the formula for the class number of an order given as [35, Thm. 12.12]:
-$h_{\Delta}=\frac{h_{\!\Delta_{0}}}{[\mathcal{O}_{\!\Delta_{0}}^{\times}:\mathcal{O}_{\!\Delta}^{\times}]}\frac{\left|\left(\mathcal{O}_{\!\Delta_{0}}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\right|}{\left|\left(\mathcal{O}_{\Delta}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\right|}.$ (6.7)
-We will estimate the size of each factor in this formula.
-For any real nonsquare discriminant $\Delta^{\prime}$, let $\varepsilon_{\!\Delta^{\prime}}$ be the smallest unit of $\mathcal{O}_{\!\Delta^{\prime}}$ such that $\varepsilon_{\!\Delta^{\prime}}>1$, and let $\varepsilon_{\!\Delta^{\prime}}^{+}$ be the smallest such totally positive unit.
-The Brauer–Siegel theorem [12]—specifically, Siegel’s part, concerning real quadratic fields [40]— gives the asymptotic formula
-$\frac{\log\!\left(h_{\!\Delta_{0}}\operatorname{Reg}(\mathbb{Q}(\Delta_{0}))\right)}{\log{\Delta_{0}}}\sim 1\mbox{ as }\Delta_{0}\to\infty.$ (6.8)
-Moreover,
-$[\mathcal{O}_{\!\Delta_{0}}^{\times}:\mathcal{O}_{\!\Delta}^{\times}]=\frac{\log\varepsilon_{\!\Delta}}{\log\varepsilon_{\!\Delta_{0}}}=\frac{\log\varepsilon_{\!\Delta}}{\log\operatorname{Reg}(\mathbb{Q}(\Delta_{0}))}.$ (6.9)
-Thus, (6.8) may be rewritten as
-$\log\!\left(\frac{h_{\!\Delta_{0}}}{[\mathcal{O}_{\!\Delta_{0}}^{\times}:\mathcal{O}_{\!\Delta}^{\times}]}\log\varepsilon_{\Delta}\right)\sim\log{\Delta_{0}}\mbox{ as }\Delta_{0}\to\infty.$ (6.10)
-Fix some $\epsilon>0$. Then, there exists some natural number $N$ such that, whenever $\Delta_{0}>N$,
-$(1-\epsilon)\log\Delta_{0}-\log\log\varepsilon_{\Delta}<\log\!\left(\frac{h_{\!\Delta_{0}}}{[\mathcal{O}_{\!\Delta_{0}}^{\times}:\mathcal{O}_{\!\Delta}^{\times}]}\right)<(1+\epsilon)\log\Delta_{0}-\log\log\varepsilon_{\Delta}.$ (6.11)
-For $\Delta_{0}\leq N$, there is some constant $B$ such that
-$B-\log\log\varepsilon_{\Delta}\leq\log\!\left(\frac{h_{\!\Delta_{0}}}{[\mathcal{O}_{\!\Delta_{0}}^{\times}:\mathcal{O}_{\!\Delta}^{\times}]}\right)\leq B-\log\log\varepsilon_{\Delta}.$ (6.12)
-We may bound $\left|\left(\mathcal{O}_{\!\Delta_{0}}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\right|$ above by $f^{2}$ and below by
-$\left|\left(\mathcal{O}_{\!\Delta_{0}}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\right|$ $=f^{2}\prod_{\mathfrak{p}|f}\left(1-\frac{1}{\operatorname{Nm}(\mathfrak{p})}\right)\geq f^{2}\prod_{p|f}\left(1-\frac{1}{p}\right)^{2}=\varphi(f)^{2},$ (6.13)
-where the first product is over prime ideals of $\mathcal{O}_{\!\Delta_{0}}$ and the second is over rational primes.
-Since $\sqrt{\Delta}=f\Delta_{0}\equiv 0\,\left(\operatorname{mod}\,f\mathcal{O}_{\!\Delta_{0}}\right)$, the multiplicative group $\left(\mathcal{O}_{\Delta}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\cong\left(\mathbb{Z}/f\mathbb{Z}\right)^{\times}$, so $\left|\left(\mathcal{O}_{\Delta}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\right|=\varphi(f)$.
-Thus,
-$\varphi(f)\leq\frac{\left|\left(\mathcal{O}_{\!\Delta_{0}}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\right|}{\left|\left(\mathcal{O}_{\Delta}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\right|}\leq\frac{f^{2}}{\varphi(f)}.$ (6.14)
-Using (on both sides) the inequality $\varphi(f)>\frac{f}{\ell(f)}$ where $\ell(f)$ is a real-valued function satisfying $\ell(n)>1$ for all $n$ and $\ell(n)=e^{\gamma}\log\log n$ as $n\to\infty$ [34, Thm.
-2.9], we have
-$\frac{f}{\ell(f)}<\frac{\left|\left(\mathcal{O}_{\!\Delta_{0}}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\right|}{\left|\left(\mathcal{O}_{\Delta}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\right|}\leq f\ell(f),$ (6.15)
-which after taking logarithms becomes
-$\log f-\log\ell(f)<\log\!\left(\frac{\left|\left(\mathcal{O}_{\!\Delta_{0}}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\right|}{\left|\left(\mathcal{O}_{\Delta}/f\mathcal{O}_{\!\Delta_{0}}\right)^{\times}\right|}\right)<\log f+\log\ell(f).$ (6.16)
-Using (6.7) and combining (6.11) and (6.16), for $\Delta_{0}>N$, we obtain
-$\left|\log h_{\Delta}-\left(\log\sqrt{\Delta}_{0}+\log f\right)\right|<\epsilon\log\sqrt{\Delta}_{0}+\left|\log\log\varepsilon_{\Delta}\right|+\log\ell(f).$ (6.17)
-On the left-hand side, $\log\sqrt{\Delta}_{0}+\log f=\log\sqrt{\Delta}\sim\log d$ as $d\to\infty$. On the right-hand side, as $d\to\infty$, $\log\sqrt{\Delta_{0}}<\log\sqrt{\Delta}\ll\log d$, while
-$\log\log\varepsilon_{\Delta}\leq\log\log\varepsilon_{\Delta}^{+}=\log\log\frac{d-1+\sqrt{\Delta}}{2}\ll\log\log d,$ (6.18)
-and $\log\ell(f)\ll\log\log\log d$. Thus,
-$\left|\log h_{\Delta}-\log d\right|\ll\epsilon\log d\mbox{ as }d\to\infty.$ (6.19)
-In the case when $\Delta_{0}\leq N$, combining (6.11) and (6.17), we instead obtain
-$\left|\log h_{\Delta}-\left(\log\sqrt{\Delta}_{0}+\log f\right)\right|<B+\left|\log\log\varepsilon_{\Delta}\right|+\log\ell(f),$ (6.20)
-which by a similar argument also implies
-$\left|\log h_{\Delta}-\log d\right|\ll\epsilon\log d\mbox{ as }d\to\infty.$ (6.21)
-Combining (6.19) and (6.21) with (6.6) proves the theorem in all cases. ∎
-Corollary 6.4.
-Assuming 5.4, as $d\to\infty$,
-$\log\left|{\rm WHSIC}(d)/\operatorname{PEC}(d)\right|=\log d+o(\log d).$ (6.22)
-Proof.
-Follows from 6.3. ∎
-7. Conjectures on counting Galois multiplets of Weyl–Heisenberg SICs
-Recall that 4.10 implies that there are a finite number of well-defined Galois multiplets of SICs in dimension $d$, each consisting of a finite number of geometric equivalence classes; we will now refine this prediction to a precise count in every dimension.
-Our Multiplet Count Conjecture (7.1) predicts the number of multiplets in dimension $d$.
-The Order-to-Multiplet Conjecture (7.2) predicts the number of geometric equivalence classes in each multiplet.
+$\Delta_{0}$, $f^{\prime}$ runs over divisors of $f$, $h_{(f^{\prime})^{2}\Delta_{0}}=\left|\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})\right|$, and $\sigma_{0}(f)=\left|\{\mathcal{O}:\mathcal{O}_{\Delta}\subseteq\mathcal{O}\subseteq{\mathcal{O}_{\Delta_{0}}}\}\right|$.
+$d$ $\Delta$ $\Delta_{0}$ $f^{\prime}$$\widetilde{\#}$ $h_{(f^{\prime})^{2}\Delta_{0}}$ $\left|\operatorname{Clm}(\mathcal{O}_{\Delta})\right|$
+$51$ $2496$ $561$,2,41241,2,41 , 2 , 4 $3$ $2,4,8$ $14$
+$52$ $2597$ $53$ $1,7$ $2$ $1,3$ $4$
+$53$ $2700$ $21$,3,$5,1$ 3 , 5 , 15 $4$ $1,1,2,6$ $10$
+$54$ $2805$ $2805$ $1$ $1$ $4$ $4$
+$55$ $2912$ $728$ $1,2$ $2$ $2,4$ $6$
+$56$ $3021$ $3021$ $1$ $1$ $6$ $6$
+$57$ $3132$ $348$ $1,3$ $1$ $2,6$ $8$
+$58$ $3245$ $3245$ $1$ $1$ $4$ $4$
+$59$ $3360$ $840$ $1,2$ $2$ $4,8$ $12$
+$60$ $3477$ $3477$ $1$ $1$ $4$ $4$
+$61$ $3596$ $3596$ $1$ $1$ $6$ $6$
+$62$ $3717$ $413$ $1,3$ $2$ $1,4$ $5$
+$63$ $3840$ $60$ $1,2,4,8$ $4$ $2,2,4,8$ $16$
+$64$ $3965$ $3965$ $1$ $1$ $4$ $4$
+$65$ $4092$ $4092$ $1$ $1$ $8$ $8$
+$66$ $4221$ $469$ $1,3$ $2$ $3,6$ $9$
+$67$ $4352$ $71$,2,4,8,161248161,2,4,8,161 , 2 , 4 , 8 , 16 $5$ $1,1,1,2,4$ $9$
+$68$ $4485$ $4485$ $1$ $1$ $4$ $4$
+$69$ $4620$ $4620$ $1$ $1$ $8$ $8$
+$70$ $4757$ $4757$ $1$ $1$ $5$ $5$
+ $d$ $\Delta$ $\Delta_{0}$ $f^{\prime}$$\widetilde{\#}$ $h_{(f^{\prime})^{2}\Delta_{0}}$ $\left|\operatorname{Clm}(\mathcal{O}_{\Delta})\right|$ $71$ $4896$ $361$,2,3,612361,2,3,61 , 2 , 3 , 6 $4$ $2,4,4,8$ $18$ $72$ $5037$ $5037$ $1$ $1$ $4$ $4$ $73$ $5180$ $5180$ $1$ $1$ $4$ $4$ $74$ $5325$ $213$ $1,5$ $2$ $1,6$ $7$ $75$ $5472$ $521$,2,3,612361,2,3,61 , 2 , 3 , 6 $4$ $1,2,4,8$ $15$ $76$ $5621$ $5621$ $1$ $1$ $6$ $6$ $77$ $5772$ $5772$ $1$ $1$ $8$ $8$ $78$ $5925$ $237$ $1,5$ $2$ $1,6$ $7$ $79$ $6080$ $380$ $1,2,4$ $3$ $2,4,8$ $14$ $80$ $6237$ $77$ $1,3,9$ $3$ $1,2,6$ $9$ $81$ $6396$ $6396$ $1$ $1$ $12$ $12$ $82$ $6557$ $6557$ $1$ $1$ $3$ $3$ $83$ $6721$ $051$,2,4,812481,2,4,81 , 2 , 4 , 8 $4$ $2,2,4,8$ $16$ $84$ $6885$ $85$ $1,3,9$ $3$ $2,2,6$ $10$ $85$ $7052$ $7052$ $1$ $1$ $4$ $4$ $86$ $7221$ $7221$ $1$ $1$ $10$ $10$ $87$ $7392$ $8481$,2121,21 , 2 $2$ $4,8$ $12$ $88$ $7565$ $7565$ $1$ $1$ $4$ $4$ $89$ $7740$ $860$ $1,3$ $2$ $2,8$ $10$ $90$ $7917$ $7917$ $1$ $1$ $4$ $4$
+Table 5.2.
+Ring class numbers and ring class monoid numbers for $51\leq d\leq 90$.
+Here $\Delta=(d+1)(d-3)=f^{2}\Delta_{0}$ for a fund.
+disc.
+$\Delta_{0}$, $f^{\prime}$ runs over divisors of $f$, $h_{(f^{\prime})^{2}\Delta_{0}}=\left|\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})\right|$, and $\widetilde{\#}$ is |$\left|\{\mathcal{O}:\mathcal{O}_{\Delta}\subseteq\mathcal{O}\subseteq{\mathcal{O}_{\Delta_{0}}}\}\right|=\sigma_{0}(f)$.
+(The values $\left|\operatorname{Clm}(\mathcal{O}_{\Delta})\right|$ agree with empirical data for $\left|\operatorname{WHSIC}(d)/\operatorname{PEC}(d)\right|$ of Scott [40]).
 5.4. Statement of the Multiplet Count Conjecture
+We showed in 4.10 that 1.1 implies that there are a finite number of well-defined Galois multiplets of WHSICs in dimensions $d\geq 4$, each consisting of a finite number of geometric equivalence classes; we now refine this prediction to a precise count in every dimension.
 […]
@@
 Let $\Delta=\Delta_{d}=(d+1)(d-3)$ and $K=\mathbb{Q}(\sqrt{\Delta})$.
-The number of Galois multiplets of Weyl-Heisenberg SICs in dimension $d$ equals the number of quadratic orders $\mathcal{O}^{\prime}$ with $\mathcal{O}_{\!\Delta}\subseteq\mathcal{O}^{\prime}\subseteq\mathcal{O}_{K}$.
+The number of Galois multiplets of Weyl–Heisenberg SICs in dimension $d$ equals the number of quadratic orders $\mathcal{O}^{\prime}$ with $\mathcal{O}_{\!\Delta}\subseteq\mathcal{O}^{\prime}\subseteq\mathcal{O}_{K}$.
 That is,
@@
 where $\sigma_{0}(f)$ denotes the number of positive divisors of $f$.)
-The content of 7.1 is that the multiplets of Weyl–Heisenberg SICs are in one-to-one correspondence with orders between $\mathcal{O}$ and $\mathcal{O}_{K}$.
-The fact that the number of such orders is simply $\sigma_{0}(f)$ uses the fact that an order in a quadratic field is determined uniquely by its discriminant, which can be $(f^{\prime})^{2}\Delta_{0}$ for a fundamental discriminant $\Delta_{0}$ and a positive integer $f^{\prime}$.
+The content of 5.8 is that the multiplets of Weyl–Heisenberg SICs are in one-to-one correspondence with orders between $\mathcal{O}_{\Delta}$ and $\mathcal{O}_{K}$.
+An order in a quadratic field is determined uniquely by its discriminant, which can be $(f^{\prime})^{2}\Delta_{0}$ for a fundamental discriminant $\Delta_{0}$ and a positive integer $f^{\prime}$.
+The order $\mathcal{O}_{\Delta}=\mathcal{O}_{f^{2}\Delta_{0}}$ is contained in $\mathcal{O}^{\prime}=\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}}$ if and only if $f^{\prime}\mid f$, so there are $\sigma_{0}(f)$ of them.
 The order $\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}}$ contains $\mathcal{O}_{f^{2}\Delta_{0}}=\mathcal{O}_{\!\Delta}$ if and only if $f^{\prime}\mid f$.
-(These facts are not true for general number fields.)
-7.2. Statement of the Extended Order-to-Multiplet Conjecture
-We next formulate a conjecture positing a correspondence between orders and multiplets with certain properties.
-The number of geometric equivalence classes $[S]$ inside a Galois multiplet $[[S]]$ should be equal to the class number of the corresponding orders, and the fields $F^{\rm trip}\!\left(S\right)$ and $F^{\rm proj}\!\left(S\right)$ associated to different multiplets are related in a natural way.
-(Recall that we have already shown these fields to be are well-defined on $\operatorname{PEC}(d)$-equivalence classes, and they are well-defined on Galois super-equivalence classes under the assumption of 4.10.)
-Conjecture 7.2 (Extended Order-to-Multiplet Conjecture).
-Fix a positive integer $d\neq 3$, let $\Delta=\Delta_{d}=(d+1)(d-3)$, and write $\Delta=f^{2}\Delta_{0}$ for a fundamental discrimiant $\Delta_{0}$ and some positive integer $f$.
-Then there is a bijection $\mathcal{M}$,
-{positive divisors of $\{\mbox{positive divisors of $f$}\}\xrightarrow{\mathcal{M}}\{[[S]]:S\mbox{ a Weyl--Heisenberg line-SIC in }\mathbb{C}^{d}\}$ (7.3)
-having the following properties:
-  • (1)
-    The number of geometric equivalence classes $[S]$ in the multiplet $\mathcal{M}(f^{\prime})=[[S]]$ is $\left|\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})\right|$;
-  • (2)
-    If $f_{1}\mid f_{2}\mid f$, $[S_{1}]\in\mathcal{M}(f_{1})$, and $[S_{2}]\in\mathcal{M}(f_{2})$, then $F^{\rm trip}\!\left(S_{1}\right)\subseteq F^{\rm trip}\!\left(S_{2}\right)$ and $F^{\rm vec}\!\left(S_{1}\right)=F^{\rm proj}\!\left(S_{1}\right)\subseteq F^{\rm vec}\!\left(S_{2}\right)=F^{\rm proj}\!\left(S_{2}\right)$.
-Part (1) of this conjecture is consistent with 5.4.
-The strengthening of this conjecture compared with 1.2 is that part (2) asks for additional compatibility under inclusion of the number fields $F^{\rm trip}\!\left(S\right)$ and $F^{\rm proj}\!\left(S\right)$ with divisibility of the associated conductors $f^{\prime}$ (i.e., reverse inclusion of the associated quadratic orders).
-Assuming 1.1, we have $F^{\rm vec}\!\left(S\right)=F^{\rm proj}\!\left(S\right)$.
-This conjecture can be tested by study of numerical invariants. One may attempt to reverse-engineer the map $\mathcal{M}$ from numerical data.
-  1. (1)
-    One first verifies that the number of multiplets matches the number of orders (7.1).
-  2. (2)
-    One can compute all the class numbers of the orders $\mathcal{O}$, and one can numerically compute the number of $\operatorname{PEC}$-equivalence classes in each multiplet.
-These multisets of numbers must be identical for a map to exist.
-  3. (3)
-    If all the class numbers of the orders $\mathcal{O}^{\prime}$ are pairwise distinct, then the putative map $\mathcal{M}$ is uniquely specified.
-Otherwise, the map exists if and only if the number of $\operatorname{PEC}(d)$-orbits in the different Galois equivalence classes are a permutation of these class numbers.
-  4. (4)
-    The reverse-inclusion requirement of the triple product and projection SIC-fields as the orders vary provides a further test of the conjecture.
-The inclusion relations between the different fields gives further constraints on existence of such a map.
-One can study weaker invariants, the degrees of the fields, which are easier to compute.
-A necessary condition for inclusion of fields is a divisibility relation between their field degrees over $\mathbb{Q}$.
-Ambiguity may remain. A case where there would be unresolvable ambiguity using numerical invariants is any case where there are two distinct multiplet $[[S_{1}]]$ and $[[S_{2}]]$ having
-  1. (1)
-    identical sizes $\left|[[S_{1}]]\right|=\left|[[S_{2}]]\right|$,
-  2. (2)
-    identical triple product fields $F^{\rm trip}\!\left(S_{1}\right)=F^{\rm trip}\!\left(S_{2}\right)$,
-  3. (3)
-    identical projection fields $F^{\rm proj}\!\left(S_{1}\right)=F^{\rm proj}\!\left(S_{2}\right)$.
-Such an ambiguity can occur, as discovered by Markus Grassl for $d=47$ and $d=67$; see Section 9.3.
-In Section 8, we present a refined 8.4 about the fields $F^{\rm trip}\!\left(S\right)$ and $F^{\rm proj}\!\left(S\right)$, identifying them with certain ray class fields of the orders.
+(These facts do not hold for general number fields.)
 5.5. Verification of cases of the Multiplet Count Conjecture
 We have verified that the two sets in the Multiplet Count Conjecture (5.8) have the same size for dimensions $d\leq 90$ ($d\neq 3$) under the assumption that the published lists of Weyl–Heisenberg SICs in these dimensions are complete.
-Scott [38] gives a putatively complete list of numerical SICs in all dimensions $d\leq 90$, building on earlier work of Scott and Grassl [39].
-Scott and Grassl find SIC fiducials by numerically minimizing a frame potential—a function whose global minima are SICs if SICs exist.
-Their search is constrained to Weyl–Heisenberg covariant sets of vectors.
-By repeating their search with a large number of seed vectors, they establish a high degree of numerical confidence in their results but do not rigorously rule out the existence of other SICs (Weyl–Heisenberg covaraint or otherwise).
-7.4. Evidence for the Extended Order-to-Multiplet Conjecture
-Some evidence for the Extended Order-to-Multiplet Conjecture (7.2) is given in Table 3 at the end of the paper, in dimensions $d\leq 15$.
-In the dimensions where there is more than one multiplet ($d\in\{7,8,11,12,15\}$), a conjectural correspondence between divisors $f^{\prime}\mid f$ and Galois multiplets (written using Scott and Grassl’s orbit labels) has been given.
-Under this correspondence, note that for divisors $f_{1},f_{2}$ of $f$ such that $f_{1}\mid f_{2}$, the corresponding degrees $\left.[F^{\rm proj}\!\left(S_{1}\right):K]\middle|[F^{\rm proj}\!\left(S_{2}\right):K]\right.$, and similarly for the triple product fields in the cases that have been computed.
-The table also supports the stronger 8.5, which implies 7.2.
+Proposition 5.9 (Empirical Proposition 2).
+The total number of constructed Galois multiplets $[[S]]$ of exact SICs for dimensions
+$d\in\{4,5,6,7,8,9,10,11,12,23,14,15,16,17,18,19,20,21,24,28,30,35,39,48\}$ (5.12)
+equals the total number of orders
+$s(\mathcal{O}_{\Delta_{d}}):=\left|\{\mathcal{O}^{\prime}:\mathcal{O}_{\Delta}\subseteq\mathcal{O}^{\prime}\subseteq\mathcal{O}_{\Delta_{0}}\}\right|,$ (5.13)
+where $\Delta=(d+1)(d-3)$.
+Equivalently, writing $\Delta=f^{2}\Delta_{0}$, where $\Delta_{0}$ is a fundamental discriminant, this number $s(\mathcal{O}_{\Delta_{d}})=\sigma_{0}(f)$, the number of positive divisors of $f$.
+(The function $\sigma_{0}(f)$ is also denoted $\tau(f)$ or $d(f)$ by some authors.)
+For the dimensions in (5.12), there is at least one bijective map
+:$\mathcal{M}:\{\mathcal{O}^{\prime}:\mathcal{O}_{\Delta}\subseteq\mathcal{O}^{\prime}\subseteq\mathcal{O}_{\Delta_{0}}\}\to\{[[S]]:S\,\mbox{a Weyl--Heisenberg line-SIC in}\,\mathbb{C}^{d}\}$ (5.14)
+such that the number of geometric equivalence classes $[S]\in\mathcal{M}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})$ is $\left|\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})\right|$.
+Proof.
+The result follows from comparison of [6, Tab.
+1 and Tab.
+2] on known exact SICs with the number theoretic data on $h_{(f^{\prime})^{2}\Delta_{0}}$ from our Table 5.1.
+Here [6, Tab.
+1] gives results of Scott and Grassl [41] for exact multiplets for $4\leq d\leq 16$, excluding $d=15$; [6, Tab.
+2] gives the remaining exact multiplets in (5.12).
+Each Galois multiplet contains a number of $\operatorname{PEC}(d)$-orbits, indicated by letters; for example, the multiplet $35bcdg$ consists of four $\operatorname{PEC}(35)$-orbits.
+In the dimensions in (5.12), the number of Galois multiplets is found to agree with the data for $s(\mathcal{O}_{\Delta_{d}})=\sigma_{0}(f)$, given in column $\sigma_{0}(f)$ of Table 5.1.
+The existence of a bijective map $\mathcal{M}$ is equivalent to the assertion that the unordered multiset of candinalities |$\left|\left\{[S]\in[[S]]\right\}\right|$ of $\operatorname{PEC}(d)$-orbits of SICs in each Galois multiplet is equal to the unordered multiset of ring class numbers $h_{(f^{\prime})^{2}\Delta_{0}}$ such that $f^{\prime}\mid f$.
+The multiset equality was verified comparing the exact multiplet data of [6] with that in column labeled $h_{(f^{\prime})^{2}\Delta_{0}}$ in Table 5.1.
+∎
+The bijective map $\mathcal{M}$ is not uniquely specified using these class number counts alone.
+However, when augmented by requiring agreement of all inclusion relations between the associated SIC fields $F^{\rm proj}\!\left(S\right)$, the truth of 1.3 in dimension $d$ uniquely determines the map $\mathcal{M}$ determined for the dimensions (5.12).
+(See Section 8.1; the smallest dimension where $\mathcal{M}$ is not determined by the inclusion relations is $d=47$.)
+Example 5.10.
+Let $d=35$. Here $\Delta=36\cdot 32=2^{7}\cdot 3^{2}$. Here $\Delta_{0}=8$ and $f=12$. The set of divisors $f^{\prime}$ of $12$ are $\{1,2,3,4,6,12\}$.
+We exhibit a bijection $\mathcal{M}$ of 1.2 from orders to multiplets in Figure 5.1.
+This bijection satisfies part (1) of 1.2 for $d=35$.
+This choice of $\mathcal{M}$ also satisfies part (2) of 1.2 for $d=35$, as shown in 6.7.
+${\mathcal{O}_{1152}}\atop{h=4}$${\mathcal{O}_{288}}\atop{h=2}$${\mathcal{O}_{72}}\atop{h=1}$${\mathcal{O}_{128}}\atop{h=1}$${\mathcal{O}_{32}}\atop{h=1}$${\mathcal{O}_{8}}\atop{h=1}$4499449
+[…]
+Figure 5.1.
+Left side: For $d=35$, over-orders of $\mathcal{O}_{1152}$, and their ring class numbers $h=h_{\Delta_{0}(f^{\prime})^{2}}=\left|\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})\right|$, with $\Delta_{0}=8$.
+The direction on edges indicate (reverse) set inclusions, with edge labels giving relative indices of orders.
+Thus $\mathcal{O}_{1152}\subset\mathcal{O}_{8}$ has index $144$ in $\mathcal{O}_{8}$.
+Right side: For $d=35$, $\mathcal{M}$-correspondence to Galois multiplets of $d=35$ SICs, with multiplet letter labels ($35a$ to $35j$) corresponding to $\operatorname{PEC}(35)$-equivalence classes of SICs.
+Class numbers match the number of orbit letter labels in each multiplet.
 6. Conjectures on precise fields of definition of Weyl–Heisenberg SICs
-In this section, we give a refinement of both 5.4 and of the “class field hypothesis” of Appleby, Flammia, McConnell, and Yard [10].
-The refined conjecture predicts the triple product field and the projector field associated to each multiplet.
+In this section, we study the number-theoretic properties of the ray class fields of orders $\mathcal{O}^{\prime}$ predicted to be fields of definition of Weyl–Heisenberg SICs by 1.3.
+We give formulas for the degrees of those ray class field over $K$, by means of studying the associated ray class groups.
+In 6.6, we test numerical predictions of relative degrees and field inclusions implied by 1.2 and 1.3.
+We also study the relative degrees of field inclusions predicted in 1.2(2), noting there exist cases where the map $\mathcal{M}$ is not uniquely determined by the conclusions of 1.2 and 1.3.
+Before studying the special ray class fields conjecturally associated to SICs, we summarize some properties of ray class fields of orders in general.
 6.1. Ray class fields of orders
@@
 In particular, the ring class field of an order $\mathcal{O}$ has Galois group over $K$ isomorphic to $\operatorname{Cl}(\mathcal{O})$ and generalizes the Hilbert class field, which is the maximal unramified extension of $K$ and has class group $\operatorname{Cl}(\mathcal{O}_{K})$.
-In [33], the authors modify the standard constructions of class field theory to describe a distinguished collection of ray class fields of orders $H^{\mathcal{O}}_{\mathfrak{m},\Sigma}$, each specified by a level datum $(\mathcal{O};\mathfrak{m},\Sigma)$, where $\mathcal{O}$ is a (possibly non-maximal) order of $K$, $\mathfrak{m}$ is an integral $\mathcal{O}$-ideal, and $\Sigma$ is any subset of real places of $K$.
-In the case that $K$ is a real quadratic field, there are two infinite places, so $\Sigma$ is one of $\emptyset$, $\{\infty_{1}\}$, $\{\infty_{2}\}$, or $\{\infty_{1},\infty_{2}\}$.
+In [34], the authors modify the standard constructions of class field theory to describe a distinguished collection of ray class fields of orders $H^{\mathcal{O}}_{\mathfrak{m},\Sigma}$, each specified by a level datum $(\mathcal{O};\mathfrak{m},\Sigma)$, where $\mathcal{O}$ is a (possibly non-maximal) order of $K$, $\mathfrak{m}$ an integral $\mathcal{O}$-ideal, and $\Sigma$ a subset of real places of $K$.
+In the case that $K$ is real quadratic, there are two infinite places, so $\Sigma$ is one of $\emptyset$, $\{\infty_{1}\}$, $\{\infty_{2}\}$, or $\{\infty_{1},\infty_{2}\}$.
 For the datum $(\mathcal{O};\mathcal{O},\emptyset)$, the associated field $H^{\mathcal{O}}_{\mathcal{O},\emptyset}$ is the ring class field associated to the order.
 For the datum $(\mathcal{O}_{K};\mathfrak{m},\Sigma)$, the associated field $H^{\mathcal{O}_{K}}_{\mathfrak{m},\Sigma}$ is the standard (Takagi) ray class field $H_{\mathfrak{m},\Sigma}$.
-For any fixed order $\mathcal{O}$, any abelian extension of $K$ as a subfield of some $H^{\mathcal{O}}_{\mathfrak{m},\Sigma}$.
-We recall [33, Thm. 1.1]:
+For a fixed order $\mathcal{O}$, any abelian extension of $K$ as a subfield of some $H^{\mathcal{O}}_{\mathfrak{m},\Sigma}$.
+We recall [34, Thm. 1.1]:
 Theorem 6.1.
@@
 There exists a unique abelian Galois extension $H_{\mathfrak{m},\Sigma}^{\mathcal{O}}/K$ with the property that a prime ideal $\mathfrak{p}$ of $\mathcal{O}_{K}$ that is coprime to the quotient ideal $\left(\mathfrak{m}:\mathcal{O}_{K}\right)$ splits completely in $H_{\mathfrak{m},\Sigma}^{\mathcal{O}}/K$ if and only if $\mathfrak{p}\cap\mathcal{O}=\pi\mathcal{O}$, a principal prime $\mathcal{O}$-ideal having $\pi\in\mathcal{O}$ with $\pi\equiv 1\,\left(\operatorname{mod}\,\mathfrak{m}\right)$ and $\rho(\pi)>0$ for $\rho\in\Sigma$.
-To understand 8.1, recall that for any integral ideal $\mathfrak{m}$ of $\mathcal{O}$, the quotient ideal $\left(\mathfrak{m}:\mathcal{O}_{K}\right)$ is the largest ideal of $\mathcal{O}_{K}$ contained in $\mathfrak{m}$.
+To understand 6.1, recall that for any integral ideal $\mathfrak{m}$ of $\mathcal{O}$, the quotient ideal $\left(\mathfrak{m}:\mathcal{O}_{K}\right)$ is the largest ideal of $\mathcal{O}_{K}$ contained in $\mathfrak{m}$.
 An important invariant of an order $\mathcal{O}$ of an algebraic number field $K$ is its (absolute) conductor $\mathfrak{f}=\mathfrak{f}(\mathcal{O})=\left(\mathcal{O}:\mathcal{O}_{K}\right)$, which is the (set-theoretically) largest integral ideal $\mathfrak{f}\subseteq\mathcal{O}$ of $\mathcal{O}$ that is also an ideal of the maximal order $\mathcal{O}_{K}$.
@@
 Moreover, if $(\mathcal{O};\mathfrak{m},\Sigma)$ and $(\mathcal{O}^{\prime};\mathfrak{m}^{\prime},\Sigma^{\prime})$ are level data satisfying $\mathcal{O}\subseteq\mathcal{O}^{\prime}$, $\mathfrak{m}\mathcal{O}^{\prime}\subseteq\mathfrak{m}^{\prime}$, and $\Sigma\supseteq\Sigma^{\prime}$, then there is a natural quotient map $\operatorname{Cl}_{\mathfrak{m},\Sigma}(\mathcal{O})\twoheadrightarrow\operatorname{Cl}_{\mathfrak{m}^{\prime},\Sigma^{\prime}}(\mathcal{O}^{\prime})$ and a corresponding natural field inclusion $H_{\mathfrak{m}^{\prime},\Sigma^{\prime}}^{\mathcal{O}^{\prime}}\subseteq H_{\mathfrak{m},\Sigma}^{\mathcal{O}}$.
-We recall [33, Thm. 1.3]:
+We recall [34, Thm. 1.3]:
 Theorem 6.2.
-For an order $\mathcal{O}$ in a number field $K$ and any level datum $(\mathcal{O};\mathfrak{m},\Sigma)$ with associated ray class field $H_{0}:=H_{\mathfrak{m},\Sigma}^{\mathcal{O}}$, there is an isomorphism $\operatorname{Art}_{\mathcal{O}}:\operatorname{Cl}_{\mathfrak{m},\Sigma}(\mathcal{O})\to\operatorname{Gal}(H_{0}/K)$, uniquely determined by its behavior on prime ideals $\mathfrak{p}$ of $\mathcal{O}$ that are coprime to $\mathfrak{f}(\mathcal{O})\cap\mathfrak{m}$, having the property that
+For an order $\mathcal{O}$ in a number field $K$ and any level datum $(\mathcal{O};\mathfrak{m},\Sigma)$ with associated ray class field $H_{0}:=H_{\mathfrak{m},\Sigma}^{\mathcal{O}}$, there is an isomorphism $\operatorname{Art}_{\mathcal{O}}:\operatorname{Cl}_{\mathfrak{m},\Sigma}(\mathcal{O})\to\operatorname{Gal}(H_{0}/K)$, uniquely determined by its behavior on prime $\mathcal{O}$-ideals $\mathfrak{p}$ coprime to $\mathfrak{f}(\mathcal{O})\cap\mathfrak{m}$, having the property that
 $\operatorname{Art}_{\mathcal{O}}([\mathfrak{p}])(\alpha)\equiv\alpha^{q}\,\left(\operatorname{mod}\,\mathfrak{P}\right),$ (6.2)
@@
 where $\operatorname{Art}:\operatorname{Cl}_{\left(\mathfrak{m}:\mathcal{O}_{K}\right),\Sigma}(\mathcal{O}_{K})\to\operatorname{Gal}(H_{1}/K)$ is the usual Artin map in class field theory, with $H_{1}=H_{\left(\mathfrak{m}:\mathcal{O}_{K}\right),\Sigma}^{\mathcal{O}_{K}}$ being a (Takagi) ray class field for the maximal order $\mathcal{O}_{K}$, and $H_{0}\subseteq H_{1}$.
-In 8.2, the set of prime ideals coprime to $\mathfrak{f}(\mathcal{O})\cap\mathfrak{m}$ includes all but finitely many of the prime ideals of $\mathcal{O}$.
-The Artin map is fully defined by (8.2), which specifies the map on prime ideals of $\mathcal{O}$ coprime to $\mathfrak{f}(\mathcal{O})\cap\mathfrak{m}$, because it turns out that any ray ideal class in $\operatorname{Cl}_{\mathfrak{m},\Sigma}(\mathcal{O})$ has a representative that is coprime to $\mathfrak{f}(\mathcal{O})\cap\mathfrak{m}$ (see [33, Lem.
+In 6.2, the set of prime ideals coprime to $\mathfrak{f}(\mathcal{O})\cap\mathfrak{m}$ includes all but finitely many of the prime ideals of $\mathcal{O}$.
+The Artin map is fully defined by (6.2), which specifies the map on prime ideals of $\mathcal{O}$ coprime to $\mathfrak{f}(\mathcal{O})\cap\mathfrak{m}$, because it turns out that any ray ideal class in $\operatorname{Cl}_{\mathfrak{m},\Sigma}(\mathcal{O})$ has a representative that is coprime to $\mathfrak{f}(\mathcal{O})\cap\mathfrak{m}$ (by [34, Lem.
 4.12]), and ideals of $\mathcal{O}$ that are coprime to $\mathfrak{f}$ enjoy unique factorization into prime ideals (even though arbitrary nonzero ideals of $\mathcal{O}$ do not).
 When $\mathcal{O}=\mathcal{O}_{K}$, (6.2) specializes to the standard definition of the Artin map in the case of a Takagi ray class field.
-For a real quadratic field $K$, the infinite places in $\Sigma$ relate to the notions of “wide” and “narrow” class fields.
-If $\mathfrak{m}=\mathcal{O}$, the (wide) ring class field is $H_{\mathcal{O},\emptyset}^{\mathcal{O}}=H_{\mathcal{O},\{\infty_{1}\}}^{\mathcal{O}}=H_{\mathcal{O},\{\infty_{2}\}}^{\mathcal{O}}$, and the narrow ring class field is $H_{\mathcal{O},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}}$.
-The narrow ring class field is either equal to the wide ring class field or is a degree $2$ extension thereof.
-More generally, there are four ray class fields associated to the ideal $\mathfrak{m}$ as well as a fifth subfield $H_{\mathfrak{m},+}^{\mathcal{O}}$ of $H_{\mathfrak{m},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}}$.
-[…]
-Each of the field inclusions in this diagram is either trivial or degree $2$.
-In the “nondegenerate” case when all have degree $2$, the Galois group $\operatorname{Gal}\!\left(H_{\mathfrak{m},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}}/H_{\mathfrak{m},\emptyset}^{\mathcal{O}}\right)\cong\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2\mathbb{Z}$.
-The fields $H_{\mathfrak{m},\emptyset}^{\mathcal{O}}$, $H_{\mathfrak{m},+}^{\mathcal{O}}$, and $H_{\mathfrak{m},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}}$ are all Galois over $\mathbb{Q}$, whereas in the nondegenerate case $H_{\mathfrak{m},\{\infty_{1}\}}^{\mathcal{O}}$ and $H_{\mathfrak{m},\{\infty_{2}\}}^{\mathcal{O}}$ are not.
-The field $H_{\mathfrak{m},\emptyset}^{\mathcal{O}}$ is totally real, $H_{\mathfrak{m},+}^{\mathcal{O}}$ is (in the nondegenerate case) totally complex and CM, and $H_{\mathfrak{m},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}}$ is (in the nondegenerate case) totally complex and not CM.
-8.2. Statement of the Extended Ray Class Fields of Orders Conjecture
-We conjecture an identification of fields of triple product SIC-fields and projection SIC-fields with certain class fields of orders.
+6.2. Special ray class fields of real quadratic orders
 We recall the fact that orders of a quadratic field $K$ are completely specified by their discriminants $\Delta^{\prime}=(f^{\prime})^{2}\Delta_{0}$, where $\Delta_{0}$ is the field discriminant of $K$ (a fundamental discriminant; see the beginning of Section 5.4) and $f^{\prime}$ is a positive integer.
 Definition 6.3.
-Fix a positive integer $d\geq 4$, write $\Delta=(d+1)(d-3)=f^{2}\Delta_{0}$ for a fundamental discriminant $\Delta_{0}$, and let $f^{\prime}$ be a positive divisor of $f$.
+Fix a positive integer $d\geq 4$, write $\Delta=(d+1)(d-3)=f^{2}\Delta_{0}$ with $\Delta_{0}$ a fundamental discriminant, and let $f^{\prime}$ be a positive divisor of $f$.
 Let $\Delta^{\prime}=(f^{\prime})^{2}\Delta_{0}$ and $\mathcal{O}^{\prime}=\mathcal{O}_{\Delta^{\prime}}$.
-We introduce two families of class fields parametrized by $d$ and $f^{\prime}$:
-$E_{d,f^{\prime}}$ $=H_{d\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}^{\prime}};$ (8.5)
-$\tilde{E}_{d,f^{\prime}}$ $=H_{d^{\prime}\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}^{\prime}}.$ (8.6)
+We introduce two families of ray class fields of the order $\mathcal{O}^{\prime}$ parametrized by $d$ and $f^{\prime}$:
+$E_{d,f^{\prime}}$ $=H_{d\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}^{\prime}}\quad\mbox{and}\quad\tilde{E}_{d,f^{\prime}}:=H_{d^{\prime}\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}^{\prime}}.$ (6.4)
 Here $d^{\prime}=d$ if $d$ is odd and $d^{\prime}=2d$ if $d$ is even.
-Note that $E_{d,f^{\prime}}=\tilde{E}_{d,f^{\prime}}$ if $d$ is odd, and $E_{d,f^{\prime}}\subseteq\tilde{E}_{d,f^{\prime}}$ if $d$ is even.
-Specifically, when $d$ is even, the index $[\tilde{E}_{d,f^{\prime}}:E_{d,f^{\prime}}]=\left|\operatorname{Cl}_{2d\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})/\operatorname{Cl}_{d\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})\right|=2$ by 10.2.
-Conjecture 8.4 (Extended Ray Class Fields of Orders Conjecture).
-Fix a positive integer $d\geq 4$, write $\Delta=(d+1)(d-3)=f^{2}\Delta_{0}$ for a fundamental discriminant $\Delta_{0}$, and let $f^{\prime}\mid f$, $f^{\prime}>0$.
-Let :{positive divisors of $\mathcal{M}:\{\mbox{positive divisors of $f$}\}\to\{[[S]]:S\mbox{ a Weyl--Heisenberg line-SIC in }\mathbb{C}^{d}\}$ be any function satisfying the conditions of 7.2.
-If $S$ is a Weyl–Heisenberg SIC with $[S]\in\mathcal{M}(f^{\prime})$, then
-$F^{\rm trip}\!\left(S\right)$ $=E_{d,f^{\prime}}$ and $F^{\rm vec}\!\left(S\right)=F^{\rm proj}\!\left(S\right)$ $=\tilde{E}_{d,f^{\prime}}.$ (8.7)
-8.4 implies that $F^{\rm proj}\!\left(S\right)=F^{\rm eproj}\!\left(S\right)$ for any Weyl–Heisenberg SIC $S$.
-This is because $\tilde{E}_{d,f^{\prime}}=H^{\mathcal{O}^{\prime}}_{d^{\prime}\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\}}$ contains the ray class field of $\mathbb{Q}$ with level datum $(\mathbb{Z};d^{\prime}\mathbb{Z},\{\infty\})$, which is simply the field $\mathbb{Q}(\zeta_{d^{\prime}})=\mathbb{Q}(\xi_{d})$ of $d^{\prime}$-th roots of unity; thus, $E_{d,f^{\prime}}(\xi_{d})=\tilde{E}_{d,f^{\prime}}$.
-8.4 is a natural extension of 1.1 of Appleby, Flammia, McConnell, and Yard.
-Under the additional (likely true) assumption that $F^{\rm vec}\!\left(S\right)$ is closed under complex conjugation (so that $F^{\rm proj}\!\left(S\right)=F^{\rm vec}\!\left(S\right)$), their conjecture becomes equivalent to the statement that $F^{\rm proj}\!\left(S\right)=\tilde{E}_{d,1}$ for any $S$ in the minimal multiplet.
-8.3. Statement of the Ideal-Class-to-SIC Conjecture
-We now formulate a conjecture that refines 5.4, 7.1, and 8.4 by insisting on compatibility with the Artin map. To state this conjecture, we decompose
-$\operatorname{Clm}(\mathcal{O}_{\!\Delta})=\bigsqcup_{f^{\prime}|f}\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})$ (8.8)
-by means of 5.3.
-Conjecture 8.5 (Ideal-Class-to-SIC Conjecture).
-Fix a positive integer $d>3$, let $\Delta=(d+1)(d-3)$, and write $\Delta=f^{2}\Delta_{0}$ where $\Delta_{0}$ is a fundamental discriminant of a quadratic field. Then there is a bijection
-$\bigsqcup_{f^{\prime}|f}\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})\xrightarrow{\Omega}\operatorname{WHSIC}(d)/\operatorname{PEC}(d)$ (8.9)
-under which an ideal class $\mathcal{A}\in\operatorname{Cl}(\mathcal{O}^{\prime})$ with $\mathcal{O}^{\prime}=\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}}$ maps to a geometric equivalence class $\Omega(\mathcal{A})=[S_{\mathcal{A}}]$ with associated fields $F^{\rm trip}\!\left(S_{\mathcal{A}}\right)=E_{d,f^{\prime}}$ and $F^{\rm proj}\!\left(S_{\mathcal{A}}\right)=\tilde{E}_{d,f^{\prime}}$.
-Furthermore, $\Omega$ can be chosen to be compatible with the Artin map (defined in 8.2)
-$Art_{\mathcal{O}^{\prime}}:\operatorname{Cl}_{d^{\prime}\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\}}(\mathcal{O}^{\prime})\to\operatorname{Gal}(\tilde{E}_{d,f^{\prime}}/\mathbb{Q}(\Delta_{0})),$ (8.10)
-in the sense that
-$[S_{\bar{\mathcal{A}}\bar{\mathfrak{B}}}]=[(\operatorname{Art}_{\mathcal{O}^{\prime}}(\mathcal{A}))(S_{\bar{\mathfrak{B}}})]\mbox{ for any }\mathcal{A},\mathfrak{B}\in\operatorname{Cl}_{d^{\prime}\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\}}(\mathcal{O}^{\prime})$ (8.11)
-and their images $\bar{\mathcal{A}},\bar{\mathfrak{B}}\in\operatorname{Cl}(\mathcal{O}^{\prime})$ under the natural quotient map.
-The required compatibility with the Artin map implies that there is exactly one Galois super-equivalence class of Weyl–Heisenberg SICs—equivalently, one Galois multiplet—for each order $\mathcal{O}$ between $\mathcal{O}_{\!\Delta_{0}}$ and $\mathcal{O}_{\!\Delta}$, which is 7.1.
-We formulate this hypothesis because its truth would provide a unifying principle explaining the structure of all the other conjectures; we present no mechanism to construct a candidate map $\Omega$.
-8.5 would define a map $\mathcal{M}$ of the type given in 7.2 by means of the following commutative diagram,
-${\displaystyle\bigsqcup_{f^{\prime}|f}\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})}$${\operatorname{WHSIC}(d)/\operatorname{PEC}(d)}${positive divisors of ${\{\mbox{positive divisors of $f$}\}}$${\{[[S]]:S\mbox{ a line-SIC in }\mathbb{C}^{d}\}}$$\scriptstyle{\Omega}$$\scriptstyle{\mathcal{M}}$ (8.12)
-where the downward maps send $\mathcal{A}\in\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})$ to $f^{\prime}$ and $[S]$ to $[[S]]$, respectively.
-8.5 is compatible with conjectures of the first author [32] and other authors [6, 11] connecting SICs to the Stark conjectures for real quadratic fields [41], particularly as formulated by Tate [43, 44], where the Stark units carry a Galois action compatible with the Artin map.
-9. Theorems on degeneration and nondegeneration of inclusions of ray class fields of orders
-In this section, we prove unconditional results about when the fields $E_{d,f^{\prime}}$ introduced in 8.3 can coincide, while fixing $d$ and varying $f^{\prime}$.
-We prove similar results about some closely related fields.
-These results allows us to explain some observed phenomena about SICs and their associated number fields.
-The fact that two distinct multiplets $[[S_{1}]]\neq[[S_{2}]]$ can have the same associated field $F^{\rm proj}\!\left(S_{1}\right)=F^{\rm proj}\!\left(S_{2}\right)$ was discovered experimentally by Markus Grassl [27].
-This observation came as a surprise, as experimental evidence in low dimensions had pointed to multiplets being determined by the associated fields.
-We determine exactly when there is an equality of class fields $E_{d,f^{\prime}}=E_{d,f^{\prime\prime}}$ (equivalently, $\tilde{E}_{d,f^{\prime}}=\tilde{E}_{d,f^{\prime\prime}}$) in the case when $f^{\prime\prime}\mid f^{\prime}$.
-We do not rule out such an equality for $f^{\prime}$, $f^{\prime\prime}$ incomparable in the divisibility order.
-(Some such cases could be ruled out by considering ramification properties of primes in these fields as extensions of $K$.) Nonetheless, our results explain the examples discovered by Grassl.
-We characterize all cases where an equality $E_{d,f^{\prime}}=E_{d,f^{\prime\prime}}$ happens for some distinct $f^{\prime},f^{\prime\prime}$ satisfying $f^{\prime\prime}\mid f^{\prime}\mid f$ (with $(d+1)(d-3)=f^{2}\Delta_{0}$), purely in terms of the dimension $d$.
-We show that the set of such $d$ has a limiting natural density of $\frac{1}{48}$.
-8.4 then predicts that, in such dimensions $d$, there will be two distinct SIC multiplets that generate the same field.
-This assertion is experimentally testable.
-A calculation of this criterion for $d\leq 500$ is compared with Grassl’s experimental data for occurrence of equality of associated fields.
-9.1. Preliminary results on class groups and class fields of orders
-We recall results from [33] needed to relate different ray class fields to each other and determine the relative Galois group (including whether or not it is the trivial group).
-We first recall a result on when there is an inclusion between ray class fields of two orders $\mathcal{O}\subseteq\mathcal{O}^{\prime}$, which is [33, Thm.
-6.5].
-Theorem 9.1.
-For two orders $\mathcal{O}\subseteq\mathcal{O}^{\prime}$ in a number field $K$ and any level datum $(\mathcal{O};\mathfrak{m},\Sigma)$ for $\mathcal{O}$, there are inclusions of ray class fields $H_{\mathfrak{m}\mathcal{O}^{\prime},\Sigma}^{\mathcal{O}^{\prime}}\subseteq H_{\mathfrak{m},\Sigma}^{\mathcal{O}}\subseteq H_{\left(\mathfrak{m}:\mathcal{O}^{\prime}\right),\Sigma}^{\mathcal{O}^{\prime}}$.
-(Here, the extended ideal $\mathfrak{m}\mathcal{O}^{\prime}$ is the smallest $\mathcal{O}^{\prime}$-ideal containing $\mathfrak{m}$, and the quotient ideal $\left(\mathfrak{m}:\mathcal{O}^{\prime}\right)$ is the maximal $\mathcal{O}^{\prime}$-ideal contained in $\mathfrak{m}$.)
-To determine cases of equality $H_{\mathfrak{m}\mathcal{O}^{\prime},\Sigma}^{\mathcal{O}^{\prime}}=H_{\mathfrak{m},\Sigma}^{\mathcal{O}}$, we will convert the problem via the Artin isomorphism (8.2) to a calculation of isomorphisms between ray class groups in the two orders.
-We next recall an exact sequence important for the calculation of ray class groups of orders and maps between them.
-The exact sequence relates class groups of two different orders in terms of the behavior of various subgroups of local and global unit groups of the two orders.
-In order to write down the exact sequence, we first need the following definition and conventions concerning the unit groups.
-Definition 9.2.
-For a commutative ring with unity $R$ and an ideal $I$ of $R$, define the group
-$\operatorname{U}_{I}(R):=\{\alpha\in R^{\times}:\alpha\equiv 1\,\left(\operatorname{mod}\,I\right)\}=(1+I)\cap R^{\times}.$ (9.1)
-If in addition $R$ has real embeddings and $\Sigma$ is a subset of the real embeddings of $R$, define
-$\operatorname{U}_{I,\Sigma}(R):=\{\alpha\in R^{\times}:\alpha\equiv 1\,\left(\operatorname{mod}\,I\right)\mbox{ and }\rho(\alpha)>0\mbox{ for }\rho\in\Sigma\}.$ (9.2)
-We also use the following extension of this notation: If there is an map $\phi:R_{1}\to R_{2}$ implicit in the discussion, and $I$ is an ideal of $R_{1}$, then we let $\operatorname{U}_{I}(R_{2}):=\operatorname{U}_{\phi(I)R_{2}}(R_{2})$ and $\operatorname{U}_{I,\Sigma}(R_{2}):=\operatorname{U}_{\phi(I)R_{2},\Sigma}(R_{2})$.
-The follow theorem is [33, Thm.
-5.4].
-Theorem 9.3.
-Let $K$ be a number field and $\mathcal{O}\subseteq\mathcal{O}^{\prime}\subseteq\mathcal{O}_{K}$ be orders of $K$, and let $\mathfrak{f}=\mathfrak{f}_{\mathcal{O}^{\prime}}(\mathcal{O})$.
-Let $\mathfrak{m}$ be an ideal of $\mathcal{O}$, $\mathfrak{m}^{\prime}$ an ideal of $\mathcal{O}^{\prime}$ such that $\mathfrak{m}\mathcal{O}^{\prime}\subseteq\mathfrak{m}^{\prime}$, and $\Sigma^{\prime}\subseteq\Sigma\subseteq\{\mbox{embeddings }K\hookrightarrow\mathbb{R}\}$.
-Let ${\mathfrak{d}}$ be any $\mathcal{O}^{\prime}$-ideal such that ${\mathfrak{d}}\subseteq\left(\mathfrak{m}:\mathcal{O}^{\prime}\right)$.
-With the $\operatorname{U}$-groups defined as in 9.2, we have the following exact sequence.
-$1\to\frac{\operatorname{U}_{\mathfrak{m}^{\prime},\Sigma^{\prime}}\!\left(\mathcal{O}^{\prime}\right)}{\operatorname{U}_{\mathfrak{m},\Sigma}\!\left(\mathcal{O}\right)}\to\frac{\operatorname{U}_{\mathfrak{m}^{\prime}}\!\left(\mathcal{O}^{\prime}/{\mathfrak{d}}\right)}{\operatorname{U}_{\mathfrak{m}}\!\left(\mathcal{O}/{\mathfrak{d}}\right)}\times\{\pm 1\}^{\left|\Sigma\setminus\Sigma^{\prime}\right|}\to\operatorname{Cl}_{\mathfrak{m},\Sigma}(\mathcal{O})\to\operatorname{Cl}_{\mathfrak{m}^{\prime},\Sigma^{\prime}}(\mathcal{O}^{\prime})\to 1$ (9.3)
-9.2. Nondegeneracy of certain ray class fields in a fixed order $\mathcal{O}$ varying infinite places
-In this subsection, we show nondegeneracy of the inclusion of subfields of $E_{d,f^{\prime}}=H_{d\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}^{\prime}}$ obtained by dropping one or both of the infinite places from the level datum.
-It follows that $E_{d,f^{\prime}}$ is totally complex and non-CM.
-The field $H_{d\mathcal{O}^{\prime},\{\infty_{1}\}}^{\mathcal{O}^{\prime}}$ appears in [10] (with the notation $\mathfrak{R}_{1}$) as a field generated by SIC overlaps after making appropriate choices.
-Nondegeneracy will be deduced from the following lemma.
-This lemma is very special to the family of ray class groups we are considering, and it is crucial to all the calculations.
-To not interrupt the exposition, we defer the proof to Appendix A.
-Lemma 9.4.
-Let $d\geq 4$, $\Delta=(d+1)(d-3)$, and $\varepsilon_{d}=\frac{d-1+\sqrt{\Delta}}{2}$.
-Let $\Sigma$ be a subset of the real embedding of $\mathbb{Q}(\sqrt{\Delta})$.
-Fix an order $\mathcal{O}$ satisfying
-$\mathcal{O}_{\!\Delta}:=\mathbb{Z}[\varepsilon_{d}]\subseteq\mathcal{O}\subseteq\mathcal{O}_{\mathbb{Q}(\sqrt{\Delta})}.$ (9.4)
-Then,
-$\operatorname{U}_{d\mathcal{O},\Sigma}(\mathcal{O})=\langle\varepsilon_{d}^{3}\rangle.$ (9.5)
-The following result shows nondegeneracy of class groups whilst varying the infinite places. It generalizes [32, Lem. 5.3].
-Theorem 9.5.
-Let $d\geq 4$, $\Delta=(d+1)(d-3)$, and $\varepsilon_{d}=\frac{d-1+\sqrt{\Delta}}{2}$. For every order
-$\mathbb{Z}[\varepsilon_{d}]\subseteq\mathcal{O}\subseteq\mathcal{O}_{\mathbb{Q}(\sqrt{\Delta})},$ (9.6)
-all the maps in the diagram
-${\operatorname{Cl}_{d\mathcal{O},\{\infty_{1},\infty_{2}\}}(\mathcal{O})}$${\operatorname{Cl}_{d\mathcal{O},\{\infty_{1}\}}(\mathcal{O})}$${\operatorname{Cl}_{d\mathcal{O},\{\infty_{2}\}}(\mathcal{O})}$${\operatorname{Cl}_{d\mathcal{O},\emptyset}(\mathcal{O})}$ (9.7)
-are surjections with kernel $\mathbb{Z}/2\mathbb{Z}$.
-Proof.
-We apply the exact sequence of 9.3, with $\mathcal{O}^{\prime}=\mathcal{O}$, $\mathfrak{m}=\mathfrak{m}^{\prime}=d\mathcal{O}$, $\Sigma^{\prime}\subseteq\Sigma$, and ${\mathfrak{d}}=(d\mathcal{O}:\mathcal{O})=d\mathcal{O}$.
-It says that there is an exact sequence
-$1\to\frac{\operatorname{U}_{d\mathcal{O},\Sigma^{\prime}}\!\left(\mathcal{O}\right)}{\operatorname{U}_{d\mathcal{O},\Sigma}\!\left(\mathcal{O}\right)}\to\frac{\operatorname{U}_{d\mathcal{O}}\!\left(\mathcal{O}/d\mathcal{O}\right)}{\operatorname{U}_{d\mathcal{O}}\!\left(\mathcal{O}/d\mathcal{O}\right)}\times\{\pm 1\}^{\left|\Sigma\setminus\Sigma^{\prime}\right|}\to\operatorname{Cl}_{d\mathcal{O},\Sigma}(\mathcal{O})\to\operatorname{Cl}_{d\mathcal{O},\Sigma^{\prime}}(\mathcal{O})\to 1$ (9.8)
-9.4 shows for any order $\mathcal{O}\subseteq\mathcal{O}_{\mathbb{Q}(\sqrt{\Delta})}$ containing $\varepsilon_{d}$ that $\frac{\operatorname{U}_{d\mathcal{O},\Sigma^{\prime}}\!\left(\mathcal{O}\right)}{\operatorname{U}_{d\mathcal{O},\Sigma}\!\left(\mathcal{O}\right)}=\frac{\langle\varepsilon_{d}^{3}\rangle}{\langle\varepsilon_{d}^{3}\rangle}=1$.
-We deduce that
-$1\to\{\pm 1\}^{\left|\Sigma\setminus\Sigma^{\prime}\right|}\to\operatorname{Cl}_{d\mathcal{O},\Sigma}(\mathcal{O})\to\operatorname{Cl}_{d\mathcal{O},\Sigma^{\prime}}(\mathcal{O})\to 1$ (9.9)
-It follows that the class groups $\operatorname{Cl}_{(d)}(\mathcal{O})$, $\operatorname{Cl}_{(d)\infty_{1}}(\mathcal{O})$, $\operatorname{Cl}_{(d)\infty_{2}}(\mathcal{O})$, and $\operatorname{Cl}_{(d)\infty_{1}\infty_{2}}(\mathcal{O})$ are all distinct and have the relationships given in the theorem statement.
-∎
-Remark 9.6.
-By 8.2, we obtain for any order $\mathcal{O}$ satisfying $\mathbb{Z}[\varepsilon_{d}]\subseteq\mathcal{O}\subseteq\mathcal{O}_{\mathbb{Q}(\sqrt{\Delta})}$ the following inclusions of ray class fields of orders,
-[…]
-where each edge in the diagram denotes an extension of degree 2.
-We also know that $\operatorname{Gal}(H_{d\mathcal{O},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}}/H_{d\mathcal{O},\emptyset}^{\mathcal{O}})\cong\{\pm 1\}^{2}\cong(\mathbb{Z}/2\mathbb{Z})^{2}$.
-9.3. Cases of equality in ray class groups and ray class fields of two orders $\mathcal{O}\subsetneq\mathcal{O}^{\prime}$
-In certain cases, the ray class group of an order does not change the under extension of the order, and consequently, the correpsonding ray class fields are identical.
-This phenomenon can be understood using the exact sequence given in 9.3: The rightmost two groups in the exact sequence are isomorphic if and only if the leftmost two groups in the sequence are isomorphic.
-We quantify exactly when this degeneration phenomenon occurs for the level data appearing in our conjectures about SICs.
-Theorem 9.7.
-Let $d\geq 4$, $\Delta=(d+1)(d-3)$, and $\varepsilon_{d}=\frac{d-1+\sqrt{\Delta}}{2}$.
-Write $\Delta=f^{2}\Delta_{0}$, where $\Delta_{0}$ is a fundamental discriminant.
-Let $\Sigma^{\prime}\subseteq\Sigma$ be subsets of $\{\infty_{1},\infty_{2}\}$.
-For every containment of orders
-$\mathbb{Z}[\varepsilon_{d}]\subseteq\mathcal{O}\subseteq\mathcal{O}^{\prime}\subseteq\mathcal{O}_{\mathbb{Q}(\sqrt{\Delta})},$ (9.11)
-the surjective map
-$\operatorname{Cl}_{d\mathcal{O},\Sigma}(\mathcal{O})\to\operatorname{Cl}_{d\mathcal{O}^{\prime},\Sigma^{\prime}}(\mathcal{O}^{\prime})$ (9.12)
-induced by extension of ideals is an isomorphism if and only if $\Sigma=\Sigma^{\prime}$ and either:
-  • (1)
-    $\mathcal{O}=\mathcal{O}^{\prime}$, or
-  • (2)
-[…]
-[…]
-Proof.
-Let $g=[\mathcal{O}^{\prime}:\mathcal{O}]$ be the relative index of the orders and let $\Delta^{\prime}=\operatorname{disc}{\mathcal{O}^{\prime}}$, so that the two orders have the following expressions as $\mathbb{Z}$-modules:
-$\mathcal{O}^{\prime}$ $=\mathbb{Z}+\frac{\Delta^{\prime}+\sqrt{\Delta^{\prime}}}{2}$ and $\mathcal{O}$ $=\mathbb{Z}+g\frac{\Delta^{\prime}+\sqrt{\Delta^{\prime}}}{2}.$ (9.13)
-A direct computation shows that the conductor ideal $\left(\mathcal{O}:\mathcal{O}^{\prime}\right)=g\mathcal{O}^{\prime}$, so the colon ideal $\left(d\mathcal{O}:\mathcal{O}^{\prime}\right)=dg\mathcal{O}^{\prime}$.
-We apply the exact sequence of 9.3, choosing $\mathfrak{m}=d\mathcal{O}$, $\mathfrak{m}^{\prime}=d\mathcal{O}^{\prime}$, $\Sigma^{\prime}\subseteq\Sigma$, and ${\mathfrak{d}}=dg\mathcal{O}^{\prime}$, to obtain
-$1\to\frac{\operatorname{U}_{d\mathcal{O}^{\prime},\Sigma^{\prime}}(\mathcal{O}^{\prime})}{\operatorname{U}_{d\mathcal{O},\Sigma}(\mathcal{O})}\to\frac{\operatorname{U}_{d\mathcal{O}^{\prime}}\!\left(\mathcal{O}^{\prime}/dg\mathcal{O}^{\prime}\right)}{\operatorname{U}_{d\mathcal{O}}\!\left(\mathcal{O}/dg\mathcal{O}^{\prime}\right)}\times\{\pm 1\}^{\left|\Sigma\setminus\Sigma^{\prime}\right|}\to\operatorname{Cl}_{d\mathcal{O},\Sigma}(\mathcal{O})\to\operatorname{Cl}_{d\mathcal{O}^{\prime},\Sigma^{\prime}}(\mathcal{O}^{\prime})\to 1$ (9.14)
-By 9.4, $\operatorname{U}_{d\mathcal{O}^{\prime},\Sigma^{\prime}}(\mathcal{O}^{\prime})=\operatorname{U}_{d\mathcal{O},\Sigma}(\mathcal{O})=\langle\varepsilon_{d}^{3}\rangle$, so the leftmost term in the sequence is the trivial group.
-Thus, the rightmost map (which is the surjection induced by extension of ideals mentioned in the theorem statement) is an isomorphism if and only if the group $\frac{\operatorname{U}_{d\mathcal{O}^{\prime}}\!\left(\mathcal{O}^{\prime}/df\mathcal{O}^{\prime}\right)}{\operatorname{U}_{d\mathcal{O}}\!\left(\mathcal{O}/df\mathcal{O}^{\prime}\right)}\times\{\pm 1\}^{\left|\Sigma\setminus\Sigma^{\prime}\right|}=1$.
-We compute the order of this group as
-$\left|\frac{\operatorname{U}_{d\mathcal{O}^{\prime}}\!\left(\mathcal{O}^{\prime}/dg\mathcal{O}^{\prime}\right)}{\operatorname{U}_{d\mathcal{O}}\!\left(\mathcal{O}/dg\mathcal{O}^{\prime}\right)}\times\{\pm 1\}^{\left|\Sigma\setminus\Sigma^{\prime}\right|}\right|=\left|{\frac{\operatorname{U}_{d\mathcal{O}^{\prime}}\!\left(\mathcal{O}^{\prime}/dg\mathcal{O}^{\prime}\right)}{\operatorname{U}_{d\mathcal{O}}\!\left(\mathcal{O}/dg\mathcal{O}^{\prime}\right)}}\right|\cdot\left|\{\pm 1\}^{\left|\Sigma\setminus\Sigma^{\prime}\right|}\right|=\left|{\frac{\operatorname{U}_{d\mathcal{O}^{\prime}}\!\left(\mathcal{O}^{\prime}/dg\mathcal{O}^{\prime}\right)}{\operatorname{U}_{d\mathcal{O}}\!\left(\mathcal{O}/dfg\mathcal{O}^{\prime}\right)}}\right|\cdot 2^{\left|\Sigma\setminus\Sigma^{\prime}\right|}.$ (9.15)
-Thus, the map is an isomorphism if and only if $\operatorname{U}_{d\mathcal{O}^{\prime}}\!\left(\mathcal{O}^{\prime}/dg\mathcal{O}^{\prime}\right)=\operatorname{U}_{d\mathcal{O}}\!\left(\mathcal{O}/dg\mathcal{O}^{\prime}\right)$ and $\Sigma=\Sigma^{\prime}$.
-We now determine the cases when the former happens.
-We wish to compare the group $U^{\prime}:=\operatorname{U}_{d\mathcal{O}^{\prime}}\!\left(\mathcal{O}^{\prime}/dg\mathcal{O}^{\prime}\right)$ to its subgroup $U:=\operatorname{U}_{d\mathcal{O}}\!\left(\mathcal{O}/dg\mathcal{O}^{\prime}\right)$ and determine when they are equal.
-They are clearly equal when $g=1$, giving (1); from now on, we assume $g>1$.
-The structure of the proof is as follows:
-  • (A)
-    If $g$ is an odd prime, we will show that $U^{\prime}\neq U$.
-  • (B)
-    If $g=2$, we will characterize exactly when $U^{\prime}=U$, giving (2).
-  • (C)
-    If $g=4$, we will use (B) to show that $U^{\prime}\neq U$ by considering intermediate orders between $\mathcal{O}$ and $\mathcal{O}^{\prime}$.
-  • (D)
-    We will use (A) and (C) to conclude that $U^{\prime}\neq U$ for all $g>2$ by considering intermediate orders between $\mathcal{O}$ and $\mathcal{O}^{\prime}$.
-Let $\omega=\frac{\Delta^{\prime}+\sqrt{\Delta^{\prime}}}{2}$, so $\mathcal{O}^{\prime}=\mathbb{Z}+\omega\mathbb{Z}$ and $\mathcal{O}=\mathbb{Z}+g\omega\mathbb{Z}$. Let
-$M^{\prime}$ =$=\{\beta\in\mathcal{O}^{\prime}/dg\mathcal{O}^{\prime}:\beta\equiv 1\,\left(\operatorname{mod}\,d\mathcal{O}^{\prime}\right)\}\mbox{ and}$ (9.16)
-$M$ =$=\{\beta\in\mathcal{O}/dg\mathcal{O}^{\prime}:\beta\equiv 1\,\left(\operatorname{mod}\,d\mathcal{O}\right)\}.$ (9.17)
-Then $U^{\prime}$ consists of the invertible elements of $M^{\prime}$, and $U$ consists of the invertible elements of $M$, that is, the elements of each that are coprime to $dg$.
-The elements of $M^{\prime}$ and $M$ may be written down as follows, where $\overline{\alpha}$ stands for the equivalence class $\,\left(\operatorname{mod}\,dg\mathcal{O}^{\prime}\right)$ represented by $\alpha$.
-$M^{\prime}$ =$=\left\{\overline{1+dm+dn\omega}:0\leq m,n<g\right\},\mbox{ and}$ (9.18)
-$M$ =$=\left\{\overline{1+dm}:0\leq m<g\right\}.$ (9.19)
-Case (A). Suppose $g=p$ is an odd prime. For $n\in\mathbb{Z}$ we have, using the norm in $\mathcal{O}_{K}$,
-$\operatorname{Nm}(1+dn\omega)$ $=\left(1+dn\frac{\Delta^{\prime}+\sqrt{\Delta^{\prime}}}{2}\right)\left(1+dn\frac{\Delta^{\prime}-\sqrt{\Delta^{\prime}}}{2}\right)$ (9.20)
-$=1+dn\Delta^{\prime}+d^{2}n^{2}\frac{(\Delta^{\prime})^{2}-\Delta^{\prime}}{4}.$ (9.21)
-The quadratic polynomial $F(x)=1+d\Delta^{\prime}x+d^{2}\frac{(\Delta^{\prime})^{2}-\Delta^{\prime}}{4}x^{2}$ has at most two roots in $\mathbb{Z}/p\mathbb{Z}$, and $F(0)=1$ (so $x=0$ is not a root), so for $p\geq 5$ there exists some $n\in\{1,\ldots,p-1\}$ such that $\operatorname{Nm}(1+dn\omega)=F(n)\not\equiv 0\,\left(\operatorname{mod}\,p\right)$.
-[…]
-Finally $\operatorname{Nm}(1+dn\omega)\equiv 1\,\left(\operatorname{mod}\,d\right)$, so we conclude $\operatorname{Nm}(1+dn\omega)$ is a unit modulo $dg$, and thus $\overline{1+dn\omega}\in U^{\prime}\setminus U$, so $U\subsetneq U^{\prime}$.
-Case (B). Suppose $g=2$. The elements of $M^{\prime}$ and $M$ may be written down as follows.
-$M^{\prime}$ =$=\{\overline{1},\overline{1+d},\overline{1+d\omega},\overline{1+d+d\omega}\},\mbox{ and}$ (9.22)
-$M$ =$=\{\overline{1},\overline{1+d}\}.$ (9.23)
-Clearly, every element of $M^{\prime}$ and $M$ are coprime to $d$.
-If $d$ is even, then they are all coprime to $2d$, so $M^{\prime}=U^{\prime}$, $M=U$, and $U^{\prime}\neq U$.
-Now suppose that $d$ is odd; then $U=\{\overline{1}\}$.
-We compute
-$\operatorname{Nm}(1+d\omega)=\begin{cases}1\,\left(\operatorname{mod}\,2d\right)&\mbox{if }\Delta^{\prime}\equiv 0\,\left(\operatorname{mod}\,8\right),\\ 1+d\,\left(\operatorname{mod}\,2d\right)&\mbox{if }\Delta^{\prime}\equiv 1\,\left(\operatorname{mod}\,8\right),\\ 1+d\,\left(\operatorname{mod}\,2d\right)&\mbox{if }\Delta^{\prime}\equiv 4\,\left(\operatorname{mod}\,8\right),\\ 1\,\left(\operatorname{mod}\,2d\right)&\mbox{if }\Delta^{\prime}\equiv 5\,\left(\operatorname{mod}\,8\right).\end{cases}$ (9.24)
-and
-$\operatorname{Nm}(1+d+d\omega)=\begin{cases}1+d\,\left(\operatorname{mod}\,2d\right)&\mbox{if }\Delta^{\prime}\equiv 0\,\left(\operatorname{mod}\,8\right),\\ 1+d\,\left(\operatorname{mod}\,2d\right)&\mbox{if }\Delta^{\prime}\equiv 1\,\left(\operatorname{mod}\,8\right),\\ 1\,\left(\operatorname{mod}\,2d\right)&\mbox{if }\Delta^{\prime}\equiv 4\,\left(\operatorname{mod}\,8\right),\\ 1\,\left(\operatorname{mod}\,2d\right)&\mbox{if }\Delta^{\prime}\equiv 5\,\left(\operatorname{mod}\,8\right).\end{cases}$ (9.25)
-Thus,
-$U^{\prime}=\begin{cases}\{\overline{1},\overline{1+d\omega}\}&\mbox{if }\Delta^{\prime}\equiv 0\,\left(\operatorname{mod}\,8\right),\\ \{\overline{1}\}&\mbox{if }\Delta^{\prime}\equiv 1\,\left(\operatorname{mod}\,8\right),\\ \{\overline{1},\overline{1+d+d\omega}\}&\mbox{if }\Delta^{\prime}\equiv 4\,\left(\operatorname{mod}\,8\right),\\ \{\overline{1},\overline{1+d\omega},\overline{1+d+d\omega}\}&\mbox{if }\Delta^{\prime}\equiv 5\,\left(\operatorname{mod}\,8\right).\end{cases}$ (9.26)
-[…]
-Case (C). Suppose $g=4$. Consider the intermediate order $\mathcal{O}_{1}=\mathbb{Z}+2\omega\mathbb{Z}$. We have
-$[\operatorname{Cl}_{d,\Sigma}(\mathcal{O}):\operatorname{Cl}_{d,\Sigma}(\mathcal{O}^{\prime})]=[\operatorname{Cl}_{d,\Sigma}(\mathcal{O}):\operatorname{Cl}_{d,\Sigma}(\mathcal{O}_{1})][\operatorname{Cl}_{d,\Sigma}(\mathcal{O}_{1}):\operatorname{Cl}_{d,\Sigma}(\mathcal{O}^{\prime})],$ (9.27)
-but $\operatorname{disc}(\mathcal{O}_{1})$ is even, and thus $[\operatorname{Cl}_{d,\Sigma}(\mathcal{O}):\operatorname{Cl}_{d,\Sigma}(\mathcal{O}_{1})]=2$ by (B).
-Thus, $\operatorname{Cl}_{d,\Sigma}(\mathcal{O})\neq\operatorname{Cl}_{d,\Sigma}(\mathcal{O}^{\prime})$ in this case (i.e., $U^{\prime}\neq U$).
-Case (D).
-Now consider any $g>2$.
-Write $g=h_{1}h_{2}$ where either $h_{1}=4$ or $h_{1}$ is an odd prime.
-There is then an intermediate order $\mathcal{O}_{1}=\mathbb{Z}+h_{1}\omega\mathbb{Z}$ with $[\mathcal{O}^{\prime}:\mathcal{O}_{1}]=h_{1}$, and by (A) and (C), $[\operatorname{Cl}_{d,\Sigma}(\mathcal{O}_{1}):\operatorname{Cl}_{d,\Sigma}(\mathcal{O}^{\prime})]>1$.
-Thus, as in (C), $\operatorname{Cl}_{d,\Sigma}(\mathcal{O})\neq\operatorname{Cl}_{d,\Sigma}(\mathcal{O}^{\prime})$ (i.e., $U^{\prime}\neq U$).
-∎
-We obtain the following criterion for equality of the associated ray class fields.
-Theorem 9.8.
-The set of $d\geq 4$ such that, for $\Delta=(d+1)(d-3)$ and $\varepsilon_{d}=\frac{d-1+\sqrt{\Delta}}{2}$, there exist distinct orders $\mathcal{O}\subseteq\mathcal{O}^{\prime}$ with
-$\mathbb{Z}[\varepsilon_{d}]\subseteq\mathcal{O}\subseteq\mathcal{O}^{\prime}\subseteq\mathcal{O}_{\mathbb{Q}(\sqrt{\Delta})},$ (9.28)
-[…]
-Proof.
-[…]
-The Galois correspondence gives an isomorphism
-$\operatorname{Gal}\!\left(H_{d\mathcal{O},\Sigma}^{\mathcal{O}}/H_{d\mathcal{O}^{\prime},\Sigma}^{\mathcal{O}^{\prime}}\right)\cong\ker\!\left(\operatorname{Cl}_{d\mathcal{O},\Sigma}(\mathcal{O})\to\operatorname{Cl}_{d\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})\right).$ (9.29)
-Thus, $H_{d\mathcal{O},\Sigma}^{\mathcal{O}}=H_{d\mathcal{O}^{\prime},\Sigma}^{\mathcal{O}^{\prime}}$ if and only if the (always surjective) map $\operatorname{Cl}_{d\mathcal{O},\Sigma}(\mathcal{O})\to\operatorname{Cl}_{d\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})$ is an isomorphism.
-The corollary now follows.
-∎
-We now summarize the degeneration phenomenon in the context and notation of the fields $E_{d,f^{\prime}}$, $\tilde{E}_{d,f^{\prime}}$ appearing in our conjectures.
-Corollary 9.9.
-Let $d\geq 4$.
-Write $\Delta=(d+1)(d-3)=f^{2}\Delta_{0}$ for $\Delta_{0}$ a fundamental discriminant, and suppose $f^{\prime},f^{\prime\prime}$ are positive integers satisfying $f^{\prime\prime}\mid f^{\prime}\mid f$ and $f^{\prime}\neq f^{\prime\prime}$.
-The following are equivalent:
-  • (1)
-    rad(Δ)1(mod $\operatorname{rad}(\Delta)\equiv 1\,\left(\operatorname{mod}\,8\right)$ and $f^{\prime}=2f^{\prime\prime}$.
-  • (2)
-[…]
-  • (3)
-    $E_{d,f^{\prime}}=E_{d,f^{\prime\prime}}$
-  • (4)
-    $\tilde{E}_{d,f^{\prime}}=\tilde{E}_{d,f^{\prime\prime}}$
-Proof.
-The equivalence of (1), (2), and (3) follows immediately from 9.7 and the proof of 9.8, taking $\mathcal{O}=\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}}$ and $\mathcal{O}^{\prime}=\mathcal{O}_{(f^{\prime\prime})^{2}\Delta_{0}}$.
-The equivalence of (3) and (4) follows from the fact that
-$\tilde{E}_{d,f^{\prime}}^{\prime}$ $=E_{d,f^{\prime}}$ and $\tilde{E}_{d,f^{\prime\prime}}$ $=E_{d,f^{\prime\prime}}$ (9.30)
-[…]
-9.4. Rarity of degeneration for ray class fields of orders
-We determine the density of $d$ for which the criterion of 9.7 holds.
-Proposition 9.10.
-We have
-[…]
-Proof.
-[…]
-Let $\mathcal{D}=\bigcup_{k\geq 0}{\mathcal{D}}_{k}$ where
-$\mathcal{D}_{k}=\{d\geq 4:2^{2k}\mid\mid\Delta_{1}\mbox{ and }\Delta_{1}\equiv 1\,\left(\operatorname{mod}\,8\right)\}.$ (9.32)
-We show that each $\mathcal{D}_{k}$ has a natural density
-${\bf d}_{k}:=\lim_{n\to\infty}\frac{1}{N}\#\{d\in\mathcal{D}_{k}:\,d\leq N\}$ (9.33)
-by showing it is a finite union of complete arithmetic progressions. The analysis below gives
-${\bf d_{0}=\bf d_{1}=\bf d_{2}=\bf d_{3}=\bf d_{4}=\bf d_{5}}=0$ (9.34)
-and ${\bf d_{k}}=\frac{2}{2^{2k}}$ for $k\geq 6$.
-  1. (1)
-[…]
-  2. (2)
-[…]
-    1. (a)
-[…]
-    2. (b)
-[…]
-    3. (c)
-[…]
-    4. (d)
-      The pattern continues for higher $k$, yielding a density set $\frac{1}{2^{2k+3}}$ set at level $k$.
-  3. (3)
-[…]
-    1. (a)
-[…]
-    2. (b)
-      The pattern continues for higher $k$, yielding a density set $\frac{1}{2^{2k+3}}$ set at level $k$.
-[…]
-$\sum_{k=1}^{\infty}\frac{1}{2^{5+2k}}=\frac{1}{96}.$
-[…]
-Remark 9.11.
-The rarity of degeneration is special to the case of SIC fields.
-Degeneration is common in ring class towers over real quadratic fields; indeed, for any nonsquare $\Delta>0$, any prime $p$, and any fixed modulus $\mathfrak{m}$, it may be shown that the tower of fields
-$H_{\mathfrak{m}}^{\mathcal{O}_{0}}\subseteq H_{\mathfrak{m}}^{\mathcal{O}_{1}}\subseteq H_{\mathfrak{m}}^{\mathcal{O}_{2}}\subseteq\cdots\subseteq H_{\mathfrak{m}}^{\mathcal{O}_{n}}\subseteq\cdots$ (9.35)
-associated to the orders $\mathcal{O}_{n}=\mathbb{Z}[p^{n}\sqrt{\Delta}]$ is eventually constant.
-The lack of degeneration for inclusions of SIC fields seems due to the particular properties of the unit group of the base field modulo $d$.
-The tendency of SIC fields to be large, and “aligned SIC” towers incorporating SIC fields in different dimensions [5] to be noncyclotomic (and thus necessarily “horizontal,” in the sense of introducing infinitely many primes into both the degree and the modulus), holds for similar reasons.
-9.5. Experimental testing of degeneration for ray class fields of orders
-One can compare the prediction of 9.8 with experimental data on algebraic SICs, which can detect equality of SIC-fields for SICs in different Galois multiplets (which conjecturally correspond to different orders.) Only three cases of equality have been found experimentally, all by Marcus Grassl in unpublished work [27].
-Those occur in dimensions
-$d\in\{47,67,259\},$ (9.36)
-and in each case, there is a secondary multiplet having the same projection SIC-field as the minimal multiplet (conjecturally corresponding to $\mathcal{O}^{\prime}=\mathcal{O}_{K}$ or $f^{\prime}=1$).
-Algebraic SICs have not been computed in every dimension up to $d=259$, but rather, Grassl has computed them is certain dimensions (including $d=259$) where the projection field has small enough expected degree to permit the calculation.
-It is much easier to write down a list of dimensions $d$ and corresponding order conductors $f^{\prime},2f^{\prime}$ where our conjectures predict degeneration of SIC-fields, that is, where the conditions of 9.8 are satisfied.
-For $d\leq 500$, equality of ray class fields occurs for
-$d\in\{47,67,83,175,211,259,303,339,431,447,467\}.$ (9.37)
-This data is consistent with Grassl’s data for SIC-fields.
-For every such $d$, the conductor pair $(1,2)$ degenerates (i.e., gives the same field ray class field in the manner of 9.8), and an additional conductor pair degenerates in three of these dimensions.
-The data on ray class field degeneration for $d\leq 500$ is given in Table 2.
-$d$ $\Delta_{0}$ $f^{\prime},2f^{\prime}$ $h$ $[E_{d,f^{\prime}}:\mathbb{Q}(\Delta_{0})]$
-$47$ $33$ $1,2$ $1$ $1472$
-$67$ $71$,2121,21 , 2 $1$ $1452$
-$83$ $051$,2121,21 , 2 $2$ $9184$
-$175$ $473$ $1,2$ $3$ $43200$
-$211$ $689$ $1,2$ $4$ $117600$
-$259$ $65$ $1,2$ $2$ $31104$
-$303$ $57$ $1,2$ $1$ $40800$
- $d$ $\Delta_{0}$ $f^{\prime},2f^{\prime}$ $h$ $[E_{d,f^{\prime}}:\mathbb{Q}(\Delta_{0})]$ $303$ $75$,105105,105 , 10 $6$ $244800$ $339$ $7851$,2121,21 , 2 $8$ $408576$ $431$ $321$ $1,2$ $3$ $371520$ $431$ $213$,6363,63 , 6 $9$ $1114560$ $447$ $777$ $1,2$ $4$ $355200$ $467$ $377$ $1,2$ $2$ $290784$ $467$ $773$,6363,63 , 6 $8$ $1163136$
-Table 2.
-SIC dims.
-$d\leq 500$ with an equality of ray class fields of orders $E_{d,f^{\prime}}=E_{d,f^{\prime\prime}}$.
-The columns show the fund.
-disc.
-$\Delta_{0}$, each pair of conductors $(f^{\prime},f^{\prime\prime})=(f^{\prime},2f^{\prime})$ exhibiting field equality, the order class number $h=h_{(f^{\prime})^{2}\Delta_{0}}=h_{(2f^{\prime})^{2}\Delta_{0}}$, and the ray class field extension degree $[E_{d,f^{\prime}}:\mathbb{Q}(\Delta_{0})]=[E_{d,2f^{\prime}}:\mathbb{Q}(\Delta_{0})]$.
-10. Comparison of degrees of ray class fields of orders and degrees of SIC-fields
-We give formulas for the degrees of the ray class fields of orders $E_{d,f^{\prime}}$ and $\tilde{E}_{d,f^{\prime}}$.
-These degrees conjecturally coincide with the SIC fields $F^{\rm trip}\!\left(S\right)$ and $F^{\rm proj}\!\left(S\right)$ for those Weyl–Heisenberg SICs $S$ assigned to these orders in 8.4, and we make a numerical comparison in Section 11.
-Definition 10.1.
-If $K$ is a field and $n\in\mathbb{N}$, define the number field totient function by
-$\varphi_{K}(n)=\left|\left(\mathcal{O}_{K}/n\mathcal{O}_{K}\right)^{\times}\right|.$ (10.1)
-This function is a natural generalization of Euler’s totient function $\varphi(n)=\varphi_{\mathbb{Q}}(n)$.
-It is easily seen to be multiplicative by the Chinese remainder theorem.
-It may be further generalized to $\varphi_{\mathcal{O}}(\mathfrak{a}):=\left|(\mathcal{O}/\mathfrak{a})^{\times}\right|$, for any order $\mathcal{O}$ and any nonzero $\mathcal{O}$-ideal $\mathfrak{a}$, which we do not need here.
-Theorem 10.2.
-Let $d\geq 4$ and let $\Delta=(d+1)(d-3)=f^{2}\Delta_{0}$ for some fundamental discriminant $\Delta_{0}$.
-Let $K=\mathbb{Q}(\sqrt{\Delta})$.
-Let $\varepsilon_{d}=\frac{(d-1)+\sqrt{\Delta}}{2}$, and let $\mathcal{O}^{\prime}$ be a $K$-order such that $\mathbb{Z}[\varepsilon_{d}]\subseteq\mathcal{O}^{\prime}\subseteq\mathcal{O}_{K}$, that is, with $\operatorname{disc}(\mathcal{O}^{\prime})=(f^{\prime})^{2}\Delta_{0}$ for some $f^{\prime}\mid f$.
-Then
-$\left|\operatorname{Cl}_{d\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})\right|=\begin{cases}\dfrac{2^{\left|\Sigma\right|}h_{K}\operatorname{Reg}_{K}\varphi_{K}(df^{\prime})}{9\log(\varepsilon_{d})\varphi(f^{\prime})},&\mbox{if}d\equiv 3\,\left(\operatorname{mod}\,9\right)\mbox{ and }3\mid f^{\prime},\vspace{5pt}\\ \dfrac{2^{\left|\Sigma\right|}h_{K}\operatorname{Reg}_{K}\varphi_{K}(df^{\prime})}{6\log(\varepsilon_{d})\varphi(f^{\prime})},&\mbox{otherwise}.\end{cases}$ (10.2)
-If $d$ is even, then $\left|\operatorname{Cl}_{2d\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})\right|=2\left|\operatorname{Cl}_{d\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})\right|$.
-Proof.
-Let $\tilde{d}=d$ if $d$ is odd and $\tilde{d}\in\{d,2d\}$ if $d$ is even. By [33, Thm. 5.6], the cardinality of the ray class group is
-$\left|\operatorname{Cl}_{\tilde{d}\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})\right|=\frac{h_{K}}{\left[\mathcal{O}_{K}^{\times}:\operatorname{U}_{\tilde{d}\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})\right]}\cdot\frac{2^{\left|\Sigma\right|}\left|\left(\mathcal{O}_{K}\middle/\left(\tilde{d}\mathcal{O}^{\prime}:\mathcal{O}_{K}\right)\right)^{\times}\right|}{\left|\operatorname{U}_{\tilde{d}\mathcal{O}^{\prime}}\!\left(\mathcal{O}^{\prime}\middle/\left(\tilde{d}\mathcal{O}^{\prime}:\mathcal{O}_{K}\right)\right)\right|}.$ (10.3)
-By 9.4, $\operatorname{U}_{d\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})=\langle\varepsilon_{d}^{3}\rangle$. On the other hand, when $\tilde{d}=2d$,
-[…]
-[…]
-$\left[\mathcal{O}_{K}^{\times}:\operatorname{U}_{\tilde{d}\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})\right]=\begin{cases}\left[\langle-1,\varepsilon_{K}\rangle:\langle\varepsilon_{d}^{3}\rangle\right]=\dfrac{2\log(\varepsilon_{d}^{3})}{\log(\varepsilon_{K})}=\dfrac{6\log(\varepsilon_{d})}{\operatorname{Reg}_{K}},&\mbox{if }\tilde{d}=d,\vspace{5pt}\\ \left[\langle-1,\varepsilon_{K}\rangle:\langle\varepsilon_{d}^{6}\rangle\right]=\dfrac{2\log(\varepsilon_{d}^{6})}{\log(\varepsilon_{K})}=\dfrac{12\log(\varepsilon_{d})}{\operatorname{Reg}_{K}},&\mbox{if }\tilde{d}=2d.\end{cases}$ (10.5)
-The quotient ideal appearing in (10.3) is $\left(\tilde{d}\mathcal{O}^{\prime}:\mathcal{O}_{K}\right)=\tilde{d}\left(\mathcal{O}^{\prime}:\mathcal{O}_{K}\right)=\tilde{d}f^{\prime}\mathcal{O}_{K}$.
-Thus,
-$\left|\left(\mathcal{O}_{K}\middle/\left(\tilde{d}\mathcal{O}^{\prime}:\mathcal{O}_{K}\right)\right)^{\times}\right|=\left|\left(\mathcal{O}_{K}\middle/\tilde{d}f^{\prime}\mathcal{O}_{K}\right)^{\times}\right|=\varphi_{K}(\tilde{d}f^{\prime}).$ (10.6)
-Now we must compute the size of $\operatorname{U}_{\tilde{d}\mathcal{O}^{\prime}}\!\left(\mathcal{O}^{\prime}\middle/\left(\tilde{d}\mathcal{O}^{\prime}:\mathcal{O}_{K}\right)\right)=\operatorname{U}_{\tilde{d}\mathcal{O}^{\prime}}\!\left(\mathcal{O}^{\prime}\middle/\tilde{d}f^{\prime}\mathcal{O}_{K}\right)$.
-Write $\mathcal{O}_{K}=\mathbb{Z}+\omega\mathbb{Z}$ and $\mathcal{O}^{\prime}=\mathbb{Z}+f^{\prime}\omega\mathbb{Z}$.
-Thus, $\mathcal{O}^{\prime}/\tilde{d}f^{\prime}\mathcal{O}_{K}=\frac{\mathbb{Z}+f^{\prime}\omega\mathbb{Z}}{\tilde{d}f^{\prime}\mathbb{Z}+\tilde{d}f^{\prime}\omega\mathbb{Z}}$ and $\tilde{d}\mathcal{O}^{\prime}=\tilde{d}\mathbb{Z}+\tilde{d}f^{\prime}\omega\mathbb{Z}$, so the multiplicative group
-$\operatorname{U}_{\tilde{d}\mathcal{O}^{\prime}}\!\left(\mathcal{O}^{\prime}\middle/\left(\tilde{d}\mathcal{O}^{\prime}:\mathcal{O}_{K}\right)\right)$ =$=\left\{u\in\frac{\mathbb{Z}+f^{\prime}\omega\mathbb{Z}}{\tilde{d}f^{\prime}\mathbb{Z}+\tilde{d}f^{\prime}\omega\mathbb{Z}}:u\equiv 1\,\left(\operatorname{mod}\,\tilde{d}\mathbb{Z}+\tilde{d}f^{\prime}\omega\mathbb{Z}\right)\right\}$ (10.7)
-$\cong\left\{u\in\mathbb{Z}/\tilde{d}f^{\prime}\mathbb{Z}:u\equiv 1\,\left(\operatorname{mod}\,\tilde{d}\right)\right\}.$ (10.8)
-[…]
-$\left|\operatorname{U}_{\tilde{d}\mathcal{O}^{\prime}}\!\left(\mathcal{O}^{\prime}\middle/\left(\tilde{d}\mathcal{O}^{\prime}:\mathcal{O}_{K}\right)\right)\right|=\begin{cases}\frac{3}{2}\varphi(f^{\prime}),&\mbox{if }d\equiv 3\,\left(\operatorname{mod}\,9\right)\mbox{ and }3\mid f^{\prime},\\ \varphi(f^{\prime}),&\mbox{otherwise}.\end{cases}$ (10.9)
-In the case $\tilde{d}=d$, plugging in (10.5), (10.6), and (10.9) to (10.3) yields (10.2).
-In the case $\tilde{d}=2d$ (with $d$ even), we have $\varphi_{K}(\tilde{d}f^{\prime})=4\varphi_{K}(df^{\prime})$, and again using plugging in (10.5), (10.6), and (10.9) to (10.3) and comparing to (10.2), we see that $\left|\operatorname{Cl}_{2d\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})\right|=2\left|\operatorname{Cl}_{d\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})\right|$.
-∎
-To use 10.2 for computation, the following elementary lemma is helpful.
+The definitions yield $E_{d,f^{\prime}}=\tilde{E}_{d,f^{\prime}}$ when $d$ is odd, and the natural field inclusion $E_{d,f^{\prime}}\subseteq\tilde{E}_{d^{\prime},f^{\prime}}$ when $d$ is even.
+We have $[\tilde{E}_{d,f^{\prime}}:E_{d,f^{\prime}}]=2$ when $d$ is even, using Theorem 6.5(2) below.
+The larger fields $\tilde{E}_{d,f^{\prime}}$ appear in 1.3, which asserts $\tilde{E}_{d,f^{\prime}}=F^{\rm proj}\!\left(S\right)$ whenever $\mathcal{M}(\mathcal{O}^{\prime})=[[S]]$.
+1.3 is a natural extension of 1.1.
+1.1(1) is equivalent to the statement that $F^{\rm proj}\!\left(S\right)=\tilde{E}_{d,1}$ for any $S$ in the minimal multiplet.
+In Appendix A the smaller fields $E_{d,f^{\prime}}$ are conjecturally associated to $F^{\rm trip}\!\left(S\right)$ for $[[S]]$, the multiplet associated to $\mathcal{O}^{\prime}$ by the map $\mathcal{M}$.
+We wish to compute the degrees of these fields.
+The Galois correspondence of ray class field theory for orders (Theorem 6.2) implies that the degree of $\tilde{E}_{d,f^{\prime}}$ (resp., $E_{d,f^{\prime}}$) over $K=\mathbb{Q}(\sqrt{\Delta})$ equals the ray class number $\left|\operatorname{Cl}_{d^{\prime}\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\}}(\mathcal{O}^{\prime})\right|$ (resp., $\left|\operatorname{Cl}_{d\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\}}(\mathcal{O}^{\prime})\right|$).
+6.5 below gives a formula for these ray class numbers, valid for ray class groups of real quadratic orders specifically of the form $\Delta=(d+3)(d-1)$.
+To state the formula, we define for the real quadratic field $K$ the number field totient function $\varphi_{K}(n)=\left|\left(\mathcal{O}_{K}/n\mathcal{O}_{K}\right)^{\times}\right|.$ It is computed by the following formula.
 Lemma 6.4.
@@
 ∎
-Remark 10.4.
-Applying the Brauer–Siegel formula to 10.2, one obtains (for $\mathbb{Z}[\varepsilon_{d}]\subseteq\mathcal{O}^{\prime}\subseteq\mathcal{O}_{\mathbb{Q}(\sqrt{\Delta_{0}})}$, $\Delta_{0}=\operatorname{disc}(\mathbb{Q}(\varepsilon_{d}))$, $\operatorname{disc}(\mathcal{O}^{\prime})=(f^{\prime})^{2}\Delta_{0}$) the asymptotic estimate
-$\log\left|\operatorname{Cl}_{d\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})\right|\sim\frac{1}{2}\log((f^{\prime})^{2}\Delta_{0})+2\log(d)\mbox{ as }d\to\infty.$ (10.11)
-In particular,
-$d^{2+o(1)}\ll\left|\operatorname{Cl}_{d\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})\right|\ll d^{3+o(1)},$ (10.12)
-with both extremes being achieved for infinitely many pairs $(d,f^{\prime})$.
-11. Numerical evidence for the conjectures
-We present formal numerical evidence for the conjectures, based on the published tables of Scott and Grassl [39, 38], a paper of Appleby, Chien, Flammia, and Waldron [7], and Flammia’s online tables [22].
-The tables of Scott and Grassl contain putatively complete lists of $\operatorname{EC}(d)$-orbits of Weyl–Heisenberg covariant SICs as numerical approximations for $d\leq 90$ and as exact algebraic solutions for $d\leq 14$.
-The paper [7] produced putatively complete lists of $\operatorname{EC}(d)$-orbits as exact algebraic solutions for all $d\leq 21$.
-(All of these works also produced many SICs in higher dimensions, without producing a complete list.)
-  • For dimensions $d\leq 90$, we check that the size of $\operatorname{Clm}(\mathbb{Z}[\varepsilon_{d}])$ is equal to the number of known $\operatorname{PEC}(d)$-orbits (or geometric equivalence classes) of Weyl–Heisenberg covariant SICs in dimension $d$, as per 5.4.
-This data was tabulated in Table 1 in Section 5.3.
-  • For dimensions $d\leq 15$, we check that the known $\operatorname{PEC}(d)$-orbits and may be assigned to orders $\mathbb{Z}[\varepsilon_{d}]\subseteq\mathcal{O}^{\prime}\subseteq\mathcal{O}_{K}$ in such a way that $\left|\operatorname{Cl}(\mathcal{O}^{\prime})\right|$ orbits are assigned to $\mathcal{O}^{\prime}$, and those orbits form a Galois multiplet, as per 7.2.
-In addition, the degrees of the class fields $E_{d,f^{\prime}}$ and $\tilde{E}_{d,f^{\prime}}$ over $K=\mathbb{Q}(\sqrt{\Delta_{d}})$ match the degrees of $F^{\rm trip}\!\left(S\right)$ and $F^{\rm proj}\!\left(S\right)$, respectively, as per 8.4.
-The data is tabulated in Table 3 below.
-Isomorphisms of fields have not been checked except in a few small dimensions, due to computational difficulties.
-Other sources of evidence also point toward the truth of these conjectures but have not been formally tabulated here.
-The connection to Stark units supports 8.5, including the compatibility with the Artin map [32, 6, 11].
-In all cases tested (including all $d\leq 15$), the cardinality of the subgroup of $\operatorname{PEC}(d)$ fixing the line $\mathbb{C}v$ spanned by a fiducial vector $v$ (called ${\rm S}$ in [39, Table I]) equals the cardinality of the image of corresponding order unit group $(\mathcal{O}^{\prime})^{\times}$ in $(\mathcal{O}^{\prime}/d\mathcal{O}^{\prime})^{\times}$.
-Finally, work in preparation of Appleby, Flammia, and Kopp will present further evidence and refinements for the conjectures [8].
-Orbit labels in the “multiplet” column are based on Flammia’s online table [22], following [39, 38, 7], specifically using the labels for exact SICs, which differ in some cases from Scott and Grassl’s labels for numerical SICs.
-We recall the definitions of the ray class fields appearing in the table:
+We now give formulas for the cardinality of various ray class groups $\left|\operatorname{Cl}_{d^{\prime}\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})\right|$.
+We recall that for $K$ a real quadratic field, the regulator $\operatorname{Reg}_{K}$ of $K$ takes the value $\log\varepsilon$, where $\varepsilon=\varepsilon_{K}>1$ is the fundamental unit of $\mathcal{O}_{K}$.
+Theorem 6.5.
+Let $d\geq 4$ and let $\Delta=(d+1)(d-3)=f^{2}\Delta_{0}$ for some fundamental discriminant $\Delta_{0}$.
+Let $K=\mathbb{Q}(\sqrt{\Delta})$.
+Let $\varepsilon_{d}=\frac{(d-1)+\sqrt{\Delta}}{2}$, and let $\mathcal{O}^{\prime}$ be a $K$-order such that $\mathbb{Z}[\varepsilon_{d}]\subseteq\mathcal{O}^{\prime}\subseteq\mathcal{O}_{K}$, that is, with $\operatorname{disc}(\mathcal{O}^{\prime})=(f^{\prime})^{2}\Delta_{0}$ for some $f^{\prime}\mid f$.
+  • (1)
+    Then
+    $\left|\operatorname{Cl}_{d\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})\right|=\begin{cases}\dfrac{2^{\left|\Sigma\right|}h_{K}\operatorname{Reg}_{K}\varphi_{K}(df^{\prime})}{9\log(\varepsilon_{d})\varphi(f^{\prime})},&\mbox{if}d\equiv 3\,\left(\operatorname{mod}\,9\right)\mbox{ and }3\mid f^{\prime},\vspace{5pt}\\ \dfrac{2^{\left|\Sigma\right|}h_{K}\operatorname{Reg}_{K}\varphi_{K}(df^{\prime})}{6\log(\varepsilon_{d})\varphi(f^{\prime})},&\mbox{otherwise}.\end{cases}$ (6.6)
+  • (2)
+    If $d$ is even, then $\left|\operatorname{Cl}_{2d\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})\right|=2\left|\operatorname{Cl}_{d\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})\right|$.
+Proof.
+A proof is given in [36], taking $d=n+1$.
+For the reader’s convenience, we include a proof in Appendix B. The proof depends on a specific congruence for the units in such orders, stated as B.1. ∎
+6.3. Evidence for the Ray Class Fields of Orders Conjecture
+Assuming the truth of 1.3, the Galois correspondence of class field theory (Theorem 6.2) requires that the degree of $F^{\rm proj}\!\left(S\right)$ over $K=\mathbb{Q}(\sqrt{(d+1)(d-3)})$ equal the class number $\left|\operatorname{Cl}_{d\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\}}(\mathcal{O}^{\prime})\right|$.
+We verify this numerical requirement.
+Proposition 6.6 (Empirical Proposition 3).
+For each dimension $d$ with $4\leq d\leq 15$, there exists a bijective map $\mathcal{M}$ from orders to (empirical) constructed Galois multiplets.
+Taking $\Delta=(d+1)(d-3)=f^{2}\Delta_{0}$ and $K=\mathbb{Q}(\sqrt{\Delta})$, the map $\mathcal{M}$ may be chosen to have the property that, for any $f^{\prime}\mid f$ and constructed SIC $S$ such that $[[S]]=\mathcal{M}(\mathcal{O}_{((f^{\prime})^{2}\Delta_{0}})$, we have
+$[F^{\rm proj}\!\left(S\right):K]=[\tilde{E}_{d,f^{\prime}}:K].$ (6.7)
+Proof.
+The numerical equalities are tabulated for $4\leq d\leq 15$ and for $d=35$ in Table 6.1 below.
+By the Galois correspondence of class field theory (6.2) the degree of $\tilde{E}_{d,f^{\prime}}$ over $K=\mathbb{Q}(\sqrt{(d+1)(d-3)})$ equals the class number $\left|\operatorname{Cl}_{d\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\}}(\mathcal{O}^{\prime})\right|$.
+We computed the relevant class numbers using the right side of (6.6).
+The empirical calculations of field degrees $[F^{\rm proj}\!\left(S\right):K]$ is based on data for these fields given in [41, Tab.
+III] and in [6, Tab.
+A7].
+The calculations of field degrees $[F^{\rm proj}\!\left(S\right):K]$ for multiplets $[[S]]$ of exact SICs for $d=15$ and $d=35$ also appear in [6, Tab.
+A7].
+The field is $\mathbb{E}_{1}$; the degree for $\mathbb{E}$ given in the table must be divided by $4$, because the degree in [6, Tab.
+A7] is given over $\mathbb{Q}$, and because $\mathbb{E}=\mathbb{E}_{1}(i)$, an extension of degree $2$ since $\mathbb{E}_{1}$ does not contain $i$ for odd $d\geq 4$.
+For inclusions of fields for $d=35$, we made use of additional calculations [2, unpublished], which determined the inclusions between the fields $F^{\rm proj}\!\left(S\right)$ in the six multiplets $[[S]]$ of exact SICs for $d=35$, correcting the picture in [6, Fig.
+1], by showing the field associated to multiplet $35e$ is contained in the field associated to multiplet $35af$.
+The inclusions of fields $\tilde{E}_{d,f^{\prime}}$ coincide exactly with the reverse ordering of the lattice of divisibility relations of the set $\{f^{\prime}:f^{\prime}\geq 1,\,f^{\prime}\mid f\}.$ On comparison we find there exists a unique bijection $\mathcal{M}$ for all the dimensions in Table 6.1 compatible with the field inclusions.
+In Table 6.1 the orbit labels in the “multiplet” column are based on Flammia’s online table [22], following [41, 40, 6], specifically using the labels for exact SICs, which differ in some cases from Scott and Grassl’s labels for numerical SICs [41].
+The correspondence between the columns labeled $f^{\prime}$ and “multiplet” in Table 6.1 specifies the map $\mathcal{M}$.
+∎
+Example 6.7.
+We illustrate in Figure 6.1 a choice of map $\mathcal{M}$ in the case $d=35$, matching inclusions of ray class fields of orders with inclusions of SIC fields $F^{\rm vec}\!\left(S(35*)\right)$ where $*\in\{a,b,c,d,e,f,g,h,i,j\}$.
+We have $\Delta=36\cdot 32=2^{7}\cdot 3^{2}$, so $\Delta_{0}=2$ and $f=12$.
+The set of divisors $f^{\prime}$ of $12$ are $\{1,2,3,4,6,12\}$.
+The SIC fields for the multiplets for $d=35$ are extracted from [6, Tab.
+9].
+The illustrated choice of $\mathcal{M}$ satisfies 1.2(2), which determines $\mathcal{M}$ uniquely.
+This choice also satisfies 1.2(1); see 5.10.
+[…]
+[…]
+Figure 6.1.
+Left side: For $d=35$, field inclusions of ray class fields $\tilde{E}_{35,f^{\prime}}$ (mod $\,\left(\operatorname{mod}\,35\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}}\right)$ of over-orders of $\mathcal{O}_{1152}$, with edge labels giving relative field degrees.
+$\tilde{E}_{35,1}$ has degree $72$ over $\mathbb{Q}(\sqrt{\Delta_{0}})$, where $\Delta_{0}=8$.
+Right side: For $d=35$, SIC multiplets with letter labels ($35a$ to $35j$) specifying $\operatorname{PEC}(35)$-orbits, and edge labels giving relative degrees of associated fields $F^{\rm vec}\!\left(S_{35*}\right)$ for $*\in\{a,b,c,d,e,f,g,h,i,j\}$.
+$F^{\rm vec}\!\left(S_{35j}\right)$ has degree $72$ over $\mathbb{Q}(\sqrt{\Delta_{0}})$.
+Inclusions of SIC fields on the right side of Figure 6.1 for $d=35$ were given in [6, Fig.
+1].
+Our figure for $d=35$ here incorporates a later correction [2, unpublished]: the ratio SIC field for $35e$ is contained in the ratio SIC field for $35af$.
+$d$ $\mathcal{O}^{\prime}$ $f^{\prime}$ $h_{\mathcal{O}^{\prime}}$ $[\tilde{E}_{d,f^{\prime}}:K]$ multiplet $[F^{\rm proj}\!\left(S\right):K]$
+$4$ $\mathbb{Z}[\frac{1+\sqrt{5}}{2}]$ $1$ $1$ $8$ $4a$ $8$
+$5$ $\mathbb{Z}[\sqrt{3}]$ $1$ $1$ $16$ $5a$ $16$
+$6$ $\mathbb{Z}[\frac{1+\sqrt{21}}{2}]$ $1$ $1$ $24$ $6a$ $24$
+$7$ $\mathbb{Z}[\sqrt{2}]$ $1$ $1$ $12$ $7b$ $12$
+$\mathbb{Z}[2\sqrt{2}]$ $2$ $1$ $24$ $7a$ $24$
+$8$ $\mathbb{Z}[\frac{1+\sqrt{5}}{2}]$ $1$ $1$ $16$ $8b$ $16$
+$\mathbb{Z}[\frac{1+3\sqrt{5}}{2}]$ $3$ $1$ $64$ $8a$ $64$
+$9$ $\mathbb{Z}[\sqrt{15}]$ $1$ $2$ $72$ $9ab$ $72$
+$10$ $\mathbb{Z}[\frac{1+\sqrt{77}}{2}]$ $1$ $1$ $96$ $10a$ $96$
+$11$ $\mathbb{Z}[\sqrt{6}]$ $1$ $1$ $80$ $11c$ $80$
+$\mathbb{Z}[2\sqrt{6}]$ $2$ $2$ $160$ $11ab$ $160$
+$12$ $\mathbb{Z}[\frac{1+\sqrt{13}}{2}]$ $1$ $1$ $32$ $12b$ $32$
+$\mathbb{Z}[\frac{1+3\sqrt{13}}{2}]$ $3$ $1$ $96$ $12a$ $96$
+$13$ $\mathbb{Z}[\sqrt{35}]$ $1$ $2$ $192$ $13ab$ $192$
+$14$ $\mathbb{Z}[\frac{1+\sqrt{165}}{2}]$ $1$ $2$ $288$ $14ab$ $288$
+$15$ $\mathbb{Z}[\sqrt{3}]$ $1$ $1$ $48$ $15d$ $48$
+$\mathbb{Z}[2\sqrt{3}]$ $2$ $1$ $96$ $15b$ $96$
+$\mathbb{Z}[4\sqrt{3}]$ $4$ $2$ $192$ $15ac$ $192$
+$35$ $\mathbb{Z}[\sqrt{2}]$ $1$ $1$ $72$ $35j$ $72$
+$\mathbb{Z}[2\sqrt{2}]$ $2$ $1$ $144$ $35i$ $144$
+$\mathbb{Z}[3\sqrt{2}]$ $3$ $1$ $288$ $35c$ $288$
+[…]
+$\mathbb{Z}[6\sqrt{2}]$ $6$ $2$ $576$ $35af$ $576$
+$\mathbb{Z}[12\sqrt{2}]$ $12$ $4$ $1152$ $35bcdg$ $1152$
+Table 6.1.
+Table of $d$, orders $\mathcal{O}^{\prime}=\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}}$ with $\Delta=(d+1)(d-3)=f^{2}\Delta_{0}$, class numbers $h_{\mathcal{O}^{\prime}}$, degrees of fields $\tilde{E}_{d,f^{\prime}}$, and index of fields $[F^{\rm proj}\!\left(S\right):K]$, $K=\mathbb{Q}(\sqrt{\Delta})$.
+6.4. Verification of cases of the Order-to-Multiplet Conjecture
+We consider whether the bijection $\mathcal{M}$ in 1.2 can be determined from numerical data.
+We have performed a test of the conjecture on numerical invariants: We did not check inclusions $F^{\rm proj}\!\left(S_{1}\right)\subseteq F^{\rm proj}\!\left(S_{2}\right)$ directly, but we did check divisibility of the degrees of the fields $\left.[F^{\rm proj}\!\left(S_{1}\right):\mathbb{Q}]\,\middle|\,[F^{\rm proj}\!\left(S_{2}\right):\mathbb{Q}]\right.$.
+However, there would be unresolvable ambiguity in any case where there are two distinct multiplets $[[S_{1}]]$ and $[[S_{2}]]$ having
   1. (1)
-    $E_{d,f^{\prime}}:=H_{d\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}^{\prime}}$, the ray class field having level datum $(\mathcal{O}^{\prime};d\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\})$, where $(d+1)(d-3)=f^{2}\Delta_{0}$ for a fundamental discrimiant $\Delta_{0}$, $f^{\prime}$ is a positive divisor of $f$, and $\mathcal{O}^{\prime}$ is the quadratic order of discriminant $(f^{\prime})^{2}\Delta_{0}$.
+    identical sizes $\left|[[S_{1}]]\right|=\left|[[S_{2}]]\right|$, and
   2. (2)
-    $\tilde{E}_{d,f^{\prime}}:=H_{d^{\prime}\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}^{\prime}}$, where $d^{\prime}=d$ if $d$ is odd and $d^{\prime}=2d$ if $d$ is even.
-8.4 predicts that $F^{\rm trip}\!\left(S\right)=E_{d,f^{\prime}}$ and $F^{\rm proj}\!\left(S\right)=\tilde{E}_{d,f^{\prime}}$.
-$d$ $\mathcal{O}^{\prime}$ $f^{\prime}$ $h_{\mathcal{O}^{\prime}}$ $[E_{d,f^{\prime}}:K]$ $[\tilde{E}_{d,f^{\prime}}:K]$ multiplet $[F^{\rm trip}\!\left(S\right):K]$ $[F^{\rm proj}\!\left(S\right):K]$
-$4$ $\mathbb{Z}[\frac{1+\sqrt{5}}{2}]$ $1$ $1$ $4$ $8$ 4a $4$ $8$
-$5$ $\mathbb{Z}[\sqrt{3}]$ $1$ $1$ $16$ $16$ 5a $16$ $16$
-$6$ $\mathbb{Z}[\frac{1+\sqrt{21}}{2}]$ $1$ $1$ $12$ $24$ 6a $12$ $24$
-$7$ $\mathbb{Z}[\sqrt{2}]$ $1$ $1$ $12$ $12$ 7b $12$ $12$
-$\mathbb{Z}[2\sqrt{2}]$ $2$ $1$ $24$ $24$ 7a $24$ $24$
-$8$ $\mathbb{Z}[\frac{1+\sqrt{5}}{2}]$ $1$ $1$ $8$ $16$ 8b $8$ $16$
-$\mathbb{Z}[\frac{1+3\sqrt{5}}{2}]$ $3$ $1$ $32$ $64$ 8a $32$ $64$
-$9$ $\mathbb{Z}[\sqrt{15}]$ $1$ $2$ $72$ $72$ 9ab $72$ $72$
-$10$ $\mathbb{Z}[\frac{1+\sqrt{77}}{2}]$ $1$ $1$ $48$ $96$ 10a $48$ $96$
-$11$ $\mathbb{Z}[\sqrt{6}]$ $1$ $1$ $80$ $80$ 11c $\ast$ $80$
-$\mathbb{Z}[2\sqrt{6}]$ $2$ $2$ $160$ $160$ 11ab $\ast$ $160$
-$12$ $\mathbb{Z}[\frac{1+\sqrt{13}}{2}]$ $1$ $1$ $16$ $32$ 12b $\ast$ $32$
-$\mathbb{Z}[\frac{1+3\sqrt{13}}{2}]$ $3$ $1$ $48$ $96$ 12a $\ast$ $96$
-$13$ $\mathbb{Z}[\sqrt{35}]$ $1$ $2$ $192$ $192$ 13ab $\ast$ $192$
-$14$ $\mathbb{Z}[\frac{1+\sqrt{165}}{2}]$ $1$ $2$ $144$ $288$ 14ab $\ast$ $288$
-$15$ $\mathbb{Z}[\sqrt{2}]$ $1$ $1$ $48$ $48$ 15d $\ast$ $48$
-$\mathbb{Z}[2\sqrt{2}]$ $2$ $1$ $96$ $96$ 15b $\ast$ $96$
-$\mathbb{Z}[4\sqrt{2}]$ $4$ $2$ $192$ $192$ 15ac $\ast$ $192$
-Table 3.
-Table of $d$, orders $\mathcal{O}^{\prime}=\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}}$ with $\Delta=(d+1)(d-3)=f^{2}\Delta_{0}$, class numbers $h_{\mathcal{O}^{\prime}}$, degrees of fields $E_{d,f^{\prime}}$ and $\tilde{E}_{d,f^{\prime}}$, and index of fields $[F^{\rm trip}\!\left(S\right):K]$ and $[F^{\rm proj}\!\left(S\right):K]$.
-Entries denoted by $\ast$ have not been computed.
+    identical projection fields $F^{\rm proj}\!\left(S_{1}\right)=F^{\rm proj}\!\left(S_{2}\right)$.
+Such an ambiguity can occur.
+It was discovered experimentally by Markus Grassl [26, unpublished] that two distinct multiplets $[[S_{1}]]\neq[[S_{2}]]$ can have the same associated projection SIC field $F^{\rm proj}\!\left(S_{1}\right)=F^{\rm proj}\!\left(S_{2}\right)$ The identified ambiguous multiplets occurred in dimensions $d\in\{47,67,259\}$, where in each case, there is a secondary multiplet having the same projection SIC field as the minimal multiplet (conjecturally corresponding to $\mathcal{O}^{\prime}=\mathcal{O}_{K}$ or $f^{\prime}=1$).
+Algebraic SICs have currently not been computed in every dimension up to $d=259$, but rather, Grassl has computed them is certain dimensions (including $d=259$) where the projection field has small enough expected degree to permit the calculation.
+[…]
+7. Theorems on sizes of class monoids of real quadratic fields
+In this section, we state rigorous results about the behavior of $\left|\operatorname{Clm}(\mathcal{O})\right|$ for orders with discriminants $\Delta_{d}=(d+1)(d-3)$.
+These results have conditional consequences for SICs via 5.4.
+Going beyond Zauner’s conjecture, one would like to know not only that SICs exist in every dimension $d$, but also how many there are.
+The connection between SICs and class numbers gives us a precise prediction for the number of SICs up to equivalence (excluding any SICs that are not Weyl–Heisenberg covariant).
+In this section, we use known results on the growth of class numbers to derive conditional results on the number of SICs as $d\to\infty$.
+The first growth result characterizes those dimensions in which our conjectures predict a unique Weyl–Heisenberg SIC, using a theorem of Byeon, Kim, and Lee [11].
+Proposition 7.1.
+One has $\left|\operatorname{Clm}(\mathcal{O}_{\!\Delta_{d}})\right|>1$ for all $d>22$.
+The values of $d\neq 3$ with $\left|\operatorname{Clm}(\mathcal{O}_{\!\Delta_{d}})\right|=1$ are $d\in\{1,2,4,5,6,10,22\}$.
+Proof.
+Set $\Delta=\Delta_{d}=(d+1)(d-3)$ and $s(d):=\left|\operatorname{Clm}(\mathcal{O}_{\!\Delta})\right|$; we are trying to determine when $s(d)=1$. By 5.3,
+$s(d)=\sum_{f^{\prime}|f}h_{(f^{\prime})^{2}\Delta_{0}}.$ (7.1)
+where $\Delta=f^{2}\Delta_{0}$ and $\Delta_{0}$ is fundamental. If $f>1$, then $s(d)>1$.
+If $f=1$, then $\Delta=\Delta_{0}$ is fundamental.
+Write $\Delta=(d-1)^{2}-4=n^{2}-4$, where $n:=d-1$.
+Byeon, Kim, and Lee solved the class number 1 problem for real quadratic fields with fundamental discriminant of the form $n^{2}-4$, a result previously known as Mollin’s conjecture [11, Thm 1.2]; whenever $n>21$ (so $d>22$), $s(d)=h_{\Delta}>1$.
+The list of dimensions where $s(d)=1$ follows from a finite calculation of class numbers.
+∎
+Corollary 7.2.
+Assuming 5.4, there is more than one $\operatorname{PEC}(d)$-class of Weyl–Heisenberg SIC in dimension $d$ for all $d>22$.
+The dimensions $d$ with a unique $\operatorname{PEC}(d)$-class of Weyl–Heisenberg SIC are $d\in\{1,2,4,5,6,10,22\}$.
+Proof.
+The corollary follows from 7.1 combined with 5.4. ∎
+The second growth result is an asymptotic formula for the size of the ray class monoid of $\mathcal{O}_{\!\Delta_{d}}$.
+The proof uses the Brauer–Siegel theorem, which is an (ineffective) estimate of the growth of the product of the class number and the regulator for maximal orders.
+The presented result for orders must account for the class numbers of non-maximal orders.
+The conductor $f=f_{\Delta_{d}}$ is an extra degree of freedom not present in the Brauer–Siegel theorem.
+Theorem 7.3.
+As $d\to\infty$, and $\Delta_{d}=(d+1)(d-3)$, the size of the class monoid of $\mathcal{O}_{\!\Delta_{d}}$ obeys the asymptotic formula
+$\log\left|\operatorname{Clm}(\mathcal{O}_{\!\Delta_{d}})\right|=\log d+o(\log d).$ (7.2)
+Proof.
+This is proved in [36], using results from [35]. ∎
+Corollary 7.4.
+Assume 5.4. Then, as $d\to\infty$,
+$\log\left|{\rm WHSIC}(d)/\operatorname{PEC}(d)\right|=\log d+o(\log d).$ (7.3)
+Proof.
+The corollary follows from 7.3 combined with the conjecture. ∎
+8. Theorems on strictness or nonstrictness of inclusions of ray class fields of orders
+In [36] we prove unconditional results determining exactly when the fields $\tilde{E}_{d,f^{\prime}}$ introduced in 6.3 can coincide, while fixing $d$ and varying $f^{\prime}$.
+Theorem 8.1.
+The set of $d\geq 4$ such that, for $\Delta=(d+1)(d-3)$ and $\varepsilon_{d}=\frac{d-1+\sqrt{\Delta}}{2}$, there exist distinct orders $\mathcal{O}\neq\mathcal{O}^{\prime}$ with
+$\mathbb{Z}[\varepsilon_{d}]\subseteq\mathcal{O}\subseteq\mathcal{O}_{\mathbb{Q}(\sqrt{\Delta})},$ $\mathbb{Z}[\varepsilon_{d}]\subseteq\mathcal{O}^{\prime}\subseteq\mathcal{O}_{\mathbb{Q}(\sqrt{\Delta})},$ (8.1)
+having $H_{d\mathcal{O},\Sigma}^{\mathcal{O}}=H_{d\mathcal{O}^{\prime},\Sigma}^{\mathcal{O}^{\prime}}$ for some $\Sigma$ (equivalently, all $\Sigma$) is exactly the set of all $d$ such that the squarefree part of $\Delta$ (called the radical $\operatorname{rad}(d)$) has rad(Δ)1(mod $\operatorname{rad}(\Delta)\equiv 1\,\left(\operatorname{mod}\,8\right)$.
+The following theorem classifies all pairs where coincidence may occur.
+Theorem 8.2.
+Let $d\geq 4$.
+Write $\Delta=(d+1)(d-3)=f^{2}\Delta_{0}$ for $\Delta_{0}$ a fundamental discriminant, and suppose $f^{\prime},f^{\prime\prime}$ are positive integers satisfying $f^{\prime}\mid f$, $f^{\prime\prime}\mid f$ and $f^{\prime\prime}<f^{\prime}$.
+The following are equivalent:
+  • (1)
+    $\tilde{E}_{d,f^{\prime}}=\tilde{E}_{d,f^{\prime\prime}}$
+  • (2)
+    $E_{d,f^{\prime}}=E_{d,f^{\prime\prime}}$
+  • (3)
+    rad(Δ)1(mod $\operatorname{rad}(\Delta)\equiv 1\,\left(\operatorname{mod}\,8\right)$, $f^{\prime}=2f^{\prime\prime}$, and $f^{\prime\prime}$ is odd.
+  • (4)
+[…]
+We also show that the condition rad(Δ)1(mod $\operatorname{rad}(\Delta)\equiv 1\,\left(\operatorname{mod}\,8\right)$ of Theorem 8.1 is satisfied for a set of $d$ having natural density $\frac{1}{48}$.
+Proposition 8.3.
+We have
+[…]
+Proof.
+This is [36, Prop. 9.6]. ∎
+8.1. Numerical data on degeneration for ray class fields of orders
+We present data on the implications of 8.2 for $4\leq d\leq 500$ In Table 8.1, we use the conditions of 8.2 to write down all cases of equality of ray class fields of relevant orders $\tilde{E}_{d,f^{\prime\prime}}=\tilde{E}_{d,f^{\prime}}$ with $(f^{\prime\prime},f^{\prime})=(f^{\prime\prime},2f^{\prime\prime})$.
+For $d\leq 500$, equality of ray class fields $\tilde{E}_{d,f^{\prime\prime}}=\tilde{E}_{d,f^{\prime}}$ occurs for
+$d\in\{47,67,83,175,211,259,303,339,431,447,467\}.$ (8.3)
+For every such $d$, the conductor pair $(1,2)$ degenerates (i.e., gives the same field ray class field in the manner of 8.2), and an additional conductor pair degenerates in three of these dimensions.
+This data is consistent with, and accounts for, Grassl’s data for collapse of SIC field inclusions in dimensions $\{47,67,259\}$.
+$d$ $\sqrt{\Delta_{0}}$ $f^{\prime\prime},f^{\prime}$ $h$ $[\tilde{E}_{d,f^{\prime\prime}}:\mathbb{Q}(\sqrt{\Delta_{0}})]$
+$47$ $33$ $1,2$ $1$ $1472$
+$67$ $71$,2121,21 , 2 $1$ $1452$
+$83$ $051$,2121,21 , 2 $2$ $9184$
+$175$ $473$ $1,2$ $3$ $43200$
+$211$ $689$ $1,2$ $4$ $117600$
+$259$ $65$ $1,2$ $2$ $31104$
+$303$ $57$ $1,2$ $1$ $40800$
+ $d$ $\sqrt{\Delta_{0}}$ $f^{\prime\prime},f^{\prime}$ $h$ $[\tilde{E}_{d,f^{\prime\prime}}:\mathbb{Q}(\sqrt{\Delta_{0}})]$ $303$ $75$,105105,105 , 10 $6$ $244800$ $339$ $7851$,2121,21 , 2 $8$ $408576$ $431$ $321$ $1,2$ $3$ $371520$ $431$ $213$,6363,63 , 6 $9$ $1114560$ $447$ $777$ $1,2$ $4$ $355200$ $467$ $377$ $1,2$ $2$ $290784$ $467$ $773$,6363,63 , 6 $8$ $1163136$
+Table 8.1.
+Dimensions  $d\leq 500$ with an equality of ray class fields of orders $\tilde{E}_{d,f^{\prime\prime}}=\tilde{E}_{d,f^{\prime}}$.
+The columns show the fund.
+disc.
+$\Delta_{0}$, each pair of conductors $(f^{\prime\prime},f^{\prime})=(f^{\prime\prime},2f^{\prime\prime})$ exhibiting field equality, the order class number $h=h_{(f^{\prime\prime})^{2}\Delta_{0}}=h_{(f^{\prime})^{2}\Delta_{0}}$, and ray class field extension degree $[\tilde{E}_{d,f^{\prime\prime}}:\mathbb{Q}(\sqrt{\Delta_{0}})]=[\tilde{E}_{d,f^{\prime}}:\mathbb{Q}(\sqrt{\Delta_{0}})]$.
 9. Concluding remarks
+We state a conjectures refining 1.2 and 1.3. We also discuss other number fields attached to SICs that have been studied in the literature.
+9.1. The Ideal-Class-to-SICs Conjecture
+We formulate a final conjecture that refines 1.2 and 1.3 by insisting on compatibility with the Artin map.
+We formulate 9.1 because its truth would provide a unifying principle explaining the structure of all the other conjectures.
+In this paper, we present no mechanism to construct a candidate map $\Omega$ made in its statement.
+To state this conjecture, we decompose
+$\operatorname{Clm}(\mathcal{O}_{\!\Delta})=\bigsqcup_{f^{\prime}|f}\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})$ (9.1)
+by means of 5.3.
+Conjecture 9.1 (Ideal-Class-to-SIC Conjecture).
+Fix a positive integer $d>3$, let $\Delta=(d+1)(d-3)$, and write $\Delta=f^{2}\Delta_{0}$ where $\Delta_{0}$ is a fundamental discriminant of a quadratic field. Then there is a bijection $\Omega$,
+$\Omega:\bigsqcup_{f^{\prime}|f}\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})\xrightarrow{\Omega}\operatorname{WHSIC}(d)/\operatorname{PEC}(d),$ (9.2)
+under which an ideal class $\mathcal{A}\in\operatorname{Cl}(\mathcal{O}^{\prime})$ with $\mathcal{O}^{\prime}=\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}}$ maps to a geometric equivalence class $\Omega(\mathcal{A})=[S_{\mathcal{A}}]$ with associated fields $F^{\rm trip}\!\left(S_{\mathcal{A}}\right)=E_{d,f^{\prime}}$ and $F^{\rm proj}\!\left(S_{\mathcal{A}}\right)=\tilde{E}_{d,f^{\prime}}$.
+Furthermore, $\Omega$ can be chosen to be compatible with the Artin map (defined in 6.2)
+$\operatorname{Art}_{\mathcal{O}^{\prime}}:\operatorname{Cl}_{d^{\prime}\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\}}(\mathcal{O}^{\prime})\to\operatorname{Gal}(\tilde{E}_{d,f^{\prime}}/\mathbb{Q}(\Delta_{0})),$ (9.3)
+in the sense that
+$[S_{\bar{\mathcal{A}}\bar{\mathfrak{B}}}]=[(\operatorname{Art}_{\mathcal{O}^{\prime}}(\mathcal{A}))(S_{\bar{\mathfrak{B}}})]\mbox{ for any }\mathcal{A},\mathfrak{B}\in\operatorname{Cl}_{d^{\prime}\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\}}(\mathcal{O}^{\prime})$ (9.4)
+and their images $\bar{\mathcal{A}},\bar{\mathfrak{B}}\in\operatorname{Cl}(\mathcal{O}^{\prime})$ under the natural quotient map.
+The required compatibility of $\Omega$ with the Artin map implies that there is exactly one Galois multiplet for each order $\mathcal{O}$ between $\mathcal{O}_{\!\Delta_{0}}$ and $\mathcal{O}_{\!\Delta}$.
+9.1 would define a map $\mathcal{M}$ of the type given in 1.2 by means of the following commutative diagram,
+${\displaystyle\bigsqcup_{f^{\prime}|f}\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})}$${\operatorname{WHSIC}(d)/\operatorname{PEC}(d)}${positive divisors of ${\{\mbox{positive divisors of $f$}\}}$${\{[[S]]:S\mbox{ a line-SIC in }\mathbb{C}^{d}\}}$$\scriptstyle{\Omega}$$\scriptstyle{\mathcal{M}}$ (9.5)
+where the downward maps send $\mathcal{A}\in\operatorname{Cl}(\mathcal{O}_{(f^{\prime})^{2}\Delta_{0}})$ to $f^{\prime}$ and $[S]$ to $[[S]]$, respectively.
+9.1 is compatible with conjectures of the first author [33] and other authors [5, 10] connecting SICs to the Stark conjectures for real quadratic fields [43, 42], which posit the existence of special algebraic units constructed from partial zeta functions and carrying a Galois action compatible with the Artin map.
+Work in preparation by Appleby, Flammia, and Kopp [7] will make the classification scheme suggested by 9.1 explicit.
+The explicit classification also yields exact formulas for the stabilizers $\operatorname{stab}(S)$ and thus the size of the set $\operatorname{WHSIC}(d)$ discussed in 5.6.
+9.2. Refined SIC fields
 This paper considered fields that are geometric invariants of a SIC.
 It can be useful in practice to define SICs using data in smaller subfields that are not geometric invariants.
-One is the field generated by the entries of a single fiducial projector $\Pi$.
+One such subfield is the field generated by the entries of a single fiducial projector $\Pi$.
 That field can be smaller than the full projector field by nearly a factor of $d$ is some special cases, and this is extremely useful for computing exact SICs in some large dimensions that would otherwise be inaccessible to numerical investigation [5, 10].
 Another is the field generated by the SIC overlaps $\operatorname{Tr}(\Pi D_{p})$ (or else the closely related SIC overlap phases $\sqrt{d+1}\operatorname{Tr}(\Pi D_{p})$), where $D_{p}$ are the displacement operators.
-The overlap field (and the overlap phase field) appear to be ray class fields with ramification at just one of the infinite places $\infty_{1}$, provided one chooses a “strongly centered” fiducial projector; see [9, 20].
+The overlap field (and the overlap phase field) both appear to be ray class fields specified by ramification at just one of the infinite places $\infty_{1}$, provided one chooses a “strongly centered” fiducial projector; see [8, 20].
 That phenomenon relates to the connection to Stark units, which live in a ray class field with ramification at the other infinite place $\infty_{2}$ and are (at least in some cases) Galois conjugate to squares of overlap phases.
-Work in preparation by Appleby, Flammia, and Kopp [8] will make the classification scheme suggested by this paper explicit, giving a conjectural construction of each presumptive equivalence class of Weyl–Heisenberg covariant SICs in dimension $d$.
-Specifically, an explicit construction of a candidate for a map $\Omega$ satisfying the conditions of 8.5 (our strongest conjecture) will be given.
 10. Acknowledgements
@@
 We are grateful to the developers of Magma and in particular David Kohel, whose software package reduced_forms (and whose package class_number, on which reduced_forms relies) we used to compute class numbers of real quadratic orders.
-We thank Markus Grassl for finding numerical examples of degeneration of SIC-fields and sharing those examples with us.
-Professor Grassl’s insights prompted the first author to correct an originally incorrect statement and proof of 9.7.
+We thank Markus Grassl for finding numerical examples of degeneration of SIC fields and sharing those examples with us.
 We thank Marcus Appleby, Steven Flammia, and Gary McConnell for valuable conversations.
 We also acknowledge the contributions of Steve Donnelly, who independently connected SICs to orders in real quadratic number fields and shared his insights with Marcus Appleby, Steven Flammia, and Markus Grassl through informal discussions.
-Appendix A Proof of 9.4
-We will now prove 9.4, which we restate as A.1.
-Lemma A.1.
-Let $d\geq 4$, $\Delta=(d+1)(d-3)$, and $\varepsilon_{d}=\frac{1}{2}(d-1+\sqrt{\Delta})$.
-Let $\Sigma$ be a subset of the two real embeddings of $\mathbb{Q}(\sqrt{\Delta})$.
+Appendix A Conjectures for triple product fields
+We formulate conjectures about the triple product fields $F^{\rm trip}\!\left(S\right)$.
+Recall that $F^{\rm trip}\!\left(S\right)\subseteq F^{\rm proj}\!\left(S\right)$.
+We have limited empirical evidence for these conjectures.
+Recall that $E_{d,f^{\prime}}=H_{d\mathcal{O}^{\prime},\{\infty_{1},\infty_{2}\}}^{\mathcal{O}^{\prime}}.$ Note that $E_{d,f^{\prime}}=\tilde{E}_{d,f^{\prime}}$ if $d$ is odd, and $E_{d,f^{\prime}}\subseteq\tilde{E}_{d,f^{\prime}}$ if $d$ is even.
+Conjecture A.1 (Extended Ray Class Fields of Orders Conjecture).
+Fix an integer $d\geq 4$ and write $\Delta=(d+1)(d-3)=f^{2}\Delta_{0}$ where $\Delta_{0}=\operatorname{disc}(\mathbb{Q}(\sqrt{\Delta}))$.
+Then the bijection $\mathcal{M}$ of 1.3 can be chosen to satisfy the properties: If $S$ is a Weyl–Heisenberg SIC with $[S]\in\mathcal{M}(f^{\prime})$, then
+$F^{\rm trip}\!\left(S\right)$ $=E_{d,f^{\prime}}$ and $F^{\rm vec}\!\left(S\right)=F^{\rm proj}\!\left(S\right)$ $=\tilde{E}_{d,f^{\prime}}.$ (A.1)
+A.1 implies part 1 of the following conjecture, which can be tested against data.
+Conjecture A.2.
+Let $d\geq 4$. Then for a Weyl–Heisenberg SIC $S$,
+  • (1)
+    The index
+    $[F^{\rm proj}\!\left(S\right):F^{\rm trip}\!\left(S\right)]=\begin{cases}1&\mbox{ if }d\mbox{ is odd},\\ 2&\mbox{ if }d\mbox{ is even}.\end{cases}$ (A.2)
+  • (2)
+    $F^{\rm proj}\!\left(S\right)=F^{\rm trip}\!\left(S\right)(\xi_{d})$.
+We have numerically verified part (1) of this conjecture, the index computation, for $4\leq d\leq 10$.
+Recall that 3.6(2) showed unconditionally that $[F^{\rm proj}\!\left(S\right):F^{\rm trip}\!\left(S\right)]$ divides $d^{2}$.
+Appendix B Proof of 6.5
+For completeness, this appendix presents results that are to appear elsewhere [36], reformulated using the variable $n=d+1$.
+Let $K=\mathbb{Q}(\sqrt{\Delta})$. Write $\Delta=f^{2}D$, $f\geq 1$, where $D$ is squarefree, so $K=\mathbb{Q}(\sqrt{D})$ and
+$\mathcal{O}_{K}=\left\{\frac{1}{2}(a+b\sqrt{D}):(a,b)\in\mathbb{Z}^{2},\,a\equiv b\,\left(\operatorname{mod}\,2\right),\,a\equiv 0\,\left(\operatorname{mod}\,2\right)\mbox{ if }D\not\equiv 1\,\left(\operatorname{mod}\,4\right)\right\}.$ (B.1)
+For the datum $(\mathfrak{m},\Sigma)$ consisting of an ideal $\mathfrak{m}$ of the order $\mathcal{O}$ in $K$ and a set of real places $\Sigma$ of $K$, recall the definition of the unit subgroup
+$\operatorname{U}_{\mathfrak{m},\Sigma}(\mathcal{O}):=\left\{\alpha\in\mathcal{O}^{\times}:\alpha\equiv 1\,\left(\operatorname{mod}\,\mathfrak{m}\right),\,\sigma(\alpha)>0\mbox{ for }\sigma\in\Sigma\right\}.$ (B.2)
+6.5 depends on the following lemma concerning unit groups of orders, which is special to the family $\Delta=(d+1)(d-3)$ of real quadratic orders, with $d\geq 4$.
+Lemma B.1.
+Let $d\geq 4$, $\Delta=(d+1)(d-3)$, and $\varepsilon_{d}=\frac{d-1+\sqrt{\Delta}}{2}$.
+Let $\Sigma$ be a subset of the real embedding of $\mathbb{Q}(\sqrt{\Delta})$.
 Fix an order $\mathcal{O}$ satisfying
-$\mathcal{O}_{\!\Delta}:=\mathbb{Z}[\varepsilon_{d}]\subseteq\mathcal{O}\subseteq\mathcal{O}_{\mathbb{Q}(\sqrt{\Delta})}.$ (A.1)
+$\mathcal{O}_{\!\Delta}:=\mathbb{Z}[\varepsilon_{d}]\subseteq\mathcal{O}\subseteq\mathcal{O}_{\mathbb{Q}(\sqrt{\Delta})}.$ (B.3)
 Then,
-$\operatorname{U}_{d\mathcal{O},\Sigma}(\mathcal{O})=\langle\varepsilon_{d}^{3}\rangle.$ (A.2)
+$\operatorname{U}_{d\mathcal{O},\Sigma}(\mathcal{O})=\langle\varepsilon_{d}^{3}\rangle.$ (B.4)
+For B.1, the unit group of interest is $\operatorname{U}_{d\mathcal{O},\Sigma}(\mathcal{O})$. The unit $\varepsilon_{d}$ is totally positive, and
+$\varepsilon_{d}^{3}=\frac{1}{2}\left((d-1)(d^{2}-2d-2)+d(d-2)f\sqrt{D}\right)\equiv 1\,\left(\operatorname{mod}\,d\mathcal{O}_{\!\Delta}\right).$ (B.5)
+Thus, we have the containment relations
+$\langle\varepsilon_{d}^{3}\rangle\subseteq\operatorname{U}_{d\mathcal{O}_{\!\Delta},\{\infty_{1},\infty_{2}\}}\subseteq\operatorname{U}_{d\mathcal{O},\Sigma}(\mathcal{O})\subseteq\operatorname{U}_{d\mathcal{O}_{K}}(\mathcal{O}_{K}).$ (B.6)
+These relations show it suffices to prove (B.4) for the special case $\mathcal{O}=\mathcal{O}_{K}$ with $\Sigma=\emptyset$.
+The following stronger assertion in the special case, about the minimal unit $\eta\in\mathcal{O}_{K}$ such that $\eta>1$ and $\eta=\pm 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$, immediately implies Lemma B.1.
+Lemma B.2.
+Let $d\geq 4$, $\Delta=(d+1)(d-3)$ and $\varepsilon_{d}=\frac{1}{2}(d-1+\sqrt{\Delta})$. Let $K=\mathbb{Q}(\sqrt{\Delta}).$ Then the smallest unit $\eta>1$ in $\mathcal{O}_{K}$ having
+$\eta=\pm 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right).$ (B.7)
+is $\eta=\varepsilon_{d}^{3}$.
 Proof.
-Let $K=\mathbb{Q}(\sqrt{\Delta})$. Write $\Delta=f^{2}D$, $f\geq 1$, where $D$ is squarefree, so $K=\mathbb{Q}(\sqrt{D})$ and
-$\mathcal{O}_{K}=\left\{\frac{1}{2}(a+b\sqrt{D}):(a,b)\in\mathbb{Z}^{2},\,a\equiv b\,\left(\operatorname{mod}\,2\right),\,a\equiv 0\,\left(\operatorname{mod}\,2\right)\mbox{ if }D\not\equiv 1\,\left(\operatorname{mod}\,4\right)\right\}.$ (A.3)
-For the datum $(\mathfrak{m},\Sigma)$ consisting of an ideal $\mathfrak{m}$ of the order $\mathcal{O}$ in $K$ and a set of real places $\Sigma$ of $K$, recall the definition of the unit subgroup
-$\operatorname{U}_{\mathfrak{m},\Sigma}(\mathcal{O}):=\left\{\alpha\in\mathcal{O}^{\times}:\alpha\equiv 1\,\left(\operatorname{mod}\,\mathfrak{m}\right),\,\sigma(\alpha)>0\mbox{ for }\sigma\in\Sigma\right\}.$ (A.4)
-For A.1, the unit group of interest is $\operatorname{U}_{d\mathcal{O},\Sigma}(\mathcal{O})$. The unit $\varepsilon_{d}$ is totally positive, and
-$\varepsilon_{d}^{3}=\frac{1}{2}\left((d-1)(d^{2}-2d-2)+d(d-2)f\sqrt{D}\right)\equiv 1\,\left(\operatorname{mod}\,d\mathcal{O}_{\!\Delta}\right).$ (A.5)
-Thus, we have the containment relations
-$\langle\varepsilon_{d}^{3}\rangle\subseteq\operatorname{U}_{d\mathcal{O}_{\!\Delta},\{\infty_{1},\infty_{2}\}}\subseteq\operatorname{U}_{d\mathcal{O},\Sigma}(\mathcal{O})\subseteq\operatorname{U}_{d\mathcal{O}_{K}}(\mathcal{O}_{K}).$ (A.6)
-These relations show it suffices to prove (A.2) for the special case $\mathcal{O}=\mathcal{O}_{K}$ with $\Sigma=\emptyset$.
-The proof studies the minimal unit $\eta>1\in\mathcal{O}_{K}$ such that
-$\eta=\pm 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right).$ (A.7)
+Let $\Delta_{0}$ be the discriminant of the maximal order $\mathcal{O}_{K}$.
+Note that $\varepsilon_{d}>1$ is a unit of $\mathcal{O}_{K}$ of positive norm and is a root of $x^{2}-(d-1)x+1=0$.
+If $\varepsilon_{d}^{\prime}>0$ is its Galois conjugate, then $\varepsilon_{d}^{\prime}=\frac{1}{\varepsilon_{d}}$ and $\operatorname{Tr}(\varepsilon_{d})=\varepsilon_{d}+\varepsilon_{d}^{\prime}=d-1$, and $\varepsilon_{d}<d-1$.
 […]
-Claim 1. We have $\operatorname{Nm}(\eta)=\eta\eta^{\prime}=1$. Consequently $\eta$ is totally positive, and $10<\eta^{\prime}<$ 1.
+Claim 1. We have $\operatorname{Nm}(\eta)=\eta\eta^{\prime}=1$. Then $\eta^{\prime}=\frac{1}{\eta}$ and $10<\eta^{\prime}<$ 1.
 Proof.
-We know $\eta\eta^{\prime}=\pm 1$.
-By hypothesis, $\eta=\pm 1+d\omega$ for some $\omega\in\mathcal{O}_{K}$, and taking algebraic conjugates yields $\eta^{\prime}=\pm 1+d\omega^{\prime}$.
-Then $\eta\eta^{\prime}=1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$, which forces $\eta\eta^{\prime}=1$.
-Total positivity and the bound for $\eta^{\prime}$ follow, proving the claim.
+By hypothesis, $\eta=\pm 1+d\omega$ for some $\omega\in\mathcal{O}_{K}$, and taking Galois conjugates yields $\eta^{\prime}=\pm 1+d\omega^{\prime}$.
+It follows that $\eta\eta^{\prime}=1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$.
+Now $\operatorname{Nm}(\eta)=\eta\eta^{\prime}=\pm 1$ and $1\not\equiv-1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$ since $d\geq 4$, so we conclude $\eta\eta^{\prime}=1$.
+Thus $\eta^{\prime}>0$ and $10<\eta^{\prime}<$ 1, proving the claim.
 Claim 2. For some integer $k\geq 1$, $\eta^{k}=\varepsilon_{d}^{3}.$
 Proof.
-We have $1<\eta\leq\varepsilon_{d}^{3}$, so for some $k\geq 1$, there holds $\eta^{k-1}<\varepsilon_{d}^{3}\leq\eta^{k}$.
-Then $\omega:=\eta^{k}\varepsilon_{d}^{-3}$ is a unit, $1\leq\omega<\eta$, and $\omega\equiv\pm 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$; hence by minimality of $\eta>1$, we must have $\omega=1$, proving the claim.
-Claim 3. $\eta^{k}=\varepsilon_{d}^{3}$ for $k\geq 3$ cannot occur.
+We have $1<\eta\leq\varepsilon_{d}^{3}$, hence there holds $\eta^{k-1}<\varepsilon_{d}^{3}\leq\eta^{k}$, for some $k\geq 1.$ Then $\omega:=\eta^{k}\varepsilon_{d}^{-3}$ is a unit, $1\leq\omega<\eta$, and $\omega\equiv\pm 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$.
+However, by minimality of $\eta>1$, we must have $\omega=1$, proving the claim.
+Claim 3. $\eta^{k}=\varepsilon_{d}^{3}$ for some $k\geq 3$ cannot occur.
 Proof.
-If $k=3$, then $\eta^{3}=\varepsilon_{d}^{3}$, which, since $\eta>1$ forces $\eta=\varepsilon_{d}$.
-We know $\eta\equiv\pm 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$ and obtain a contradiction by showing that $\varepsilon_{d}\not\equiv\pm 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$.
-To show the latter, we argue by contradiction and suppose $\varepsilon_{d}\equiv\pm 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$.
-We know $\varepsilon_{d}\varepsilon_{d}^{\prime}=\eta\eta^{\prime}=1$, which then holds $\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$ and thus implies $\varepsilon_{d}^{\prime}\equiv\pm 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$; hence, $\varepsilon_{d}+\varepsilon_{d}^{\prime}\equiv\pm 2\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$.
-However, $\varepsilon_{d}+\varepsilon_{d}^{\prime}=\frac{1}{2}(d-1+\sqrt{\Delta})+\frac{1}{2}(d-1+\sqrt{\Delta})=d-1\equiv-1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$, giving a contradiction for $d\geq 4$.
-Thus $\varepsilon_{d}\not\equiv\pm 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$ holds, and $\eta\neq\varepsilon_{d}$.
-If $k\geq 4$, then $\eta=(\varepsilon_{d})^{3/k}<\varepsilon_{d}$ (taking the positive $k$-th root.) Now Claim 1 gives (since $x+\frac{1}{x}$ is an increasing function for $x\geq 1$) that
-$11<\eta+\eta^{\prime}=\operatorname{Tr}(\eta)<\varepsilon_{d}+\varepsilon_{d}^{\prime}=\operatorname{Tr}(\varepsilon_{d})=d-$ 1 . (A.8)
-The trace is an integer, and
-$\operatorname{Tr}(\eta)$ $\equiv\eta+\eta^{\prime}\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)\equiv\pm 2\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right).$ (A.9)
-Since $d\geq 4$ and $\operatorname{Tr}(\eta)\in\mathbb{Z}$, the bounds on $\operatorname{Tr}(\eta)$ give
-$\operatorname{Tr}(\eta)=\begin{cases}2&\mbox{if }\eta\equiv 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right),\\ d-2&\mbox{if }\eta\equiv-1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right).\end{cases}$ (A.10)
+If $k=3$, then we have $\eta^{3}=\varepsilon_{d}^{3}$, and, since $\eta>1$, we have $\eta=\varepsilon_{d}$.
+By hypothesis $\eta\equiv\pm 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$, so we deduce $\varepsilon_{d}\equiv\pm 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$.
+We rule out this possibility, arguing by contradiction.
+So suppose $\varepsilon_{d}\equiv\pm 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$.
+We know $\varepsilon_{d}\varepsilon_{d}^{\prime}=\eta\eta^{\prime}=1$, which viewed $\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$ implies $\varepsilon_{d}^{\prime}\equiv\varepsilon_{d}\equiv\pm 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right).$ We conclude that $\varepsilon_{d}+\varepsilon_{d}^{\prime}\equiv\pm 2\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$.
+On the other hand
+$\varepsilon_{d}+\varepsilon_{d}^{\prime}=\frac{1}{2}(d-1+\sqrt{\Delta})+\frac{1}{2}(d-1+\sqrt{\Delta})=d-1\equiv-1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right).$ (B.8)
+Since ±$\pm 2\not\equiv-1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$ for any $d\geq 4$, we have a contradiction, whence it always holds that $\varepsilon_{d}\not\equiv\pm 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)$.
+If $k\geq 4$, then $\eta=(\varepsilon_{d})^{3/k}<\varepsilon_{d}$ (taking the positive $k$-th root.) We study the trace $\operatorname{Tr}(\eta)=\eta+\eta^{\prime}$ and note, by Claim 1, $\operatorname{Tr}(\eta)=\eta+\frac{1}{\eta}.$
+Now, since $x+\frac{1}{x}$ is an increasing function for $x\geq 1$, and $1<\eta<e_{d}$, we have
+$11<\eta+\eta^{\prime}=\operatorname{Tr}(\eta)<\varepsilon_{d}+\varepsilon_{d}^{\prime}=\operatorname{Tr}(\varepsilon_{d})=d-$ 1 . (B.9)
+Now $\operatorname{Tr}(\eta)\in\mathbb{Z}$, and
+$\operatorname{Tr}(\eta)$ $\equiv\eta+\eta^{\prime}\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right)\equiv\pm 2\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right).$ (B.10)
+Since $d\geq 4$, the bounds (B.9) on $\operatorname{Tr}(\eta)$ give
+$\operatorname{Tr}(\eta)=\begin{cases}2&\mbox{if }\eta\equiv 1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right),\\ d-2&\mbox{if }\eta\equiv-1\,\left(\operatorname{mod}\,d\mathcal{O}_{K}\right).\end{cases}$ (B.11)
 We have two cases.
@@
 Then $\eta+\eta^{\prime}=2$ and $\eta\eta^{\prime}=1$, so $\eta$ is a root of the polynomial $x^{2}-2x+1=(x-1)^{2}$, so $\eta=1$, contradicting $\eta>1$.
-  • Case 2: Suppose $\operatorname{Tr}(\eta)=d-2$. We may assume $d\geq 5$, since the case $d=4$ is covered by Case 1. Now
-    $d-3<\eta=\operatorname{Tr}(\eta)-\eta^{\prime}<\varepsilon_{d}<\operatorname{Tr}(\varepsilon_{d})=d-1.$ (A.11)
-    We have for $d\geq 6$ that
-    $(d-3)^{3}<\eta^{3}<\varepsilon_{d}^{3}<(d-1)^{3}<(d-3)^{5}<\eta^{5},$ (A.12)
-    since $(d-1)^{3}<(d-3)^{5}$ for $d\geq 6$.
-Thus $\eta^{k}=\varepsilon_{d}^{3}$ cannot hold.
-For $d=5$, we have $\eta+\eta^{\prime}=3$ and $\eta\eta^{\prime}=1$, so $\eta$ is a root of the polynomial $x^{3}-3x+1$, so $\eta=\frac{3+\sqrt{5}}{2}$.
-However, for $d=5$, $\Delta=12$, so $\varepsilon_{5}\in\mathbb{Q}(\sqrt{3})$ cannot be an integer power of $\eta$.
-This completes the proof of Claim 3. It remains to rule out $k=2$, where $\varepsilon_{d}^{3}$ is the square of a unit $\tilde{\eta}$. We first classify all $d$ where such a unit exists.
+  • Case 2: Suppose $\operatorname{Tr}(\eta)=d-2$.
+In this case $\eta$ satisfies $x^{2}-(d-2)x+1=0$, hence $\eta=d-2-\eta^{\prime}>d-3$.
+For $d\geq 7$, noting that $(d-1)^{3}<(d-3)^{4}$ holds, and recalling $\varepsilon_{d}<d-1$, we have
+    $(d-3)^{3}<\eta^{3}<\varepsilon_{d}^{3}<(d-1)^{3}<(d-3)^{4}<\eta^{4}\leq\eta^{k},$ (B.12)
+    so there are no solutions $\eta^{k}=\varepsilon_{d}^{3}$ for $k\geq 4$ in this case.
+    For the remaining values $d=4,5,6$, case $d=4$ is already ruled out in Case 1; $d=5$ has $\eta=\frac{3+\sqrt{5}}{2}$, while $\varepsilon_{5}=2+\sqrt{3}$; $d=6$ has $\eta=2+\sqrt{3}$ while $\varepsilon_{6}=\frac{5+\sqrt{21}}{2}$, so $\eta^{k}=\varepsilon_{d}^{3}$ has no solution since $\eta$ and $\varepsilon_{d}$ are in different number fields.
+This completes the proof of Claim 3.
+It remains to rule out $k=2$, where $\varepsilon_{d}^{3}$ is the square of a unit $\tilde{\eta}$. We first classify all $d$ where such a unit exists.
 Claim 4. The set of all $d\geq 4$ having a unit $\tilde{\eta}\in\mathcal{O}_{K}$ that satisfies $\tilde{\eta}^{2}=\varepsilon_{d}^{3}$ is:
@@
     $d=a^{2}-1$ for any $a\geq 3$. Then, $\tilde{\eta}=\pm\eta\varepsilon_{d}$, where
-    $\eta=\varepsilon_{a+1}=\frac{a+\sqrt{a^{2}-4}}{2}$ (A.13)
+    $\eta=\varepsilon_{a+1}=\frac{a+\sqrt{a^{2}-4}}{2}$ (B.13)
     has $\operatorname{Nm}(\eta)=+1$. Here $\sqrt{\Delta}=a\sqrt{d-3}=a\sqrt{a^{2}-4}$.
@@
 […]
-$\eta^{2}=\frac{1}{4}\left((a^{2}+b^{2}D)+2ab\sqrt{D}\right),$ (A.14)
-while $\varepsilon_{d}=\frac{1}{2}(d-1+f\sqrt{D})$. To obtain a solution, we must have
-$ab$ $=f;$ (A.15)
-$a^{2}+b^{2}D$ $=2(d-1).$ (A.16)
-We have $a\neq 0$ since $f\neq 0$, so the first equation gives $b^{2}=\frac{f^{2}}{a^{2}}$, and the second equation can be rewritten (since $\Delta=Df^{2}$) as
-$a^{2}+\frac{\Delta}{a^{2}}=2(d-1).$ (A.17)
+$\eta^{2}=\frac{1}{4}\left((a^{2}+b^{2}\Delta_{0})+2ab\sqrt{\Delta_{0}}\right),$ (B.14)
+while $\varepsilon_{d}=\frac{1}{2}(d-1+f\sqrt{\Delta_{0}})$. To obtain a solution, we must have
+$ab$ $=f;$ (B.15)
+$a^{2}+b^{2}\Delta_{0}$ $=2(d-1).$ (B.16)
+We have $a\neq 0$ since $f\neq 0$, so the first equation gives $b^{2}=\frac{f^{2}}{a^{2}}$, and the second equation can be rewritten (since $\Delta=\Delta_{0}f^{2}$) as
+$a^{2}+\frac{\Delta}{a^{2}}=2(d-1).$ (B.17)
 Introducing the new variable $y=a^{2}$, we find
-$y^{2}-2(d-1)y+(d+1)(d-3)=(y-(d+1))(y-(d-3))=0,$ (A.18)
+$y^{2}-2(d-1)y+(d+1)(d-3)=(y-(d+1))(y-(d-3))=0,$ (B.18)
 which yields two cases. The first is $y=a^{2}=d+1$, which gives solutions $d=a^{2}-1$ for $a\geq 3$, to make $d\geq 4$. The second is $y=a^{2}-3$, which gives solutions $d=a^{2}+3$ for $a\geq 1$.
-  • Case 1: Suppose $d=a^{2}+1$ (so $a\geq 3$). Then $\Delta=a^{2}(a^{2}-4)$, $\sqrt{\Delta}=a\sqrt{a^{2}-4}$, and $a^{2}-4=Db^{2}$. Now
-    $\eta$ $=\varepsilon_{a+1}=\frac{1}{2}(a+\sqrt{a^{2}-4});$ (A.19)
-    $\varepsilon_{d}$ $=\eta^{2}=\frac{1}{2}(a^{2}-2+a\sqrt{a^{2}-4}).$ (A.20)
+  • Case 1: Suppose $d=a^{2}+1$ (so $a\geq 3$). Then $\Delta=a^{2}(a^{2}-4)$, $\sqrt{\Delta}=a\sqrt{a^{2}-4}$, and $a^{2}-4=\Delta_{0}b^{2}$. Now
+    $\eta$ $=\varepsilon_{a+1}=\frac{1}{2}(a+\sqrt{a^{2}-4});$ (B.19)
+    $\varepsilon_{d}$ $=\eta^{2}=\frac{1}{2}(a^{2}-2+a\sqrt{a^{2}-4}).$ (B.20)
     Here $\operatorname{Nm}(\eta)=+1$, and the unit $\tilde{\eta}=\eta\varepsilon_{d}>1$ has $\operatorname{Nm}(\tilde{\eta})=1$.
   • Case 2: Suppose $d=a^{2}-3$, ($a\geq 1$). Then $\Delta=a^{2}(a^{2}+4)$ and $\sqrt{\Delta}=a\sqrt{a^{2}+4}$. So
-    $\eta$ $=\frac{1}{2}(a+\sqrt{a^{2}+4});$ (A.21)
-    $\varepsilon_{d}$ $=\eta^{2}=\frac{1}{2}(a^{2}+2+a\sqrt{a^{2}+4}).$ (A.22)
+    $\eta$ $=\frac{1}{2}(a+\sqrt{a^{2}+4});$ (B.21)
+    $\varepsilon_{d}$ $=\eta^{2}=\frac{1}{2}(a^{2}+2+a\sqrt{a^{2}+4}).$ (B.22)
     Here $\operatorname{Nm}(\eta)=-1$, and the unit $\tilde{\eta}=\eta\varepsilon_{d}>1$ has $\operatorname{Nm}(\tilde{\eta})=-1$.
-This proves Claim 4. To complete the proof, we must rule out the case $k=2$.
+This proves Claim 4.
+To complete the proof, we rule out the final case $k=2$.
 Claim 5. $\eta^{k}=\varepsilon_{d}^{3}$ for $k=2$ does not occur.
-We must show that all the $\tilde{\eta}$ described by Claim 4 have
-$\tilde{\eta}\not\equiv\pm 1\,\left(\operatorname{mod}\,(a^{2}-1)\mathcal{O}_{K}\right).$ (A.23)
+Proof. We must show that all the units $\tilde{\eta}$ described by Claim 4 have
+$\tilde{\eta}\not\equiv\pm 1\,\left(\operatorname{mod}\,(a^{2}-1)\mathcal{O}_{K}\right).$ (B.23)
 The Claim 4 solutions for $d=a^{2}+1$ are ruled out by Claim 1, since $\operatorname{Nm}(\tilde{\eta})=-1$.
@@
 We have
-$\tilde{\eta}=\frac{1}{2}\left(a(a^{2}-3)+(a^{2}-1)\sqrt{a^{2}-4}\right).$ (A.24)
-Now $\Delta=(d+1)(d-3)=a^{2}(a^{2}-4)$; hence, $\Delta=f^{2}D$ has $D\mid(a^{2}-4)$ and $K=\mathbb{Q}(\sqrt{a^{2}-4})$. We assert for $d=a^{2}-1$ that
-$\tilde{\eta}\equiv-a\,\left(\operatorname{mod}\,(a^{2}-1)\mathcal{O}_{K}\right).$ (A.25)
+$\tilde{\eta}=\frac{1}{2}\left(a(a^{2}-3)+(a^{2}-1)\sqrt{a^{2}-4}\right).$ (B.24)
+Now $\Delta=(d+1)(d-3)=a^{2}(a^{2}-4)$; hence, $\Delta=f^{2}\Delta_{0}$ has $\Delta_{0}\mid(a^{2}-4)$ and $K=\mathbb{Q}(\sqrt{a^{2}-4})$. We assert for $d=a^{2}-1$ that
+$\tilde{\eta}\equiv-a\,\left(\operatorname{mod}\,(a^{2}-1)\mathcal{O}_{K}\right).$ (B.25)
 The assertion implies $a\pm 1\not\in d\mathcal{O}_{K}$ holds for $d=a^{2}-1$, with $d\geq 3$, which implies Claim 5. It is proved in two cases.
 […]
-    $\tilde{\eta}$ $=\frac{1}{2}(a(a^{2}-3)+(a^{2}-1)\sqrt{a^{2}-4}))$ (A.26)
-    $\equiv\frac{1}{2}a(a^{2}-3)-\frac{1}{2}(a^{2}-1)$ $\,\left(\operatorname{mod}\,(a^{2}-1)\mathcal{O}_{K}\right)$ (A.27)
-    $\equiv\frac{a-1}{2}(a^{2}-1)-a$ $\,\left(\operatorname{mod}\,(a^{2}-1)\mathcal{O}_{K}\right)$ (A.28)
-    $\equiv-a$ $\,\left(\operatorname{mod}\,(a^{2}-1)\mathcal{O}_{K}\right).$ (A.29)
+    $\tilde{\eta}$ $=\frac{1}{2}(a(a^{2}-3)+(a^{2}-1)\sqrt{a^{2}-4}))$ (B.26)
+    $\equiv\frac{1}{2}a(a^{2}-3)-\frac{1}{2}(a^{2}-1)$ $\,\left(\operatorname{mod}\,(a^{2}-1)\mathcal{O}_{K}\right)$ (B.27)
+    $\equiv\frac{a-1}{2}(a^{2}-1)-a$ $\,\left(\operatorname{mod}\,(a^{2}-1)\mathcal{O}_{K}\right)$ (B.28)
+    $\equiv-a$ $\,\left(\operatorname{mod}\,(a^{2}-1)\mathcal{O}_{K}\right).$ (B.29)
   • Case 2. Suppose $a$ is even. Let $a=2a^{\prime}$. Then, $\Delta=a^{2}\Delta_{0}$ with $\Delta_{0}=a^{2}-4=4((a^{\prime})^{2}-1)$, so $K=\mathbb{Q}(\sqrt{(a^{\prime})^{2}-1})$. Now
-    $\tilde{\eta}$ $=\frac{a}{2}(a^{2}-3)+(a^{2}-1)\sqrt{(a^{\prime})^{2}-1}$ (A.30)
-    $\equiv\frac{a}{2}(a^{2}-1)-a$ $\,\left(\operatorname{mod}\,(a^{2}-1)\mathcal{O}_{K}\right)$ (A.31)
-    $\equiv-a$ $\,\left(\operatorname{mod}\,(a^{2}-1)\mathcal{O}_{K}\right).$ (A.32)
-We have proved the assertion, hence Claim 5, which completes the proof. ∎
+    $\tilde{\eta}$ $=\frac{a}{2}(a^{2}-3)+(a^{2}-1)\sqrt{(a^{\prime})^{2}-1}$ (B.30)
+    $\equiv\frac{a}{2}(a^{2}-1)-a$ $\,\left(\operatorname{mod}\,(a^{2}-1)\mathcal{O}_{K}\right)$ (B.31)
+    $\equiv-a$ $\,\left(\operatorname{mod}\,(a^{2}-1)\mathcal{O}_{K}\right).$ (B.32)
+We have proved the assertion, hence Claim 5 is proved. Thus Lemma B.2 is proved. ∎
+Now we prove 6.5.
+Proof of 6.5.
+Let $\tilde{d}=d$ if $d$ is odd and $\tilde{d}\in\{d,2d\}$ if $d$ is even. By [34, Thm. 5.6], the cardinality of the ray class group is
+$\left|\operatorname{Cl}_{\tilde{d}\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})\right|=\frac{h_{K}}{\left[\mathcal{O}_{K}^{\times}:\operatorname{U}_{\tilde{d}\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})\right]}\cdot\frac{2^{\left|\Sigma\right|}\left|\left(\mathcal{O}_{K}\middle/\left(\tilde{d}\mathcal{O}^{\prime}:\mathcal{O}_{K}\right)\right)^{\times}\right|}{\left|\operatorname{U}_{\tilde{d}\mathcal{O}^{\prime}}\!\left(\mathcal{O}^{\prime}\middle/\left(\tilde{d}\mathcal{O}^{\prime}:\mathcal{O}_{K}\right)\right)\right|}.$ (B.33)
+By B.1, $\operatorname{U}_{d\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})=\langle\varepsilon_{d}^{3}\rangle$. On the other hand, when $\tilde{d}=2d$,
+[…]
+[…]
+$\left[\mathcal{O}_{K}^{\times}:\operatorname{U}_{\tilde{d}\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})\right]=\begin{cases}\left[\langle-1,\varepsilon_{K}\rangle:\langle\varepsilon_{d}^{3}\rangle\right]=\dfrac{2\log(\varepsilon_{d}^{3})}{\log(\varepsilon_{K})}=\dfrac{6\log(\varepsilon_{d})}{\operatorname{Reg}_{K}},&\mbox{if }\tilde{d}=d,\vspace{5pt}\\ \left[\langle-1,\varepsilon_{K}\rangle:\langle\varepsilon_{d}^{6}\rangle\right]=\dfrac{2\log(\varepsilon_{d}^{6})}{\log(\varepsilon_{K})}=\dfrac{12\log(\varepsilon_{d})}{\operatorname{Reg}_{K}},&\mbox{if }\tilde{d}=2d.\end{cases}$ (B.35)
+The quotient ideal appearing in (B.33) is $\left(\tilde{d}\mathcal{O}^{\prime}:\mathcal{O}_{K}\right)=\tilde{d}\left(\mathcal{O}^{\prime}:\mathcal{O}_{K}\right)=\tilde{d}f^{\prime}\mathcal{O}_{K}$.
+Thus,
+$\left|\left(\mathcal{O}_{K}\middle/\left(\tilde{d}\mathcal{O}^{\prime}:\mathcal{O}_{K}\right)\right)^{\times}\right|=\left|\left(\mathcal{O}_{K}\middle/\tilde{d}f^{\prime}\mathcal{O}_{K}\right)^{\times}\right|=\varphi_{K}(\tilde{d}f^{\prime}).$ (B.36)
+Now we must compute the size of $\operatorname{U}_{\tilde{d}\mathcal{O}^{\prime}}\!\left(\mathcal{O}^{\prime}\middle/\left(\tilde{d}\mathcal{O}^{\prime}:\mathcal{O}_{K}\right)\right)=\operatorname{U}_{\tilde{d}\mathcal{O}^{\prime}}\!\left(\mathcal{O}^{\prime}\middle/\tilde{d}f^{\prime}\mathcal{O}_{K}\right)$.
+Write $\mathcal{O}_{K}=\mathbb{Z}+\omega\mathbb{Z}$ and $\mathcal{O}^{\prime}=\mathbb{Z}+f^{\prime}\omega\mathbb{Z}$.
+Thus, $\mathcal{O}^{\prime}/\tilde{d}f^{\prime}\mathcal{O}_{K}=\frac{\mathbb{Z}+f^{\prime}\omega\mathbb{Z}}{\tilde{d}f^{\prime}\mathbb{Z}+\tilde{d}f^{\prime}\omega\mathbb{Z}}$ and $\tilde{d}\mathcal{O}^{\prime}=\tilde{d}\mathbb{Z}+\tilde{d}f^{\prime}\omega\mathbb{Z}$, so the multiplicative group
+$\operatorname{U}_{\tilde{d}\mathcal{O}^{\prime}}\!\left(\mathcal{O}^{\prime}\middle/\left(\tilde{d}\mathcal{O}^{\prime}:\mathcal{O}_{K}\right)\right)$ =$=\left\{u\in\frac{\mathbb{Z}+f^{\prime}\omega\mathbb{Z}}{\tilde{d}f^{\prime}\mathbb{Z}+\tilde{d}f^{\prime}\omega\mathbb{Z}}:u\equiv 1\,\left(\operatorname{mod}\,\tilde{d}\mathbb{Z}+\tilde{d}f^{\prime}\omega\mathbb{Z}\right)\right\}$ (B.37)
+$\cong\left\{u\in\mathbb{Z}/\tilde{d}f^{\prime}\mathbb{Z}:u\equiv 1\,\left(\operatorname{mod}\,\tilde{d}\right)\right\}.$ (B.38)
+[…]
+$\left|\operatorname{U}_{\tilde{d}\mathcal{O}^{\prime}}\!\left(\mathcal{O}^{\prime}\middle/\left(\tilde{d}\mathcal{O}^{\prime}:\mathcal{O}_{K}\right)\right)\right|=\begin{cases}\frac{3}{2}\varphi(f^{\prime}),&\mbox{if }d\equiv 3\,\left(\operatorname{mod}\,9\right)\mbox{ and }3\mid f^{\prime},\\ \varphi(f^{\prime}),&\mbox{otherwise}.\end{cases}$ (B.39)
+In the case $\tilde{d}=d$, plugging in (B.35), (B.36), and (B.39) to (B.33) yields (6.6).
+In the case $\tilde{d}=2d$ (with $d$ even), we have $\varphi_{K}(\tilde{d}f^{\prime})=4\varphi_{K}(df^{\prime})$, and again using plugging in (B.35), (B.36), and (B.39) to (B.33) and comparing to (6.6), we see that $\left|\operatorname{Cl}_{2d\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})\right|=2\left|\operatorname{Cl}_{d\mathcal{O}^{\prime},\Sigma}(\mathcal{O}^{\prime})\right|$.
+∎